A Complete Bibliography of Publications in Physics in Perspective (PIP)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

15 December 2023 Version 1.71

Title word cross-reference

$105.00 [Ryd13]. $11.95 [RS10d]. $110.00 [Gre09]. $124.50 [Hec07]. $126.00 [Seg10a]. $13.95 [Tay05]. $134.50/£ [Gre06]. $14.95 [Hol01, Sim07, Wes10]. $15.95 [Bed09a]. $16.00 [Hil07, RS06d]. $16.50 [Col00]. $16.95 [Hob02, PP08, RS06d]. $17.95 [Kah03a, Noe09]. $17.99 [But19]. $18.00 [Ger99a, PCGD02]. $18.95 [RS10d, RS11e]. $19.95 [Ehr04, Hun01]. $20.00 [Ham11]. $20.95 [Oha08b]. $20.99 [RS10e]. $21.00 [Gol99a, Roe10]. $21.95 [Sil07, Sim12a]. $22.00 [Can01, Dom05]. $22.50 [Car07, Pri99, Rio00a]. $22.95 [Bed09c, For01, Kle11, vB02a]. $23.95 [But00]. $24.95 [Bed01a, Bed08, Bri07, Com05, Eme11, Fel11, Fra01b, Gre07a, Hal09a, Hob02, Jac04, New04, Pas12a, Pes11, Ram05, RS07a, RS07b, RS07c, RS08c, RS08d, RS09b, RS09d, RS12h, RS12d, Seg11a]. $25 [Gea99]. $25.00 [Bro04, Dah06a, Fel08, Ger06a, Goo12, Hie99, MG21, RS09c, Tay10]. $25.95 [KP02, RS07a, vB11]. $26.00 [Bro06a, For02, Fri08, RS09d, Yeh05]. $26.95 [Mil10]. $26.99 [RS11e]. $27.00 [Bed09b, For08, Lus10, Tri08].

$27.50 [PCGD02]. $27.95 [DeW99, Dor06, Ger06a, Gri06, vB01, vB10a]. $27.98 [Lan10]. $27.99 [Fra12a]. $28.00 [Fra04, Lar05c, Sha02]. $28.95 [Cre12]. $28.99 [Pes15]. $29.00 [Har05, Hob04, Pip99b]. $29.95 [Blo06, Bor11, Cus00, Das18, Fra09a, Goo09a, Gre07b, Hol09, Kah03a, Mer01, Mod18, Nal20, Oha08a, PCGD02, Pas04, RS10c, Roe12, Ros01, Ste03b, vB06, vB08b, Bee15]. $29.99 [Hol19, Rob18, Seg15]. $30 [RS11b]. $30.00 [Bru03, Coh11, Eve00, Fra07, Ger06b, Gib05, Jam00, Kle07, Lep07, PCGD02, Per01, Sch00a, Sch05]. $32.00 [Car11, RS07b]. $32.50 [vB99]. $34.50 [Gre05, Rig07, Sti08, Wal07]. $34.95 [Bed09a, Fra10, Fri11, Goo09b, Gou01, RS09c, Tei02]. $35.00 [Ahe02, Aur04, Bed06, Bri99, Eve08a, For09, Fri09, Hol07, Kob02, Lus04, Mas07, Neu99, Par00, Par06, Ryc10, Shi06, Sob08, Sty05, Tor02, Wat03, Wei04, vB08a, vB10b]. $35.95 [Ger99a, Jon07]. $37.50 [Eve04, Nyc09, Wei04]. $38.00 [Fre00, Phi13, Rio00a]. $39.95 [Bla00b, Fra01a]. $40 [RS10e]. $40.00 [Cas10, RS08d, Ros08]. $42.00 [Eve08b]. $45.00 [Che11, Lin10, New99, Pri99, RS10b, Sch01, Seg09, Seg13]. $49.50 [Ber08b, Hil11, Ryn99]. $49.95 [Alf09, Bla00a, Ger03, Stu06]. $50 [Smi20]. $50.00 [DeV00, Hub06]. $55.00 [Bru02, Dom05, Hie99, Noe11, Tri07]. $59.00 [Pas10]. $59.50 [Fre04, LM06]. $59.95 [Jac03, PP10]. $60 [Gea99]. $60.00 [Bru99a, For01, Tat13, Wat01, vB10c]. $61 [Kra16]. $65.00 [Oha10, PCGD02, vB12]. $69.95 [Mar03b]. $70.00 [Eme09, Goo08, Per10]. $74.50 [Gre05]. $75 [vB02b]. $75.00 [Fre05a, Pip99b, Van03]. $77.50 [Hun01]. $79.95 [vB00]. $80.00 [Cas11, Kro10]. $84.99 [For20]. $85.00 [Bru06, Har05, Tri09]. $90.00 [Sha06]. $95.00 [Han04, Sim04]. $99.99 [Ano18a]. α_{1−} [VMR19]. β_{2} [AR19]. g − 2 [Cra00]. N [Fri99].

-6 [FEVvdB19]. -glycoprotein [AR19].

. [Nal20].

0 [RS13d]. 0-19-982807-5 [RS13d]. 0-231-50959-6 [Hug09].

13q [FEVvdB19]. 1703 [Wei04]. 1773 [Sch01]. 1848 [Per10]. 1910 [Smi20]. 1913 [RA13]. 1919 [CdL16]. 1926 [Car11]. 1939 [Van03]. 1944 [Sch21]. 1945 [Cas05a, HHH03]. 1949 [Ber08b]. 1960s [Gau18, Wei17]. 1964 [Lem07]. 1972 [Rio00a]. 1980s [Bad09b, Wes18, Cas10]. 1997 [Rig11]. 19th [Sta06a, Hen05b].

2 [Cas12, Wes18]. 2001 [Ber02b]. 2002 [RS02d]. 2007 [Hug09]. 2009 [Seg10b]. 2016 [Hof17]. 20th [Hug09, vB10c, Ren08]. 21st [RS07c]. 24 [MDK19]. 24-Parganas [MDK19]. 29 [Ber02b].

3 [Wes19]. 36 [Pro16]. 362 [Tri08]. 362-364 [Tri08].

46 [FEVvdB19].

5 [Nye08, RS13d].

6 [Tri08]. 6q [FEVvdB19].

978 [Cas12]. 978-0-670-02276-2 [Cas12].

A. [Sta07]. Aarau [Hun15]. Aaserud [Kra16]. Abdus [Alf09, Fra08b]. Abel [New04, Pes03]. aberrations [CFL19]. aborted [KKG^{+}19]. Abraham [Kle07]. Absorbing [Joh06a]. Absorption [DMA07, Fra02b]. Academic [RS09a, GR08a]. Accelerating [Goo08, Kir02, Pas04, Tri09, Kra07]. Accelerator [HC11, Kap08, Wie17b]. Accepted [Bru99b]. accommodation [Hof05]. Account [LoS16, Wes17, Wil17, Sta10]. Achievements [RA15]. acides [PMAS19]. Acoustics [Ban14, Fix15]. Act [PA21]. activity [GR08a, RS09a]. AdA [Ber04a]. Adam [Liu20, Sal20]. adaptation [Fla98, New99]. Addison [Ger99a, RS06b]. Addison-Wesley [Ger99a]. Addison-Wesley/Perseus [Ger99a]. adjustment [HW08, Sim08]. Administration [RS07a]. Ado [Eme11, Fis10]. Adrienne [Seg09]. Adult [Cas12]. adulthood [Gav01]. Adventure [Hu19a, Hu19b, Jac10, Wes15]. Adventures [Bri99, FL06, Pes11, RS06b, Tis09a, Tis09b, Tow99, Tri05, FL88, Hug10, FLH85]. Adversaries [For02, Wei01]. advice [FGL^{+}12]. Advising [Her00, vB02a]. Aether [Mul01]. Affair [Seg13, Tur12]. Affordable [Ber16]. After [Ste23]. Aftermath [Ber08b, Hen07b]. Afterward [Tay10]. Again [Coo11, Eis06, Oha08a]. Against [Gol16, OO20]. Age [Aur04, Cas11, Cha02b, Ehr04, Lin10, RS10e, Car10, Coe07, Hol08a, Hol08b, Hol08c, Mag09, Pan03, Bri07, Seg10b]. Aged [DFAB19]. Ages [Goo12, Osl10]. agrestis [PS19]. Aid [Bed05]. Aiken [Gou01, Coh00]. Akademgorodok [KA14, KA15]. Akademie [Gre09]. Aksant [Sim10a]. Al [Col00, Sil07]. Al-Khalili [Col00]. Alamos [Bed01b, Sei00]. Alan [Goo09b, Hol01, RS09a]. Albert [Ham11, RD05, Die05, Goe03, Hun10, Hun15, Kle10, Ren05b, Ren05a, RD05, RS06a, van12]. Alberto [Gri06, Oha10, Seg11a]. Alchemy [Ram05, Mor03]. Alenush [TT12]. Alex [Hof17]. Alexander [vB10b, Roe10, Rup05, Wal09, Wal11]. Alexandre [dP07]. Alferov [Bor11, Jos10]. Alfred [Bed06, Fol05, JP99]. Algebra [Gau18]. Algebraical [Ano00a]. Alien [Pas12a, Jay11]. Alisa [Kro10]. Allan [Jac03, Eve04]. allele [VMR19]. Allen [Hol19]. alliance [Ber02a]. along [Gab00]. Alpher [PCGD02, Alp12]. Alvin [Goo03b, Goo03c]. Always [Fra16]. Amateur [Har01, Man22]. Amateurs [Ano22]. Amazonia [CdL16]. Ambition [Bed09c, Jon08]. amble [SS08]. America [RS11a, Tis09b, Wal09, Wal11, Wei08, vB10b, von09]. American [Ger06a, Bad05, Bed06, BS05, Cas05b, Gor09, Gre21, Ham00, Hu16, Moo08, Pas12b]. Amherst [Fra04, Hob04, Lan10, Lar05c, PP08]. amino [PMAS19]. amino-acides [PMAS19]. among [SB19]. Ampère [Fre05a, Hof17, Ste05, Dar00, Ste05, SL16]. Amsterdam [Gre09, Sim10a]. Analog [Joh06a]. Anastopoulos [For09]. Ancient [Bar03, Gib05]. Anderson [MG21, Phi13, ZZ18]. Andrei [Eve06]. Andrew [Bru02, Har05, MG21, Sch05]. Andriesse [Sha06]. Angeles [Bed01a, Neu99]. Angeles/London [Neu99]. Anger [Alf09, Fra08b]. Animal [Man23]. Ann [Ber08b, vB12]. Anna [Cas05a]. Annihilation [OR21]. Anniversary [Bon18]. Anomalous [LoS16]. Anpassung [HW08, Sim08]. Answers [RS04d, LR03]. Anthony [DeV00]. anthropic [Lar07]. Anti [RS12a]. Anti-science [RS12a]. Antimatter [QN08, Fra09a]. Antiproton [OR21]. antiquiora [Sal17]. Antiquity [Bru99a]. Antonino [LPR04]. Antony [Seg10a]. Apart [Lep07]. Apparatus [Gre21]. Apple [Bal08, RS09d]. application [Bad09c]. Appointment [Bad09a]. Appreciation [RS12g, Bed05, Day04, Gol99b, Har02b, Har01, Hol03, JS02, Kah03b, Mil07, NW08, Rei01a, TF07]. Approach [Bro05, RS06c]. April [RS02d]. Archibald [DeW99]. Archimedean [Sal17]. Archimedes [Hir10, RS09a]. Archipelago [Wei17]. Arena [Bad01]. Argument [Gra11]. Ariadne [Tri08]. Arianrhod [RS06d]. Aristotelian [Sal17]. Aristotle [Noe09, Cro07]. Arkhangelsk [SE14]. Arms [RS06d, McM05]. Arnold [Car11, Set10]. Art [Dah09, Eve00, Fle21a, Fle21b, RS12d, Bar99a, HW11, Lei08, SR06, RS08c]. Arthur [PCGD02, Fri08, Wie17a, vB10a]. Artificial [Buc20, GLR06, GR09, GLR12, GLR20, AGR04]. Arts [RS13e, Pet11, Cre12]. Arturo [Mer01, Oha08a]. Aspects [RS09a, GR08a]. Astray [But19]. Astrid [Bed09a]. Astrology [Ano01a, Liu20]. Astronomers [Sal20]. Astronomy [Kos13, Liu20, RS07b, Smi20, Cla09, Sol06]. Astrophysicist [Tri05]. Astrophysics [Wie18]. Atmospheric [LoS16]. Atom [Bal08, Far09b, Kra11a, Kra16, Lal19, RS09d, vB99, Pul98, RS10c]. Atomic [Bed08, Bør12, Can00, Cas11, Cas17, For01, Her00, Jen11, Neu99, Nye99, Pal00, Rob18, vB02a, Car10, Kel07, Ros98, Bed01a]. Atoms [Sal19, Cer98, Bla00a]. attempts [Nav06]. Audretsch [Jon07]. Aufrecht [Lem07]. August [Rig11]. Ausstellung [RD05]. Authors [RS06a, Ren05b, Ren05a]. Autonomie [HW08, Sim08]. autonomy [Hof05, HW08, Sim08]. Autoren [Ren05b, Ren05a]. Axel [vB99]. Axiomatics [Mil99].

B [Hil07, Mar03b, GORS09, Sta02]. b&w [Ano18a]. B-VIII [GORS09]. Bacciagaluppi [Seg10a]. Back [Coo11]. Background [Alp12]. Bacon [Cre17]. Bad [Van07]. Badash [Cas10]. Baggott [Gre05]. Bais [Oha08b]. Bakeout [Cre09a]. Balancing [PA21]. Balíbar [RS09d]. Baltimore [Fel08, Goo12, Gre07b, Ham11, Oha10, RS08c, Tat13]. Band [Hu19a, Hu19b]. banding [Seh19]. Bands [Har09, Seh19]. Bang [Cas12, RS09h, AH01, PCGD02, Seg11b]. bangs [Sol06]. Bantu [VMR19]. Barbara [Bad09a, Sim07]. Barcelona [RRR13]. Barr [Gib05]. Barri [Coh11]. barrier [NHW08, Bed09a]. Barrow [Eve00]. Based [RS12g]. Basel [PP10, vB00, Mar03b]. Basic [Ano05a, Bro06a, But19, Hil07, PCGD02, RS10c, Seg10b, Seg15, Wei08, vB11]. basics [RS06f]. basis [Loe03]. Battle [Sal20]. Be [CMS20a, RS09h, RS13h]. Beam [Hec10, Blo06]. Beaten [Bro06a, Fey05]. Beauty [But19, Joh99a, New99, PCGD02, Per01, RS10e, Fla98, Goo02, Hol08a, Hol08b, Hol08c, Mil01]. Became [Cha18, Dah06a, Mor05]. Beck [dP07]. Becquerel [Kip00]. Bedfellows [NMS22]. Before [Ste23]. began [Cla09, RS07b]. Begin [RS06e]. Beginner [Wes10, Joh09]. Beginnings [CMS21b]. begins [RS01d, WWH01]. Behind [Bru03, Sha06, AM05, Fri01, Goo09b]. Beijing [Hu19a, Hu19b]. Bekenstein [Bel14]. Belief [Jac00, Tei02]. Beliefs [RS10b, Dav10]. Believing [Fra17]. Belknap [vB08b]. Bell [Ber08c, JS02, Liu20, Sal20]. Beller [Par00]. Bellevue [Bed09b, Lus10]. Bengal [MDK19]. Bennett [Wei04]. Beranek [Hal09a]. Bergeron [Jac04]. Bergmann [Hal05]. Berkeley [Bed01a, Bla00b, Neu99, Dah06b, Wes17]. Berkeley/Los [Neu99]. Berlin [Mar03b, Stu06, Hof99, Hof00, van12]. Bern [Hen05a]. Bernard [Dom05, Gou01, vB99]. Bernardini [Stu06]. Bernstein [For08, Ger06a, RS10d]. Bertoloni [Gre07b]. Beta [Fra02b, GR09]. Bethe [Fra01b, RS08b, BL06, Goo99, Jac09, Sch00b, Sch14]. Betrayal [Fri08, Mil05]. better [Ruh20]. Bettyann [Ger99a]. Between [Bar99a, Hof05, Man23, PA21, RA15, Wes17, Fix15, Haa03, HW08, Seh19, Sim08, dWtH20, Eve00]. Beyond [Bag04, Cas09, HW11, Kro10, Pas12a, Sok08, Bok08, Jay11, Mil07, Bed09b, Gre05, Lus10, RS12d]. Biagioli [Mas07]. Bieri [Pas10]. Big [Cas12, PCGD02, Sol06, AH01, Giu12, Seg11b, Wes14]. Biggest [OM18]. Bild [D’A04]. biographical [Kur02]. Biographie [RS03e, Stu04, Tri08]. Biography [Cre99, CMS19a, Ger03, Ger06b, Rio00a, Sch14, Tri08, Bri05, Lar05a, MM00, RS03e, SRD06, PP08]. Biology [JFEH15]. Bird [Bed06]. Birkhäuser [Mar03b, PP10, vB00]. Birth [Bed08, RS06d, She13, Tay10, vB00, Ber02b, Bru02, CFL19, Kel07, MW99, McM05, Mel09]. BIS [Ste03a, Cas05a, HHH03]. BIS-Verlag [Ste03a]. Black [AK99, AK12, Bed08, DeW99, Fri08, ORSSBO^{+}17, Tay10, Loe03, Mel09, Mil05, WF98, Col00, Sei00]. black-body [Loe03]. Blackett [Fre04, Hor03, Nye99]. Blau [RS03e, Stu04, Tri08, RS03e, Sim13, SRD06]. Blocks [Noe11, Coo10]. Blueprint [RS09b, Hoo08]. Blunder [OM18]. body [Loe03]. Böhlau [Stu04, Tri08]. Bohm [For20]. Bohr [Kra16, Par06, RS07b, Sch11, Sch21, Ano04a, Hal12, Hol05, JFEH15, Kle10, Kra11a]. Bokulich [Kro10]. Bologna [Stu06, DS13]. Boltzmann [Bla00a, Bad09c, Boy12, Cer98, Dah09, Rei07]. Bomb [Bed08, Bla00b, Fra01b, Her00, Neu99, vB02a, Cas09, Kel07, Per99, Ros98, RS07i, Sch00b, Bed09b, For01, Hol09, Lus10]. Bombardment [Hon03]. Bombs [Can00, Gor09]. Bone [Ger99a, Kev97]. Bonebrake [SS10]. bones [Hen07a]. Bonn [Pip02]. Bonolis [Stu06]. Book [Ahe02, AHSB09, Alf09, Ano11a, Ano11b, Ano12a, Ano12b, Ano13a, Ano13b, Ano13c, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15b, Ano16d, Ano16a, Ano16b, Ano16c, Ano17, Ano18a, Ano18b, Ano18c, Aur04, Bad21, Bed01a, Bed06, Bed08, BHF^{+}09, Bed09a, Bed09b, Bed09c, Bee15, BEF^{+}08, Ber08b, Bla00a, Bla00b, Blo06, Bor11, Bri99, BHS07, Bri07, Bro04, Bro06a, Bru99a, Bru02, Bru03, Bru06, But00, But19, CvB01, Can01, Car07, Car11, CGYC05, Cas05a, Cas10, Cas11, Cas12, CFH^{+}11, Che11, Coh11, Col00, Com05, Cre12, Cus00, Cus02, DBBG06, Dah06a, Das18, Das21, Day02, DeV00, DeW99, DSFG05, Dom05, Dor06, Eck23, Ehr04, Eme09, Eme11, Eve00, Eve04, EvBF^{+}08, Eve08a, Eve08b, Fel08, Fel11, For01, For02, For08, For09, For20, FFF^{+}01, Fra01a]. Book [Fra04, Fra07, Fra09a, Fra10, Fra12a, Fra01b, Fre00, Fre04, Fre05a, Fri08, Fri09, Fri11, Gea99, Ger03, Ger06a, Ger99a, GBS^{+}06, Ger06b, Gib05, Gol99a, GEBF09, Goo09a, Goo12, Goo02, GWL03, Goo08, Goo09b, Gou01, Gre05, Gre07a, Gre06, Gre07b, Gre09, Gri06, Hal09a, Ham11, Han04, HTR05, Har05, Hec07, Hie99, Hil07, Hil11, Hob02, Hob04, Hol01, Hol09, Hol07, Hol19, Hub06, Hug09, HMHB01, Hun01, Jac03, Jac04, Jam00, Jon07, KBS03, Kah03a, KP02, KMG^{+}07, Kle07, KHC11, Kle11, KSAF02, Kob02, Kra16, KPP^{+}10, Kro10, Lal19, Lan10, Lar05b, Lar05c, Lep07, Lin10, LM06, LEJ04, Lus04, Lyo20, Mar03b, Mas07, McC22, Mer01, Mil10, Mod18, MG21, Nal20, Neu99, New99, New04, NNT^{+}09, Noe09, Noe11, Nyc09, OTRF08, Oha08a, Oha08b]. Book [Oha10, Par00, Par06, PCGD02, PFW04, Pas04, PP08, PFC^{+}10, PP10, Pas10, Pas12a, PORL10, Per10, Per01, Pes11, Pes15, PSTR13, Phi13, Pip99b, Pri99, Ram05, RS06a, RS06b, RS06c, RS06d, RS07a, RS07b, RS07c, RS07d, Rig07, RS08a, RS08b, RS08c, RS08d, RS09a, RS09b, RS09c, RS09d, RS10b, RS10c, RS10d, RS10e, RS11b, RS11c, RS11d, RS11e, RS12h, RS12d, RS12b, RS12c, RS13c, RS13d, RS13a, RS13b, Rio00a, Rob18, Roe10, RPF12, Roe12, Ros01, Ros08, Ryc10, Ryd13, Ryn99, Sch00a, Sch01, Sch23, Sch05, Seg09, Seg10a, Seg10b, Seg11a, Seg13, Seg15, Sha06, Sha02, SGvB06, Shi06, Sil07, SEA^{+}04, Sim04, SHL^{+}07, Sim07, Sim08, Sim10a, SvBC^{+}12, Sim12a, Smi20, Sob08, Sta06a, SGJ03, Ste03a, Ste03b, SvBG08, Sti08, SHH04, Stu04, SBLM^{+}06, Stu06]. Book [Sty05, Tat13, Tay05, Tay10, TKP^{+}02, Tei02, Tor02, Tri07, Tri08, Tri09, Van03, WRJH07, Wal07, Wat01, WMV03, Wat03, Wei04, Wes10, Yeh05, vB99, vB00, vB01, vBBH02, vB02a, vB02b, vB06, vB08a, vB08b, vBSTM10, vB10a, vB10b, vB10c, vB11, vB12]. Books [Bro06a, Bru03, But19, DeV00, Fra04, Ger99a, Hil07, Hob04, Lan10, Lar05c, PCGD02, PP08, RS06d, RS10c, RS10e, Seg10b, Seg15, vB11, Mor90]. Books/Henry [Bru03]. Bose [Ban16]. Boson [Ber16, Hen07a]. Boston [Mar03b]. Boston/Basel [Mar03b]. Both [Hon00]. Bottle [Gre07a]. Bottom [Hil07]. Box [Sei00]. Boyd [RS12d]. Boys [Foc22]. bra [D’A02]. Bradbury [Sei00]. Bragg [Eme09, Hun04, Jen08, LM06]. Brazilian [CdL16]. Breadth [JS02]. Breakfast [Joh99b]. Brecht [Kra22]. Brian [Ger06b, RS10b]. Brief [Kos13]. Briefe [Cas05a, HHH03]. Briggs [NHKL14]. Bright [Cre08]. Brigitte [Stu04, Tri08, Tri08]. Brilliance [RS04c]. Brilliant [Seg10b, Mag09]. Bristol [Bru99a, Fre00, Ryn99, Van03, Wat01, vB02b, BP08]. Britain [Gea99, JP07, Smi98]. British [Kra11a]. Broke [Bed09a, NHW08]. Broken [Bri07]. Bromley [Jac03]. Brookhaven [Rio00a, Bon18, Cre99, Wil17]. Brown [Fre00, RS06c, RS08b]. Bruce [RS07d, Bed09a, Ham11]. Bruno [Seg15, Bon11, Clo15]. Brunswick [Ger99a, Hie99, Yeh05]. Bryan [Boy12]. Buchwald [Bru02, Dom05]. Budapest [Kov03]. Building [Ano18a, Wil17]. Bush [RS07a]. Byers [Eve08a].

C [Alp12, Bed08, Bed09b, Ger06a, Hol03, Hol19, Lus10, Rob18, Sim12a]. C. [Ban14, Sha06, Sin02]. C3 [MPFP19, SB19, MPFP19]. ca [Sal14, Sal17]. Calculating [NGZ23]. Calendar [Liu20]. California [Bad09a, Bed01a, Bla00b, Neu99, RS07a, Tri08, Hol11]. Calle [Lan10]. Callendar [RA15]. Cambridge [Ano18a, Bor11, Bro04, Bru02, Bru06, But00, Car11, Cas10, Cas11, Coh11, Com05, Cre12, Dom05, Eve08a, For02, For08, Fra01a, Fra12a, Goo09a, Gou01, Gre07a, Hal09a, Han04, Har85, Har05, Hob02, Hol19, Jac04, Kah03a, Kro10, Mil10, Mod18, New99, New04, Oha08b, Par06, Pas10, Pri99, RS06c, RS09d, RS10e, RS11b, Ros08, Ryd13, Sch05, Seg10a, Sha06, Sim04, Ste03b, Sty05, Tei02, Tri07, Wat03, vB08b, Nye08, Pip99a, Por09, War03]. Camille [Kel22]. Campaign [Bør12]. Can [RS10f, Sil07]. Canaday [For01]. captivates [RS08h]. Cardinal [Ano10a, Her00, vB02a]. Cards [Ghi05, Shi06]. Care [RS06b, FL88]. Career [Huf12, Eic08, Sim10a]. Carl [van12, vB10a, Cas15, Dav00a, Jam00, Mil09, Pou99, Sch00a, Sch04]. Carlo [Bla00a, Stu06]. Carlos [Lan10]. Carried [VV14]. carrière [Eic07]. Carrington [RS07b, Cla09]. Carson [Cas11]. Case [Das15, Fre99, GLR20, Maa07, Man22, Jan02, LPR04]. Cass [Fre04]. Cassidy [Bed09b, Ger06a, Hol19, Lus10, Rob18]. Catalyzed [Jac10]. Catapult [NHKL14]. Catherine [Seg09]. Cathode [Ryn99, Dah97]. Catholic [Hel05, Hub06]. Cathryn [Cas11]. Causality [Sch09b]. Causes [PCGD02, Goo02, Mil01]. Cautionary [Kle11, Rus15, Goo10]. Cavendish [Nav06, Pip99a]. Celebrate [RS01a]. Celebration [Jac11]. Celebrity [Hal18]. Celestial [Sal20]. cell [KKG^{+}19, PS19]. centenario [Ber02b]. centenary [Ber02b]. Centennial [FJ21]. Center [Hof99, KG19, Lal19]. Central [Wes14]. Centrifugation [Ree09]. Centrifuge [RS03b]. Centuries [SE14, SS08]. Century [Bru06, Eve08a, GMS^{+}03, Gre20, Hug09, Jac03, Mou20, Per01, Ren08, RS07c, Ros08, Rus15, Sta06a, Wat03, vB10c, Bro02, Cas05b, Har06, Kra99, BW06, Cus00, Dar04, Fra09c, Ger06a, Ger99a, Gre07b, Gre21, Han04, Hen05b, Hie99, Hob04, Hu16, Jac06, Joh99a, Kev97, LR03, Liu20, Pas12b, Pur97, RS06c, Sti08, TS01, Wee10, Har85]. Cercignani [Bla00a]. cérébro [PMAS19]. cérébro-spinal [PMAS19]. CERN [Kri05, Lac07, PA21]. CERN’s [PA21]. Cesar [VV14]. Cham [For20, Rob18]. Chance [Kip00, Par07, vB06]. Chang [RS08b]. Chang-Hwan [RS08b]. Change [Mar15, RS11f]. Changed [Fel08, Sti08, Den07, Har06]. Changes [Per04]. Chaos [New99, Fla98]. Character [FLH85, FL06, FL88, RS06b]. Charap [Ehr04]. Charis [For09]. charisma [RS04c]. Charles [Yeh05, Bri99, Hol03]. Charlottenburg [Hof00].

Charting [RS07c]. Chemical [Joh04, Sim04, LR03]. chemins [Eis02]. Chemists [Sim02]. chez [PMAS19]. Chicago [Ahe02, Car07, Dah06a, Fra07, Gea99, Ger06a, Har05, Lep07, Lin10, Mas07, Nyc09, Par00, PCGD02, RS07d, Rio00a, Roe10, Seg09, Seg13, Sha02, Tay10, Tor02, vB10b]. Chief [Ren05b, Ren05a, RS06a, RS11c, RS11d]. Chill [Wes14]. China [Hu19a, Hu19b]. Chinese [Hu16, Liu20]. Choices [Her00, vB02a]. Chongqing [Hu19a, Hu19b]. Christian [Ste03a, Kra08b]. Christy [Lip06a, Lip06b]. Chromosome [CFL19]. chromosomes [Seh19]. Church [Wil99]. Cinematic [Waz10]. Circuit [Hal10]. Citadel [Sic06]. City [Hec04, Hof99, Bed03, Ros01]. Claim [Buc20, Sim02]. Clark [RS07b]. Classic [FL06, Hug99, RS06b]. Classical [Kro10, Bok08]. classification [Hof02]. Climbing [MM00, Ger03]. clocks [Mar04a]. Clockwork [vB08b, New07]. Close [RS10d, RS11e, Seg15]. Closely [CP14d]. Cloth [Ros01, Sch01, Ahe02, Alf09, Aur04, Bed01a, Bed06, Bed08, Bed09b, Ber08b, Bla00a, Bla00b, Blo06, Bor11, Bri99, Bri07, Bro04, Bro06a, Bru99a, Bru02, Bru03, Bru06, But00, Can01, Car11, Cas05a, Cas10, Cas11, Che11, Coh11, Col00, Com05, Cre12, Cus00, DeW99, Dom05, Dor06, Ehr04, Eme09, Eme11, Eve00, Eve04, Eve08a, Eve08b, Fel08, Fel11, For01, For02, For08, For09, Fra01a, Fra04, Fra07, Fra09a, Fra10, Fra01b, Fre00, Fre04, Fri08, Fri09, Fri11, Gea99, Ger03, Ger06a, Ger99a, Ger06b, Gib05, Gol99a, Goo09a, Goo08, Goo09b, Gre05, Gre07a, Gre06, Gre09, Gri06, Hal09a, Han04, Har05, Hec07, Hie99, Hil11, Hob02, Hob04, Hol09, Hol07, Hub06, Hun01, Jac03, Jac04, Jam00, Jon07, Kah03a, KP02, Kle07, Kle11]. cloth [Kro10, Lan10, Lar05c, Lin10, LM06, Lus04, Lus10, Mar03b, Mas07, Mer01, Mil10, Neu99, New99, New04, Noe11, Nyc09, Oha08a, Oha08b, Oha10, Par00, Par06, PCGD02, Pas04, PP10, Pas10, Pas12a, Per10, Pes11, Pip99b, Pri99, Ram05, Rig07, Rio00a, Roe12, Ros08, Ryc10, Ryd13, Ryn99, Sch00a, Sch05, Seg09, Seg10a, Seg11a, Seg13, Seg15, Sha06, Shi06, Sim04, Sim08, Sim12a, Sob08, Sta06a, Ste03b, Sti08, Stu04, Stu06, Tat13, Tay10, Tei02, Tor02, Tri07, Tri09, Van03, Wal07, Wat03, Wei04, vB99, vB00, vB01, vB02b, vB06, vB08a, vB08b, vB10a, vB10b, vB10c, vB11, vB12]. Club [For22]. CNO [Wie18]. Code [Tay10, Mel09]. Coen [Lin10]. Cohen [Dom05, Gou01]. Coherence [NW08]. Cold [Ruh20, Sal19, Seg13, Yeh05, Cha02a, MW03, Tur12, Kle23]. Colin [Can01]. Collaboration [Ber10b, Eme09, Jen08]. Colleagues [Kle10]. Collected [Fey05]. Collecting [Man22]. College [Gre06, RS08a, FG05]. Collide [CMP18d]. Collider [Ber04a, Giu12, Lin09, RS09c, Rio00b]. Collins [PCGD02]. Colonial [Ban14]. Color [Pes13]. Colors [Rei19]. colour [vB02b]. Columbia [Hug09, vB10c]. combines [RS06b]. Comes [RS12a]. Commentariolus [Rom99]. commentary [Hug99]. commitment [Gra08]. Committee [Bed05, RS07c]. Common [Ano99a, Hie00, Mic06]. Communist [Bor11, Jos10]. Community [Das19, ORSSBO^{+}17]. Company [Jam00, KP02, RS09a, Sim07, vB10a]. Comparative [MPFP19]. Comparison [dWtH20]. competed [SR06]. complement [SB19]. Complete [Gol99b, BV09]. Complex [New99, Ree14, Fla98]. Complexity [MPFP19]. Component [MPFP19, Ree14, SB19]. Compton [Jen02]. Computation [DeV00, Hey99, Jon07]. Computational [New99, Fla98]. Computer [Gou01, KP02, New99, Coh00, Dav00b, Fla98]. Computers [Hey99, DeV00]. Computing [Par11]. Concept [For09, McM02, Noe11, dWtH20, Ana08, Coo10]. conception [D’A04]. Conceptions [Kra07, Goo08, Tri09]. Concepts [Jam07]. Conceptual [Bro13]. Concerning [RS11d]. Concerns [ZZ18]. Concrete [Wes18, Wes19]. Condensed [Pav10, Wil17]. Condon [Day08]. Conference [BV09, Seg10a]. Conferences [Ano16e, Sch16]. Confirmed [Nal20]. Conflict [Bru99a]. conformist [Gav01]. Congress [Man23]. Connected [RS08c, Lei08]. Connection [Kra00, Seh19]. Connections [Har02b]. Conoscere [Ber02b]. Conscience [vB12, LH11]. consciousness [RK11, Bed09a, RS07d]. Constancy [RS11f]. Constant [Hol99]. Constants [Kra19]. Constructing [JP07]. Construction [Ryd13, HDS12]. Contact [Hon00]. Contemporary [Bal08, Mer01, Omn99a, RS09d, Hug99]. Contestation [IT23]. Context [DS22, Ren05c, Wal03b, RS06a]. Contextual [Sch14]. Contra [Gra11]. contraction [RS02b]. Contraption [vB06, Par07]. Contributing [Hon03]. Contribution [Nau05]. Contributions [BW06, Eve08a, ORSSBO^{+}17, Sch04, Loe03]. Controversial [Ger06b, Bri05]. Controversy [Hun01, Kra96, Waz10, PFD08]. Converging [Mic06]. Conversation [Goo99, Goo03b, Goo04, GF05a, GF05b, Lip06a, Lip06b, Lip07, LG08, PCGD02, Day02, Goo01, Goo03c, LC01]. Cool [Car07]. Cooper [Gol99a, Wei04]. Cooperation [Sal19]. Coopersmith [Noe11]. Copenhagen [RS07a, Cas02, Kra22, PA01, Seg07]. Copernican [Gra08]. Copernicus [Lal19, Rom99]. Coriolis [Gra11]. Coriolis-Force [Gra11]. Correction [Hu19a, Ill20]. Corridors [Nye99]. Cosmic [Fra08b, Fri89, Gre12, Kah03a, Mil09, vB10a, Fri00, Alf09, Alp12]. Cosmologies [Wie17a]. Cosmology [Kra96, HDS12, Kra07, Seg11b, Goo08, Hun01, Ryd13, Tri07, Tri09, Cas12]. Cosmos [Goo08, Kir02, Tri09, vB10b, Kra07, Pou99, Jam00, Pas04, Wal09, Wal11]. Coulomb [Hee06]. Councils [vB00, MW99]. Countryside [Gre00]. couple [Bad09c]. Course [DE07, RS07c]. Court [NGZ23, Sal20]. cover [Ste03a]. Cracking [Mel09, Tay10]. Cracow [Str04]. Craft [Jac00, Tei02]. Crafting [Set10, Car11]. Crane [Hol11]. Crapshoot [vB08b, New07]. Crazy [Ano05b]. Crease [Kle07, Rio00a]. Creator [Bri07]. Creators [Bed08, Kel07]. Credit [Mas07]. Crelinsten [Hol07]. Crimes [Ano00b]. Crisis [RS02c]. critical [Cas05a, HHH03]. Cropper [Lus04]. Crosbie [Gea99]. Crossroads [Row01, Seg10a, BV09]. Crowe [Noe09]. Crystals [Rei19]. Cultural [DS22, For02, Gea99, HKT02, RS08a, Smi98, Wei01]. Culture [Aur04, Gau18, Hen99a, Kov03, PCGD02, ZZ18, Day02, Day04, LC01, Pan03, PS19, Ros98, Sok08, Stö02, Goo09b, Neu99]. Cultures [Mon00, Sha02, Day04]. Curie [Sim07]. Curies [Ger06b, Bri05]. Curiosity [Fra12a, Huf11]. Curious [Eis06, FLH85, FL06, Oha08a, RS06b, FL88]. curmudgeon [And11, Phi13]. Current [Par11, Lar07]. Curt [RS10e]. Curvature [Tor00]. Cycles [Wie18]. Cyclotron [Aus16]. cynicism [Haa03]. Cynthia [Bed08].

D [Eve00, Hie99, Jac04, RS08a, Sha06]. D. [Jac03]. Daedalus [O’C14]. Dahl [Ryn99, Van03, Wat01]. Dahlem [Hof00]. Dame [Gib05, Hub06]. Dan [Gol99a, RS09b]. dangerous [Ber02a]. Daniel [Ahe02, Nal20, vB08a]. dans [PMAS19]. Dardo [Bru06]. Dark [Kir02, Pas04, RS01d]. Darkness [Seg10b, Mag09]. Darrigol [Fre05a]. Darwin [Seg11a, Mar11]. Dassow [vB10b]. Data [RS12e, Wes16, Wes18, Wes19]. Daughter [Bed01a, Pal00]. David [Ger06a, Hol19, Bed09b, Dor06, Eme11, For20, Hol09, Kle11, Lep07, Lus10, RS07b, Rob18, Roe12, Sch03, vB06]. Davidson [Sch00a]. Davies [RS10b, RS11b]. Davis [KP02]. Day [Lan12, Pip99b]. Days [Eck99, Hol11]. DC [Cas11, Dor06, Ram05, Hal10]. Dear [Cas05a, Gre09, RS07d, HHH03]. Death [GR13, Kel22]. Debate [FJ21]. Deborah [Lin10]. Debye [Sim10a, Eic07, Eic08]. Decade [Sch16]. Decay [GR09]. December [Sch21]. Deciphering [Mil09, vB10a]. Declassification [Kle23]. Decline [RS11a]. Decoding [RS10c]. Decolonizing [CMS19b]. Dedicated [For20]. Dee [Ger06a]. Defection [Seg13, Tur12]. Defending [Haa03, Haa07, Fra04]. defiant [Gav01]. Define [Gre12]. Degrees [Dor06, Lin04]. Delbrück [Cas12, JFEH15, Seg11b]. Della [RS09a, Ber02b]. Demise [Rio00b]. Democracy [RS11e, Fer10]. demystified [Wol03, Tay05]. Denis [Ger06b]. Denmark [Sch11]. Dennis [vB01]. Denny [Fel08]. deny [Bør12]. Department [BS11]. dependent [Seh19]. Depth [JS02]. Description [VMR19]. Design [Ehr06, Cal09, Lan10]. Desire [Wes16, Wes18, Wes19]. Dessauer [Goe03]. DESY [Hab10, HHH17]. detect [Nav06]. Detecting [Arn01]. Detection [Lev04]. Deutsche [HW08, Sim08]. Deutsches [Sim10b]. Development [Hun01, Kra96, Rus15]. Deviations [Bro06a, Fey05]. Diagrams [Lep07]. Dialog [Jha11]. Dialogue [Par00, Bel99, RS11d]. Diary [Lin03]. Dibner [Dom05]. Dick [Hol11]. Dictionary [HKT02, Hol01, RS08a, Isa00]. Did [Nor16, ORSSBO^{+}17, Jen11]. Dieter [Sim08]. Different [Hil07, Phi13, Wes06, And11]. Difficulties [D’A00]. Diffusion [Ree11]. Dilemma [IT23]. Dilemmas [Vey13]. Dimensional [Hal07]. Dimensionality [Hal04]. Dimming [RS13f]. diploid [PS19, SE19]. Dirac [RS10c, Bro06b, D’A02, Far09b, Sim02]. direction [WWH01]. Directs [RS11f]. Disappearance [GR13, Seg10b, Mag09]. Disaster [Kap08]. Discord [Eve04, Fra02a]. Discover [Nor16]. Discovered [O’C10, RS10e, Hol08a, Hol08b, Hol08c]. Discoverer [WJS12]. discoveries [Die05]. Discovering [NB09, Pas10]. Discovery [Bro04, Buc20, DE07, Fri99, Gab00, Ger99b, GLR06, GR09, GLR12, GLR20, Kip00, Noe11, OR21, Pav10, Per04, Pes15, Sim00, Sin02, VV14, AGR04, Coo10, LPR04, Wei83, Wei90, Wei03, Wei15]. discussing [FR11]. Disintegration [Tri08, RS03e, SRD06]. disparate [Mul99]. Dispersal [NHKL14]. Dispersion [Lep07, Pip01]. Displaced [Bed05]. Display [Sim10b]. Disposition [NGZ23]. Disrupting [Moo08, Fri09]. distinguished [Mul99]. Distortion [RS07a]. district [MDK19]. Disunity [PA21]. Diurnal [Gra11]. Diverse [Das19]. diversity [Cha19]. Divided [Seg15, Clo15]. Divine [GLR20]. Do [Ano16e, RS05b, RS06b, Sil07, FL88]. Documents [RD05, RS06a]. Does [RS12e]. Dog [Bed08]. Dokumente [RD05]. Dolejsek [Fri99]. Doll [Pip02]. Domenico [Gre07b]. Dominique [Lal19]. Don [RS09c]. Don’t [CMS20a, FF05, Fra16, RS13e]. Door [Eve12]. Dorion [Car07]. Double [Bed12]. Doubleday [RS07b]. Doubt [Nal20]. Down [Hil07]. Drag [Fre05b]. Drama [Sch22]. Dramatic [Rob18, Cas17]. Drawing [Ber03, Lep07]. Dreams [Kle23]. Dresden [Kah03b]. Dritten [HW08, Sim08, MPFP19]. Drives [Fir12, RS13d]. Dry [Bee15]. Dublin [O’C14]. DuBridge [Goo03b, Goo03c]. Duhem [Cro99, Kra08b]. Duitsland [Eic07]. duplication [SE19]. During [NGZ23, Hof05, Ree09, Ree11]. Dutch [Sha06]. Dynamics [Nau05].

E-36 [Pro16]. E. [Day08, RS08a, RS10b]. Early [Aus16, Eck99, Goo12, Gre09, Hu16, Hub06, Huf12, Hug09, Kle23, Kra11a, Lev04, Nav06, ORSSBO^{+}17, Ram99, RSD07, Sal17, vB10c, Hel05, Osl10, Ren08, Sch06, Sch03, Ruh20, Sal19]. Earth [Gra11, Kra15]. East [Hof09]. Eclipse [CdL16, Nal20]. Eclipses [RS09c, LWE09]. economics [RS08h]. ed [Bed08, Bro06a, Bru02, DeV00, Dom05, Eve08a, Fre04, Gre06, Gre09, Han04, Hob04, Hol01, Hol09, Jon07, PCGD02, Pip99b, RS06c, RS09d, Sim04, Sim08, Ste03a, Stu04, Stu06, Tri08]. Eddington [Dur03, Nyc09, Sta07]. Edinburgh [Hen07a]. Edited [Cas05a, Ryc10, RS11b]. Edition [Mal12, RS09c, Shi06, Hol01, vB02a]. Editorial [Ano16f, Ano20, CP14a, CP15a, CMP17a, CMP18b, Fan16, Mar20, RS02c, RS02d, RS03a, RS03b, RS03c, RS05b, RS06g, RS06f, RS07f, RS07g, RS07h, RS07e, RS08g, RS08h, RS08e, RS08f, RS09g, RS09h, RS09e, RS09f, RS10g, RS11g, RS12e, RS12f]. Editors [RS08b]. Edizioni [RS09a]. eds [RS12d, vB00]. Education [Hal05, Hen99a, Huf12, Hun15]. Educational [CS18]. Edward [Ano18a, Coo11, Rig11, Wat03]. Effect [Jac14, Jen02, Sin02]. Ehrenfest [Hal04, HK07, Kle10]. Ehrlich [Jon07]. Eickhoff [Sim10a]. Eighteenth [Mou20, Han04]. Eighteenth-Century [Mou20, Han04]. Eighth [RS06e]. Eighty [RS12f]. Eighty-Six [RS12f]. Einstein [Nal20, Oha08a, Oha10, Par06, Ren05b, Ren05a, RS06a, RS07b, RS11c, Seg11a, Tat13, Tay05, Tay10, vB08a, Ari03, Ari05, Coo10, Cre06, Cro07, Dar00, Die05, Eis02, Eis06, Fre05a, FR11, Goe03, Goo09a, Hal07, Hal18, Ham11, Hol09, Hol05, Hol07, Hol19, Hug99, Hun10, Hun15, Ill12, Ill19, Ill20, Jan02, Ken07, Kle10, Lis16, Maa07, Mar03b, Mar04a, Mar04b, Mar09, Mar11, Mel09, Mil99, Mil01, Noe09, Noe11, Nor16, OM18, Ove00, Ove01, PCGD02, Pri99, Ren05b, Ren05a, RD05, Ren05c, RS06a, RS06d, Row01, RS07i, Sch08, Sta02, Van07, Waz10, Wol03, dWtH20, van12, vB01, Goo02]. Einstein-Marić [Hol19]. Einsteins [Ren05c]. Eisenhower [RS06h]. Eisenstaedt [Oha08a]. Elderly [DFAB19]. Electrical [Kra18]. Electricity [Cha01, Man23]. Electrodynamics [Dar00, Sch04, Mil07, Ste05, SL16, Fre05a, Hof17]. Electron [Ber04a, D’A00, Ros12, Bru02, Dah97, Ryn99]. Electron-Interference [Ros12]. Electronic [Bri07]. Elegant [Lar05c, Lai04]. Elektrodynamik [Ste05]. Elementary [Per04, RS07c, LR03]. Elementary-Particle [Per04]. Elements [Sim00, LR03]. Elisabeth [Pip02]. Ellen [LKK04]. Eltern [Cas05a, HHH03]. Eludes [RS12f]. Emanation [GLR20]. Emanative [Jac14]. Embattled [Sal19]. Emergence [Sil07]. Emergency [Bed05]. Emerging [Hon03]. Emil [Fre99]. Emission [Mar04b]. Emotion [RS11g]. Emperors [Sal20]. Empire [Mil05, Fri08]. empirical [Loe03]. Emulsions [Sim13, VV14]. Enabler [Ber16]. Encounter [Rig07]. Encountered [Liu20]. Encounters [Bed09a, RS07d, RK11]. encyclopedia [Kur02]. end [RS08h]. Endings [CMS21b]. ends [WWH01]. Energy [Car07, Coh11, Coo10, Dah99, Fel11, Gea99, Har09, Jen11, Kir02, Pas04, Gol10, Gol12, Smi98, Sta10, Noe11, Wat01]. Engine [Esp21]. Engineer [Maa07, Ren05b, Ren05a, RS06a]. Engineering [Wes14]. England [Esp21]. English [AR19, Sch06]. Englishman [Bor08]. Enigma [Bed09a, RS07d, Ber04b, RK11, Ger06a]. Enlightenment [Per10, RS13f, Sch01, Lev09, Sch97, Aur04, Lac07, Pan03]. Enrico [AGR04, BB04, Coo99, Gam03, Gol99b, Gol99a, GLR06, GR09, GLR20, Stu06]. Entangled [Jon07]. Entanglement [Rig07]. Entropy [Kel22, Kra08b]. Eötvös [Stu05]. Epistemology [Jha11]. Eponymy [Cre09b]. Equations [Das15]. era [RS01c]. Eric [Car07, Das18, Sob08]. Erosion [Gre01, Ahe02]. Erratum [KA15]. Escape [RS12e]. espace [Eis02]. espace-temps [Eis02]. Espenak [RS09c]. Essay [Fra09b, Hof17, Lus10, New04, Pes03]. Essays [Eve00, Bar99a]. Essen [Ess10]. Essential [Hil11, Mul10]. Esterson [Hol19]. Estrangement [Fix15]. Ether [Fre05b, Pip01]. Ethical [Ahe02, Gre01]. Ettore [Seg10b, DE07, GR08a, GR08b, GR13, Mag09, RS09a]. Eugen [VJ10]. Eugene [Stu05]. EUR [Lal19, Sta06a]. Eureka [Hir10, RS09a]. Europe [Goo12, IT23, Osl10, Tis09a, vD20]. European [RS08d, Sal07]. euros [Ste03a, Stu04]. Events [LoS16]. Ever [Fra05]. Every [Hil11, Mul10]. Everyone [ZZ18, For04, Com05]. Everywhen [Pes11, Hug10]. Everywhere [Hug10, Pes11]. Evolution [For09, Ana08]. Evolved [Wes16, Wes18, Wes19]. Exact [Hie00]. Excellence [Bru03, Fri01]. Excursions [Hof00]. exhibition [RD05]. Exile [Bed12]. Exiles [Fra05]. Exner [Stö02]. expanded [Tri08, vB02a]. Expanding [Kra15, NB09, Pas10]. Experiment [Fra99, Fre05b, Hon00, Pro16, Ros12, Fra02a, Sta01, Eve04]. Experimental [Hon03, Hug09, RA04, She13, Sic06, Wes16, Wes18, Wes19, vB10c, Ren08]. Experimentalist [RA10]. Experimentalists [Gea14]. Experimentalpraxis [HRS00, Ste03a]. Experimente [Ste05]. Experiments [DMA07, Eck99, Gea14, Hee06, Hof17, Lev04, Mil99, NHKL14, PA21, Tat13, VJ10, HRS00, Ill12, RS01b, Ste03a, Ste05, SL16]. Explain [Pes15, Wei15]. Explaining [Cha02b, Ehr04]. Exploding [Kir02, Pas04]. Explorations [New99, Fla98]. Explorative [Ste05]. Exploratory [SL16, Hof17, Ste05]. Explorers [CP15b, JFEH15]. Exploring [DeV00, Hey99]. Extraordinary [Seg10b, Eme09, Jen08, Mag09]. Extravagant [Kir02, Pas04]. Eye [Che01, Lan12]. Eyewitnesses [Bed08, Kel07].

F [Cas05a, Hec07, Lip06a, Lip06b, RS08a, Ryn99, Van03, Wat01, vB02b]. Facility [Rus15]. Facing [For02, Wei01]. Fact [dWtH20, Goo10, Kle11]. Factor [Hon03]. Facts [ZZ18, RS01a]. Fair [Man22]. Faith [Bar03, Kra08b, RS03d, RS05c, Rus00, Can01, Gib05]. Falk [Ste03a]. Fall [Bri07, DFAB19, Sal14, Ste23]. Fame [Van07]. Family [Ger06b, Bri05]. FAPP [RS07f]. Fara [Fra10]. Faraday [Can01, Hof17, Pip99b, Ste05, Day99, Rus00, Ste05, SL16]. Farm [Cas17, Rob18]. Farmelo [RS10c]. Fascination [Jon07]. Fascist [Bon11]. Father [Eme09, WJS12, Jen08]. Fathers [RS05c]. Faust [RS07a, Seg07]. Fe [Noe09]. Features [RS06e]. February [Sch21]. Federico [Sch09a]. Feed [Ree14]. Ferdinand [NGZ23, Sal20]. Fermi [Gol99a, Stu06, AGR04, Ber02b, BB04, Coo99, Gam03, Gol99b, GLR06, GR08b, GR09, GLR20]. Fermilab [HK03, KW08, Seg09, Wes17]. Ferris [RS11e]. fetus [KKG^{+}19]. few [Ano01b]. Feynman [Bro06a, Lep07, Noe11, RS06b, RS12h, Ano05a, Bro05, Bro06a, Coo10, CP14b, DeV00, Fey05, FL06, FGL^{+}06, FGL^{+}12, FR11, GG00, Hey99, Kra11b, RS06b, RS06c]. Fiber [Hec04, Ros01]. Field [D’A00, Hal18, McM02, Pes12]. Fiftieth [Bon18]. Fifty [Cas15, CP14b, Wes06]. Fifty-Year [Cas15]. Film [MM23, WJS12, Waz10]. Finches [Seg11a, Mar11]. Finding [Har09]. Finite [RS12d, HW11, Lin03]. Finn [Kra16]. Firestein [RS13d]. First [Alf09, Ber04a, Can00, For01, GMS^{+}03, Hob04, LoS16, Mil99, PA21, Pro16, Rig07, Wes17, von09, AGR04, Fra08b, RS06c, Fra09c, Man23]. First-Person [Wes17]. fish [SE19]. Fisher [Eme11]. Fisica [Stu06]. Fission [Sim00, Sim10b, Sim12b, Van03, Dah02]. Five [Fel08, Hal07, Sti08, Den07, Har06]. Five-Dimensional [Hal07]. Fizeau [Fre05b]. Flake [New99]. Flammarion [Kel22]. Flash [Dah97, Ryn99]. Flawed [Bør12]. Flight [Boy12]. Flow [Car07]. Flower [NHKL14]. Flying [Dah09]. Foam [DeW99, WF98]. Foldy [TF07]. follow [RS07h]. Following [Nye99]. Force [Gra11, Sal14, Hoo08, Sch06, RS09b]. Ford [Com05, DeW99]. Forecasting [NGZ23]. Foreign [Bed05]. foreword [Cas05a]. Forget [RS05d, RS13e, Ruh20]. Forgetting [Sim12b]. Forging [Smi20]. Forgotten [GR08b, RS10h]. Form [Mul01, Her03]. Forman [Sch14]. Formation [PS19]. Forty [Per04]. Found [Eis06, Oha08a]. Foundational [Ree14]. Foundations [DS22, PFD08]. Four [Fra10, Wie17b, ZZ18, Far09a, Far10]. fourth [Hol01]. Fowler [GF05a, GF05b]. Fox [Gre06]. Fractals [New99, Fla98]. Fragments [Bed01a, Kra19, Pal00]. France [Per10, Lev09]. Francesco [Che11, RS09a]. Francis [Cre17]. Francisco [RS06b]. Franck [Lem07, Gea14, Lem07, LH11, vB12]. Franco [Goo01]. Frank [Fre04, RS11e, Goo04, Seg15, SS07]. Franklin [Eve04]. Franz [Stö02]. Fraser [Alf09, RS06c]. Fraud [Kle11, Goo10]. Fraunhofer [Tei02, Jac00]. Fred [Bed09a, RS07d, RS09c]. Frederick [RS02d]. Free [Sal14]. freedom [Eve06]. Freeman [Bru03]. Freeman/Times [Bru03]. Freiherr [Cas15]. Freire [For20]. French [Mer01, Oha08a, Yeh05, vB99, Eis02]. frequencies [AR19]. Friedlander [Kah03a]. Friedman [Bru03]. Friedrich [Cas15, Goe03, Hof17]. Friends [Kle10]. Friendship [Cas15, Fri08, Mil09, vB10a, Mil05]. Fritz [Tra08, Bed05, Hof09]. Fritzsch [RS11c]. Front [Kle11, Goo10]. Frontier [CMP17d, KW08, Lin09, RS09c, Seg09, HK03]. Frontiers [Hie00]. Frontiersman [Wil00]. Fuchs [Ber10b, Hof09]. Fueling [CS18]. full [CFL19]. full-term [CFL19]. Fuller [Tor02]. Fulvio [Tay10]. Fundamental [Fle21a, Sch16]. Further [FL88]. Fusion [Fel11, Jac10, Wei08, Sta10]. Future [Hil11, NGZ23, RS08d, Mul10].

G [DeV00, vB08b, Seh19]. G. [Jen02]. Gaetano [Man22]. Galilei [Gra10]. Galileo [Lus04, RS11d, Sal14, Cre12, Cro01, DFAB19, Fri11, Gra08, Gra10, Gra11, Hei10a, Kra22, Mas07, OT10, Pet11, RS11c, RS11d, Sal17, Sch06, Wal03b, Wil99]. Galison [Mar04a]. Gamma [Ger99b]. Gamow [Cas12, Hob02, Har01, Seg11b]. Gans [von09]. garden [Sch03]. Garrett [Mod18]. Gary [Eve08a, New99]. Gases [BF13, Fis10, Eme11]. Gauss [Sta06a, Hen05b]. Gaußens [Hen05b, Sta06a]. Gdańsk [Jan11]. Gears [Wes19]. Gedanken [Mil99]. Gell [Per01, Joh99a]. Gell-Mann [Per01, Joh99a]. Gendered [Hug09, vB10c, Ren08]. gene [Cha19, SE19]. General [CdL16, Eis02, Mil99]. générale [Eis02]. Generation [Fra12b, RS10e, Hol08a, Hol08b, Hol08c, Sta01]. Generations [Cus00, Kra99]. Genesis [AH01, PCGD02, Unn00]. Genetic [SB19]. Geneva [Lac07]. Genius [Bri07, Dor06, Seg10b, Sim07, Lin04, Mag09, Sch08, Goo09a]. Geniuses [Cas12, Seg11b]. Genomics [Cas12, Seg11b]. Geochemistry [Kra00]. Geography [Ste23]. Geometrical [Ano00a]. Geons [WF98, DeW99]. Geophysics [Ham00]. George [Cas12, RS08d, Bla00b, Boy12, Har01, Hob02, Per01, Seg11b, Stu05]. Georges [Lal19, Yeh05]. Gerald [Shi06, Par06, RS08b]. Gerjuoy [Coo11]. German [Bed09a, Ber08b, Cas05a, Cas11, Jon07, Neu99, Rob18, Sim08, Ste03a, Tri08, Cas02, Cas17, For22, HRS00, HHH03, Hen05b, Hen07b, Hof05, HW08, Lem07, MPFP19, Ren05b, Ren05a, RD05, Ren05c, Ros98, RS03e, Sim12b, Ste05, Van07, Wal03a, dWtH20]. Germans [SB19]. Germany [Cas05a, Bed12, Bør12, CS18, Eic08, Hel05, Hof09, Hub06, Jac06, Ros08, Sim10a]. Gerry [Jac11]. Gerson [Pav11]. Gertrude [Gol16]. Gesellschaft [HW08, Sim08]. Ghirardi [Shi06]. Gian [Che11]. GianCarlo [Shi06]. Giere [Fra07]. Gillin [Ano18a]. Gino [RS07a, Cas12]. Giudice [Che11]. Giuliano [Aur04]. Glasgow [Joh06b]. Glashausser [Yeh05]. Gleditsch [LKK04]. Global [Ano11c, Ano16h, Bla00b, Fra12a, Sil07, Huf11, Per99]. Globalization [RS03a]. glycoprotein [AR19]. Go [Fra05]. God [Goo12, Shi06, Ghi05, Jac14, Osl10, Sam10]. Gold [Coh11]. Golden [Hab10]. Goldhaber [Gol16, Pav11]. Goldsmith [Sim07]. Golub [Ste03b]. Gonzo [CMP17b]. Good [Ano02, RS01b]. Gooday [Gre06]. Goodstein [Kle11]. Gordon [RS06c, Alf09]. Gore [Sil07]. got [Lin03]. Gothard [VJ10]. Göttingen [Sta06a, Hen99b, HR01, Sch03]. Graeme [Gre06, LM06]. Graham [RS10c]. Grand [vB06, Par07]. Gravitational [DMA07, vB08a, Ken07]. Gravity [Eis06, Kra15, Lev04, Tor00]. Gravity-Wave [Lev04]. Great [Cro01, FJ21, JP07, Lus04]. Greatest [vB11]. Green [Noe09]. Greenberg [Ahe02]. Gregersen [RS11b]. Gregg [vB02a]. Gregoire [vB00]. Groping [BF13]. Group [Sim07, MDK19]. Guerra [RS09a]. Guest [Fan16]. Guide [Bai07, RS09i, Sim07, Joh09, Oha08b, Wes10]. guides [LWE09]. Guido [Seg10a, dP07]. Guimarães [Gri06]. Günther [Bed09a]. Gutenberg [Hug09]. Gyros [Ill19, Ill20].

H [Bri99, Bru03, Lus04, Tri08, Gor09]. H-Bombs [Gor09]. Haack [Fra04]. Haas [Wie17a]. Hadron [Giu12, Lin09, RS09c]. Hahn [Wal03a, Sim06, Sim12b, Wal03a, Wal06]. Haibel [Bed09a]. Haigerloch [GORS09]. Half [Buc20, Clo15, Rus15, Seg15]. Half-Century [Rus15]. Half-Life [Buc20, Clo15, Seg15]. Hall [Rob18, Cas17]. Hamburg [Hab10]. Hand [Ano11d, Gre09, Hen05b, RSD07, Sta06a]. Hans [Bed09a, BL06, Goo99, Jac09, RS08b, Sch14]. Hans-Joachim [Bed09a]. Happened [Bri99, Cra00, Tow99]. Harald [RS11c]. hardback [Ano18a]. Hardbound [DeV00, Gou01, Kob02, Per01, Sha02, Wat01, PCGD02]. Hardcover [MG21, But19, Das18, For20, Hol19, Kra16, Lal19, Mod18, Nal20, Rob18, Smi20]. Hargittai [Sti08]. Harmonious [Jac06, Ros08]. Harmony [BF13, RS08c, Lei08]. HarperCollins [Pes15, RS09b, RS11e]. Harriot [Sch06]. Harry [PCGD02, Pas10]. Hartley [Boy12]. Harvard [Com05, Cre12, For02, Goo09a, Kah03a, Mil10, Oha08b, Par06, RS10d, Ste03b, vB08b, Gau18]. havoc [Mil01, PCGD02, Goo02]. Haw [RS07c]. Hawking [PP08, Cro01, Lar05a, Lus04]. Haxel [Joh04]. Hazard [Hei10b]. Head [RS07g]. Hearing [Lam14]. Heart [RS07c, RS07g]. Heat [RA10]. Heaven [Eve12]. Heavens [NGZ23, Ryd13, HDS12]. Heavy [Bør12, Dah99, Wat01]. Hebrew [Unn00]. Hecht [Blo06, Ros01]. Heering [Ste03a]. Heidelberg [Stu06]. Heike [RA04, RA13]. Heilbron [Fri11, Kra16, RS11d]. Heinrich [Mul99, Mul01, Pip02]. Heisenberg [Bed09b, Cas05a, Lus10, Par06, RS07b, Car10, Cas09, Cas11, Cas15, FR11, Hol05, Lus10, Neu99, Ros98, Sch22]. Helen [Fra09a]. Helge [Cus00, Goo08, Hun01, PP10, Tri09]. Helium [Wes14, Gav01]. Hellyer [Hub06]. Helmholtz [D’A04, Pes13]. Helmut [Cas05a]. Help [RS10f]. Helping [Ano11d]. Helpless [Hon00]. Helsinki [Hol04]. Henri [RA10]. Henrik [RS11b]. Henry [Bru03, Dor06, Jam00, Ram05]. Hentschel [Ber08b, Sta06a, vB12]. Her [Buc20]. Herbert [Har02b]. Herbig [Cas05a]. Heritage [Dah06b, Joh06b]. Herken [vB02a]. Herman [PCGD02, Alp12]. Hermann [Pes09, Ryc10]. Heroes [RS06d, Ari03, Ari05]. Herschel [Ryd13, HDS12]. Hertz [D’A00, D’A04, Gea14, Mul99, Mul01, Pip02].

Hevesy [Stu05]. Hevesy [?]LorándStuewer:2005:GHL. Hey [DeV00, Sty05]. Hidden [Far09b, Lei08, RS07c, RS10c, RS08c]. Hide [Mal12, PCGD02, Cus02]. Higgs [Ber16, Hen07a]. Highest [RA13]. Highlights [Hen05a]. Hilbert [Row01, Sch03]. Hill [OV22]. Hirsch [Cas05a, RS08a]. Hirsch-Heisenberg [Cas05a]. Hirshfeld [RS09a]. Historian [Hu12, Huf12]. Historians [Gea02, Kel07, Bed08]. Historical [Bro13, Cas11, Cha02a, Hen99b, Hof99, Hun01, Kra96, PA01, HRS00, Ste03a]. Histories [Nye19, Bru02]. historischer [HRS00, Ste03a]. History [Ano22, BS11, Ber03, CP15c, CMP17b, CMP18a, CMS19a, CMS20b, Cro99, Cus00, Dom05, Eis06, Eme11, Fle21a, Fre00, Gea99, GG00, Goo08, Han04, Hec07, Joh04, Joh09, Kos13, Kra19, LoS16, Mar03a, Oha08a, Pas12b, Ram99, Ryn99, Sch14, Sim04, Sim10b, Sim12a, Tor02, Tri09, VV14, Wie18, Wil99, Yeh05, vB99, vB08b, Bro99, Cas17, Dah97, Far09a, Far10, Fis10, Ful00, HRS00, Kra99, Kra07, Mal11, Mar04a, MW03, New07, Nye08, Por09, Pul98, Smi98, Sol06, Ste03a, vB02b, DS22, Fra10, Rob18, Wes10]. Hit [Ano04b]. Hoax [Goo09b, Sok08]. Hoboken [Ger06b]. Hoddeson [Mod18, Seg09]. Hoffmann [RS12d, Sim08]. Hofmann [Ste17]. Hole [ORSSBO^{+}17, Tay10, Mel09]. Holes [AK99, Col00, DeW99, AK12, Mil05, WF98, Fri08]. Holmes [RS10e]. Holographic [Hec07]. Holography [Joh06a]. Holt [Bru03, Jam00]. Holton [Par06]. Holtzmann [Ger99a]. Honor [RA13]. Hooke [Wei04, BCHJ03, Inw02, Nau05, Wei04]. Hooper [RS09b]. Hopkins [Fel08, Goo12, Gre07b, Ham11, Oha10, RS08c, RS09c, Tat13]. Hore [Fre04]. Hoskin [Ryd13]. Hossenfelder [But19]. Houghton [Fri08]. Houses [Ano18a]. Howard [Coh00, Gou01]. Hubble [OO20, RS09g]. Huff [Fra12a]. Huggett [Pes11, Pri99]. Hugh [RA15]. Human [Goo12, MPFP19, vB99, KKG^{+}19, Osl10, Pul98, Seh19, SB19]. Humanities [Hie00, RS10f]. Humankomplements [MPFP19]. Humboldt [Roe10, vB10b, Rup05, Wal09, Wal11]. Humor [Hal12]. Hundert [Ren05b, Ren05a]. Hundred [RS06a, Mor90, Ren05b, Ren05a]. Hungarian [Bed12, Stu05]. Hunt [Ham11, Pes12, vB11, Sam10]. Hunter [LM06, Wei04]. Huygens [AM05, Sha06]. Hwan [RS08b]. Hydrostatics [Sal17]. Hypothesis [Mar04b].

I. [Day04, Gou01]. Iain [RS12d]. Ian [vB11]. ice [Ber02a, Jac04]. Icy [Rei10]. Idea [Cre08, Ruh20]. Identical [Ber08c]. Identity [van12]. Ideology [RS04a]. Ignorance [Fir12, RS13d]. II [Fre00, Rob18, Sch09a, Che11, Bad05, Bro99, Cas17, Cre05b, Cre09a, Goo03c, GF05b, Lip06b, LG08, OW13b, Pou01, Sim12b, Tis09b]. Ike [CMS20a]. ills [Ano18a]. Illuminate [Eve12]. Illustrated [Bai07, Oha08b]. Illy [Tat13]. im [Hen05b, HW08, Lem07, Ren05c, Sim08, Sta06a, HRS00, Ste03a]. Images [Mas07, van12]. imaginary [Mar04a]. Imaging [Ger99a, Kev97]. Imagining [RS06d, Ari03, Ari05]. Immigrant [CMP17c]. Impact [Bla00b, Wie18, Per99]. imperatives [Bro99, Fre00]. Importers [Gre21].

Inconvenient [Sil07, Sim10b]. indeterminist [Stö02]. India [Ban16, Bla00b, Per99]. Indian [PW19]. Induced [GLR06, GR09, GLR20, AGR04]. Industrialisation [RSD07, Gre09]. Industry [Ber08a, Ber10a, Hal09a]. Inevitable [Fra08a]. Infinity [Pes12]. Influence [GR09, HC11, Hu16]. Information [DG10, Jon07, RS10c, RS11b]. Informational [Bel14]. Informing [Bel14]. Ingenieur [Ren05b, Ren05a]. Ingenium [Den07, Fel08]. Inner [Eve00, Sim07, Bar99a]. Innovation [Buc16, Waz10]. Inquiry [Gre09, RSD07]. insider [Sta10]. insights [FGL^{+}12]. Inspector [Hen05b, Sta06a]. Instead [For22]. Institute [Bru99a, Cas11, Col00, Dom05, Fre00, Pip99b, Ryn99, Van03, Wat01, vB02b, Foc22, Hal12, Sch11, Sch21]. Institution [JP07]. Instruction [Hen99a]. Instrument [Ros08, Hen05b, Jac06, Sta06a]. Instrumentenbauer [Hen05b, Sta06a]. Instruments [Man22, Mas07]. Intellectual [Huf11, Fra12a]. intellectuals [RS08g]. Intelligence [Bør12]. Intelligent [Ehr06]. Intelligibility [RS07d]. Interest [Hen99b]. Interference [Rig07, Ros12]. Interpretations [Man23]. Interpreting [Mer01, Omn99a]. Interrogating [OM18]. Interview [Ano05a]. INTOR [Sta10]. Intriguing [Esp21]. Introduced [CdL16]. Introducing [Wal07]. Introduction [Hil11, Clo09, Mul10, Ryc10, RS10d]. Inventing [Kni06]. Invention [Dor06, Gre09, RSD07, Lin04]. Inventions [Mod18, Tat13, Ill12]. Inventive [Wei04, Inw02]. Investigation [KKG^{+}19]. invisible [Hen05b, Sta06a]. Inwood [Wei04]. Iranian [TT12]. Irish [Gre00]. Isaac [Bee15, BC01, Dom05, Dry14]. Isaacs [Hol01]. ISABELLE [Cre05a, Cre05b]. ISBN [Cas12, RS13d]. ISBN-13 [Cas12]. Isidor [Kri05]. Islamic [RS08d, Sal07]. Island [Wes15, Sch16]. Isolation [Ham00]. Isospondyli [SE19]. Issues [Par11, Lar07]. István [Sti08]. Italian [Man23, Shi06, Ber02b, Man22]. Italiana [Stu06]. Italy [Bon11, Esp21]. iv [Tri08]. Ivan [Ger06a]. Iwan [Dah06a]. ix [Eve08b, Ger06b, Goo09a, Hob02, Oha08a, Phi13, Tri07].

J [Ano18a, Bed06, Coh11, Dah97, DeV00, Eic07, Eic08, Eve08b, Goo12, Ham11, Hob04, Hu12, Lar05c, NHKL14, Noe09, RS06d, Ryn99, Sim10a]. J. [BS05, Cas05b, Dah97, Ger06a, Kle07, McM05, ORSSBO^{+}17, RS04c, RS08c]. Jackson [Ros08, Tei02]. Jacob [Bel14]. Jagdish [Ger03]. Jahrhundert [Hen05b, Sta06a]. James [Ham11, Lem07, RS08a, vB12, Hun10, Lem07, LH11, Ste17]. Jardine [Wei04]. JASON [IT23]. Jauncey [Jen02]. Jay [Ste03b]. Jayawardhana [Pas12a]. Jean [Oha08a, Yeh05]. Jed [Bru02, Dom05]. Jeff [Blo06, Ros01]. Jeffrey [Hol07]. Jenkin [Eme09]. Jennifer [Noe11]. Jensen [Joh04]. Jeremy [For08, Ger06a, RS10d]. Jersey [Ger06b]. Jerusalem [Unn00]. Jesuit [Hub06, Hel05]. Jewish [Bed12]. Jim [Wei04, Col00, Gre05]. JLab [Wes18, Wes19]. JLab’s [Wes14, Wes16, Wes19]. Jo [Sim04]. Joachim [Bed09a]. João [Seg10b]. Joel [Bri07]. Johann [Liu20, Sal20]. Johannes [LPR04]. John [Fel08, Ger06b, Kra16, Sch00a, Ber08c, Ber10b, DeW99, Ehr04, Eme09, Eve00, For01, Fra01a, Fri11, JS02, Mar03b, O’C10, Rei10, RS11d, Stu18, Tra08]. Johns [Goo12, Gre07b, Ham11, Oha10, RS08c, RS09c, Tat13]. Johnson [Per01, Wes10]. Johnston [Hec07, vB02b]. Join [Das19]. Joking [FLH85, RS06b]. Jones [Bed09c]. Jordan [JFEH15, Kra15]. Jose [Tri07]. Joseph [Dor06, Ram05, RS08a, Sch01, Tei02, Fix15, Jac00, Sch97, Vey13, WJS12]. Josephson [Bor11]. Joshua [Smi20]. Jost [vB12]. Journalist [FJ21]. Journals [Ste23]. Journey [Che11, TS01, Wat03, Giu10]. József [Tat13]. Jr [RS08a]. Jubilee [Hab10]. Judith [Hob04, Wat03]. Julian [Ger03, Gau18, MM00, Mil07]. Jung [Mil09, vB10a]. Junior [For20]. Jürgen [RS06a, Jon07]. Jury [Hol07, Cre06]. Just [Foc22].

Kai [Bed06]. Kaiser [Lep07]. Kamerlingh [RA04, RA13]. Karl [Rei19]. Karyotype [FEVvdB19]. Keay [Sch00a]. Keith [Lar05c]. Kelly [Bed08, Fri09]. Kelvin [Dor06, Lin04, Tra08]. Ken [RS09c]. Kennefick [Nal20, vB08a]. Kenneth [DeW99, Com05, Jac04]. Kerr [Tay10]. ket [D’A02]. Kett [RS08a]. Kevles [Ger99a]. Khalili [Col00]. Kimball [Ger03]. Kinematics [Mar09, Oha10]. king [Mor05, Dah06a, Esp21]. Kingdom [Hob02]. Kings [RS07b, Cla09]. Kirshner [Pas04]. Kiss [Lam14]. Klaus [Ber08b, Ber10b, Hof09, Sta06a]. Klein [Hal07, Hu12]. KLMN [Fri99]. Knew [Wei04, Inw02]. Knocking [Eve12]. Knopf [Bed06, Per01]. Know [Ber08d, ZZ18, Mul10, For08, Hil11]. Knowledge [Ano18a, Jha11, Kle23, Mon00, RS12g, Sha02, Sch06]. Kolb [Seg09]. Köln [Stu04, Tri08]. Komponente [MPFP19]. Koninkliijke [Gre09]. Kontext [Ren05c]. Kragh [Cus00, Goo08, Hun01, PP10, Tri09]. Krakow [Lal19]. Krauss [RS12h]. Kristine [PP08]. kritischer [Cas05a, HHH03]. Kuhn [Tor02, Ful00, Huf12]. Kuttner [Bed09a, RS07d].

L [Fri11, Lip07, LG08, Sha02]. Labinger [PCGD02]. Labor [HRS00, Ste03a]. Laboratories [Gre06, FG05]. Laboratory [For22, Hen99a, Rio00a, Wes06, AGR04, HRS00, Ste03a, Aus16, Bon18, Cre99, Nav06, Pip99a, Wil17]. Labyrinth [O’C14, Wal07]. Laidler [Lar05c]. Lake [Wei09, Mil10, Rhe02]. Lambert [Lal19]. Landmarks [Gab00]. Lange [Hof09]. Langen [Cas05a]. Langley [Loe03]. Language [RS06d, Ari03, Ari05]. Large [RS09c, Giu12, Lin09]. Lars [Car00]. Larsen [PP08]. Laser [Bri99, Hec10, Tow99, Blo06]. Last [Ram05, Rus15, Mor03]. Late [Gre09, Sal14, RSD07]. later [RS06h]. Latin [von09]. Lattes [VV14]. Laughlin [Hil07]. Laura [vB10b]. Laureate [Bor11, Jos10]. Laureates [Bru06, Dar04]. Laurel [NHKL14]. Laurie [RS06c]. Lauritsen [Hol03]. Law [Coo11, Cre09b, OO20]. Lawrence [Eme09, Hun04, Jen08, LM06, Neu99, Cas10, RS12h]. Laws [Bon11, DFAB19, Fra08a, RS11e, D’A02, Fer10]. lead [RS07h]. Leader [Hil11, Mul10]. Leading [Lus04, Ste23, Cro01]. Leads [But19]. leaps [Ber09, Sol06, WWH01, RS10d]. Learn [Ber02b]. learned [Ano11e]. Learning [CMS19b, Gre06, FG05]. Leben [Ren05c]. Lebensweges [RD05]. Lectures [CP14b, RS06b, FGL^{+}06]. Lederman [HK03]. Lee [RS08b, Goo03b, Goo03c]. Legacy [RS09a, BB04, Hir10, Stu06]. Legend [OM18]. Leibniz [KP02, Noe11, Coo10, Dav00b]. Leibowitz [RS08c]. Leiden [HK07, van12]. Leipzig [SS08]. Lemaître [Lal19]. Lemmerich [vB12]. Lenard [Mul99]. Length [Ess10]. Lenin [Bor11, Jos10]. Lens [Ban16]. Leo [Fra05, Hal09a]. Leon [HK03, Ste03b]. Leonardo [Wei04, BCHJ03]. Lepton [Per04]. Leslie [TF07]. Less [Ehr06]. Lesson [RS09g]. Lessons [Hab10]. Lest [RS05d]. Letter [Sch22]. Letters [Bro06a, Cas05a, Fey05, HHH03]. Leventhal [Bed08]. Levine [Hof17]. Levitt [Per10]. LHC [Che11, Giu10]. Liberal [RS13e]. Liberalism [Lin10, Coe07]. Liberty [Hal09b, RS11e, Fer10]. libres [PMAS19]. Liebe [Cas05a, HHH03]. Life [Ber08a, Ber10a, Bor11, Buc20, DeW99, Far09b, For20, For22, Hal09a, Hal09b, Hin12, Hon00, Jam00, Kra11b, Lal19, LM06, Lus04, Lus10, Man23, Mod18, NW08, Pas12a, Pav11, Pip99b, RA15, Rei07, Ren05c, RS06a, RS09a, RS10c, RS12h, Sch01, Seg10b, Seg15, WJS12, Wei04, Wie17a, vB12, BCHJ03, Clo15, Coe07, Cro01, Dav00a, FG05, Hir10, Hun04, Inw02, Jay11, Jos10, LH11, Mag09, Pou99, RD05, Sch97, WF98, Car07, Gre06, Kle07, Lin10, RS07c, Sch00a, Seg15]. Lifetime [Har02b]. Light [Bed09a, Cas02, Cre08, Cre09a, DMA07, Ess10, Ham11, Hec04, Hun04, Kra18, Mou20, Pip01, Ros01, Hun10, NHW08, vB02b, LM06]. Like [CMS20a, FJ21, RS13h]. Lillian [Mod18, Seg09]. Limerick [Ano99b]. Limit [Tri05]. Limits [DeV00, Hey99]. Lincoln [RS09c]. Lindley [Dor06, RS07b]. Lines [Kle11, Wes17, Goo10]. Linus [Bad09a]. Lion [Noe09]. Liquefier [Wes14]. Liquid [Ree11, Gav01]. liquide [PMAS19]. Lisa [Wei04]. Lisbon [SDC12]. Lissa [Gre09]. Literacy [GMS^{+}03, HKT02, Hob04, RS08a]. Literary [Bed09b, Lus10]. Literature [Can00, Coh11, For01, Kra16, Gol10, Gol12, Mor90]. Little [BF13, Maa07]. Littmann [RS09c]. Liverpool [Hu19a, Hu19b]. Lives [Wie17b]. LLC [Jam00]. Local [Ano16h]. Lockwood [Wal07]. Lodestone [Gri06, Gui05]. Logic [Kip00, Lar05c, Lai04]. Lomonosov [CS13, CS18]. London [Ahe02, Car11, Col00, Com05, Dah06a, DeW99, Fra07, Fre04, Goo09a, Gre07a, Lin10, Neu99, Par06, Pip99b, Tei02, Wei04, vB08b, BCHJ03, Cre17, Wei04]. Long [Fre05b, Mor90]. Longbourne [RA15]. Look [Ghi05, Shi06, Mor90]. Looks [FJ21]. Loránd [Stu05]. Lord [Tra08]. Lorentz [D’A00, Jan02]. Lorenz [Kra18]. Loschmidt [Cre09b]. Loss [Wes17]. Lost [Eis06, Oha08a, Oha10, Sta01, Mar09, But19]. Louis [Ess10, Fre00]. Love [vB01, Ove00, Ove01, Kra16]. Loves [Mal12, PCGD02, Cus02]. low [CFL19]. Lowest [RA13]. Ludvig [Kra18]. Ludwig [Bla00a, Boy12, Cer98, Rei07]. Luisa [Stu06]. lure [RS02a]. Lydia [Pas10]. Lyman [NHKL14].

M [Ber08b, Bro06b, Ehr04, Fel11, Gib05, HK03, Jen02, Nye99, Ste03b, vB12].

M. [Das15]. MA [Hol19, Mod18]. Mach [dWtH20]. Machine [Maa07, Sal14, Hen05b, Sta06a]. Machines [AK99, Fel08, AK12, Den07, Col00]. Macmillan [Bri07, Wei04]. Macmillian [RS07c]. Madison [For01]. Madrid [SR06]. Magic [Hun15]. Magnetic [Gri06, Pav10, Ram99, Gui05]. Magnets [Wil17]. Magueijo [Seg10b]. Mahishyas [MDK19]. Majid [RS09d]. Majorana [RS09a, Seg10b, DE07, DMA07, GR08a, GR08b, GR13, Mag09]. Make [Blo06, Hec10, RS01b]. maker [Hen05b, Sta06a]. Makers [Ros08, Jac06]. Makes [Ehr06, RS07d]. Making [Cre99, Das19, Man22, Par00, RS08d, Bel99, Sal07, Rio00a]. Malin [PCGD02]. Malley [Sim12a]. Malsbary [Shi06]. Man [Ano99a, Bla00a, Fac82, Far09b, Kra11b, Mod18, RS09a, RS10c, RS12h, Sha06, Wei04, AM05, Cer98, Hir10, Hol03, Inw02]. Manhattan [Bed08, Kel07, Ree09, Ree11, Ree14, SS10, Vey13]. Manipulated [RS12e]. Mann [Per01, Joh99a]. Manufacturers [Gre21]. Manuscript [O’C10]. Manuscripts [Sal17, Dry14, Bee15]. Many [Sil05, MG21]. Mapping [NGZ23]. Mara [Par00]. Maracineanu [Buc20]. Marage [vB00]. Marc [Bru03]. March [Rig11]. Marcus [Hub06]. Margaret [Goo12]. Maria [Cas05a, Hug09, vB10c]. Marić [Hol19]. Marie [Sim07]. Marietta [RS03e, Sim13, SRD06, Stu04, Tri08]. Marine [Ham00]. Mario [Mas07]. Marius [Gra10]. Marjorie [Sim12a]. Mark [Sim08, Cre12, Fel08, RS07c, RS09c]. Mars [Smi20]. Marshall [Hob04]. Martians [Sti08, Har06]. Martijn [Sim10a]. Martin [Bed06, Hu12, KP02]. Martinez [Oha10, Seg11a]. Mary [Bed01a, Sim04]. Maschinen [Hen05b, Sta06a]. Maschinen-Inspector [Hen05b, Sta06a]. Mason [Sim10a]. Mass [Bor11, Car11, Cas10, Coh11, Com05, Dom05, Goo09a, Gre07a, Jac04, Kah03a, Mil10, New99, New04, Oha08b, Par06, Pri99, Ros08, Smi20, Ste03b, Tei02, vB08b]. Massachusetts [Bru02, For02, Wat03]. Massive [Sam10, vB11]. Master [RA04]. Masters [War03, Har05]. Material [Gau18, Mar04a, Kle07]. Materials [Hug09, Ree14, vB10c, Ren08]. Math [But19]. Mathematical [Har05, New04, Sic06, Nye08, Pes03, War03]. Mathematics [Ber03, Cre12, Eck99, Hon00, Row01, Ryc10, Ari03, Ari05, Mic06, Pes09, Pet11, RS06d]. Matricon [Yeh05]. Matter [Ano16e, For09, MG21, Pav10, RS07c, RS09b, RS10b, Ana08, Dav10, Hoo08]. Matthew [Nyc09]. Mauro [Bru06]. Max [Cas12, Kah03b, Seg11b]. Maxwell [Cha01, D’A00]. Maxwellian [Kra18]. May [CdL16]. Mayer [Joh04]. McClelland [O’C10]. McMillan [RS06d]. Mean [Jen11]. Meaning [EG05, Goo09a, Gre05, Hol99, New04, RS06a, Bag04, Pes03, Sch08]. means [Zim02, Fra01a]. Measure [Cra00, Gre05, Bag04, Cha19]. Measurement [Gau18, Jha11, Kni06]. Measurements [Buc20, vB02b]. Measuring [Che01]. Mechanics [Cro07, Dir09, Gau18, Ghi05, Gre07b, Omn99b, RS12f, Sal17, Fol05, FR11, Hen05b, Her03, PFD08, Sch06, Sta06a, Kob02, Noe09, Shi06, Bro06b, Mul01]. Mechanicus [Hen05b, Sta06a]. Mechanism [Cha01]. Medal [Tra08].

Media [RS10h, Smi20]. Medical [Ger99a, Kev97]. meek [RS06g]. Meeting [Van07]. Meets [Row01]. Megascience [Ano16g, KW08, Pro16, Seg09]. Mehra [Ger03]. Melba [NW08]. Melbourne [Hom05]. Meli [Gre07b]. Melia [Tay10]. Members [Das19]. Memoir [Bed01b]. Memoirs [TS01, Wat03]. memorable [Mor90]. Memoriam [TT12, Fre08, JP99, Pai99, Rei07, Stu18, Tra08]. memory [Sim06]. men [Mul99]. Mental [Ber08b, Hen07b]. Mentality [Ber08b, Hen07b]. Merli [Ros12]. meson [VV14]. Messenger [LM06, Hun04]. metabiography [Rup05, Roe10]. metaphysical [Cha02a]. Metaphysics [Jac14, DG10, RS11b]. method [Seh19]. mew [VMR19]. Meyer [Rei01a]. Meyerstein [Hen05b, Sta06a, Hen05b]. Michael [Hob04, Pip99b, Wei04, Can01, Day99, Fra12b, Kah03a, Noe09, Rus00, Ryd13, Wal07]. Michelle [Ano05a, Bro06a]. Michigan [Aus16]. microphysics [Bru02]. Microtus [PS19]. Microwave [Alp12]. Mid [Wes18, Wes19]. Mid-1980s [Wes18]. mid-1990 [Wes19]. Middle [DFAB19, Goo12, Osl10, RS07c]. Middle-Aged [DFAB19]. Miedema [Sha06]. Mifflin [Fri08]. Migrations [Bed12]. Mikhail [CS13, CS18]. Miles [SE14]. Mileva [Hol19]. Military [Fre00, Moo08, Bro99]. Mill [CS18]. Millennium [Bru99a]. Miller [Fri08, PCGD02, vB10a]. Millikan [Hol99]. Mills [Rig11]. Milton [Ger03]. Mind [MG21, Pes09, Ryc10]. Mindful [Gre09, RSD07]. Minnesota [Rhe02]. Minor [CMP18c]. Mirror [van12]. Mis [Kel22]. Mises [SS07]. Misrepresentation [Buc16]. Miss [ORSSBO^{+}17]. Missing [Fra09a, QN08, vB11]. Missionary [Liu20]. Missiroli [Ros12]. Mistaken [van12]. Mistakes [Ano00b]. Mix [Ano01c]. Model [GR08b, Joh04]. Modern [Bar03, Eve12, For09, Gol99a, Goo12, Gre05, Hub06, Nye19, Pes15, RS06d, RS07b, Sim04, vB00, Ana08, Bag04, Cla09, Coo99, Die05, Hel05, MW99, McM05, Nye08, Osl10, Sch06, Wei15, Gib05]. modernen [RS03e, Stu04, Tri08]. Molecule [RS13g]. Møller [Kra22]. Money [Ahe02, Gre01]. Montgomery [Sha02]. Moon [PP10, Kra08a]. Moonshine [Jen11]. Moore [Fri09]. Moral [Fra01b, Vey13, Sch00b]. Morality [Buc16]. Moritz [Hen05b, Sta06a, Hen05b]. Morris [Ram05]. Morrison [Fre08, Mor90]. Morus [Dah06a]. Mössbauer [Wes06]. Most [Ger06b, Bri05, Eme09, Jen08]. Mother [TT12]. Motion [Sal14]. motu [Sal17]. Mountain [Ger03, Hun15, NHKL14, MM00]. Movements [Mon00, Moo08, Sha02]. moves [Gra08]. Movies [OT10]. Mr [FLH85, Hob02, RS06b, RS11c]. Mr. [FR11]. MRI [Ber16]. Much [BF13, Fis10, Wei04, Inw02, Eme11]. Muller [Hil11, RS08d, Cas05a]. Multiple [Fra05, Kel22, TF07]. Munich [Cas05a, TEW02]. Muon [Jac10]. Muon-Catalyzed [Jac10]. Murray [Per01, Joh99a]. Muse [Can00, Cre12, For01, Pet11]. Museum [Sch09a, Sim10b, Pip99a, Sei00]. Museums [Rhe02]. Music [CP14c, Fix15]. Musicians [Ros08, Jac06]. Muslim [Alf09, Fra08b]. Muslims [MDK19]. My [Hin12, Jac09, RS01d]. Myles [Ros08, Tei02]. Mysteries [Ghi05, Shi06]. Mysterious [Seg10b, Sim12a, Mal11]. Mystery [Fra09a, QN08]. Mystic [Far09b, Nyc09, RS10c, Sta07]. Myth [vB06, Par07]. Myths [Goo08, Tri09, Kra07, Mar11, Seg11a].

N [Bri07, Das15]. naam [Eic07]. Nadia [RS09a]. Naked [Kev97, Lan12, Ger99a]. Naked-Eye [Lan12]. NAL [Pro16]. Nall [Smi20]. Name [Mar15, Eic08, Sim10a]. Napier [Hen07a]. Naples [Sch09a, Esp21]. nascita [Ber02b]. National [Bon18, Cre99, Rio00a, Wil17, Cre08, Cre09a]. Nationalism [Hol09, RS07i]. Natural [CMS19c, Dom05, Gre20, Hub06, Joh04, BC01, Hel05]. Nature [Cha01, Goo12, Hoo08, Kat05, Kra19, Mal12, New99, RS04d, RS07c, RS10b, RS11b, Ryc10, DG10, Dav10, Fer10, Fla98, Osl10, Pes09, PCGD02, RS09b, RS11e, Cus02]. Nauheim [Van07]. Nazi [Neu99, Sim10a, Eic07, Eic08, Ros98]. nazi-Duitsland [Eic07]. Near [Bad09a]. Near-Appointment [Bad09a]. Nearest [GP01, Ste03b]. Nebula [Wee10]. Nederlandse [Gre09]. née [Pip02]. Need [Ber08d, For08]. needed [RS08g]. Needs [Hil11, Mul10]. Nei [Cha19]. neighbor [Sch03]. Neil [Bru99a]. nel [Ber02b]. neonates [CFL19]. Nerds [MM23]. Netherlands [Kle23]. Neumann [Ber10b, Sch04]. Neutrino [Arn01, Fel12, LoS16, Nav06, RS11e]. Neutron [GLR06, GR09, GLR20, Rus15, AGR04]. Neutron-Induced [GLR06, GR09, GLR20, AGR04]. Neutronics [GORS09]. Neutrons [GLR06]. Newell [NW08]. News [Smi20]. Newton [Bee15, Dom05, Dry14, vB08b, BC01, Dry14, FR11, Jac14]. Nick [Pes11, Pri99]. Nicolaas [Roe10]. Niels [RS11b, Kle10, Sch11, Sch21]. Night [Lan12]. Nimtz [Bed09a]. Nina [Eve08a]. Nineteenth [Gre20, Hie99, Ros08, Har85, Gre21, Jac06, LR03, Pur97, Wee10]. Nineteenth-Century [Ros08, Gre21, Jac06, LR03, Wee10]. Niobium [Ber16]. Nir [Fra09a]. NJ [Hie99, Nal20, Yeh05]. No [RS04b, Nal20]. Nobel [Alf09, Bru03, Bru06, Dar04, Fra08b, Fri01, Goo03a, Har02a, Kur02, Nye19, PA21, RA13]. Noble [Eme11, Fis10]. Non [Kra18]. Nordström [Hal04]. Normale [RS09a]. North [Fol05, CdL16, SB19]. North-Brazilian [CdL16]. Norton [DeW99, KP02, RS06b, RS08d, RS12h, Sim07, Tay05, vB10a]. Norway [Kle23, Wit07]. Norwegian [Bor08, Bør12]. note [Cha19]. notebook [AGR04]. Notes [Phi13, RS06a, RS06b, RS06c, RS06d, RS07a, RS07b, RS07c, RS07d, RS08a, RS08b, RS08c, RS08d, RS09a, RS09b, RS09c, RS09d, RS10b, RS10c, RS10d, RS10e, RS11b, RS11c, RS11d, RS11e, RS12h, RS12d, RS12b, RS12c, RS13c, RS13d, RS13a, RS13b, Sal14, And11]. Nothing [Clo09, Eme11, Fis10, RS10d]. Notice [RS12e]. Notre [Gib05, Hub06]. nourrissons [PMAS19]. Novosibirsk [KA15, KA14]. NSTA [RS08a]. Nuclear [Bad01, Ber08d, Bla00b, Bør12, Can00, Cas02, Cas10, Dah99, For01, For22, Gol16, GORS09, Hin12, Joh04, Kle23, Kra00, Pav10, Ree14, RS03a, Ruh20, Seg10b, Seg13, Sim00, Sim10b, Sim12b, Sim13, Van03, VV14, Wat01, Wie17b, Wie18, Wit07, Bad05, Bad09b, Ber02a, Dah02, Mag09, Mil07, Per99, Tur12, For08]. Nuclear-Fission [Sim10b, Sim12b].

Number [Mil09, vB06, vB10a, Par07]. Numbers [Gre12, D’A02]. Nussbaumer [Pas10]. NY [Lan10]. Nye [Sim04].

Object [Lam14]. Objectivity [Fra09b]. Objects [Gre07b]. Observation [OR21, Fra17]. Observations [Gra10]. Obsession [Fri08, Mil05]. Obsessive [Sim07]. Occupation [Sch21]. Occupational [Hei10b]. Occupied [Sch11]. odd [Bad09c]. Odds [Gol16]. Odyssey [Bee15, Che11, Dry14, Giu10]. Old [Man23, Roe12, Wei11]. Oldenburg [Ste03a]. Olival [For20]. Olivier [Fre05a]. Omnès [Kob02, Mer01]. Once [Hon00]. One [PCGD02, Ren05b, Ren05a, RS13g, Day02, LC01, RS06a]. Onion [FJ21]. Onnes [RA04, RA13]. Opening [Sei00]. Opera [Hin12, Lam14]. Oporto [dP07]. Oppenheimer [Fra01b, Ger06a, Goo09a, Kle07, RS06d, Bed06, Ber04b, BS05, Cas05b, Ger06a, McM05, ORSSBO^{+}17, RS04c, Sch00b, Sch08]. Opportunity [Kip00, ORSSBO^{+}17]. Opposition [Kra11a, Nye99]. Optical [Per10, Lev09]. Optics [Hec04, Jac00, Ros01, Tei02]. Oracle [Car00]. Orbital [Nau05]. Order [Lan10, Cal09]. Orderly [Pes12]. orders [SE19]. Ordinary [Seg11b, Cas12]. original [BV09]. originally [PS19]. Origins [Cas12, Hof17, Kat05, McM02, Oha10, Mar09, Seg11b, Sta01, Ste05, SL16]. Osler [Goo12]. Oslo [Ped11]. Ostariophysi [SE19]. Other [RS06b, Seg11a, FL88, Mar11]. Otto [BF13, Har02b, Sim06, Sim12b, Wal03a, Wal06]. Ought [ZZ18]. Our [Gre12, Pas12a, RS05c, Tor02, Ful00, GP01, Jay11, Ste03b]. Outer [Eve00, Kah03a, Bar99a, Fri00]. Overbye [vB01]. Ovshinsky [Mod18]. own [Pip99b]. Oxford [Alf09, Bed09a, Bed09c, Bee15, Ber08b, Bla00a, Blo06, Bri99, Can01, Che11, Das18, Eme09, Eme11, Eve00, Fel11, Fra10, Fre05a, Fri11, Ger03, Gol99a, Goo08, Goo09b, Gre05, Gre06, Hec07, Hol01, Kle07, Kra16, LM06, Lus04, MG21, Noe11, PCGD02, Per10, Pes11, RS06d, RS07d, Rig07, RS09c, RS10b, RS10c, RS10d, RS11e, RS12d, RS13d, Ros01, Shi06, Sim12a, Sob08, Sti08, Tri09, Wal07, Wei04, Wes10, vB99, FG05].

P [Bro06a, Fey05, Gri06, Kle07, Pas04, Rio00a, Sim10a, Stu05]. P. [Eic07, Eic08, Nye99]. Pace [Hob04]. Pacifist [Sal19]. Pages [MG21, Ahe02, Alf09, Ano18a, Aur04, Bed01a, Bed06, Bed08, Bed09a, Bed09b, Bed09c, Bee15, Ber08b, Bla00a, Bla00b, Blo06, Bor11, Bri99, Bri07, Bro06a, Bru99a, Bru02, Bru03, Bru06, But00, But19, Can01, Car07, Car11, Cas05a, Cas10, Cas11, Cas12, Che11, Coh11, Col00, Com05, Cre12, Cus00, Dah06a, Das18, DeV00, DeW99, Dom05, Eme09, Eme11, Eve04, Eve08a, Eve08b, Fel08, Fel11, For01, For02, For08, For09, Fra01a, Fra04, Fra07, Fra09a, Fra10, Fra12a, Fra01b, Fre00, Fre04, Fre05a, Fri08, Fri09, Fri11, Gea99, Ger03, Ger06a, Ger06b, Gib05, Gol99a, Goo09a, Goo12, Goo08, Goo09b, Gre07a, Gre06, Gre07b, Gre09, Gri06, Hal09a, Ham11, Han04, Har05, Hec07, Hie99, Hil07, Hil11, Hob02, Hob04, Hol01, Hol09]. pages [Hol07, Hol19, Hub06, Hun01, Jac03, Jac04, Jam00, Jon07, Kah03a, KP02, Kle07, Kle11, Kob02, Kra16, Kro10, Lal19, Lan10, Lar05c, Lep07, Lin10, LM06, Lus04, Lus10, Mar03b, Mas07, Mer01, Mil10, Mod18, Nal20, Neu99, New04, Noe09, Noe11, Nyc09, Oha08a, Oha08b, Oha10, Par00, Par06, PCGD02, Pas04, PP08, PP10, Pas10, Pas12a, Per10, Per01, Pes11, Pes15, Phi13, Pip99b, Pri99, RS07b, Rig07, RS09a, RS09b, RS09d, RS10b, RS10c, RS10d, RS10e, RS11b, RS11e, RS12d, RS13d, Rio00a, Rob18, Roe10, Roe12, Ros01, Ros08, Ryc10, Ryd13, Ryn99, Sch00a, Sch01, Sch05, Seg09, Seg10a, Seg11a, Seg13, Seg15, Sha06, Sha02, Shi06, Sil07, Sim04, Sim07, Sim08, Sim10a, Sim12a, Smi20, Sob08, Sta06a, Ste03a, Ste03b, Sti08, Stu04, Stu06]. pages [Sty05, Tat13, Tay05, Tay10, Tei02, Tor02, Tri07, Tri08, Tri09, Van03, Wal07, Wat01, Wat03, Wei04, Wes10, Yeh05, vB99, vB00, vB01, vB02a, vB02b, vB06, vB08a, vB08b, vB10a, vB10b, vB10c, vB11, vB12]. Painstaking [Wes18]. Pais [Kle07]. Palace [Ano18a]. Palevsky [Bed01a]. Palmisano [Hob04]. Pancaldi [Aur04]. Panisperna [Foc22]. Pantheon [RS10e]. paper [Bed09a, Bed09c, Car07, Dah06a, Dom05, For01, Fra12a, Fre05a, Gea99, Ger99a, Goo12, Gre05, Gre07b, Ham11, Har05, Hie99, Hil07, Hob02, Hol01, Hun01, Kah03a, Lep07, Noe09, PCGD02, PP08, Phi13, Pip99b, Pri99, RS06d, Rio00a, Roe10, Sil07, Sim07, Sim10a, Sty05, Tay05, Tri08, Wes10, Yeh05, vB02a]. Papers [Bee15, Dry14]. Paradigm [OO20]. Paradoxes [RS03c]. parallax [Gra08]. parents [Cas05a, HHH03]. Parganas [MDK19]. Paris [Lam14]. Parisian [Gab00]. Park [Sch01, vB06]. Parliament [Ano18a]. Part [Cre08, Cre09a, Goo03b, Goo03c, GF05a, GF05b, Lip06a, Lip06b, Lip07, LG08, OW13a, OW13b, Tis09a, Tis09b, Wes18, Wes19]. Partial [FEVvdB19]. Particle [Ana08, Per04, RS07c, Sim13, Tri08, vB11, RS03e, Sam10, SRD06, For09]. Particles [Kah03a, RS13g, Fri00, Wei83, Wei90, Wei03, Bro04]. Partisans [Sal19]. Pasachoff [Ste03b]. Pascual [Kra15]. Passage [vB10b, Wal09, Wal11]. Passion [Bed09c, Jon08, Rei07]. Past [FJ21, Hab10]. Patents [Tat13, Ill12]. Paternity [Gor09]. Path [Ram05, Eve06, Mor03]. pathway [RD05, RS06a]. Patricia [Fra10]. Patrick [Sty05, Fre04, Hor03]. Paul [RS10c, RS11b, Bor11, Bro06b, Far09b, Ger99b, HK07, Kle10, Neu99, Sch14]. Pauli [vB10a, Mil09]. Pauling [Bad09a]. Peace [Hol09, Sal19, Day08, RS07i]. Peaceful [Sal19]. Peaks [Rei10]. Pearson [RS06b]. Pedagogy [Gau18]. Penguin [RS06d]. Pennsylvania [Sch01]. People [RS06b, FL88]. Percussion [Sal14]. Perfectly [Fey05, Bro06a]. period [HHH03]. Periodic [Ram05, Sce07, Sob08, Mor03]. Peripatetic [Hen05a]. Peripheral [Das15, Das19]. Periphery [CMS19b, KG19]. Perkovich [Bla00b]. Perseus [DeV00, Ger99a, Wat03]. Person [LoS16, Wes17]. Personal [Bed01b, Jac10, NW08]. personalities [Ano04c]. Perspective [For22, Mal12, Huf11, Fra12a, PCGD02, Tri08, RS03d, RS06e, Cus02]. Perspectives [Gor09, Lar07, Par11]. Perspectivism [Fra07]. Pesic [Gre07a, New04, Ryc10]. Peter [Gre09, Mod18, Ryc10, Sim10a, CS18, Fre04, Gre07a, Hal05, New04, Pip99b, RS07d, Ste03a]. Peterson [Cre12]. pH [Seh19]. pH-dependent [Seh19]. Phenomena [Wei08, Gui05, Gri06]. phenotype [AR19]. Philadelphia [Bru99a, Fre00, Hal09b, Ryn99, Van03, Wat01, vB02b]. Philip [MG21, Fre08, Mor90, Phi13, ZZ18]. Philipp [Mul99, SS07]. Phillips [NW08]. Philosopher [Pip99b, Day99]. Philosophical [Gre21, Tor02, Ful00]. Philosophy [But00, CMS19c, Cro99, Das18, Goo09b, Gre05, Gre20, Hag12, Hub06, Mer01, NMS22, Omn99a, Ryc10, Tor99, Bag04, Bar99a, BC01, Fol05, Hel05, Hug10, Mic06, Pes09, PFD08, Sok08, Sta01, Dom05, Eve00, Pes11]. Philosophysics [DS22]. Photographic [Sim13]. Photography [Hin12, Joh06a]. Photons [Pou00, Pou01]. Physical [Bed03, BP08, Buc16, DS13, Gre00, Hen07a, Hof00, Hof05, HW08, Hol04, Hom05, Hun15, Joh06b, KW10, Kov03, Lac07, Ped11, PA01, Rei01b, Rhe02, RS09i, RRR13, SR06, SS08, Sim04, Sim08, SDC12, Str04, D’A02, D’A04, Nye08, Sta06b, KG19]. Physicist [BS04, Eve08b, Gol99b, Gol16, Hal05, Hal18, Hu12, IT23, Nye99, O’C14, OW13a, OW13b, Seg15, Tis09b, Wil00, von09, Clo15, Eve06, HK07, Lem07, Wei07, Tis09a]. Physicists [Ber08b, HR01, Jan11, Lus04, RS05b, RS07g, RS09g, RS10a, RS12f, Ros08, Sal19, Sim08, Sti08, TEW02, Bad05, Cro01, Har06, Har85, Hen07b, Jac06, Mul99, Bru99a, HW08]. Physics [Ano16e, Ano16h, Ano22, Bal08, Ban14, Ban16, Bar03, Bed03, Bed09a, Bed09b, BS11, Bel14, BP08, Bør12, Bru99a, But19, BW06, Can00, Can01, Cha02b, Che11, Col00, Com05, Coo11, Cre99, CP14c, CP14d, CMP17c, CMP17d, CMP18a, CMS19a, CMS19b, CMS19c, CMS19d, CMS20b, Cro99, Cus00, Dah06b, Dah06a, DS22, DE07, Eck99, Eve12, Fle21a, Foc22, For01, FG05, Fra08a, Fra16, Fra17, Fra09c, Fre00, Gea99, Gib05, Gin10, Gre05, Hag12, Hal04, Hal09b, Hal12, Ham11, HC11, Hen07a, Hil07, Hil11, Hin12, Hof99, Hof00, Hol04, Hom05, Hon03, Hon00, Hu16, Hub06, Huf12, JS02, Jam07, Joh99a, Joh06b, Kat05, KG19, KW08, KA14, Lus10, Mal12, Mar03a, MNS22, McM02, Mul10, MG21, NMS22, NMS23, Nye19, OV22, PCGD02, Pas12b, Per04, Pes11]. Physics [PFD08, Pip99b, Pip02, Pur97, Rei10, RA13, Rhe02, RS01c, RS02c, RS03b, RS04b, RS05b, RS06b, RS06c, RS07c, RS07d, Rig07, RS08c, RS10h, RS11b, Rio00a, Row01, Ryn99, SR06, SS08, Sch09b, Sch16, Seg09, Sim13, Ste23, Str04, TEW02, Tor99, Unn00, Van03, Wat01, Wei08, Wei17, Wit07, dWtH20, dP07, vB02b, Ana08, Bad09c, Bag04, Bro02, BL06, Cas09, Coo99, Cus02, Dar04, DG10, FGL^{+}06, FGL^{+}12, For04, Giu10, Har85, HRS00, Hel05, Hug10, Hun10, Isa00, Kra99, Lei08, MW99, Mel09, Mic06, Mil07, Mor05, New07, Pes09, RS02b, RS06h, RS08h, RK11, RS03e, Ruh20, Rus00, Sch03, Seg07, Smi98, Ste03a, SRD06, TF07, Tur12, War03, WF98, Bru06, But00, DeW99, Ehr04, Eve08a, For09, Gol99a, Gre06, Hal12]. Physics [Har05, Hie99, Hil11, Hol01, Jac03, KA15, Lep07, Per01, RS05a, RS06b, RS07a, RS08b, RS08d, RS09d, Ryc10, Sch09a, Seg13, Tay10, Tri08, vB00, vB08b, RS07c, RS03d, RS06e]. Physikalische [HW08, Sim08]. Physiker [Lem07, HW08, Sim08]. Physikgeschichte [HRS00, Ste03a]. Pi [Ger06a, VV14]. Pi-meson [VV14]. Picasso [Goo02, PCGD02, Mil01]. Pierpont [Loe03]. Pierre [Cro99, Kra08b, vB00]. PiL [VMR19]. Pioneer [LKK04, Rei01a, Sim13, Tri08, Coh00, RS03e, SRD06, Gou01]. Pioneering [Hof02]. Pisa [RS09a, Gam03]. Pittsburgh [Eve04, Hof17, Seg11a, Smi20]. Planck [Bad09c, Gea02, Hol99]. Planetary [Kos13, Sil07]. Planets [Pas12a, Jay11]. Plans [Wes18, Wes19]. Planting [Sch03]. Plasma [RS10e, Wei08, Wei17, SBD09]. Plateau [WJS12]. Playful [BS04, Hal12]. Playing [Gau18]. pluralism [Bok08, Kro10]. Pluribus [CMS21a]. Poincaré [Kat05, Mar04a]. Point [CdL16]. Points [Hen99b, Kos13]. Polanyi [Fra12b, Jha11]. Polanyian [Jha11]. Pole [Ano04b]. Poli [Esp21]. Policy [Ano10a, Mar14]. Political [Ahe02, Bad01, Gre01, Per10, Goe03, Lev09]. Politics [Ahe02, Bru03, Buc16, Cas10, CMP18d, Gre01, Hol09, Moo08, Sim12b, TS01, Bad09b, Day08, Fri01, PFD08, RS07i, Sim06, Wat03]. Pollen [NHKL14]. Polonium [Buc20]. Polonium-210 [Buc20]. Polymath [CS13, Har01]. Polymorphism [MPFP19, SB19, VMR19]. Polymorphismus [MPFP19]. Polyphony [BF13]. Polyploidization [PS19]. Pomor [CS13]. Pomorie [SE14]. Pontecorvo [Seg13, Seg15, Clo15, Tur12]. population [AR19, Cha19, PS19, VMR19]. Porter [Bru99a, Han04]. Portrait [Ger06a, Gou01, Ber04b, Coh00]. position [HK07]. Positron [Ber04a]. Possibility [Par11]. Post [Sal17, Ste23]. Post-Aristotelian [Sal17]. Post-Soviet [Ste23]. Posttransferrin [MPFP19]. Posttransferrins [MPFP19]. Postwar [Lep07, Wei08]. Potential [Hon00]. Pound [Pav10]. Pounds [Ano01b]. Poundstone [Jam00]. Power [Che01, Ham11, Nye99, Ber02a, Hun10]. Pozzi [Ros12]. pp [Eve00, For20, Ger99a, Gou01, Gre05, New99]. Practical [Par11, Sta07, Tat13, Ill12, Nyc09]. Practically [Eme11, Fis10]. Practice [Car11, Sic06, FGL^{+}12, Goe03, Set10]. Practices [Hug09, vB10c, Ren08]. Pre [DS22]. Pre- [DS22]. Precision [Jac00, Tei02]. Predictable [RS09g]. predominantly [PS19]. Presented [Mul01, Her03]. Preservation [Van07]. Presidential [Her00, vB02a]. Presidents [Hil11, Mul10, RS08d]. Press [Ahe02, Alf09, Ano18a, Aur04, Bed01a, Bed09a, Bed09b, Bed09c, Bee15, Ber08b, Bla00a, Bla00b, Blo06, Bor11, Bri99, Bro04, Bru02, Bru06, But00, Can01, Car07, Car11, Cas10, Cas11, Che11, Coh11, Com05, Cre12, Cus00, Dah06a, Das18, Dom05, Dor06, Ehr04, Eme09, Eme11, Eve00, Eve04, Eve08a, Fel08, Fel11, For01, For02, For08, For09, Fra01a, Fra07, Fra09a, Fra10, Fra12a, Fra01b, Fre05a, Fri09, Fri11, Gea99, Ger03, Ger06a, Ger99a, Gib05, Gol99a, Goo09a, Goo12, Goo08, Goo09b, Gou01, Gre05, Gre07a, Gre06, Gre07b, Hal09a, Hal18, Ham11, Han04, Har05, Hec07, Hie99, Hil11, Hob02, Hof17, Hol01, Hol09, Hol07, Hol19, Hub06, Hug09, Hun01, Jac04, Kah03a, Kle07, Kle11, Kob02, Kra16, Kro10, Lal19, Lep07, Lin10]. Press [LM06, Lus04, Lus10, Mas07, Mer01, Mil10, Mod18, MG21, Nal20, Neu99, New99, New04, Noe09, Noe11, Nyc09, Oha08a, Oha08b, Oha10, Par00, Par06, PCGD02, Pas04, Pas10, Pas12a, Per10, Pes11, Pri99, Ram05, RS06a, RS06c, RS06d, RS07a, RS07b, RS07d, Rig07, RS08a, RS08c, RS08d, RS09c, RS09d, RS10b, RS10c, RS10d, RS10e, RS11b, RS12d, RS13d, Rio00a, Roe10, Roe12, Ros01, Ros08, Ryc10, Ryd13, Sch01, Sch05, Seg09, Seg10a, Seg11a, Seg13, Sha06, Sha02, Shi06, Sim04, Sim12a, Smi20, Sob08, Ste03b, Sti08, Sty05, Tat13, Tay10, Tei02, Tor02, Tri07, Tri08, Tri09, Wal07, Wei04, Wes10, Yeh05, vB99, vB02a, vB06, vB08a, vB08b, vB10b, vB10c, vB12, Goo04]. Pressure [Cha18]. Priestley [Sch01, Sch97]. Princeton [Aur04, Cus00, Ehr04, For09, Fra09a, Fra01b, Fri09, Hil11, Hol09, Hol07, Hun01, Kle11, Kob02, Mer01, Nal20, Oha08a, Pas04, Pas12a, RS06a, RS07b, RS09d, Roe12, Ryc10, Shi06, vB06, vB08a]. principle [AM05, Sch09b, Sha06]. Principles [Bro06b, Dir09, Mul01, Her03]. Priscilla [RS06d]. Private [Hol09, Lin10, Coe07, RS07i]. Prize [Bru03, RA13, Fri01, PA21]. prizes [Goo03a, Har02a]. Problem [Jha11, Mou20, Pes13, RS06b, FGL^{+}06, SE19]. Problem-Solving [RS06b, FGL^{+}06]. Problems [Eve04, Fra02a]. Proca [dP07]. proceedings [BV09]. Professional [NW08]. Professionals [Bed12]. Professor [Pip02]. Professors [Gre20]. Program [Fre05b, Ree14, Wes16, Wes18, Wes19]. progress [Kni06]. Project [Bed08, Neu99, Rob18, Ros98, Cas02, Cas17, Kel07, Ree09, Ree11, Ree14, Sim12b, SS10, Vey13]. Projectile [Mou20, Sal14]. proliferation [Per99, Bla00b]. Prometheus [Bed06, Fra04, Hob04, Lan10, Lar05c, PP08, BS05]. Proof [New04, Pes03]. Propagation [Hon03]. Properties [Wei08]. Props [RS10a]. Proto [Pro16]. Proto-Megascience [Pro16]. Providence [GLR20]. Przibram [Rei19]. pseudoscience [WWH01]. PSSC [RS06h]. Psychology [She13]. Pt [MPFP19]. Public [Cas11, Hol09, RS08g, RS11h, Waz10, Car10, RS08h, RS07i]. Publication [GR08b]. Publishable [Fre05b]. Publishers [Bed08]. Publishing [Bru99a, Fre00, Ryn99, Van03, Wat03, vB02b]. Pullman [vB99]. Purcell [Rig11]. Pure [D’A00, Van07, RS02b]. Pure-Field [D’A00]. Purrington [Hie99]. Pursuing [Hun10, Ham11]. Pursuit [Hal09b]. Puzzle [Pes12]. Pygmalion [Lam14].

Qing [NGZ23]. Quantitative [RA04]. Quantum [BF13, Ban14, Ban16, Bed09a, Bed09b, Bed09c, Bel99, Ber09, Bro05, Bro06b, Car11, Com05, Cus02, DeW99, DS22, Dir09, Fol05, For20, Gau18, Gea02, Ghi05, Gre05, Hal12, HW87, HW03, Hon03, JFEH15, Jon07, Jon08, Kob02, Kra99, Kra16, Kra11b, Kro10, Lin09, Lus10, Mal12, Omn99a, Omn99b, Par11, PCGD02, Pes12, RS06c, RS09c, RS10c, RS12f, RK11, Sch05, Shi06, Sil05, Sol06, Sty05, Wat04, WWH01, von09, Bag04, Bok08, Cas09, For04, FR11, Mil07, NHW08, PFD08, Sch03, Set10, WF98, BV09, Bed09a, Cus00, Mer01, Par00, RS07d, Rig07, RS10d, RS12h, Seg10a]. Quark [Wat04, Sch05]. quarks [Sta01]. Quarries [Rei10]. Quenched [Cre05a, Cre05b]. Quest [Fel11, Fre05b, Fri08, Kle23, vB08a, Ken07, Mil05, Sta10]. Questions [Pal00, RS04d, Bed01a]. Quinn [Fra09a]. Quirino [DMA07].

R [Fra09a, Ger06a, HK03, Lin10, Mod18, RS08c, Sob08, Ste17, Seh19]. R. [Hol99]. Rabbi [Ano04a]. Rabbits [Sil05]. Rabi [Day04, Kri05]. Race [Blo06, Dah99, Hec10, Hol07, Wat01, Cre06, McM05, RS06d]. Racial [Bon11]. Radar [Ess10, Fre00, Bro99]. radiant [Cha02a]. radiation [Loe03, Alp12]. Radioactivity [Buc20, Gab00, GLR06, GR09, GLR12, GLR20, Kip00, Mal11, Rei19, AGR04, Rei01a, Sim12a]. Radiochemistry [LKK04]. Radium [Hug09, vB10c, Ren08]. Rain [Kah03a, Fri00]. Ralph [Alp12, PCGD02]. Raman [Ban14, Sin02]. Random [Kov03, Rei01b]. Rasetti [Goo01]. rational [D’A02]. Ray [Fri99, VJ10, Pas12a]. Rays [Fra02b, Ger99b, Ryn99, Dah97, Fri89]. Reactionaries [Van07]. Reactor [GORS09, Sta10, Fel11]. Reading [CP14d, DeV00, Ger99a, Sim07]. readings [Hug99]. Ready [Fra05]. Real [Hol19, Zim02, WWH01, Fra01a]. Realism [RS02b]. realities [Mic06]. Reality [Mal12, PCGD02, RS10c, RS11b, Wes19, Cus02, DG10, HK03, Rig07]. Really [Jen11]. Reason [RS03d, RS11e, RS11g, Fer10, Haa03, Haa07, Fra04]. Reasonable [Bro06a, Fey05, Hol03]. Recentering [CMS20b]. Recently [O’C10]. Reception [Rom99]. Rechenberg [Cas05a]. recipient [Tra08]. Recognition [Nye19]. Reconfiguring [Osl10, Goo12]. Reconsideration [Buc20]. Reconsidering [Jan02, Seg10a, BV09]. recovery [AGR04]. Redshifts [OO20]. Reductionism [Kro10, Bok08]. Reexamining [Bok08, Kro10]. Reflection [Mou20, Cha02a]. Reflections [RS10b, Cha02a, Dav10, FGL^{+}12]. Reflective [Che01]. reflexes [Cha02a]. Reform [Liu20]. regional [SE19]. Regnault [RA10]. Regular [Hee06]. Reich [Sim08, Hof05, HW08, Sim06, Sim08]. Reiche [Bed05]. Reinventing [Hil07]. Reisinger [vB99]. Relation [Kro10, Bok08]. Relativistic [Kat05]. relativité [Eis02]. Relativity [Bai07, CdL16, Oha08a, Oha08b, Tay05, Tay10, Tri07, Cre06, Mar09, Bru99b, EG05, Eis02, Eis06, Hol07, Mel09, Mil99, Nal20, Oha10, RS06a, Van07, Wol03, van12]. Religion [Nyc09, Wil99, Day08, Sta07]. remarkable [Die05]. Remarks [Bro13]. Remember [RS06f]. Remembering [Kah03b]. remembers [Wei07, Eve08b]. Remigration [Hof09]. Reminiscences [Bar99b]. Renaissance [Cre12, Gre09, Sal07, Pet11, RS08d, RSD07]. Renaming [OO20]. Renn [RS06a]. Rentetzi [Hug09, vB10c]. Replicating [Hee06]. Reply [Ste17]. Reporting [Mar21]. representation [D’A02]. Representations [Hen99a]. Repression [Wal06, Wal03a]. Research [Das19, Fre05b, Ham00, HC11, HRS00, Hug09, RA04, RS11g, Ste03a, Wei08, Wes19, vB10c, Ren08]. Resisting [Kra11a]. Resonance [Pav10, Ram99]. Responsibility [Fra01b, Wal06, Bad05, Sch00b, Wal03a]. Restless [RS07c]. Rethinking [Wil99]. Revealing [RS07c]. Revenge [MM23]. Review [Ahe02, Alf09, Ano18a, Aur04, Bad21, Bed01a, Bed06, Bed08, Bed09a, Bed09b, Bed09c, Bee15, Ber08b, Bla00a, Bla00b, Blo06, Bor11, Bri99, Bri07, Bro04, Bro06a, Bru99a, Bru02, Bru03, Bru06, But00, But19, Can01, Car07, Car11, Cas05a, Cas10, Cas11, Cas12, Che11, Coh11, Col00, Com05, Cre12, Cus00, Cus02, Dah06a, Das21, Day02, DeV00, DeW99, Dom05, Dor06, Eck23, Ehr04, Eme09, Eme11, Eve00, Eve04, Eve08a, Eve08b, Fel08, Fel11, For01, For02, For08, For09, For20, Fra01a, Fra04, Fra07, Fra09a, Fra09b, Fra10, Fra12a, Fra01b, Fre00, Fre04, Fre05a, Fri08, Fri09, Fri11, Gea99, Ger03, Ger06a, Ger99a, Ger06b, Gib05, Gol99a, Goo09a, Goo12, Goo02, Goo08, Goo09b, Gou01, Gre05, Gre07a, Gre06, Gre07b, Gre09]. Review [Gri06, Hal09a, Ham11, Han04, Har05, Hec07, Hie99, Hil07, Hil11, Hob02, Hob04, Hof17, Hol01, Hol09, Hol07, Hol19, Hub06, Hug09, Hun01, Jac03, Jac04, Jam00, Jon07, Kah03a, KP02, Kle07, Kle11, Kob02, Kra16, Kro10, Lal19, Lan10, Lar05b, Lar05c, Lep07, Lin10, LM06, Lus04, Lus10, Lyo20, Mar03b, Mas07, McC22, MM23, Mer01, Mil10, MG21, Nal20, Neu99, New99, New04, Noe09, Noe11, Nyc09, Oha08a, Oha08b, Oha10, Par00, Par06, Pas04, PP08, PP10, Pas10, Pas12a, Per10, Per01, Pes11, Pes15, Phi13, Pip99b, Pri99, Ram05, Rig07, Rio00a, Roe10, Roe12, Ros01, Ros08, Ryc10, Ryd13, Ryn99, Sch00a, Sch01, Sch23, Sch05, Seg09, Seg10a, Seg10b, Seg11a, Seg13, Seg15, Sha06, Sha02, Shi06, Sil07, Sim04, Sim07, Sim08, Sim10a]. Review [Sim12a, Smi20, Sob08, Sta06a, Ste03a, Ste03b, Sti08, Stu04, Stu06, Sty05, Tat13, Tay05, Tay10, Tei02, Tor02, Tri07, Tri08, Tri09, Van03, Wal07, Wat01, Wat03, Wes10, Yeh05, vB99, vB00, vB01, vB02a, vB02b, vB06, vB08a, vB08b, vB10a, vB10b, vB10c, vB11, vB12, KG19]. Reviewed [Tri08]. Reviews [AHSB09, Ano11a, Ano11b, Ano12a, Ano12b, Ano13a, Ano13b, Ano13c, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15b, Ano16d, Ano16a, Ano16b, Ano16c, Ano17, Ano18b, Ano18c, BHF^{+}09, BEF^{+}08, BHS07, CvB01, CGYC05, CFH^{+}11, DBBG06, Das18, DSFG05, EvBF^{+}08, FFF^{+}01, GBS^{+}06, GEBF09, GWL03, HTR05, HMHB01, KBS03, KMG^{+}07, KHC11, KSAF02, KPP^{+}10, LEJ04, Mod18, NNT^{+}09, OTRF08, PCGD02, PFW04, PFC^{+}10, PORL10, PSTR13, Rob18, RPF12, SGvB06, SEA^{+}04, SHL^{+}07, SvBC^{+}12, SGJ03, SvBG08, SHH04, SBLM^{+}06, TKP^{+}02, WRJH07, WMV03, Wei04, vBBH02, vBSTM10, Mor90]. Revised [Mal12, Shi06, vB02a]. Revisited [Sch16, Kra16]. Revolution [Fle21b, Fra12a, Gol99a, Gra10, Joh99a, Per01, van12, Bel99, Huf11, Jan02, Par00]. Revolutionibus [RS11d, Rom99]. revolutions [Coo99]. Rhys [Dah06a]. Riccioli [Gra11]. Richard [Bro06a, RS07b, RS12h, Cla09, Eve08b, Fey05, GG00, Hil11, Kra11b, Ram05, RS08d, RS10e, SS07, Tay05, von09]. Riding [Ber08a, Ber10a, Hal09a]. Riemann [Pes13]. Riess [Ste03a]. Rigden [Stu18]. Rise [Bri07, Bro13, Har05, War03]. Risks [Kap08]. Ritz [Mar04b]. Riverside [Tri08]. Road [Hei10c, KP02, Dav00b, Goo03a, Har02a, HK07, TF07]. Roald [RS12d]. Robert [Bed06, BS05, Cas05b, Ger06a, Hol09, Kle07, McM05, ORSSBO^{+}17, PCGD02, RS04c, RS06d, Tri08, Alp12, BCHJ03, Bru03, Gre06, Hie99, Hil07, HK03, Inw02, Lip06a, Lip06b, Nau05, Pai99, Pas04, Pav10, Rio00a, Sch01, Stu04]. Roberto [But00]. Roberts [Gre09]. Robotti [RS09a]. Robyn [RS06d]. Rock [Rei10]. Rodale [Sil07]. Roger [Tri08, vB08b]. Roland [Kob02, Mer01]. Role [Hal04]. Roles [Fra99]. romance [Ove00, Ove01, vB01]. Romantic [RS10e, Hol08a, Hol08b, Hol08c]. Rome [Foc22]. Romer [JP99]. Ronald [Fra07]. Rose [Neu99]. Rosenblum [Bed09a, RS07d]. Rosettes [Kos13]. Rosner [Stu04, Tri08]. Rossi [Bon11].

Rotation [Gra11]. Rotblat [Vey13]. rough [HK07]. Rowe [Hol09]. Roy [Tay10, Han04]. Royal [Foc22, JP07]. Rudolf [Pip02]. Rudolph [Jon07]. Ruin [RS06d, McM05]. Rule [Ano02]. Rumford [Cha02a]. Rupke [Roe10]. Rupp [Fre99]. Ruprecht [Sta06a]. Russell [Can01, Hob02]. Russia [CS18]. Russian [Eve06]. Rutgers [Ger99a, Hie99, Yeh05]. Rutherford [Ano05b, Jen11].

S [Ahe02, Fra01b, Goo09a, Goo08, Huf12, Nyc09, Nye99, RS08a, Sta07, Stu18, Tri09, Hol19, Nal20, Sal19]. S. [Fra01b]. Sabine [But19]. Sackur [BF13]. sad [Gav01]. Saga [Cre05a, Cre05b, PP10, Kra08a]. Sagan [Car07, Jam00, Sch00a, Dav00a, Pou99]. Saha [Das15]. Sailor [Fre04, Hor03]. Sakharov [Eve06]. Salam [Alf09, Fra08b]. Saliba [RS08d]. Sally [Sha06]. Sample [vB11]. Samuel [Loe03]. San [RS06b]. Sander [Oha08b]. Sangalli [Mer01, Oha08a]. Santa [Bad09a, Noe09]. Sap [Bro13]. Sarah [Bee15]. satellite [Kra08a, PP10]. Satyendranath [Ban16]. Sauveur [Fix15]. Saw [Mod18]. Scalar [Cha18]. Scarani [Rig07]. scattering [TF07]. Scenarios [Kap08]. Scerri [Das18, Sob08]. Schaffer [Gre09]. Schall [Liu20, Sal20]. Scharff [Gol16]. Scharff-Goldhaber [Gol16]. Scheppler [Hob04]. Schmitt [Har02b]. Schneider [Car07]. Schofield [Sch01]. Scholars [Bed05]. School [Eck99]. Schrödinger [D’A04, Sil05]. Schulmann [Hol09]. Schweber [Fra01b, Goo09a]. Schwinger [Gau18, Ger03, MM00, Mil07]. Science [Ano01a, Ano18a, Aur04, Ber08a, Ber10a, Buc16, Cas10, Cas11, CP15a, CMP18d, Day08, Dom05, DS13, Eve00, Far09a, Far10, Fer02, Fir12, Fix15, For02, Fra01a, Fra04, Gea99, GP01, Goo09b, GMS^{+}03, Gre01, Hal09a, Ham00, Han04, HC11, Her00, Huf12, Joh99b, Jon08, Kle11, Kov03, Kra11b, Lai04, LH11, Lin10, LM06, Mar14, Mar21, MNS23, Mar11, Mon00, Moo08, Nyc09, Omn99a, Par06, Pav11, Ped11, PA01, RA10, Rei01b, Rhe02, RS04a, RS04d, RS07a, RS07d, RS08a, RS08d, RS09i, RS10f, RS11e, RRR13, Sch11, Sei00, Sim04, Sok08, Sta06b, Ste03b, Sti08, TS01, Van07, Wat03, Wes10, Wil99, ZZ18, vB02a, vB02b, Bad09b, Bar99a, Bri05, Car10, Coe07, Dav10, Day04, Die05, Eic08, Fer10, Fri01, Gol10, Gol12, Goo03a, Goo10, Haa03]. science [Haa07, Har02a, HW11, Hol08a, Hol08b, Hol08c, Hol05, Hun04, JP07, Jen08, Joh09, Jos10, LC01, Lac07, Lar07, Mal11, Mor90, Nye08, Pan03, Por09, RS12a, Sal07, SR06, Sim10a, Smi98, Sta07, Tre08, Wei01, Wei15, WWH01, Zim02, Ahe02, Bed09c, Bor11, Bru03, Coh11, Das18, Fra10, Fri09, Ger06b, Giu12, Han04, Har06, Hec07, Hob04, Kle11, Lar05c, Mer01, PCGD02, Pes15, RS07b, RS10b, RS10e, RS12h, RS12d, RS13d, Seg11a, Sha02, Sim12a, vB11, vB12, Eme09, Day02]. Sciences [Hie00, SDC12, Nye08, RS11c, Sim04]. Scientific [Ano18a, Das19, Ehr06, Eve08b, Eve12, Fle21b, Fra12a, Ger03, Gre00, Har01, Kap08, Lam14, O’C10, ORSSBO^{+}17, Phi13, RA15, RS06c, RS08b, RS09a, RS11g, SE14, vB01, GR08a, Huf11, Jan02, Kni06, MM00, Ove00, Ove01, Fra07]. scientism [Haa03]. Scientist [Fre04, Man22, Stu05, Tow99, Fra08b, Hor03, Sch00b, Alf09, Bri99, Fra01b].

Scientists [Bad01, CMP18c, Das15, Das18, Hof09, Moo08, RS11h, Van07, Goo03a, Har02a, Kur02, Man23]. Scott [OV22, Sha02]. SDI [Her00, vB02a]. Sean [Hec07, Wes10, vB02b]. Search [Ber16, DFAB19, Gea14, Hal18, Pas12a, RS09b, Sim00, Hoo08, HK07, Jay11]. Seasons [OW13a, OW13b]. Sébastíen [RS09d]. Secrecy [Mas07]. Secret [MNS23, SS10]. Secrets [Seg11a, Mar11]. SEDs [Bed01b]. See [Lan12]. Seeds [Gra10]. Seeing [Fra17]. Segrè [Cas12, RS07a]. Selected [Ryc10, Pes09]. selection [Day99, Lar07]. Selectivity [Fra02a, Eve04]. Seminal [Nau05]. Seminar [Bet00, Sch09b, dP07]. Seminary [SS10]. Sensationalized [Hal18]. Sense [RS07d]. Sensibility [Lar05c, Lai04]. September [Ber02b, RS02d]. Serafin [Stö02]. Serber [Pai99]. Serendipity [Cra00]. Series [Fri99]. Serum [AR19, MDK19]. Seth [Car11, RS07a]. settembre [Ber02b]. Settlers [CP15b]. Seven [Das18]. Seventeen [RS13g]. Seventeenth [Gre07b, Liu20]. Shadow [Fra01b, Nal20, Per10, Lev09, Sch00b]. Shadows [BW06, Eve08a, vB02b]. Shahn [RS09d]. Shall [RS13h]. shaped [Die05]. Shaping [vB10b, Wal09, Wal11]. shared [Sch06]. Sheilla [Bed09c]. Shell [Joh04]. Shelter [Sch16]. Sherwin [Bed06]. Shifting [Nye19]. Shifts [OO20]. Shimon [PCGD02]. Ship [HHH17]. Shockley [Bri07]. Shoolery [Wat03]. Short [Pas12b, Clo09, RS10d]. showing [Seh19]. Shulman [RS07a]. Shurkin [Bri07]. Sichau [Ste03a]. Side [Hal12]. sights [Sta06b]. Significance [Sce07, Sob08]. Silicon [Har09]. Silvan [Goo09a]. Simon [Gre09, Gra10]. Simone [Seg13]. Simple [RS04b, RS02a, Seh19]. Simply [Wol03, Tay05]. simultaneity [Mar04a]. simultaneous [LPR04]. Singapore [Eve08b, Phi13, RS08b]. Sirens [Pes13]. Sister [PA21]. Site [SS10]. Sites [PA01]. Six [RS12f]. Sky [Lan12, Tri05, Gre07a]. Slit [Ros12]. Small [MNS22]. Smith [Gea99]. Smoke [Lis16]. Sneaking [Ghi05, Shi06]. Social [Moo08, Bad05]. socialist [Hor03, Fre04]. Società [Stu06]. Society [Sim08, Hof05, HW08]. soft [Ste03a]. Sokal [Goo09b]. Solar [CdL16, Pas12a, TT12, Jay11]. Solvay [BV09, Seg10a, vB00, MW99]. Solving [RS06b, FGL^{+}06]. Some [Bro13, Hen99b, MDK19]. Sommerfeld [Car11, Bet00, Eck99, Sch09b, Set10]. Son [Eme09, Jen08]. Sons [Ger06b]. Sorcerers [Ram05, Mor03]. Soul [RS07a, RS07b, Seg07]. Sound [Ber08a, Ber10a, Hal09a, Pes13]. source [Mil07, Cre08, Cre09a]. Sources [GLR06, New04, Pes03]. Soviet [Gor09, Ste23]. Soviet-American [Gor09]. Space [Bed09a, Eve00, Goo02, Hug99, Jac14, PCGD02, Pes13, RS07c, RS09d, Bar99a, Eis02, Fri00, JP07, Lin03, Maj08, Maj12, Mil01, NHW08, Kah03a, Pri99, Tri07]. space-time [Eis02, Tri07]. Spacetime [Tor00]. Spandri [Man22]. Sparked [vB11]. Special [Bai07, Oha08b]. spectral [Hof02]. Spectroscopic [Hen99a]. spectroscopical [LR03]. Spectroscopy [Fri99, Wes06]. Spectrum [Jac00, Tei02]. Speed [Bed09a, DMA07, vB08a, Ken07, NHW08]. sphere [Car10, Cas11]. spinal [PMAS19]. Spiral [Wee10]. spontaneously [KKG^{+}19]. spots [Lin03]. Springer [For20, Jac03, Rob18, Stu06].

Springer-Verlag [Jac03]. Spurious [PP10, Kra08a]. spy [Clo15, Seg15]. SSC [RS09g]. Stability [Boy12]. Stacey [Fel11]. Stachel [Mar03b]. Stages [Ano10b]. Stands [Hol09, RS07i]. Stanford [Mod18, vB02a, vB12]. Stanley [Nyc09]. Stannard [Hob02]. Star [GP01, Ste03b]. Stark [LPR04]. Stars [Das15, Fri08, Gra10, Kir02, Pas04, Tri08, Hof02, Mil05, RS03e, SRD06]. State [Aus16, Sch01]. States [Bed12, RS13f]. statistics [Bad09c]. Stefan [Cre09b, Rei01a]. Stefania [Buc20]. Steinle [Hof17]. stellar [Gra08]. Stephanie [Hob04]. Stephen [Gib05, Lar05a, PP08, Wei04]. Steps [Mil99]. Sterne [RS03e, Stu04, Tri08]. Steve [Tor02]. Steven [Bro04, For02, Mil10, Pes15]. Stigler [Cre09b]. Still [Ano16e, Gra08]. Stockholm [Goo03a, Har02a]. Stop [RS03b]. storm [Lem07]. Storming [Sic06]. Story [Bed09c, Esp21, Hec04, Hol19, Jon08, Ros01, Sce07, Sob08, Wes14, Gav01, RS02a]. strain [KKG^{+}19]. Strange [Bee15, Fre99, Hon00, Jay11, Joh99a, Mil09, Wei04, vB10a, Dry14, Inw02, Pas12a, Per01]. Strangest [Far09b, RS10c]. String [Ehr06]. Strohmaier [Stu04, Tri08]. strong [Ano04c]. Structure [Hon03, Kra00]. Struggle [Hol99, RS07a, RS07b, Seg07]. Stuart [RS07b, RS13d]. Student [Huf12]. Studies [Dom05, MPFP19, Har85]. Study [GORS09, Hu16, MDK19, Neu99, Sch01, Ros98, Sch97]. Stuewer [Tri08]. Sturm [Lem07]. Subatomic [Bro04, Wei83, Wei90, Wei03]. subcontinent [PW19]. Subjects [Joh06a]. Sublime [RS12d, HW11]. substructured [Cha19]. Subtle [Noe11, Coo10]. Suess [Joh04]. Suman [Car11]. Sun [RS07b, LWE09, Cla09, GP01, RS09c, Ste03b]. Super [Rio00b]. Supercollider [Mar14]. Superconducting [Ber16, Mar14, Rio00b]. Superconductivity [GG00, MW03, Yeh05]. Superior [Fix15]. Supermagnets [Gri06, Gui05]. Supersymmetry [RS09b, Hoo08]. Suplee [RS10e]. Supplement [RS06b, FGL^{+}06]. supplemental [Kle07]. Suppression [RS07a]. Surdo [LPR04]. Surely [FLH85, RS06b]. Surface [WJS12]. Surprising [GP01, Ste03b]. Surveyed [Ste23]. Susan [Fra04]. Sweets [Ano01c]. Symbolic [D’A02]. Synchrotron [Cre08, Cre09a]. Synopsis [Jam07]. system [Jay11, Pas12a]. Systems [New99, Fla98, MDK19, RS11c, RS11d]. Szilard [Fra05].

t [FEVvdB19]. T. [Huf12]. Table [Joh99b, Sce07, Sob08, Mor03, Ram05]. Tacit [Jha11]. Tactical [Ruh20]. Tale [Cas10, Cra00, Das18, Dor06, RS07b, Rus15, Wes06, Bad09b, Cla09, Lin04, RS01d]. Tales [Bal08, Kle11, RS09d, Goo10]. Talk [CMS19d]. Talks [RS10a]. Tau [Per04]. Tausk [PFD08]. taux [PMAS19]. Teacher [Jac09, Stu05]. Teacher/Eugene [Stu05]. Teachers [RS08a, Stu05]. Teaching [Cro99]. Technical [Fre00, Bro99]. Technique [Hon03, RA04]. Technology [Buc16, Dom05, Ham11, Hil11, Rhe02, Sch11, Hun10, Mul10]. Teilchenphysik [RS03e, Stu04, Tri08]. Telegdi [Lip07, LG08]. Telephone [Ano04b]. Telescopes [Mas07]. Telescopic [Gra10]. Tell [Fra16]. Teller [Wat03]. temperature [Kni06]. Temperatures [RA13]. temps [Eis02]. Ten [Bed09c, Jon08]. Tension [WJS12]. Terian [TT12]. term [CFL19]. Termination [NGZ23]. Terror [RS10e, Hol08a, Hol08b, Hol08c]. Test [Hol07, Cre06]. Testable [Ehr06]. Testimonies [FJ21]. Textbooks [Fra16]. Their [RS10a]. Themes [Fle21a]. Theological [SS10]. Theologies [Kel22]. Theoretical [Bel14, DE07, Eck99, Hal12, Par11, Tis09b, dP07, Tis09a]. Theories [D’A00, Hun01, Kra96, Lep07, Mou20, RS01b]. Theory [BF13, Bro05, Bro13, Car11, Ehr06, Eis06, Gea14, GR09, Hal18, Har05, Kra18, Nal20, Oha08a, Pes12, RS09b, Seg10a, Sic06, Sta01, BV09, Bag04, D’A04, Gra08, Hoo08, Mil07, Set10, Stö02, War03, Gre05, RS06c]. Theresa [Per10]. Thermal [Ree11]. Thermodynamics [Car07, RA15]. Thermometry [RA15]. ThermoPoetics [Coh11, Gol10, Gol12]. these [RS01d]. Thesis [Bro05, RS06c]. Thin [Kah03a, Fri00]. Things [CMS21b]. Think [RS06b, FL88, FF05]. Thinking [Ano00a, Eve12, Joh04, Gre07b]. Third [MPFP19, RS09c, Sim08, Sta01, Hof05, HW08, Sim06]. Thomas [Ful00, GR08b, Sch06, Tor02]. Thomson [Ryn99, Dah97]. Thought [vB08a, Ken07, Pul98, vB99]. Thoughtful [Phi13, And11]. Thoughts [Hol09, RS07i]. Thousand [Fra10, Far09a, Far10]. Thousands [SE14]. Three [Rhe02]. Tides [Gra08]. Time [AK99, Bed09a, Bon18, Col00, Hon00, Jam07, Mon00, PCGD02, RS07c, Wal07, AK12, Cas05a, Eis02, FF05, Ful00, Goo02, Lem07, Lin03, Maj08, Maj12, Mil01, NHW08, RS09d, Sha02, Tri07]. Times [Bru03, Lus04, Lus10, Cro01, Tor02]. Timothy [RS11e]. tips [FGL^{+}06, FGL^{+}12, RS06b]. tissue [PS19]. Titanium [Ber16]. Toby [Fra12a]. today [RS06h]. Told [DFAB19]. Tomorrow [Mod18]. Tompkins [Hob02]. Tony [Sty05]. Too [Wei04, Inw02]. Top [RS12a, SS10]. Top-Secret [SS10]. Torretti [But00]. Torsion [Hee06]. total [LWE09]. Totality [LWE09, RS09c]. Tour [Hal10, Joh06b, Lan12, Rhe02]. Tourist [Bed03, BP08, Gre00, Hen07a, Hof00, Hol04, Hom05, Hun15, Joh06b, KW10, Kov03, Lac07, PA01, Rei01b, Rhe02, RS09i, SR06, SS08, Str04, Sta06b]. tours [Hof00]. Townes [Bri99]. Toys [Gau18]. Track [Bro06a, Fey05]. Trafficking [Ren08, Hug09, vB10c]. Tragedy [Bed06, Bed09c, Jon08, RS07b, BS05, Cla09, Lin04, Dor06]. Training [CS18]. Transferrin [PW19]. Transformation [Gin10, Gre07b, HHH17]. Transition [D’A00, Wes18]. Transits [She13]. Translated [Bed09a, Oha08a, Sim10a, Tri08, vB12, Ber08b, Hof17, Jon07, Mer01, Sha06, Shi06, Yeh05, vB99]. Translating [Kel22]. Translation [Mon00, Sha02, BV09]. Transmutation [Van03, Dah02]. Transnational [Ban16]. Transparency [Per10, Lev09]. Transuranium [Sim00]. Traveler [RS09i]. Traveling [Ken07, vB08a]. Travels [Gre00]. Tree [Pip99b, Day99]. Trees [Bro13, RS13h]. Trefil [RS08a]. Trends [Nye19]. Triads [Ros08, Jac06]. Trilogy [Kra16]. triploid [PS19]. trisomiques [PMAS19]. trisomy [FEVvdB19, KKG^{+}19]. Tritium [Ber02a, Jac04]. Triumph [Ahe02, Bed06, Gre01, BS05]. Troubled [Seg10b, Mag09]. True [Bee15, Dry14]. Trusted [Bla00a, Cer98]. Truth [Fra16, Seg11a, Sil07, Mar11]. Tunneling [Bed09a, NHW08]. Turbulent [Bon18]. Turchetti [Seg13]. Turing [Dav00b, KP02]. Turn [Bel14].

Turning [HHH17]. Twelve [Ano10b, Bal08, RS09d]. Twentieth [Bru06, Cus00, Eve08a, Ger99a, Per01, Sti08, Wat03, Har06, Kra99, BW06, Dar04, Hu16, Joh99a, Kev97, TS01]. Twentieth-Century [Bru06, Eve08a, Wat03, BW06, Dar04, Joh99a, TS01]. Twenty [GMS^{+}03, Hob04, RS06c, Fra09c]. Twenty-First [GMS^{+}03, Hob04, RS06c, Fra09c]. Twin [RS03c]. Twins [Ber08c]. Twists [Hee06]. Two [Ano04c, Eve04, Hun01, Kra96, RS03c, RS11d, Ros12, ZZ18, Day04, Fra02a, Mul99, RS11c]. Two-Culture [ZZ18]. Two-Slit [Ros12]. Tycho [Wes15]. Tychonic [Gra10]. Tyndall [Rei10].

U [Day08]. UA1 [PA21]. UA2 [PA21]. Unabomber [Kap08]. Uncertainty [Bed09b, Dur03, Lin10, Lus10, Cas09, Coe07, RS07b]. Undermining [RS07a]. Understand [RS05b]. Understanding [Buc20, For20, Goo12, Gri06, Mer01, Omn99a, Omn99b, RS12f, RS12g, Gui05, Osl10, Kob02]. Uneasy [NMS22]. Unexpected [RS07b, Cla09]. Unification [Hal07]. Unified [Hal18, RS09b, Hoo08]. Unit [Ess10]. United [Hob02, Bed12, RS13f]. Unity [PA21]. Universal [KP02, Dav00b, Day04]. Universe [Cha02b, Fac82, Ehr04, Eve12, Goo08, Gre12, HW87, HW03, Hil07, Kir02, Kra96, Lal19, Lan10, NB09, Pas04, Pes12, Ren05b, Ren05a, RS10c, Roe12, Tri09, Wei11, Cal09, Kra07, Lai04, Lin03, SBD09, Wei09, Hun01, Kel22, Kra12, Lar05c, Pas10, RS06a, RS10e, Sty05, Wal07]. Universitäts [Hen05b, Sta06a]. Universitäts-Mechanicus [Hen05b, Sta06a]. University [Ahe02, Alf09, Ano18a, Aur04, Aus16, Bed01a, Bed09a, Bed09c, BS11, Bee15, Ber08b, Bla00a, Bla00b, Blo06, Bri99, Bro04, Bru06, But00, Can01, Car07, Cas11, Che11, Com05, Cre12, Cus00, Dah06a, Das18, Ehr04, Eme09, Eme11, Eve00, Eve04, Eve08a, Fel08, Fel11, For01, For02, For08, For09, Fra01a, Fra07, Fra09a, Fra10, Fra12a, Fra01b, Fre05a, Fri09, Fri11, Gea99, Ger03, Ger99a, Gib05, Gol99a, Goo09a, Goo12, Goo08, Goo09b, Gre05, Gre06, Gre07b, Ham11, Han04, Har05, Hec07, Hie99, Hil11, Hob02, Hof17, Hol01, Hol09, Hol07, Hub06, Hug09, Hun01, Kah03a, Kle07, Kle11, Kob02, Kra16, Kro10, Lep07, Lin10, LM06, Lus04, Mas07, Mer01, Mil10, MG21, Nal20, Neu99, Noe11, Nyc09, Oha08a, Oha08b, Oha10, Par00]. University [Par06, PCGD02, Pas04, Pas10, Pas12a, Per10, Pes11, Pip02, RS06a, RS06c, RS06d, RS07a, RS07b, RS07d, Rig07, RS08c, RS09c, RS09d, RS10b, RS10c, RS10d, RS10e, RS11b, RS12d, RS13d, Rio00a, Roe10, Roe12, Ros01, Ryc10, Ryd13, Sch09a, Sch01, Sch05, Seg09, Seg10a, Seg11a, Seg13, Sha06, Sha02, Shi06, Sim04, Sim12a, Smi20, Sob08, Ste03b, Sti08, Sty05, Tat13, Tay10, Tor02, Tri07, Tri09, Unn00, Wal07, Wei04, Wes10, Yeh05, vB99, vB02a, vB06, vB08a, vB08b, vB10b, vB10c, vB12, Hen05b, Sta06a, Bad09a, DS22, HR01, Mar03a]. Universums [Ren05b, Ren05a]. Unlikely [Ber10b, Kra00]. Unraveling [Ghi05, Shi06]. unreasonable [Hol03]. unsichtbare [Hen05b, Sta06a]. unsolvability [Pes03, New04]. Untersuchung [HRS00, Ste03a]. Untersuchungen [MPFP19]. Upbringing [CS13]. updated [LWE09]. Upon [Hon00]. Upright [Lem07]. Uranium [For22]. Urban [Tri05].

Ursprünge [Ste05]. Uruguayan [OV22]. Utrecht [Sta06b].

V [Ban14, Sin02]. Valentine [Lip07, LG08]. Valentini [Seg10a]. Valerio [Rig07]. Vandenhoeck [Sta06a]. variants [PW19]. Varying [Kra15, Kra19]. Vasilyevich [CS13]. VCH [Bed09a, Gri06, Jon07, RS06a, Sim08]. Vector [Cha18]. Vedral [RS10c]. Velocity [Ess10]. Ven [Wes15]. Venturing [JFEH15]. Venu [PP10]. Venus [Kra08a, She13]. Verantwortung [Wal03a]. Verbiest [NGZ23, Sal20]. Verdrängung [Wal03a]. Vergleichende [MPFP19]. Verlag [Jac03, PP10, Sim08, Ste03a, Stu04, Stu06, Tri08, vB00]. Verlagsbuchhandlung [Cas05a]. Verona [Man22]. Versus [Wil99, Ano00a, Ano00b, OO20, Jan02, Par11]. Vertical [Rei10]. Very [Bai07, MG21, Clo09, RS10d, Oha08b]. Vexation [Par06, Hol05]. vi [Ber08b, Cre12, Hol01, RS10d]. via [Ste23, Foc22]. Vicinity [Hof00]. Victor [RA10, RS02d]. Victorian [Ano18a, Coh11, Gea99, Gol10, Gol12, Smi98]. Victory [Hol05, Par06]. Vienna [Coe07, DS22, Hug09, Lin10, Rei01b, Ren08, vB10c]. Viennese [DS22]. Vietnam [IT23]. View [CdL16, Kra22]. Views [Ano11c, Goe03, Wei09, Mil10]. Vignette [Pip02]. vii [RS10b, Wal07]. viii [Ahe02, Dor06, New04, PCGD02, RS09a, RS13d, Ryd13, Tri09, Wes10, GORS09]. Viking [Cas12, RS07a, vB01]. Villard [Ger99b]. Vision [HK03, Die05]. Visions [Hec07]. Visiting [Wes15]. Visual [Hen99a]. Vital [NMS23]. vitro [PS19]. Vivian [Pav10]. Vlatko [RS10c]. Void [Kos13]. Vol [Han04, Sim04, RS06a]. Vol. [Nye08]. Volta [Aur04, Pan03]. Voltaic [Hon00].

W [Com05, DeW99, Eic07, Eic08, Kah03a, KP02, Phi13, RS06b, RS08d, RS12h, Ros08, Seg09, Sim07, Sim10a, Tay05, Tei02, ZZ18, vB10a]. Walk [Gab00, Hof99, Jan11, Kov03, Rei01b, SE14]. Walker [RS09a, Sim08]. Walking [Hof00]. Wallenborn [vB00]. Walls [vB10b]. Walter [OV22]. Walters [Sty05]. War [Bad05, Bro99, Cas17, Fre00, Hol09, Rob18, Seg13, Sim12b, RS07i, Tur12, IT23, Kle23, Nye99, Ruh20, Sal19]. Warming [Sil07]. Wars [Yeh05, MW03]. Wartime [Dah99, Ess10, Hu19a, Hu19b, Sch22, Wat01]. Warwick [Bru02, Har05]. Was [Eis06, Bru99b, Oha08a]. Washington [Cas11, Dor06, Ram05, Hal10]. Wasn’t [PP10, Kra08a]. Water [Bør12, Dah99, Pip01, Wat01]. Watson [Sch05]. Watt [Ham11, Hun10]. Wave [For09, Lev04, Ana08]. Waves [Ber08a, Ber10a, Hal09a, Ken07, vB08a]. ways [Eis02]. Waysand [Yeh05]. weak [RS06g]. Weakened [RS10a]. Weapon [Ruh20]. Weapons [Ber08d, For08, For22, Nye99, Ree14, RS03a, Bad05, Ber02a]. Weather [NGZ23]. Wegbereiterin [RS03e, Stu04, Tri08]. Weighing [Pou00, Pou01]. weight [CFL19]. Weimar [Stu04, Tri08, Sch09b]. Weinberg [Bro04, For02, Mil10, Pes15]. Weinheim [Bed09a, Gri06, Jon07, Sim08]. Weintraub [Roe12]. Weiss [Eve08b]. Weisskopf [RS02d]. Weizsäcker [Cas15]. Welcoming [Rom99]. Were [Hon00]. Werk [Ren05c]. Werner [Cas05a, Cas15, Lus10]. Wesley [RS06b]. Wesley/Perseus [Ger99a]. West [MDK19]. Western [Cus02, Mal12, PCGD02]. Westfall [Seg09]. Weston [Fel11]. wetenschap [Eic07]. Wetenschappen [Gre09]. Weyl [Ryc10, Pes09]. Wheeler [DeW99]. Where [CMP17d, SR06, WWH01]. Whimper [RS09h]. Whipple [Pip99a]. Whitehead [Fol05]. Who [Ano11e, Bla00a, Mod18, Sti08, Wei04, Cer98, Har06, Inw02]. Whodunit [Wee10]. Whose [CP15c]. Whyte [RS12d]. Wide [Hei10c]. Widow [Pip02]. Wien [Stu04, Tri08]. Wife [Hol19, Seg11a, Mar11]. Wigner [Jha11, Stu05]. Wiley [Bed09a, Ger06b, Gri06, Jon07, RS06a, Sch00a, Sim08]. Wiley-VCH [Bed09a, Gri06, Jon07, RS06a, Sim08]. Will [RS07h, RS08h, RS09h]. Willcox [RS09c]. William [Hu19a, LM06, New99, Ryd13, Eme09, Fra02b, GF05a, GF05b, HDS12, Hu19b, Hun04, Jam00, Jen08, Lus04]. Williams [Eve08a]. Wilson [Fra02b, HK03]. winding [TF07]. Window [Kip00]. Winter [Bad01, Cas10, Bad09b]. Wire [Hee06]. Wire-Torsion [Hee06]. Wisconsin [For01, Mar03a]. within [Fra04, Haa03, Haa07]. Without [Lan10, Cal09]. Wobegon [Rhe02]. Wolfgang [vB10a, Mil09]. Wolfson [Tay05]. Woman [LKK04]. Women [BW06, Eve08a, Foc22, Hof02]. Wonder [RS10e, Hol08a, Hol08b, Hol08c]. Words [Bed08, RS01d, Pip99b]. Work [Joh04, Lal19, NW08, O’C10, Ren05c, RS06a, Sch01, Stu06, WJS12, Wei04, BCHJ03, BB04, Pip99b, Sch97]. Workshop [Sta10]. World [Com05, Eve08b, Eve12, Fre00, Goo12, Hil11, Hob02, Jon07, Pes15, Phi13, RS06c, RS06d, RS07c, RS08b, RS11c, RS11d, Rob18, Sim07, vB06, Ari03, Ari05, Day08, Den07, Eve06, For04, Hol03, Mul10, Osl10, Par07, Wei09, Wei15, Bad05, Bro99, Cas17, Fel08, For20, RS05a, RS07d, Sim12b]. Worlds [Pas12a, RS08c, Sil05, Jay11, Lei08]. Wormholes [AK99, Col00, AK12]. Wranglers [Har85]. Write [Nye19]. Writing [CMS21b, Fer02, Sch14]. Writings [Ryc10, Day99, Pes09]. Wrong [RS11c, FR11, WWH01]. Wudka [Tri07].

x [Blo06, Bri07, Coh11, Fri09, Hal09a, Jac03, Kra16, Kro10, Mil10, Nyc09, PCGD02, Per01, RS09d, RS10c, RS11e, Ros08, Sch05, Tei02, Fri99, VJ10]. X-Ray [Fri99, VJ10]. xi [Bee15, Bru02, Bru06, Eme11, For20, Gou01, Hil11, Lin10, Mar03b, Pri99, Ram05, Sha02, Tat13, vB02b, vB11]. xii [Bed09c, Com05, Dah06a, Ehr04, Eve04, For08, Fra01a, Fra09a, Gea99, Gib05, Gre07b, Gri06, Hub06, Jac04, Jam00, Jon07, KP02, Lus04, Pas04, RS07b, RS09a, RS10e, Rio00a, Ros01, Sch01, Sim08, Ste03b, Stu06, Van03, Wat03, Wei04, vB99, vB08a]. xiii [Alf09, Bed06, Bla00b, Cas10, Fra12a, Fre04, Hun01, Kob02, Yeh05, vB00]. xiv [Bed01a, Cus00, For02, Fri08, Fri11, Ger99a, Har05, Kle11, Noe11, Par06, Pes15, Rob18, vB06]. xix [Car07, Car11, Fre05a, Lal19, Lep07, Oha10, Seg15]. xv [Aur04, Bru99a, Bru03, Eme09, Fra10, Par00, PCGD02, Pip99b, Sty05, vB01, vB02a, vB10b]. xvi [Bed09a, Bro04, But00, Cas11, Fel08, RS11b, Wat01]. xvii [Bla00a, Rig07, RS12h, RS12d, Ryn99]. xviii [For01, Fra01b, Hil07, New99, Seg11a, Tor02, vB12]. xx [Dom05, For09, Hie99, Neu99, RS09d, Sch00a]. xxi [Cas12, Goo09b, Gre06, LM06, RS10e, Sim12a]. xxii [Col00, Kle07, Noe09, Shi06, Sob08]. xxiii [Bro06a, DeV00, Mer01, vB10c]. xxiv [Sti08]. xxix [Hol07, Wei04]. xxv [Seg10a, vB10a]. xxvi [Sha06]. xxvii [Gre09, Sim04]. xxx [Han04]. xxxvi [Hol09]. XY [FEVvdB19].

Year [Cas15, Fra10, RS05a, RS06e, Far09a, Far10]. Years [Aus16, CP14b, Per04, RS12f, Wes06, RS06h]. Yes [RS04b]. York [Bed03, Bed06, Bed08, Bed09a, Bed09b, Bed09c, BS11, Ber08b, Blo06, Bri99, Bri07, Bro06a, Bru03, Can01, Cas11, Cas12, Che11, DeW99, Eme09, Eve00, Eve08a, Fel11, Fra04, Fri08, Fri11, Ger03, Ger06a, Gol99a, Hil07, Hob04, Jac03, Jam00, KP02, Lar05c, LM06, Lus04, Lus10, PCGD02, PP08, Per10, Per01, Pes11, RS06b, RS06d, RS07b, RS07c, RS09a, Ros01, Sch00a, Seg10a, Seg10b, Seg15, Sil07, Sim07, Sim12a, Stu06, Tay05, Wal07, vB99, vB01, vB10a, vB10c, vB11]. Yossi [Fra09a]. You’re [FLH85, RS06b]. youth [Gav01].

Z [Bru02, Dom05, Sta02, Mar03b]. Zangwill [MG21]. Zeit [Cas05a, HHH03, Lem07]. Zeno [Pri99, Hug99]. Zeptospace [Che11, Giu10]. Zero [NHW08, Bed09a]. Zertrümmerung [RS03e, Stu04, Tri08]. Zhores [Bor11, Jos10]. zijn [Eic07]. Ziman [Fra01a]. Zionism [Hol09, RS07i]. Zöllner [Kra12]. zum [MPFP19]. Zur [Ste03a, HRS00, RD05]. zwischen [HW08, Sim08].

References

Acocella:2004:EFD

[AGR04] Giovanni Acocella, Francesco Guerra, and Nadia Robotti. Enrico Fermi’s discovery of neutron-induced artificial radioactivity: The recovery of his first laboratory notebook. Physics in Perspective (PIP), 6(1):29–41, April 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

             link.springer.com/article/10.1007/s00016-003-0175-x;
              http://www.springerlink.com/content/5qgmvqy95d3t33ma/

.

Alpher:2001:GBB

[AH01] Ralph Alpher and Robert Herman. Genesis of the Big Bang. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-511182-6. 256 pp. LCCN ????

Ahearne:2002:BRD

[Ahe02] John F. Ahearne. Book review: Daniel S. Greenberg, Science, Money, and Politics: Political Triumph and Ethical Erosion.

Chicago and London: The University of Chicago Press, 2001, viii + 530 pages. $35.00 (cloth). Physics in Perspective (PIP), 4(3):362–365, August 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Alford:2009:BR

[AHSB09] Mark Alford, Jeff Hughes, Gino Segré, and Benjamin Bederson. Book reviews. Physics in Perspective (PIP), 11(3):347– 353, September 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-009-0421-y.

Al-Khalili:1999:BHW

[AK99] Jim Al-Khalili. Black Holes, Wormholes and Time Machines. Institute of Physics Publishers, Bristol, UK, 1999. ISBN 0-7503-0560-6 (paperback), 0-585-26795-2 (e-book). xxii + 265 pp. LCCN QC173.59.S65 A4 1999. URL http://www.loc.

gov/catdir/enhancements/fy0745/99046327-d.html.

Al-Khalili:2012:BHW

[AK12] Jim Al-Khalili. Black holes, wormholes, and time machines. Taylor and Francis, Boca Raton, FL, USA, second edition, 2012. ISBN 1-4398-8559-1 (paperback), 1-4398-8560-5 (e-book). xxii + 184 pp. LCCN QC173.59.S65 A4 2012.

Alford:2009:BRG

[Alf09] Mark Alford. Book review: Gordon Fraser, Cosmic Anger: Abdus Salam — The First Muslim Nobel Scientist. Oxford: Oxford University Press, 2008, xiii + 305 pages. $49.95 (cloth). Physics in Perspective (PIP), 11(3):347–348, September 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Alpher:2012:RAR

[Alp12] Victor S. Alpher. Ralph A. Alpher, Robert C. Herman, and the Cosmic Microwave Background Radiation. Physics in Perspective (PIP), 14(3):300–334, September 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-012-0088-7.

Andriesse:2005:HMB

[AM05] Cornelis Dirk Andriesse and Sally Miedema. Huygens: the man behind the principle. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-85090-8 (hardcover). xxvi + 440 + 8 pp. LCCN Q143.H96 A6413 2005. URL http://

www.loc.gov/catdir/enhancements/fy0632/2005006464-d.
html;   http://www.loc.gov/catdir/enhancements/fy0733/
2005006464-b.html;   http://www.loc.gov/catdir/toc/ecip059/
2005006464.html.

Anastopoulos:2008:PWE

[Ana08] Charis Anastopoulos. Particle or wave: the evolution of the concept of matter in modern physics. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-13512-6 (hardcover). xx + 410 pp. LCCN QC171.2 .A53 2008.

Anderson:2011:MDN

[And11] P. W. (Philip W.) Anderson. More and different: notes from a thoughtful curmudgeon. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2011. ISBN 981-4350-12-5 (hardcover), 981-4350-13-3 (paperback). ix + 412 pp. LCCN Q171 .A527 2011.

Anonymous:1999:CM

[Ano99a] Anonymous. The common man. Physics in Perspective (PIP), 1(3):336, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050024;   http:/
/www.springerlink.com/content/y1h02gr3rv7wurrl/.

Anonymous:1999:L

[Ano99b] Anonymous. Ein limerick. Physics in Perspective (PIP), 1(2): 214, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2000:GVA

[Ano00a] Anonymous. Geometrical versus algebraical thinking. Physics in Perspective (PIP), 2(4):453, December 2000.

Anonymous:2000:MVC

[Ano00b] Anonymous. Mistakes versus crimes. Physics in Perspective (PIP), 2(1):3–5, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2001:AS

[Ano01a] Anonymous. Astrology and science. Physics in Perspective (PIP), 3(4):424, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2001:FMP

[Ano01b] Anonymous. A few more pounds. Physics in Perspective (PIP), 3(3):313, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/PL00000534.

Anonymous:2001:HMS

[Ano01c] Anonymous. How to mix sweets. Physics in Perspective (PIP), 3(1):132, March 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2002:GR

[Ano02] Anonymous. Good rule. Physics in Perspective (PIP), 4(4):495, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2004:BR

[Ano04a] Anonymous. Bohr and the Rabbi. Physics in Perspective (PIP), 6(3):367, September 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2004:HHT

[Ano04b] Anonymous. How to hit a telephone pole. Physics in Perspective (PIP), 6(4):486, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2004:TSP

[Ano04c] Anonymous. Two strong personalities. Physics in Perspective (PIP), 6(2):248, June 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2005:BFI

[Ano05a] Anonymous. Basic Feynman: An interview with Michelle Feynman. Web site, 2005. Comments about the preparation of [Fey05].

Anonymous:2005:RC

[Ano05b] Anonymous. Rutherford is crazy. Physics in Perspective (PIP), 7(2):263, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2010:CP

[Ano10a] Anonymous. Cardinal policy. Physics in Perspective (PIP), 12(1):118, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-010-0016-7.pdf.

Anonymous:2010:TS

[Ano10b] Anonymous. The twelve stages. Physics in Perspective (PIP), 12(4):506, December 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-010-0044-3.pdf.

Anonymous:2011:BRc

[Ano11a] Anonymous. Book reviews. Physics in Perspective (PIP), 13(3): 376–382, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-011-0070-9.

Anonymous:2011:BRd

[Ano11b] Anonymous. Book reviews. Physics in Perspective (PIP), 13(4): 484–497, December 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-011-0073-6.

Anonymous:2011:GV

[Ano11c] Anonymous. Global views. Physics in Perspective (PIP), 13 (1):124, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-011-0054-9.pdf.

Anonymous:2011:HH

[Ano11d] Anonymous. The helping hand. Physics in Perspective (PIP), 13(2):256, June 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-011-0063-8.pdf.

Anonymous:2011:WL

[Ano11e] Anonymous. Who is learned? Physics in Perspective (PIP), 13 (3):383, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-011-0069-2.pdf.

Anonymous:2012:BRc

[Ano12a] Anonymous. Book reviews. Physics in Perspective (PIP), 14(3): 371–387, September 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-012-0100-2.

Anonymous:2012:BRd

[Ano12b] Anonymous. Book reviews. Physics in Perspective (PIP), 14(4): 512–520, December 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-012-0104-y.

Anonymous:2013:BRb

[Ano13a] Anonymous. Book reviews. Physics in Perspective (PIP), 15 (2):244–251, June 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-013-0114-4.

Anonymous:2013:BRc

[Ano13b] Anonymous. Book reviews. Physics in Perspective (PIP), 15(3): 361–370, September 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-013-0121-5.

Anonymous:2013:BRd

[Ano13c] Anonymous. Book reviews. Physics in Perspective (PIP), 15(4): 502–511, December 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-013-0126-0.

Anonymous:2014:BRa

[Ano14a] Anonymous. Book reviews. Physics in Perspective (PIP), 16 (1):129–141, March 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-014-0129-5.

Anonymous:2014:BRb

[Ano14b] Anonymous. Book reviews. Physics in Perspective (PIP), 16 (2):277–289, June 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-014-0135-7.

Anonymous:2014:BRc

[Ano14c] Anonymous. Book reviews. Physics in Perspective (PIP), 16(3): 406–413, September 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-014-0144-6.

Anonymous:2014:BRd

[Ano14d] Anonymous. Book reviews. Physics in Perspective (PIP), 16(4): 480–490, December 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-014-0150-8.

Anonymous:2015:BRa

[Ano15a] Anonymous. Book reviews. Physics in Perspective (PIP), 17 (2):161–168, June 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-015-0163-y.

Anonymous:2015:BRb

[Ano15b] Anonymous. Book reviews. Physics in Perspective (PIP), 17(3): 259–263, September 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-015-0170-z.

Anonymous:2016:BRa

[Ano16a] Anonymous. Book reviews. Physics in Perspective (PIP), 17 (4):361–364, January 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-015-0173-9.

Anonymous:2016:BRb

[Ano16b] Anonymous. Book reviews. Physics in Perspective (PIP), 18 (1):148–154, April 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-016-0177-0.

Anonymous:2016:BRc

[Ano16c] Anonymous. Book reviews. Physics in Perspective (PIP), 18 (2):242–246, August 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-016-0183-2.

Anonymous:2016:BR

[Ano16d] Anonymous. Book reviews. Physics in Perspective (PIP), 18(4): 395–403, December 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-016-0194-z.

Anonymous:2016:DPC

[Ano16e] Anonymous. Do physics conferences still matter? Physics in Perspective (PIP), 18(1):1–2, April 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-016-0179-y;
 http://link.springer.com/content/pdf/10.1007/s00016-
016-0179-y.pdf.

Anonymous:2016:E

[Ano16f] Anonymous. Editorial. Physics in Perspective (PIP), 18(3): 247–248, September 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-016-0191-2.

Anonymous:2016:M

[Ano16g] Anonymous. Megascience. Physics in Perspective (PIP), 18(4): 355–356, December 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-016-0193-0.

Anonymous:2016:PGL

[Ano16h] Anonymous. Physics global and local. Physics in Perspective (PIP), 18(2):155–156, August 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-016-0184-1.

Anonymous:2017:BR

[Ano17] Anonymous. Book reviews. Physics in Perspective (PIP), 19 (1):76–87, April 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-017-0196-5.

Anonymous:2018:BR

[Ano18a] Anonymous. Book review: Edward J. Gillin, The Victorian Palace of Science: Scientific Knowledge and the Building of the Houses of Parliament, Cambridge: Cambridge University Press, 2017, 325 pages, 38 b&w ills., $99.99 (hardback). Physics in Perspective (PIP), 20(4):370–384, December 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2018:BRa

[Ano18b] Anonymous. Book reviews. Physics in Perspective (PIP), 20 (1):159–162, March 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2018:BRb

[Ano18c] Anonymous. Book reviews. Physics in Perspective (PIP), 20(2): 208–217, June 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2020:E

[Ano20] Anonymous. Editorial. Physics in Perspective (PIP), 22(4): 189–190, December 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-020-00268-8.pdf.

Anonymous:2022:AHP

[Ano22] Anonymous. Amateurs in the history of physics. Physics in Perspective (PIP), 24(1):1–2, March 2022. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL https://

link.springer.com/article/10.1007/s00016-022-00288-6.

Atkin:2019:SGP

[AR19] J. Atkin and A. T. Rundle. Serum β_{2}-glycoprotein 1 phenotype frequencies in an English population. Physics in Perspective (PIP), 21(1):81–84, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/content/pdf/10.1007/BF00278570.pdf.

Arianrhod:2003:EHI

[Ari03] Robyn Arianrhod. Einstein’s heroes: imagining the world through the language of mathematics. University of Queensland Press, St Lucia, Queensland, Australia, 2003. ISBN 0-7022-3408-7. viii + 323 pp. LCCN QC19.6 .A75 2003. US$24.00.

Arianrhod:2005:EHI

[Ari05] Robyn Arianrhod. Einstein’s heroes: imagining the world through the language of mathematics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-518370-3 (hardcover), 0-19-530890-5 (paperback). xiii + 323 pp. LCCN QC19.6 .A75 2005. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/toc/ecip053/2004026055.html.

Arns:2001:DN

[Arn01] Robert G. Arns. Detecting the neutrino. Physics in Perspective (PIP), 3(3):314–334, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/PL00000535.

Aurilia:2004:BRG

[Aur04] Antonio Aurilia. Book review: Giuliano Pancaldi, Volta: Science and Culture in the Age of Enlightenment. Princeton: Princeton University Press, 2003, xv + 384 pages. $35.00 (cloth). Physics in Perspective (PIP), 6(4):478–482, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Austin:2016:MSU

[Aus16] Sam M. Austin. The Michigan State University Cyclotron Laboratory: Its early years. Physics in Perspective (PIP), 17(4): 298–333, January 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-015-0175-7.

Badash:2001:NWS

[Bad01] Lawrence Badash. Nuclear winter: Scientists in the political arena. Physics in Perspective (PIP), 3(1):76–105, March 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160050057.

Badash:2005:APN

[Bad05] Lawrence Badash. American physicists, nuclear weapons in World War II, and social responsibility. Physics in Perspective (PIP), 7(2):138–149, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-003-0215-6;

 http://www.springerlink.com/content/u2638016081u076h/

.

Badash:2009:NAL

[Bad09a] Lawrence Badash. The near-appointment of Linus Pauling at the University of California, Santa Barbara. Physics in Perspective (PIP), 11(1):4–14, March 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-008-0387-1.

Badash:2009:NWT

[Bad09b] Lawrence Badash. A nuclear winter’s tale: science and politics in the 1980s. Transformations: studies in the history of science. MIT Press, Cambridge, MA, USA, 2009. ISBN 0-262-01272-3 (hardcover). xiii + 403 pp. LCCN

QC794.98 .B325 2009. URL http://www.loc.gov/catdir/
toc/ecip0827/2008039040.html.

Badino:2009:OCB

[Bad09c] Massimiliano Badino. The odd couple: Boltzmann, Planck and the application of statistics to physics (1900–1913). Annalen der Physik 8 (Berlin, Germany), 18(2–3):81–101, ???? 2009. CODEN ANPYA2. ISSN 0003-3804. See also [Gea02].

Badino:2021:BR

[Bad21] Massimiliano Badino. Book review. Physics in Perspective (PIP), 23(2–3):176–178, October 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Baggott:2004:BMM

[Bag04] J. E. Baggott. Beyond measure: modern physics, philosophy, and the meaning of quantum theory. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-852927-9 (hardcover), 0-19-852536-2 (paperback). xvi + 379 pp. LCCN

QC174.12 .B33 2004. URL http://catdir.loc.gov/catdir/
enhancements/fy0615/2003066207-d.html.

Bais:2007:VSR

[Bai07] Sander Bais. Very Special Relativity: an Illustrated Guide. Harvard University Press, Cambridge, MA, USA, 2007. ISBN 0-674-02611-X (hardcover), 0-674-01967-9 (paperback). 120 pp. LCCN

QC173.65 .B35 2007. URL http://www.loc.gov/catdir/toc/
ecip0712/2007009144.html.

Balibar:2008:AAT

[Bal08] Sébastien Balibar. The Atom and the Apple: Twelve Tales from Contemporary Physics. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-13108-2 (hardcover). x + 190 pp. LCCN QC24.5 .B3513 2008. URL http://www.loc.gov/

catdir/toc/ecip0816/2008018027.html.

Banerjee:2014:CVR

[Ban14] Somaditya Banerjee. C. V. Raman and colonial physics: Acoustics and the quantum. Physics in Perspective (PIP), 16(2): 146–178, June 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-014-0134-8.

Banerjee:2016:TQQ

[Ban16] Somaditya Banerjee. Transnational quantum: Quantum physics in India through the lens of Satyendranath Bose. Physics in Perspective (PIP), 18(2):157–181, August 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-016-0182-3.

Barrow:1999:BIS

[Bar99a] John D. Barrow. Between inner space and outer space: essays on science, art, and philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-850254-0. xii + 274 pp. LCCN Q162 .B364 1999.

Barschall:1999:R

[Bar99b] H. H. Barschall. Reminiscences. Physics in Perspective (PIP), 1(4):390–444, December 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050030.

Barr:2003:MPA

[Bar03] Stephen M. Barr. Modern Physics and Ancient Faith. University of Notre Dame Press, Notre Dame, IN, USA, 2003. ISBN 0-268-03471-0 (hardcover). ix + 312 pp. LCCN BL265.P4 B37 2003.

Bernardini:2004:EFH

[BB04] C. (Carlo) Bernardini and Luisa Bonolis, editors. Enrico Fermi: his work and legacy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 88-7438-015-1 (SIF, Bologna, Italy), 3-540-22141-7 (Springer-Verlag). xii + 410 pp. LCCN QC774.F4 C6613 2004. URL http://

www.loc.gov/catdir/enhancements/fy0818/2004108212-d.
html;   http://www.loc.gov/catdir/enhancements/fy0818/
2004108212-t.html.

Buchwald:2001:INN

[BC01] Jed Z. Buchwald and I. Bernard Cohen. Isaac Newton’s natural philosophy. Dibner Institute studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-02477-2 (hardcover). xx + 354 pp. LCCN Q143.N495 I8 2001.

Bennett:2003:LLL

[BCHJ03] Jim Bennett, Michael Cooper, Michael Hunter, and Lisa Jardine, editors. London’s Leonardo: the life and work of Robert Hooke. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-852579-6. xii + 224 pp. LCCN Q143.H7 L66 2003. URL http://www.e-

streams.com/es0612/es0612_2943.html;   http://www.h-
net.org/review/hrev-a0e0c8-aa;   http://www.loc.gov/
catdir/enhancements/fy0616/2003545106-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0616/2003545106-
t.html;   http://www.loc.gov/catdir/enhancements/fy0723/
2003545106-b.html.

Bederson:2001:BRM

[Bed01a] Benjamin Bederson. Book review: Mary Palevsky, Atomic Fragments: A Daughter’s Questions. Berkeley and Los Angeles: University of California Press, 2000, xiv + 289 pages. $24.95 (cloth). Physics in Perspective (PIP), 3(2):253–254, June 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bederson:2001:SAP

[Bed01b] Benjamin Bederson. SEDs at Los Alamos: a personal memoir. Physics in Perspective (PIP), 3(1):52–75, March 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160050056.

Bederson:2003:PTP

[Bed03] Benjamin Bederson. The physical tourist: Physics and New York City. Physics in Perspective (PIP), 5(1):87–121, April 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s000160300004.

Bederson:2005:AFR

[Bed05] Benjamin Bederson. In appreciation: Fritz Reiche and the Emergency Committee in Aid of Displaced Foreign Scholars. Physics in Perspective (PIP), 7(4):453–472, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-005-0245-3.

Bederson:2006:BRK

[Bed06] B. Bederson. Book review: Kai Bird and Martin J. Sherwin, American Prometheus: The Triumph and Tragedy of J. Robert Oppenheimer. New York: Alfred A Knopf, 2005, xiii + 721 pages. $35.00. (cloth). Physics in Perspective (PIP), 8(2):226– 229, May 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bederson:2008:BRC

[Bed08] Benjamin Bederson. Book review: Cynthia C. Kelly, ed., The Manhattan Project: The Birth of the Atomic Bomb in the Words of its Creators, Eyewitnesses, and Historians. New York: Black Dog and Leventhal Publishers, 2007, 495 pages. $24.95 (cloth). Physics in Perspective (PIP), 10(3):372–374, September 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bederson:2009:BRB

[Bed09a] Benjamin Bederson. Book review: Bruce Rosenblum and Fred Kuttner, Quantum Enigma: Physics Encounters Consciousness. New York: Oxford University Press, 2008, 224 pages. $15.95 (paper). Günther Nimtz and Astrid Haibel, Zero Time Space: How Quantum Tunneling Broke the Light Speed Barrier. Translated from the German by Hans-Joachim Nimtz. Weinheim: Wiley-VCH Press, 2008, xvi + 150 pages. $34.95 (paper). Physics in Perspective (PIP), 11(4):462–464, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bederson:2009:BRD

[Bed09b] Benjamin Bederson. Book review: David C. Cassidy, Beyond Uncertainty: Heisenberg, Quantum Physics, and the Bomb. New York: Bellevue Literary Press, 2009, 456 pages. $27.00 (cloth).

Physics in Perspective (PIP), 11(3):351–353, September 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bederson:2009:BRS

[Bed09c] Benjamin Bederson. Book review: Sheilla Jones, The Quantum Ten: A Story of Passion, Tragedy, Ambition, and Science. New York: Oxford University Press, 2008. xii + 340 pages. $22.95 (paper). Physics in Perspective (PIP), 11(2):231–232, June 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bederson:2012:DEM

[Bed12] Benjamin Bederson. Double exile: Migrations of Jewish– Hungarian professionals through Germany to the United States, 1919–1945. Physics in Perspective (PIP), 14(3):376–384, September 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Beeley:2015:BRS

[Bee15] Philip Beeley. Book review: Sarah Dry, The Newton Papers: The Strange and True Odyssey of Isaac Newton’s Manuscripts. Oxford: Oxford University Press, 2014, xi + 238 pages. $29.95. Physics in Perspective (PIP), 17(1):74–79, March 2015. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
015-0156-x.

Berez:2008:BR

[BEF^{+}08] Thomas M. Berez, William E. Evenson, Kenneth W. Ford, Ruth Lewin Sime, Jay M. Pasachoff, and Naomi Pasachoff. Book reviews. Physics in Perspective (PIP), 10(4):484–491, December 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-006-0364-5.

Beller:1999:QDM

[Bel99] Mara Beller. Quantum dialogue: the making of a revolution. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, 1999. ISBN 0-226-04181-6 (hardcover), 0-226-04182-4. xv + 365 + 8 pp. LCCN QC174.13 .B45 1999. URL http://www.loc.

gov/catdir/description/uchi052/99035499.htm;   http:
//www.loc.gov/catdir/toc/uchi052/99035499.htm.

Belfer:2014:IPJ

[Bel14] Israel Belfer. Informing physics: Jacob Bekenstein and the informational turn in theoretical physics. Physics in Perspective (PIP), 16(1):69–97, March 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-014-0132-x.

Bergeron:2002:TID

[Ber02a] Kenneth D. Bergeron. Tritium on ice: the dangerous new alliance of nuclear weapons and nuclear power. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-02527-2 (hardcover). x + 234 pp. LCCN TK9023 .B47 2002. URL http://www.loc.

gov/catdir/toc/fy036/2002066016.html;   https://www.
math.utah.edu/pub/tex/bib/physperspect.bib.

Bernardini:2002:CFN

[Ber02b] Carlo Bernardini. Conoscere Fermi nel centenario della nascita: 29 settembre 1901–2001. (Italian) [Learn about Fermi in the centenary of his birth: 29 September 1901–2001]. Edizioni scientifiche SIF, Bologna, Italy, second edition, 2002. ISBN 88-7438-000-3. viii + 383 pp. LCCN QC774.F4. 10.33 EUR. URL http:

//www.gbv.de/dms/casalini/02/02220733.pdf;   http://
www.sif.it/fermiindex.html; http://www.sif.it/libri/
conoscere_fermi;   http://www.zentralblatt-math.org/
zmath/en/search/?an=1073.01512. English translation in

[BB04].

Bernstein:2003:DWH

[Ber03] Jeremy Bernstein. The drawing or why history is not mathematics. Physics in Perspective (PIP), 5(3):243–261, September 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-003-0177-8.

Bernardini:2004:AFE

[Ber04a] Carlo Bernardini. AdA: The first electron–positron collider. Physics in Perspective (PIP), 6(2):156–183, June 2004. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
003-0202-y.

Bernstein:2004:OPE

[Ber04b] Jeremy Bernstein. Oppenheimer: portrait of an enigma. Ivan R. Dee, Chicago, IL, USA, 2004. ISBN 1-56663-569-1. xi + 223 pp. LCCN QC16.O62 B43 2004.

Beranek:2008:RWL

[Ber08a] Leo Leroy Beranek. Riding the Waves: a Life in Sound, Science, and Industry. MIT Press, Cambridge, MA, USA, 2008. ISBN 0-262-02629-5 (hardcover). x + 235 pp. LCCN TA140.B385 A3 2008.

Berez:2008:BRK

[Ber08b] Thomas M. Berez. Book review: Klaus Hentschel, The Mental Aftermath: The Mentality of German Physicists, 1945–1949, translated from the German by Ann M. Hentschel. New York: Oxford University Press, 2007, vi + 205 pages. $49.50 (cloth). Physics in Perspective (PIP), 10(4):484–485, December 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bernstein:2008:JBI

[Ber08c] Jeremy Bernstein. John Bell and the identical twins. Physics in Perspective (PIP), 10(3):269–286, September 2008. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
008-0380-8.

Bernstein:2008:NWW

[Ber08d] Jeremy Bernstein. Nuclear Weapons: What You Need to Know. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-88408-X. xi + 299 pp. LCCN

U264 .B453 2008.  URL http://www.loc.gov/catdir/
enhancements/fy0743/2007016625-b.html;   http://www.
loc.gov/catdir/enhancements/fy0743/2007016625-d.html;
 http://www.loc.gov/catdir/toc/ecip0716/2007016625.html.

Bernstein:2009:QL

[Ber09] Jeremy Bernstein. Quantum leaps. Belknap Press of Harvard University Press, Cambridge, MA, USA; London, UK, 2009. ISBN 0-674-03541-0. vi + 230 pp. LCCN QC174.13 .B47 2009.

Beranek:2010:RWL

[Ber10a] Leo Leroy Beranek. Riding the Waves: a Life in Sound, Science, and Industry. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-51399-4 (paperback). ???? pp. LCCN ????

Bernstein:2010:JNK

[Ber10b] Jeremy Bernstein. John von Neumann and Klaus Fuchs: an unlikely collaboration. Physics in Perspective (PIP), 12(1): 36–50, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-009-0001-1;   http://www.
springerlink.com/content/y54804274h88v546/.

Berlincourt:2016:SNT

[Ber16] T. G. Berlincourt. Superconducting niobium–titanium: Enabler for affordable MRI and the search for the Higgs boson. Physics in Perspective (PIP), 17(4):334–353, January 2016. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
015-0172-x.

Bethe:2000:SS

[Bet00] Hans A. Bethe. Sommerfeld’s Seminar. Physics in Perspective (PIP), 2(1):3–5, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050033;   http:/
/www.springerlink.com/content/2h5uxxgpn6pvngx7/fulltext.
pdf.

Badino:2013:MPL

[BF13] Massimiliano Badino and Bretislav Friedrich. Much polyphony but little harmony: Otto Sackur’s groping for a quantum theory of gases. Physics in Perspective (PIP), 15(3):295–319, September 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-013-0110-8.

Bederson:2009:BR

[BHF^{+}09] Benjamin Bederson, Donald E. Hall, Allan Franklin, David Goodstein, and Kenneth W. Ford. Book reviews. Physics in Perspective (PIP), 11(2):231–239, June 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-008-0394-2.

Brinkman:2007:BR

[BHS07] William Brinkman, Gerald Holton, and Mark P. Silverman. Book reviews. Physics in Perspective (PIP), 9(2):256–262, June 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-006-0323-1.

Brown:2006:HBHa

[BL06] G. E. (Gerald Edward) Brown and Chang-Hwan Lee, editors. Hans Bethe and his physics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2006. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). xiii + 314 pp. LCCN QC16.B46 H36 2006. URL http:

//www.worldscibooks.com/physics/5989.html;   http://
www.worldscientific.com/worldscibooks/10.1142/5989.

Blackmore:2000:BRC

[Bla00a] John Blackmore. Book review: Carlo Cercignani, Ludwig Boltzmann: The Man Who Trusted Atoms. Oxford: Oxford University Press, 1998, xvii, 327 pages. $49.95 (cloth). Physics in Perspective (PIP), 2(1):108–111, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Blanpied:2000:BRG

[Bla00b] William A. Blanpied. Book review: George Perkovich, India’s Nuclear Bomb: The Impact on Global Proliferation. Berkeley: University of California Press, 1999, xiii + 597 pages $39.95 (cloth). Physics in Perspective (PIP), 2(4):447–450, December 2000.

Bloembergen:2006:BRJ

[Blo06] N. Bloembergen. Book review: Jeff Hecht, Beam: The Race to Make the Laser. New York: Oxford University Press, 2005, x + 274 pages. $29.95 (cloth). Physics in Perspective (PIP), 8 (2):226, May 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bokulich:2008:RQC

[Bok08] Alisa Bokulich. Reexamining the quantum–classical relation: beyond reductionism and pluralism. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-85720-1. x + 195 pp.

LCCN QA805 .B685 2008. URL http://www.loc.gov/catdir/
enhancements/fy0834/2008018902-b.html;   http://www.
loc.gov/catdir/enhancements/fy0834/2008018902-d.html;
 http://www.loc.gov/catdir/enhancements/fy0834/2008018902-
t.html.

Bonolis:2011:BRR

[Bon11] Luisa Bonolis. Bruno Rossi and the racial laws of Fascist Italy. Physics in Perspective (PIP), 13(1):58–90, March 2011. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
010-0035-4.

Bond:2018:FAB

[Bon18] Peter D. Bond. The fiftieth anniversary of Brookhaven National Laboratory: A turbulent time. Physics in Perspective (PIP), 20 (2):180–207, June 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Borowitz:2008:NE

[Bor08] Sidney Borowitz. The Norwegian and the Englishman. Physics in Perspective (PIP), 10(3):287–294, September 2008. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
007-0372-0.

Borowitz:2011:BRP

[Bor11] Sidney Borowitz. Book review: Paul Josephson, Lenin’s Laureate: Zhores Alferov’s Life in Communist Science. Cambridge, Mass.: The MIT Press, 2010, 307 pages. $29.95 (cloth). Physics in Perspective (PIP), 13(4):486–493, December 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Borresen:2012:FNP

[Bør12] Hans Christofer Børresen. Flawed nuclear physics and atomic intelligence in the campaign to deny Norwegian heavy water to Germany, 1942–1944. Physics in Perspective (PIP), 14(4): 471–497, December 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-012-0094-9.

Boyd:2012:GHB

[Boy12] T. James M. Boyd. George Hartley Bryan, Ludwig Boltzmann, and the stability of flight. Physics in Perspective (PIP), 14 (1):4–32, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-011-0077-2.

Berry:2008:PTP

[BP08] Michael Berry and Brian Pollard. The physical tourist: Physics in Bristol. Physics in Perspective (PIP), 10(4):468–480, December 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-008-0398-y.

Brinkman:1999:BRC

[Bri99] William F. Brinkman. Book review: Charles H. Townes, How the Laser Happened: Adventures of a Scientist. New York: Oxford University Press, 1999. 200 pages. $35.00 (cloth). Physics in Perspective (PIP), 1(4):459–460, December 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Brian:2005:CBM

[Bri05] Denis Brian. The Curies: a biography of the most controversial family in science. John Wiley, New York, NY, USA, 2005. ISBN 0-471-27391-0. ix + 438 pp. LCCN ????

Brinkman:2007:BRJ

[Bri07] William E. Brinkman. Book review: Joel N. Shurkin, Broken Genius: The Rise and Fall of William Shockley, Creator of the Electronic Age. New York: Macmillan, 2006, x + 297 pages. $24.95 (cloth). Physics in Perspective (PIP), 9(2):256, June 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Brown:1999:RHW

[Bro99] Louis Brown. A radar history of World War II: technical and military imperatives. Institute of Physics Publishers, Bristol, UK, 1999. ISBN 0-7503-0659-9. xvi + 563 pp. LCCN D810.R33 B77 1999. URL http://

www.h-net.org/review/hrev-a0c4o5-aa;   http://www.h-
net.org/review/hrev-a0c4p3-aa;   http://www.loc.gov/
catdir/enhancements/fy0668/99038978-d.html.

Bromley:2002:CP

[Bro02] D. Allan (David Allan) Bromley. A century of physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95247-0. x + 114 pp.

LCCN QC7 .B68 2002. URL http://www.loc.gov/catdir/
enhancements/fy0816/2001018410-d.html;   http://www.
loc.gov/catdir/enhancements/fy0816/2001018410-t.html.

Brown:2004:BRS

[Bro04] Laurie M. Brown. Book review: Steven Weinberg, The Discovery of Subatomic Particles. Cambridge: Cambridge University Press, 2003, xvi + 206. $25.00 (cloth). Physics in Perspective (PIP), 6(4):485–486, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Brown:2005:FTN

[Bro05] Laurie M. Brown. Feynman’s Thesis — a New Approach To Quantum Theory. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2005. ISBN 981-256-366-0 (print), 981-256-763-1 (e-book). 142 pp.

LCCN QC174.12 .F49 2005. URL http://public.eblib.com/
EBLPublic/PublicView.do?ptiID=244532;   http://www.worldscibooks.
com/promotion/feynman.html#5852.

Brown:2006:BRM

[Bro06a] Laurie M. Brown. Book review: Michelle Feynman, ed., Perfectly Reasonable Deviations from the Beaten Track: The Letters of Richard P. Feynman. New York, Basic Books, 2005, xxiii + 486 pages. $26.00 (cloth). Physics in Perspective (PIP), 8(4): 473–474, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Brown:2006:PMD

[Bro06b] Laurie M. Brown. Paul A. M. Dirac’s The Principles of Quantum Mechanics. Physics in Perspective (PIP), 8(4):381– 407, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0276-4;   http://www.
springerlink.com/content/n5354k7513351301/.

Brown:2013:TRS

[Bro13] Harvey R. Brown. The theory of the rise of sap in trees: Some historical and conceptual remarks. Physics in Perspective (PIP), 15(3):320–358, September 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-013-0117-1.

Brush:1999:BRN

[Bru99a] Stephen G. Brush. Book review: Neil A. Porter, Physicists in Conflict: From Antiquity to the new Millennium. Bristol and Philadelphia: Institute of Physics Publishing, 1998. xv + 275 pages. $60.00 (cloth). Physics in Perspective (PIP), 1(3):339– 341, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Brush:1999:WWR

[Bru99b] Stephen G. Brush. Why was Relativity accepted? Physics in Perspective (PIP), 1(2):184–214, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s000160050015.

Brush:2002:BRJ

[Bru02] Stephen G. Brush. Book review: Jed Z. Buchwald and Andrew Warwick, ed., Histories of the electron: The birth of microphysics. Cambridge, Massachusetts: The MIT Press, 2001, xi + 514 pages. $55.00 (cloth). Physics in Perspective (PIP), 4(4): 492–493, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Brush:2003:BRR

[Bru03] Stephen G. Brush. Book review: Robert Marc Friedman, The Politics of Excellence: Behind the Nobel Prize in Science. New York: A. H. Freeman/Times Books/Henry Holt, 2001, xv + 379 pages. $30.00 (cloth). Physics in Perspective (PIP), 5(2):235– 238, May 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Brush:2006:BRM

[Bru06] Stephen G. Brush. Book review: Mauro Dardo, Nobel Laureates and Twentieth-Century Physics. Cambridge: Cambridge University Press, 2004, xi + 533 pages. $85.00 (cloth). Physics in Perspective (PIP), 8(1):105–106, March 2006. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bromberg:2004:PP

[BS04] J. L. Bromberg and A. L. Schawlow. The playful physicist. Physics in Perspective (PIP), 6(3):310–343, September 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/

s00016-003-0173-x. This interview with Arthur L. Schawlow is adapted from an interview conducted by Joan Lisa Bromberg on January 19, 1984, at Stanford University.

Bird:2005:APT

[BS05] Kai Bird and Martin J. Sherwin. American Prometheus: the triumph and tragedy of J. Robert Oppenheimer. Alfred A. Knopf, New York, NY, USA, 2005. ISBN 0-375-72626-8, 0-375-41202-6. xiii + 721 + 32 pp. LCCN QC16.O62 B57 2005. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
http://site.ebrary.com/lib/ucsc/Doc?id=10078784.

Bederson:2011:HNY

[BS11] Benjamin Bederson and H. Henry Stroke. History of the New York University Physics Department. Physics in Perspective (PIP), 13(3):260–328, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-011-0056-7.

Buchwald:2016:PMI

[Buc16] Jed Z. Buchwald. Politics, morality, innovation, and misrepresentation in physical science and technology. Physics in Perspective (PIP), 18(3):283–300, September 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-016-0187-y.

Bucurescu:2020:RSM

[Buc20] Dorel Bucurescu. A reconsideration of Stefania Maracineanu’s measurements of polonium-210’s half-life: Understanding her claim to the discovery of artificial radioactivity. Physics in Perspective (PIP), 22(3):162–181, September 2020. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Butterfield:2000:BRR

[But00] Jeremy Butterfield. Book review: Roberto Torretti, The Philosophy of Physics. Cambridge: Cambridge University Press, 1999, xvi + 512 pages. $23.95 (cloth). Physics in Perspective (PIP), 2(3):327–328, September 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Butterfield:2019:SHL

[But19] Jeremy Butterfield. Book review: Sabine Hossenfelder, Lost in Math: How Beauty Leads Physics Astray, Basic Books, 2018, 304 pages, $17.99 (hardcover). Physics in Perspective (PIP), 21(1):63–86, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bacciagaluppi:2009:QTCb

[BV09] Guido Bacciagaluppi and Antony Valentini, editors. Quantum theory at the crossroads: reconsidering the 1927 Solvay Conference; [with a complete translation of the original proceedings]. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-81421-9. LCCN QC173.96. URL http://www.loc.gov/

catdir/toc/ecip0817/2008019585.html.

Byers:2006:SCT

[BW06] Nina Byers and Gary A. Williams, editors. Out of the Shadows: Contributions of Twentieth-Century Women to Physics. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-82197-5 (hardcover). xxv + 471 pp. LCCN QC15 .O94

2006. URL http://catdir.loc.gov/catdir/enhancements/
fy0703/2006299534-d.html;   http://catdir.loc.gov/catdir/
enhancements/fy0703/2006299534-t.html;   http://catdir.
loc.gov/catdir/enhancements/fy0732/2006299534-b.html.

Calle:2009:UOD

[Cal09] Carlos I. Calle. The universe: order without design. Prometheus Books, Amherst, NY, USA, 2009. ISBN 1-59102-714-4 (hardcover). 304 + 8 pp. LCCN QB981 .C345 2009.

Canaday:2000:NML

[Can00] John Canaday. The Nuclear Muse: Literature, Physics, and the First Atomic Bombs. University of Wisconsin Press, Madison, WI, USA, 2000. ISBN 0-299-16850-6 (hardcover), 0-299-16854-9 (paperback). xviii + 310 pp. LCCN QC791.96 .C36 2000. URL

http://www.e-streams.com/es0409/es0409_1493.htm.

Cantor:2001:BRC

[Can01] Geoffrey Cantor. Book review: Colin A. Russell, Michael Faraday: Physics and Faith. New York: Oxford University Press, 2000, 128 pages. $22.00 (cloth). Physics in Perspective (PIP), 3(3):373–375, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Careri:2000:LO

[Car00] Giorgio Careri. Lars, the oracle. Physics in Perspective (PIP), 2(2):204–210, June 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s000160050042.

Carlsson:2007:BRE

[Car07] Anders Carlsson. Book review: Eric Schneider and Dorion Sagan, Into the Cool: Energy Flow, Thermodynamics, and Life. Chicago: University of Chicago Press, 2005, xix + 362 pages. $22.50 (paper). Physics in Perspective (PIP), 9(4):514–515, December 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Carson:2010:HAA

[Car10] Cathryn Carson. Heisenberg in the atomic age: science and the public sphere. Publications of the German Historical Institute. German Historical Institute, Washington, DC, USA, 2010. ISBN 0-521-82170-3 (hardcover). xvi + 541 + 8 pp. LCCN QC16.H45 C37 2010.

Carson:2011:BRS

[Car11] Cathryn Carson. Book review: Suman Seth, Crafting the Quantum: Arnold Sommerfeld and the Practice of Theory, 1890– 1926. Cambridge, Mass. and London: The MIT Press, 2010, xix + 378 pages. $32.00 (cloth). Physics in Perspective (PIP), 13(1):118–120, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Cassidy:2002:NLC

[Cas02] David C. Cassidy. New light on “Copenhagen” and the German Nuclear Project. Physics in Perspective (PIP), 4(4):447– 455, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s000160200004.

Cassidy:2005:BRW

[Cas05a] D. C. Cassidy. Book review: Werner Heisenberg, Liebe Eltern! Briefe aus kritischer Zeit 1918 bis 1945. [(German) [Dear parents! Letters from a critical time 1918 to 1945]]. Edited by Anna Maria Hirsch-Heisenberg. With a foreword by Helmut Rechenberg. Munich, Germany: Langen Müller in der F. A. Herbig Verlagsbuchhandlung, 2003, 400 pages. EUR26.00 (cloth). Physics in Perspective (PIP), 7(2):253–256, June 2005. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Cassidy:2005:JRO

[Cas05b] David C. Cassidy. J. Robert Oppenheimer and the American century. Pi Press, New York, NY, USA, 2005. ISBN 0-13-147996-2. xviii + 462 + 16 pp. LCCN QC16.O62 C37 2005.

Cassidy:2009:BUH

[Cas09] David C. Cassidy. Beyond uncertainty: Heisenberg, quantum physics, and the bomb. Bellevue Literary Press, New York, NY, USA, 2009. ISBN 1-934137-13-8. 480 pp. LCCN QC16.W518 C37 2008.

Cassidy:2010:BRL

[Cas10] David C. Cassidy. Book review: Lawrence Badash, A Nuclear Winter’s Tale: Science and Politics in the 1980s. Cambridge, Mass: The MIT Press, 2009, xiii + 403 pages. $40.00 (cloth). Physics in Perspective (PIP), 12(3):358–361, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Cassidy:2011:BRC

[Cas11] David C. Cassidy. Book review: Cathryn Carson, Heisenberg in the Atomic Age: Science and the Public Sphere. Washington, DC: German Historical Institute and New York: Cambridge University Press, 2010, xvi + 541 pages. $80.00 (cloth). Physics in Perspective (PIP), 13(2):250–255, June 2011. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Cassidy:2012:BRG

[Cas12] David C. Cassidy. Book review: Gino Segrè, Ordinary Geniuses: Max Delbrück, George Gamow, and the Origins of Genomics and Big Bang Cosmology, Viking Adult, New York, 2011, xxi + 330 pages, ISBN-13 978-0-670-02276-2. Physics in Perspective (PIP), 14(2):248–250, June 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Cassidy:2015:WHC

[Cas15] David C. Cassidy. Werner Heisenberg and Carl Friedrich Freiherr von Weizsäcker: A fifty-year friendship. Physics in Perspective (PIP), 17(1):33–54, March 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-014-0152-6.

Cassidy:2017:FHG

[Cas17] David C. In Cassidy. Farm Hall and the German Atomic Project of World War II: a dramatic history. Springer International Publishing, Cham, Switzerland, 2017. ISBN 3-319-59577-6 (print), 3-319-59578-4 (e-book). xiv + 125 + 32 pp. LCCN PS3553.A87 F37 2017.

Crispino:2016:AIG

[CdL16] Luís C. B. Crispino and Marcelo C. de Lima. Amazonia introduced to general relativity: The May 29, 1919, solar eclipse from a North-Brazilian point of view. Physics in Perspective (PIP), 18(4):379–394, December 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-016-0190-3.

Cercignani:1998:LBM

[Cer98] Carlo Cercignani. Ludwig Boltzmann: the man who trusted atoms. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-850154-4 (hardcover), 0-19-857064-3 (paperback). xvi + 329 pp. LCCN QC16.B64

C47 1998. URL http://www.loc.gov/catdir/enhancements/
fy0605/98017743-d.html;   http://www.loc.gov/catdir/
enhancements/fy0605/98017743-t.html. Foreword by Roger

Penrose.

Cheng:2011:BR

[CFH^{+}11] Ta-Pei Cheng, Bernard J. Feldman, Jacob Darwin Hamblin, Cathryn Carson, and Elisha Cohn. Book reviews. Physics in Perspective (PIP), 13(1):114–123, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
011-0052-y.

Chen:2019:CAF

[CFL19] Andrew T. L. Chen, Arthur Falek, and William Lester. Chromosome aberrations in full-term low birth weight neonates. Physics in Perspective (PIP), 21(1):13–16, March 2019. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/BF00278559.pdf.

Cassidy:2005:BR

[CGYC05] David C. Cassidy, Karl Giberson, Nai-Chang Yeh, and Robert N. Compton. Book reviews. Physics in Perspective (PIP), 7(2):253–263, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-003-0219-2.

Chalmers:2001:MMN

[Cha01] Alan Chalmers. Maxwell, mechanism, and the nature of electricity. Physics in Perspective (PIP), 3(4):425–438, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/PL00000539.

Chang:2002:RRR

[Cha02a] Hasok Chang. Rumford and the reflection of radiant cold: Historical reflections and metaphysical reflexes. Physics in Perspective (PIP), 4(2):127–169, May 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-002-8362-8.

Charap:2002:EUN

[Cha02b] John M. Charap. Explaining the Universe: the New Age of Physics. Princeton University Press, Princeton, NJ, USA, 2002. ISBN 0-691-00663-6. xii + 226 pp. LCCN QC21.3 .C48 2002. URL http://www.loc.gov/

catdir/description/prin021/2001058840.html;   http:/
/www.loc.gov/catdir/samples/prin031/2001058840.html;
 http://www.loc.gov/catdir/toc/prin031/2001058840.html.

Chalmers:2018:HPB

[Cha18] Alan Chalmers. How pressure became a scalar, not a vector. Physics in Perspective (PIP), 20(2):165–179, June 2018. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Chakraborty:2019:NNM

[Cha19] Ranajit Chakraborty. A note on Nei’s measure of gene diversity in a substructured population. Physics in Perspective (PIP), 21(1):85–88, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/BF00278571.pdf.

Chen:2001:MRP

[Che01] Xiang Chen. Measuring reflective power with the eye. Physics in Perspective (PIP), 3(4):439–461, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://link.springer.com/article/10.1007/PL00000540.

Cheng:2011:BRG

[Che11] Ta-Pei Cheng. Book review: Gian Francesco Giudice, A Zeptospace Odyssey: A Journey into the Physics of the LHC. Oxford and New York: Oxford University Press, 2010, ii + 276 pages. $45.00 (cloth). Physics in Perspective (PIP), 13(1):114– 116, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Clark:2009:SKU

[Cla09] Stuart (Stuart G.) Clark. The Sun Kings: the unexpected tragedy of Richard Carrington and the tale of how modern astronomy began. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-14126-6 (paperback). xii + 211

pp. LCCN M09.E06300. URL http://press.princeton.edu/
titles/8370.html.

Close:2009:NVS

[Clo09] Frank E. Close. Nothing: a very short introduction. Very short introductions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-922586-9 (paperback). 157 pp. LCCN QC6 .C588 2009.

Close:2015:HLD

[Clo15] Frank E. Close. Half-life: the divided life of Bruno Pontecorvo, physicist or spy. Basic Books, New York, NY, USA, 2015. ISBN 0-465-06998-3 (hardcover), 0-465-04487-5 (e-book), 1-78074-582-6 (e-book). xix + 378 pp. LCCN QC774.P66 C56 2014. US$29.99.

Crease:2017:E

[CMP17a] Robert P. Crease, Joseph D. Martin, and Peter Pesic. Editorial. Physics in Perspective (PIP), 19(4):319–320, December 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00016-017-0211-x.pdf.

Crease:2017:GH

[CMP17b] Robert P. Crease, Joseph D. Martin, and Peter Pesic. Gonzo history. Physics in Perspective (PIP), 19(2):89–90, June 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL https://link.springer.com/content/pdf/10.
1007/s00016-017-0201-z.pdf.

Crease:2017:IP

[CMP17c] Robert P. Crease, Joseph D. Martin, and Peter Pesic. Immigrant physics. Physics in Perspective (PIP), 19(1):1–2, April 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00016-017-0198-3.pdf.

Crease:2017:WPF

[CMP17d] Robert P. Crease, Joseph D. Martin, and Peter Pesic. Where is the physics frontier? Physics in Perspective (PIP), 19(3): 181–182, September 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-017-0208-5.pdf.

Crease:2018:PH

[CMP18a] Robert Crease, Joseph D. Martin, and Peter Pesic. Physics is its history. Physics in Perspective (PIP), 20(4):315–317, December 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00016-018-0231-1.pdf.

Crease:2018:E

[CMP18b] Robert P. Crease, Joseph D. Martin, and Peter Pesic. Editorial. Physics in Perspective (PIP), 20(1):1–3, March 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00016-018-0217-z.pdf.

Crease:2018:MS

[CMP18c] Robert P. Crease, Joseph D. Martin, and Peter Pesic. On ”minor” scientists. Physics in Perspective (PIP), 20(3):219– 220, September 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-018-0226-y.pdf.

Crease:2018:WSP

[CMP18d] Robert P. Crease, Joseph D. Martin, and Peter Pesic. When science and politics collide. Physics in Perspective (PIP), 20 (2):163–164, June 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-018-0222-2.pdf.

Crease:2019:BHP

[CMS19a] Robert P. Crease, Joseph D. Martin, and Richard Staley. Biography and the history of physics. Physics in Perspective (PIP), 21(3):161–162, September 2019. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00016-019-00245-w.pdf.

Crease:2019:DPL

[CMS19b] Robert P. Crease, Joseph D. Martin, and Richard Staley. Decolonizing physics: Learning from the periphery. Physics in Perspective (PIP), 21(2):91–92, June 2019. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00016-019-00240-1.pdf.

Crease:2019:PNP

[CMS19c] Robert P. Crease, Joseph D. Martin, and Richard Staley. Physics and (natural) philosophy. Physics in Perspective (PIP), 21(4):255–256, December 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00016-019-00249-6.
pdf.

Crease:2019:WWT

[CMS19d] Robert P. Crease, Joseph D. Martin, and Richard Staley. What we talk about when we talk about physics. Physics in Perspective (PIP), 21(1):1–2, March 2019. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00016-019-00237-w.pdf.

Crease:2020:DLI

[CMS20a] Robert P. Crease, Joseph D. Martin, and Richard Staley. Don’t be like Ike. Physics in Perspective (PIP), 22(2):53–54, June 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00016-020-00257-x.pdf.

Crease:2020:RHP

[CMS20b] Robert P. Crease, Joseph D. Martin, and Richard Staley. Recentering the history of physics. Physics in Perspective (PIP), 22(1):1–2, March 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-020-00253-1.pdf.

Crease:2021:PL

[CMS21a] Robert P. Crease, Joseph D. Martin, and Richard Staley. E pluribus … ? Physics in Perspective (PIP), 23(4):179–180, December 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00016-021-00282-4.pdf.

Crease:2021:WTE

[CMS21b] Robert P. Crease, Joseph D. Martin, and Richard Staley. Writing things up: Endings and beginnings. Physics in Perspective (PIP), 23(1):1–2, March 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00016-021-00272-6.
pdf.

Coen:2007:VAU

[Coe07] Deborah R. Coen. Vienna in the age of uncertainty: science, liberalism, and private life. University of Chicago Press, Chicago, IL, USA and London, UK, 2007. ISBN 0-226-11172-5 (cloth). xi + 380 pp. LCCN DB844.E98 C64 2007. URL http://

www.loc.gov/catdir/enhancements/fy0707/2007003723-b.
html;   http://www.loc.gov/catdir/enhancements/fy0707/
2007003723-d.html;   http://www.loc.gov/catdir/toc/ecip0710/
2007003723.html.

Cohen:2000:HAP

[Coh00] I. Bernard Cohen. Howard Aiken: portrait of a computer pioneer. History of computing. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-03262-7 (hardcover), 0-262-53179-8 (paperback). xx + 329 + 24 pp. LCCN QA76.2.A35 C65 2000.

Cohn:2011:BRB

[Coh11] Elisha Cohn. Book review: Barri J. Gold, ThermoPoetics: Energy in Victorian Literature and Science, Cambridge, Mass.: The MIT Press, 2010, x + 343 pages. $30.00 (cloth). Physics in Perspective (PIP), 13(1):120–123, March 2011. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Coleman:2000:BRJ

[Col00] Lawrence A. Coleman. Book review: Jim Al-Khalili, Black Holes, Wormholes and Time Machines. London: Institute of Physics, 1999, xxii + 265 pages. $16.50 (cloth). Physics in Perspective (PIP), 2(3):328–330, September 2000. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Compton:2005:BRK

[Com05] R. N. Compton. Book review: Kenneth W. Ford, The Quantum World: Quantum Physics for Everyone. Cambridge, Mass. and London: Harvard University Press, 2004, xii + 270 pages. $24.95 (cloth). Physics in Perspective (PIP), 7(2):262–263, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Cooper:1999:EFR

[Coo99] Dan Cooper. Enrico Fermi and the revolutions in modern physics. Oxford portraits in science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-511762-X. 117 pp. LCCN QC16.F46 C66 1999. URL http://www.loc.

gov/catdir/enhancements/fy0603/98034471-d.html.

Coopersmith:2010:ESC

[Coo10] Jennifer Coopersmith. Energy, the subtle concept: the discovery of Feynman’s blocks from Leibniz to Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-954650-9 (hardcover). xiv + 400 pp. LCCN QC72 .C66 2010.

Cooper:2011:EGP

[Coo11] David K. C. Cooper. Edward Gerjuoy: From physics to law and back again. Physics in Perspective (PIP), 13(4):433–455, December 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-011-0072-7.

Crease:2014:E

[CP14a] Robert P. Crease and Peter Pesic. Editorial. Physics in Perspective (PIP), 16(1):1–2, March 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s00016-
014-0130-z;   http://link.springer.com/article/10.1007/
s00016-014-0130-z.

Crease:2014:FLF

[CP14b] Robert P. Crease and Peter Pesic. The Feynman lectures, fifty years on. Physics in Perspective (PIP), 16(2):143–145, June 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-014-0137-5.

Crease:2014:PM

[CP14c] Robert P. Crease and Peter Pesic. Physics and music. Physics in Perspective (PIP), 16(4):415–416, December 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://link.springer.com/article/10.1007/s00016-014-
0147-3;   http://link.springer.com/content/pdf/10.1007/
s00016-014-0147-3.pdf.

Crease:2014:RPC

[CP14d] Robert P. Crease and Peter Pesic. Reading physics closely. Physics in Perspective (PIP), 16(3):291–292, September 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-014-0143-7;   http://link.springer.com/content/
pdf/10.1007/s00016-014-0143-7.pdf.

Crease:2015:EB

[CP15a] Robert P. Crease and Peter Pesic. Editorial: “But Is It Science?”. Physics in Perspective (PIP), 17(1):1–2, March 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-015-0154-z;   http://link.springer.com/content/
pdf/10.1007/s00016-015-0154-z.pdf.

Crease:2015:ES

[CP15b] Robert P. Crease and Peter Pesic. Explorers and settlers. Physics in Perspective (PIP), 17(3):171–172, September 2015.

CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-015-0169-5;   http://link.springer.com/content/
pdf/10.1007/s00016-015-0169-5.pdf.

Crease:2015:WHI

[CP15c] Robert P. Crease and Peter Pesic. Whose history is it? Physics in Perspective (PIP), 17(2):81–82, June 2015. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://link.springer.com/article/10.1007/s00016-015-
0161-0;   http://link.springer.com/content/pdf/10.1007/
s00016-015-0161-0.pdf.

Crane:2000:HWH

[Cra00] H. Richard Crane. How we happened to measure g − 2: a tale of serendipity. Physics in Perspective (PIP), 2(2):135–140, June 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s000160050040.

Crease:1999:MPB

[Cre99] Robert P. Crease. Making Physics: a Biography of Brookhaven National Laboratory, 1946–1972. University of Chicago Press, Chicago, IL, USA and London, UK, 1999. ISBN 0-226-12017-1. xii + 434 pp. LCCN QC789.2.U62 B763 1999.

URL http://www.loc.gov/catdir/description/uchi052/
98030327.html;   http://www.loc.gov/catdir/enhancements/
fy0608/98030327-t.html;   http://www.loc.gov/catdir/
enhancements/fy0609/98030327-b.html.

Crease:2005:QISa

[Cre05a] Robert P. Crease. Quenched! The ISABELLE saga, I. Physics in Perspective (PIP), 7(3):330–376, September 2005. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
004-0246-7.

Crease:2005:QISb

[Cre05b] Robert P. Crease. Quenched! The ISABELLE saga, II. Physics in Perspective (PIP), 7(4):404–452, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-005-0247-1.

Crelinsten:2006:EJR

[Cre06] Jeffrey Crelinsten. Einstein’s jury: the race to test relativity. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-12310-1. xxix + 397 pp. LCCN

QC173.585 .C74 2006. URL http://www.loc.gov/catdir/
enhancements/fy0654/2005032681-d.html;   http://www.
loc.gov/catdir/enhancements/fy0734/2005032681-b.html;
 http://www.loc.gov/catdir/toc/ecip063/2005032681.html.

Crease:2008:NSL

[Cre08] Robert P. Crease. The National Synchrotron Light Source, Part I: Bright idea. Physics in Perspective (PIP), 10(4):438–467, December 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-007-0357-z.

Crease:2009:NSL

[Cre09a] Robert P. Crease. The National Synchrotron Light Source, Part II: The bakeout. Physics in Perspective (PIP), 11(1):15–45, March 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-007-0358-y.

Crepeau:2009:LSS

[Cre09b] John Crepeau. Loschmidt, Stefan, and Stigler’s Law of Eponymy. Physics in Perspective (PIP), 11(4):357–378, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-009-0420-z.

Crease:2012:BRM

[Cre12] Robert P. Crease. Book review: Mark A. Peterson, Galileo’s Muse: Renaissance Mathematics and the Arts. Cambridge: Harvard University Press, 2011, vi + 336 pages. $28.95 (cloth). Physics in Perspective (PIP), 14(2):250–252, June 2012. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Crease:2017:FBL

[Cre17] Robert P. Crease. Francis Bacon’s London. Physics in Perspective (PIP), 19(3):291–306, September 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Crowe:1999:PDH

[Cro99] Michael J. Crowe. Pierre Duhem, the history and philosophy of physics, and the teaching of physics. Physics in Perspective (PIP), 1(1):54–64, March 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050005.

Cropper:2001:GPL

[Cro01] William H. Cropper. Great physicists: the life and times of leading physicists from Galileo to Hawking. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-513748-5. xii + 500 pp. LCCN QC15 .C76 2001. URL http://

www.loc.gov/catdir/enhancements/fy0637/2001021611-d.
html;   http://www.loc.gov/catdir/enhancements/fy0724/
2001021611-b.html.

Crowe:2007:MAE

[Cro07] Michael J. Crowe. Mechanics from Aristotle to Einstein. Green Lion Press, Santa Fe, NM, USA, 2007. ISBN 1-888009-32-2 (paperback). xxii + 331 pp. LCCN QA802 .C76 2007. URL

http://www.loc.gov/catdir/toc/fy0803/2007925664.html.

Crease:2013:PPU

[CS13] Robert P. Crease and Vladimir Shiltsev. Pomor polymath: The upbringing of Mikhail Vasilyevich Lomonosov, 1711–1730. Physics in Perspective (PIP), 15(4):391–414, December 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-013-0113-5.

Crease:2018:FPM

[CS18] Robert P. Crease and Vladimir Shiltsev. Fueling Peter’s Mill: Mikhail Lomonosov’s educational training in Russia and Germany, 1731–1741. Physics in Perspective (PIP), 20(3):272–304, September 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Cushing:2000:BRH

[Cus00] James T. Cushing. Book review: Helge Kragh, Quantum Generations: A History of Physics in the Twentieth Century. Princeton: Princeton University Press, 1999, xiv + 494 pages. $29.95 (cloth). Physics in Perspective (PIP), 2(2):217–218, June 2000.

CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Cushing:2002:BRB

[Cus02] James T. Cushing. Book review: Nature loves to hide: Quantum physics and reality, a Western perspective. Physics in Perspective (PIP), 4(2):245–247, May 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Cantor:2001:BR

[CvB01] Geoffrey Cantor and Hans Christian von Baeyer. Book reviews. Physics in Perspective (PIP), 3(3):373–376, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
PL00000537.

DAgostino:2000:DTM

[D’A00] Salvo D’Agostino. On the difficulties of the transition from Maxwell’s and Hertz’s pure-field theories to Lorentz’s electron. Physics in Perspective (PIP), 2(4):398–410, December 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160050052.

DAgostino:2002:RND

[D’A02] Salvo D’Agostino. From rational numbers to Dirac’s bra and ket: Symbolic representation of physical laws. Physics in Perspective (PIP), 4(2):216–229, May 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-002-8364-6.

DAgostino:2004:BCP

[D’A04] Salvo D’Agostino. The Bild conception of physical theory: Helmholtz, Hertz, and Schrödinger. Physics in Perspective (PIP), 6(4):372–389, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-003-0205-8.

Dahl:1997:FCR

[Dah97] Per F. Dahl. Flash of the cathode rays: a history of J. J. Thomson’s electron. Institute of Physics Publishers, Bristol, UK, 1997. ISBN 0-7503-0453-7 (hardcover), 0-585-20853-0 (e-book). xvii + 526 pp. LCCN 99.E01389.

Dahl:1999:HWW

[Dah99] Per F. Dahl. Heavy Water and the Wartime Race for Nuclear Energy. Institute of Physics Publishers, Bristol, UK, 1999. ISBN 0-7503-0633-5 (hardcover). xvi + 399 pp. LCCN QD169.W3

D25 1999. URL http://www.loc.gov/catdir/enhancements/
fy0661/99033672-d.html.

Dahl:2002:NTN

[Dah02] Per F. Dahl. From nuclear transmutation to nuclear fission, 1932–1939. Institute of Physics Publishers, Bristol, UK, 2002. ISBN 0-7503-0865-6. xii + 304 pp. LCCN QC790

.D33 2002. URL http://www.loc.gov/catdir/enhancements/
fy0806/2007298521-d.html.

Dahl:2006:BRI

[Dah06a] P. F. Dahl. Book review: Iwan Rhys Morus, When Physics Became King. Chicago and London: The University of Chicago Press, 2005, xii + 303 pages. $25.00 (paper). Physics in Perspective (PIP), 8(2):225, May 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Dahl:2006:BPH

[Dah06b] Per F. Dahl. Berkeley and its physics heritage. Physics in Perspective (PIP), 8(1):90–101, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-005-0252-4.

Dahmen:2009:BAF

[Dah09] Silvio R. Dahmen. Boltzmann and the art of flying. Physics in Perspective (PIP), 11(3):244–260, September 2009. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
008-0395-1.

Darrigol:2000:EAE

[Dar00] Olivier Darrigol. Electrodynamics from Ampère to Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850594-9. xix + 532 pp. LCCN 0.2dar a0165 a4100.

Dardo:2004:NLT

[Dar04] M. (Mauro) Dardo. Nobel laureates and Twentieth-Century physics. Cambridge University Press, Cambridge, UK, 2004.

ISBN 0-521-83247-0, 0-521-54008-9 (paperback). xi + 533 pp. LCCN QC7 .D27 2004. URL http://www.loc.gov/

catdir/description/cam041/2004049240.html;   http://
www.loc.gov/catdir/toc/cam041/2004049240.html.

Dasgupta:2015:SPS

[Das15] Deepanwita Dasgupta. Stars, peripheral scientists, and equations: The case of M. N. Saha. Physics in Perspective (PIP), 17(2):83–106, June 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-015-0159-7.

Dasgupta:2018:BRE

[Das18] Deepanwita Dasgupta. Book reviews: Eric Scerri, A Tale of Seven Scientists and a New Philosophy of Science, Oxford: Oxford University Press, 2016, 264 pages, $29.95 (hardcover). Physics in Perspective (PIP), 20(3):310–313, September 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Dasgupta:2019:MRM

[Das19] Deepanwita Dasgupta. Making research more diverse: How peripheral members join a scientific community. Physics in Perspective (PIP), 21(2):93–107, June 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Dasgupta:2021:BR

[Das21] Deepanwita Dasgupta. Book review. Physics in Perspective (PIP), 23(4):231–234, December 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Davidson:2000:CSL

[Dav00a] Keay Davidson. Carl Sagan: a life. John Wiley, New York, NY, USA, 2000. ISBN 0-471-39536-6. xx + 540 pp. LCCN QB36.S15 D38 2000.

Davis:2000:UCR

[Dav00b] Martin Davis. The universal computer: the road from Leibniz to Turing. W. W. Norton & Co., New York, NY, USA, 2000. ISBN 0-393-04785-7. xii + 257 pp. LCCN QA76.17. D38 2000; QA76.17 .D38 2000.

Davies:2010:WBM

[Dav10] Edward Brian Davies. Why beliefs matter: reflections on the nature of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-958620-9. x + 250 pp.

LCCN BL240.3. URL http://www.gbv.de/dms/bowker/toc/
9780199586202.pdf.

Day:1999:PTS

[Day99] Peter Day, editor. The philosopher’s tree: a selection of Michael Faraday’s writings. Institute of Physics Publishers, Bristol, UK, 1999. ISBN 0-7503-0571-1 (paperback), 0-7503-0570-3 (hardcover). xv + 211 pp. LCCN QC16.F2

A3 1999. URL http://www.loc.gov/catdir/enhancements/
fy0668/98054683-d.html.

Day:2002:BRB

[Day02] Michael A. Day. Book review: The one culture? A conversation about science. Physics in Perspective (PIP), 4(2):248–250, May 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Day:2004:ART

[Day04] Michael A. Day. In appreciation: I. I. Rabi: The two cultures and the universal culture of science. Physics in Perspective (PIP), 6(4):428–476, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-003-0206-7.

Day:2008:UCS

[Day08] Michael A. Day. E. U. Condon: Science, religion, and the politics of world peace. Physics in Perspective (PIP), 10(1):4–55, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-007-0337-3.

Dahl:2006:BR

[DBBG06] Per F. Dahl, Nicolaas Bloembergen, Benjamin Bederson, and Thomas B. Greenslade, Jr. Book reviews. Physics in Perspective (PIP), 8(2):225–231, May 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-005-0282-y.

Drago:2007:EMC

[DE07] Antonino Drago and Salvatore Esposito. Ettore Majorana’s course on theoretical physics: a recent discovery. Physics in Perspective (PIP), 9(3):329–345, September 2007. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
006-0305-3.

Denny:2007:IFM

[Den07] Mark Denny. Ingenium: five machines that changed the world. The Johns Hopkins University Press, Baltimore, MD, USA, 2007. ISBN 0-8018-8586-8. xvi + 176 pp.

LCCN T15 .D3417 2007. URL http://www.loc.gov/catdir/
enhancements/fy0701/2006026085-b.html;   http://www.
loc.gov/catdir/enhancements/fy0701/2006026085-d.html;
 http://www.loc.gov/catdir/toc/ecip0619/2006026085.html.

DeVries:2000:BRA

[DeV00] Paul L. DeVries. Book review: Anthony J. G. Hey, ed., Feynman and Computation: Exploring the Limits of Computers. Reading: Perseus Books, 1999, xxiii + 438 pages. $50.00 (Hardbound). Physics in Perspective (PIP), 2(2):220, June 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

DeWitt:1999:BRJ

[DeW99] Bryce DeWitt. Book review: John Archibald Wheeler with Kenneth Ford, Geons, Black Holes and Quantum Foam, A Life in Physics. New York and London: W. W. Norton, 1998, 380 pages. $27.95 (cloth). Physics in Perspective (PIP), 1(2):224–225, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Dias:2019:WMA

[DFAB19] Penha Maria Cardozo Dias, Mariana Faria Brito Francisquini, Carlos Eduardo Aguiar, and Marta Feijó Barroso. What the middle-aged Galileo told the elderly Galileo: Galileo’s search for the laws of fall. Physics in Perspective (PIP), 21(3):194–221, September 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Davies:2010:INR

[DG10] P. C. W. Davies and Niels Henrik Gregersen, editors. Information and the nature of reality: from physics to metaphysics.

Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-76225-1 (hardback). xvi + 382 pp. LCCN BD701 .I49

2010. URL http://assets.cambridge.org/97805217/62250/
cover/9780521762250.jpg.

Diehl:2005:AEV

[Die05] Renee D. Diehl. Albert Einstein’s vision: remarkable discoveries that shaped modern science. Physics in Perspective (PIP), 7(4): 491–492, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Dirac:2009:PQM

[Dir09] Paul Adrien Maurice Dirac. The Principles of Quantum Mechanics, volume 27 of The international series of monographs on physics. Clarendon Press, Oxford, UK, fourth revised reprinted edition, 2009. ISBN 0-19-852011-5. xii + 314 pp. LCCN QC6 .D55 2009.

Dragoni:2007:QME

[DMA07] Giorgio Dragoni, Giulio Maltese, and Luisa Atti. Quirino Majorana’s experiments on the speed of light and gravitational absorption. Physics in Perspective (PIP), 9(3):281–304, September 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-006-0304-4.

Domski:2005:BRJ

[Dom05] Mary Domski. Book review: Jed Z. Buchwald and I. Bernard Cohen, ed., Isaac Newton’s Natural Philosophy [Dibner Institute Studies in the History of Science and Technology]. Cambridge, Mass.: The MIT Press, 2001, xx + 354 pages. $55.00 (cloth), $22.00 (paper). Physics in Perspective (PIP), 7(3):377–380, September 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Dorfman:2006:BRD

[Dor06] J. R. Dorfman. Book review: David Lindley, Degrees Kelvin: A Tale of Genius, Invention, and Tragedy. Washington, DC: Joseph Henry Press, 2004. viii + 366, $27.95 (cloth). Physics in Perspective (PIP), 8(1):107–109, March 2006. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

dosSantosFitas:2007:GBA

[dP07] Augusto José dos Santos Fitas and António Augusto Passos Videira. Guido Beck, Alexandre Proca, and the Oporto Theoretical Physics Seminar. Physics in Perspective (PIP), 9(1):4–25, January 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-006-0278-2.

Dry:2014:NPS

[Dry14] Sarah Dry. The Newton papers: the strange and true odyssey of Isaac Newton’s manuscripts. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-995104-7 (hardcover). xi + 238 pp. LCCN QC16.N7 D79 2014.

Dragoni:2013:PSB

[DS13] Giorgio Dragoni and Ivana Stojanovic. Physical science in Bologna. Physics in Perspective (PIP), 15(1):92–115, March 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-013-0108-2.

DelSanto:2022:PUV

[DS22] Flavio Del Santo and Emanuel Schwarzhans. “Philosophysics” at the University of Vienna: The (pre-)History of foundations of quantum physics in the Viennese cultural context. Physics in Perspective (PIP), 24(2–3):125–153, October 2022. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL https://link.springer.com/article/10.
1007/s00016-022-00290-y.

Domski:2005:BR

[DSFG05] Mary Domski, Daniel F. Styer, A. P. French, and Daniel M. Greenberger. Book reviews. Physics in Perspective (PIP), 7(3): 377–386, September 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-003-0243-2.

Durham:2003:EU

[Dur03] Ian T. Durham. Eddington and uncertainty. Physics in Perspective (PIP), 5(4):398–418, December 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-003-0172-0.

deWaal:2020:CFG

[dWtH20] Elske de Waal and Sjang L. ten Hagen. The concept of fact in German physics around 1900: A comparison between Mach and Einstein. Physics in Perspective (PIP), 22(2):55–80, June 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00016-020-00256-y.pdf.

Eckert:1999:MET

[Eck99] Michael Eckert. Mathematics, experiments, and theoretical physics: The early days of the Sommerfeld School. Physics in Perspective (PIP), 1(3):238–252, October 1999. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160050021.

Eckert:2023:BR

[Eck23] Michael Eckert. Book review. Physics in Perspective (PIP), 25 (1–2):76–78, June 2023. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL https://link.springer.

com/article/10.1007/s00016-023-00300-7.

Einstein:2005:MRb

[EG05] Albert Einstein and Brian Greene. The Meaning of Relativity, volume 1921 of The Stafford Little lectures. Princeton University Press, Princeton, NJ, USA, expanded Princeton science library edition, 2005. ISBN 0-691-12027-7. xxiv + 166 pp. LCCN QC6.

URL http://press.princeton.edu/titles/484.html.

Ehrlich:2004:BRJ

[Ehr04] Robert Ehrlich. Book review: John M. Charap, Explaining the Universe: The New Age of Physics. Princeton: Princeton University Press, 2002, xii + 226, $19.95 (cloth). Physics in Perspective (PIP), 6(4):478, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Ehrlich:2006:WMT

[Ehr06] Robert Ehrlich. What makes a theory testable, or is intelligent design less scientific than string theory? Physics in Perspective (PIP), 8(1):83–89, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-005-0279-6.

Eickhoff:2007:NWP

[Eic07] Martijn Eickhoff. In naam der wetenschap?: P. J. W. Debye en zijn carrière in nazi-Duitsland. Nederlands Instituut voor Oorlogsdocumentatie [== Dutch Institute for War Documentation], Amsterdam, The Netherlands, 2007. ISBN ???? 195 pp. LCCN ????

Eickhoff:2008:NSP

[Eic08] Martijn Eickhoff. In the name of science?: P. J. W. Debye and his career in Nazi Germany, volume 2 of Studies of the Netherlands Institute for War Documentation. Aksant, Amsterdam, The Netherlands, 2008. ISBN 90-5260-327-8. 184 pp. LCCN QC16.D43 E36 2008.

Eisenstaedt:2002:ERG

[Eis02] Jean Eisenstaedt. Einstein et la relativité générale: les chemins de l’espace-temps. (French) [Einstein and General Relativity: the ways of space-time]. CNRS histoire des sciences. CNRS éditions, Paris, France, 2002. ISBN 2-271-05880-5. ISSN 1292-4296. 344 pp. LCCN QC173.55. EUR 29.

Eisenstaedt:2006:CHR

[Eis06] Jean Eisenstaedt. The Curious History of Relativity: How Einstein’s Theory of Gravity Was Lost and Found Again, volume 15 of CNRS editions. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-11865-5. ix + 363 pp. LCCN QC173.6 EIS; 06.E06047.

Emery:2009:BRJ

[Eme09] Guy Emery. Book review: John Jenkin, William and Lawrence Bragg, Father and Son: the most extraordinary collaboration in science. New York: Oxford University Press, 2008, xv + 458 pages. $70.00 (cloth). Physics in Perspective (PIP), 11(4): 457–462, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Emery:2011:BRD

[Eme11] Guy Emery. Book review: David E. Fisher, Much Ado About (Practically) Nothing: The History of the Noble Gases. Oxford: Oxford University Press, 2010, xi + 264 pages. $24.95 (cloth). Physics in Perspective (PIP), 13(4):484–486, December 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Esposito:2021:EII

[Esp21] Salvatore Esposito. From England to Italy: The intriguing story of Poli’s engine for the King of Naples. Physics in Perspective (PIP), 23(2–3):104–138, October 2021. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00016-021-00277-1.pdf.

Essen:2010:LEV

[Ess10] Ray Essen. Louis Essen and the velocity of light: From wartime radar to unit of length. Physics in Perspective (PIP), 12 (1):51–73, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-009-0004-y.

Evenson:2008:BR

[EvBF^{+}08] William E. Evenson, Hans Christian von Baeyer, Bernard J. Feldman, Benjamin Bederson, Lee G. Sobotka, and Hans C. Ohanian. Book reviews. Physics in Perspective (PIP), 10(3): 368–376, September 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-008-0393-3.

Evenson:2000:BRJ

[Eve00] William E. Evenson. Book review: John D. Barrow, Between Inner Space and Outer Space: Essays on Science, Art, and Philosophy, New York: Oxford University Press, 1999, 304 pp, $30.00 (cloth). Physics in Perspective (PIP), 2(1):112–113, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Evenson:2004:BRA

[Eve04] W. E. Evenson. Book review: Allan Franklin, Selectivity and Discord: Two Problems of Experiment. Pittsburgh: University of Pittsburgh Press, 2002, xii + 290 pages. $37.50 (cloth). Physics in Perspective (PIP), 6(1):119–121, April 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Evenson:2006:WAS

[Eve06] William E. Evenson. The world of Andrei Sakharov: a Russian physicist’s path to freedom. Physics in Perspective (PIP), 8(4): 480–481, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Evenson:2008:BRN

[Eve08a] William E. Evenson. Book review: Nina Byers and Gary Williams, ed., Out of the Shadows: Contributions of Twentieth-Century Women to Physics. Cambridge and New York: Cambridge University Press, 2006, 498 pages. $35.00 (cloth). Physics in Perspective (PIP), 10(3):368–369, September 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Evenson:2008:BRR

[Eve08b] William E. Evenson. Book review: Richard J. Weiss, A Physicist Remembers. Singapore: World Scientific, 2007, ix + 261 pages. $42.00 (cloth). Physics in Perspective (PIP), 10(4):485–486, December 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Evenson:2012:KHD

[Eve12] William Evenson. Knocking on heaven’s door: How physics and scientific thinking illuminate the universe and the modern world. Physics in Perspective (PIP), 14(3):371–373, September 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

College:1982:MU

[Fac82] Faculty of Lynchburg College, editor. Man and the Universe, volume 1 (series 2) of Classical selections on great issues. University Press of America, Washington, DC, USA, 1982. ISBN 0-8191-2295-5, 0-8191-2252-1 (paperback). xvi + 370 pp. LCCN Q125 .M316 1982.

Fang:2016:GE

[Fan16] Zaiqing Fang. Guest editorial. Physics in Perspective (PIP), 17(4):265–267, January 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-015-0176-6;
 http://link.springer.com/content/pdf/10.1007/s00016-
015-0176-6.pdf.

Fara:2009:SFT

[Far09a] Patricia Fara. Science: a four thousand year history. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-922689-X. xv + 408 pp. LCCN Q125 .F27 2009.

URL http://www.loc.gov/catdir/enhancements/fy0909/
2008050975-b.html;   http://www.loc.gov/catdir/enhancements/
fy0909/2008050975-d.html;   http://www.loc.gov/catdir/
toc/fy0904/2008050975.html.

Farmelo:2009:SMH

[Far09b] Graham Farmelo. The Strangest Man: The Hidden Life of Paul Dirac, Mystic of the Atom. Basic Books, New York, NY, USA, 2009. ISBN 0-465-01827-0. 539 + 8 pp. LCCN QC16.D57; QC16.D57 F37 2009. US$29.95. URL http://bnreview.

barnesandnoble.com/t5/The-Thinking-Read/The-Strangest-
Man-The-Hidden-Life-of-Paul-Dirac-Mystic-of-the/ba-p/
1243.

Fara:2010:SFT

[Far10] Patricia Fara. Science: a four thousand year history. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-922689-X, 0-19-958027-8. xviii + 482 pp.

LCCN Q125. URL http://www.gbv.de/dms/bowker/toc/
9780199580279.pdf;   http://www.gbv.de/dms/faz-rez/FD1200910282479995.
pdf.

Feldman:2008:BRM

[Fel08] Bernard J. Feldman. Book review: Mark Denny, Ingenium: Five Machines that Changed the World. Baltimore: The John Hopkins University Press, 2007, xvi + 176 pages. $25.00 (cloth). Physics in Perspective (PIP), 10(3):371–372, September 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Feldman:2011:BRW

[Fel11] Bernard J. Feldman. Book review: Weston M. Stacey, The Quest for a Fusion Energy Reactor. New York: Oxford University Press, 2010, 208 pages. $24.95 (cloth). Physics in Perspective (PIP), 13(1):116–117, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Feldman:2012:N

[Fel12] Bernard J. Feldman. Neutrino. Physics in Perspective (PIP), 14(4):519–520, December 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Ferris:2002:SW

[Fer02] Timothy Ferris. On science writing. Physics in Perspective (PIP), 4(1):3–12, February 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-002-8356-6.

Ferris:2010:SLD

[Fer10] Timothy Ferris. The science of liberty: democracy, reason, and the laws of nature. Harper, New York, NY, USA, 2010. ISBN 0-06-078150-5. 368 pp. LCCN Q175.5 .F477 2010.

Fryns:2019:PTK

[FEVvdB19] J. P. Fryns, E. Eggermont, H. Verresen, and H. van den Berghe. Partial trisomy 13: Karyotype 46,XY,-6,+t(13q,6q). Physics in Perspective (PIP), 21(1):47–54, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/content/pdf/10.1007/BF00278564.pdf.

Feynman:2005:PRD

[Fey05] Michelle Feynman, editor. Perfectly Reasonable Deviations (from the Beaten Track): the Collected Letters of Richard P. Feynman. Basic Books, New York, NY, USA, 2005. ISBN 0-7382-0636-9. xxiii + 486 pp. LCCN QC16 .F49

A4 2005. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/toc/ecip055/

2005000049.html. Foreword by Timothy Ferris. See also interview with the editor [Ano05a].

Feynman:2005:DYT

[FF05] Richard P. (Richard Phillips) Feynman and Michelle Feynman. Don’t you have time to think? Allen Lane, London, UK, 2005. ISBN 0-7139-9847-4 (hardcover). xxi + 486 pp. LCCN 06.E02646. US$20.00.

Franklin:2001:BR

[FFF^{+}01] Allan Franklin, Kenneth W. Ford, Hans Frauenfelder, A. Truman Schwartz, and Thomas D. Rossing. Book reviews. Physics in Perspective (PIP), 3(4):490–497, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://link.springer.com/article/10.1007/PL00000542.

Fox:2005:POL

[FG05] Robert Fox and Graeme Gooday, editors. Physics in Oxford, 1839–1939: laboratories, learning, and college life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-856792-8. xix + 363 pp. LCCN

QC47.G73 O947 2005. URL http://www.loc.gov/catdir/
enhancements/fy0639/2005296514-d.html;   http://www.
loc.gov/catdir/enhancements/fy0725/2005296514-b.html;
 http://www.loc.gov/catdir/toc/fy0601/2005296514.html.

Feynman:2006:FTP

[FGL^{+}06] Richard P. (Richard Phillips) Feynman, Michael A. Gottlieb, Ralph Leighton, Matthew L. (Matthew Linzee) Sands, Robert B. Leighton, and Rochus Vogt. Feynman’s tips on physics: a problem-solving supplement to the Feynman lectures on physics. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-8053-9063-4 (paperback). xiii + 162 pp. LCCN QC23

.F47 1989 Suppl. URL http://www.loc.gov/catdir/toc/
ecip0512/2005013077.html.

Feynman:2012:FTP

[FGL^{+}12] Richard P. (Richard Phillips) Feynman, Michael A. Gottlieb, Ralph Leighton, Matthew L. (Matthew Linzee) Sands, Robert B. Leighton, and Rochus Vogt, editors. Feynman’s tips on physics: reflections, advice, insights, practice. Basic Books, New York, NY, USA, second edition, 2012. ISBN 0-465-02797-0 (hardcover), 0-465-02921-3 (e-book). xiv + 182 pp. LCCN ????

Firestein:2012:IHI

[Fir12] Stuart Firestein. Ignorance: How It Drives Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-982807-5 (hardcover), 0-19-982808-3 (e-book), 1-280-59536-1. viii + 195 pp. LCCN Q175.32.K45 F57 2012.

Fisher:2010:MAA

[Fis10] David E. Fisher. Much ado about (practically) nothing: a history of the noble gases. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-539396-1 (hardcover). x + 264 pp. LCCN QD162 .F57 2010.

Fix:2015:SSM

[Fix15] Adam Fix. A science superior to music: Joseph Sauveur and the estrangement between music and acoustics. Physics in Perspective (PIP), 17(3):173–197, September 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-015-0164-x.

Florio:2021:PLL

[FJ21] Victória Flório and Olival Freire Júnior. The past looks like an onion: The centennial ”Great Debate” through journalists’ testimonies. Physics in Perspective (PIP), 23(2–3):85–103, October 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Feynman:1988:WDY

[FL88] Richard Phillips Feynman and Ralph Leighton. What do YOU care what other people think?: Further adventures of a curious character. W. W. Norton & Co., New York, NY, USA, 1988. ISBN 0-393-02659-0. 255 pp. LCCN QC16.F49 A3 1988. US$17.95.

Feynman:2006:CFA

[FL06] Richard Phillips Feynman and Ralph Leighton. Classic Feynman: All the Adventures of a Curious Character. W. W. Norton & Co., New York, NY, USA, 2006. ISBN 0-393-06132-9. x + 511 pp. LCCN QC16.F49 A3 2006. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/;   http://www.loc.
gov/catdir/toc/ecip0515/2005018928.html.

Flake:1998:CBN

[Fla98] Gary William Flake. The computational beauty of nature: computer explorations of fractals, chaos, complex systems, and adaptation. MIT Press, Cambridge, MA, USA, 1998. ISBN 0-262-06200-3 (hardcover). xviii + 493 pp. LCCN QA76.6 .F557 1998.

Fleck:2021:FTP

[Fle21a] Robert Fleck. Fundamental themes in physics from the history of art. Physics in Perspective (PIP), 23(1):25–48, March 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Fleck:2021:SRA

[Fle21b] Robert Fleck. The scientific revolution in art. Physics in Perspective (PIP), 23(2–3):139–169, October 2021. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Feynman:1985:SYJc

[FLH85] Richard P. Feynman, Ralph Leighton, and Edward Hutchings. “Surely You’re Joking, Mr. Feynman!”: Adventures of a Curious Character. Bantam Books, New York, NY, USA, 1985. ISBN 0-553-25649-1. xi + 322 pp. LCCN QC16.F49A37 1986.

Focaccia:2022:JBP

[Foc22] Miriam Focaccia. Not just boys at Via Panisperna: Women at the Royal Physics Institute in Rome. Physics in Perspective (PIP), 24(2–3):154–177, October 2022. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL https://link.

springer.com/article/10.1007/s00016-022-00291-x.

Folse:2005:QMP

[Fol05] Henry J. Folse. Quantum mechanics and the philosophy of Alfred North Whitehead. Physics in Perspective (PIP), 7(4): 494–496, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Ford:2001:BRJ

[For01] Kenneth W. Ford. Book review: John Canaday, The Nuclear Muse: Literature, Physics and the First Atomic Bomb. Madison: University of Wisconsin Press, 2000, xviii + 303 pages. $22.95 (paper), $60.00 (cloth). Physics in Perspective (PIP), 3 (4):492–493, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Ford:2002:BRS

[For02] Kenneth W. Ford. Book review: Steven Weinberg, Facing Up: Science and Its Cultural Adversaries. Cambridge, Massachusetts: Harvard University Press, 2001, xiv + 283 pages. $26.00 (cloth). Physics in Perspective (PIP), 4(3):365–366, August 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Ford:2004:QWQ

[For04] Kenneth William Ford. The quantum world: quantum physics for everyone. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01342-5 (cased), 0-674-01832-X (paperback). ix + 270 (hardcover), ix + 294 (paperback) pp. LCCN QC174.12 .F68 2004.

Ford:2008:BRJ

[For08] Kenneth W. Ford. Book review: Jeremy Bernstein, Nuclear Weapons: What You Need to Know. Cambridge: Cambridge University Press, 2007, xii + 299 pages. $27.00 (cloth). Physics in Perspective (PIP), 10(4):486–487, December 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Ford:2009:BRC

[For09] Kenneth W. Ford. Book review: Charis Anastopoulos, Particle or Wave: The Evolution of the Concept of Matter in Modern Physics. Princeton: Princeton University Press, 2008, xx + 410 pages. $35.00 (cloth). Physics in Perspective (PIP), 11(2):238– 239, June 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Forstner:2020:BRO

[For20] Christian Forstner. Book review: Olival Freire Junior, David Bohm: A Life Dedicated to Understanding of the Quantum World, Cham: Springer, 2019, xi + 250 pp., $84.99 (hardcover). Physics in Perspective (PIP), 22(3):185–187, September 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00016-020-00260-2.pdf.

Forstner:2022:LLI

[For22] Christian Forstner. Laboratory life instead of nuclear weapons: a new perspective on the German Uranium Club. Physics in Perspective (PIP), 24(4):181–207, December 2022. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s00016-022-00294-8.

Fritzsch:2011:YWM

[FR11] Harald Fritzsch and Jeanne Rostant. You are wrong, Mr. Einstein!: Newton, Einstein, Heisenberg and Feynman discussing quantum mechanics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2011. ISBN 981-4324-99-X (hardcover). xxi + 1 + 178 pp. LCCN ????

Franklin:1999:RE

[Fra99] A. Franklin. The roles of experiment. Physics in Perspective (PIP), 1(1):35–53, March 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050004.

Franklin:2001:BRJ

[Fra01a] Allan Franklin. Book review: John Ziman, Real Science: What it Is and What it Means. Cambridge: Cambridge University Press, 2000, xii + 399 pages. $39.95 (cloth). Physics in Perspective (PIP), 3(4):490–491, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Frauenfelder:2001:BRS

[Fra01b] Hans Frauenfelder. Book review: S. S. Schweber, In the Shadow of the Bomb: Bethe, Oppenheimer, and the Moral Responsibility of the Scientist. Princeton: Princeton University Press, 2000, xviii + 260 pages. $24.95 (cloth). Physics in Perspective (PIP), 3(4):493–495, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Franklin:2002:SDT

[Fra02a] Allan Franklin. Selectivity and discord: two problems of experiment. University of Pittsburgh Press, Pittsburgh, PA, USA, 2002. ISBN 0-8229-4191-0. 290 pp. LCCN Q175 .F795

2002. URL http://www.loc.gov/catdir/enhancements/
fy0708/2002011054-d.html.

Franklin:2002:WWA

[Fra02b] Allan Franklin. William Wilson and the absorption of beta rays. Physics in Perspective (PIP), 4(1):40–77, February 2002. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
002-8358-4.

Franklin:2004:BRS

[Fra04] Allan Franklin. Book review: Susan Haack, Defending Science — within reason. Amherst, New York: Prometheus Books, 2003, 411 pages, $28.00 (cloth). Physics in Perspective (PIP), 6(4): 484–485, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Frank:2005:ERG

[Fra05] Tibor Frank. Ever ready to go: The multiple exiles of Leo Szilard. Physics in Perspective (PIP), 7(2):204– 252, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://adsabs.harvard.

edu/abs/2005PhP.....7..204F;   http://link.springer.
com/article/10.1007/s00016-004-0236-9;   http://www.
springerlink.com/content/1422-6960/.

Franklin:2007:BRR

[Fra07] Allan Franklin. Book review: Ronald Giere, Scientific Perspectivism. Chicago and London: University of Chicago Press, 2006, 151 pages. $30.00 (cloth). Physics in Perspective (PIP), 9(4): 512–514, December 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Franklin:2008:LPI

[Fra08a] Allan Franklin. Are the laws of physics inevitable? Physics in Perspective (PIP), 10(2):182–211, June 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
006-0309-z.

Fraser:2008:CAA

[Fra08b] Gordon Fraser. Cosmic anger: Abdus Salam — the first Muslim Nobel scientist. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-920846-8. xiii + 305 + 16 pp. LCCN QC16.S26 F73 2008. URL http://

www.loc.gov/catdir/enhancements/fy0907/2008000391-b.
html;   http://www.loc.gov/catdir/enhancements/fy0907/
2008000391-d.html;   http://www.loc.gov/catdir/enhancements/
fy0907/2008000391-t.html.

Franklin:2009:BRH

[Fra09a] Allan Franklin. Book review: Helen R. Quinn and Yossi Nir, The Mystery of the Missing Antimatter. Princeton: Princeton University Press, 2008, xii + 278 pages. $29.95 (cloth). Physics in Perspective (PIP), 11(2):235–237, June 2009. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Franklin:2009:ERO

[Fra09b] Allan Franklin. Essay review: Objectivity. Physics in Perspective (PIP), 11(3):338–346, September 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-009-0428-4.

Fraser:2009:NPT

[Fra09c] Gordon Fraser, editor. The New Physics for the Twenty-First Century. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-14002-1. viii + 548 pp. LCCN QC7.5 .N49

2009. URL http://catdir.loc.gov/catdir/enhancements/
fy1009/2010287037-b.html;   http://catdir.loc.gov/catdir/
enhancements/fy1009/2010287037-d.html;   http://catdir.
loc.gov/catdir/enhancements/fy1009/2010287037-t.html.

Franklin:2010:BRP

[Fra10] Allan Franklin. Book review: Patricia Fara, Science: A Four Thousand Year History. Oxford: Oxford University Press, 2009, xv + 408 pages. $34.95 (cloth). Physics in Perspective (PIP), 12(3):355–358, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Franklin:2012:BRT

[Fra12a] Allan Franklin. Book review: Toby E. Huff, Intellectual Curiosity and the Scientific Revolution: A Global Perspective. Cambridge: Cambridge University Press, 2011, xiii + 354 pages. $27.99 (paper). Physics in Perspective (PIP), 14(1):120–122, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Franklin:2012:MPH

[Fra12b] Allan Franklin. Michael Polanyi and his generation. Physics in Perspective (PIP), 14(4):514–517, December 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Franklin:2016:PTD

[Fra16] Allan Franklin. Physics textbooks don’t always tell the truth. Physics in Perspective (PIP), 18(1):3–57, April 2016. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
016-0178-z.

Franklin:2017:SBO

[Fra17] Allan David Franklin. Is seeing believing?: Observation in physics. Physics in Perspective (PIP), 19(4):321–423, December 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

French:1999:SCE

[Fre99] A. P. French. The strange case of Emil Rupp. Physics in Perspective (PIP), 1(1):3–21, March 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050002.

French:2000:BRL

[Fre00] A. P. French. Book review: Louis Brown, A Radar History of World War II: Technical and Military Imperatives. Bristol and Philadelphia: Institute of Physics Publishing, 1999, 563 pages. $38.00 (cloth). Physics in Perspective (PIP), 2(4):450– 453, December 2000.

French:2004:BRP

[Fre04] A. P. French. Book review: Peter Hore, ed., Patrick Blackett: Sailor, Scientist and Socialist. London: Frank Cass, 2003, xiii + 330 pages. $59.50 (cloth). Physics in Perspective (PIP), 6(2): 244–246, June 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

French:2005:BRO

[Fre05a] A. P. French. Book review: Olivier Darrigol, Electrodynamics from Ampère to Einstein. Oxford: Oxford University Press, 2003, xix + 532 pages. $75.00 (paper). Physics in Perspective (PIP), 7(3):382–383, September 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Frercks:2005:FRP

[Fre05b] Jan Frercks. Fizeau’s research program on ether drag: a long quest for a publishable experiment. Physics in Perspective (PIP), 7(1):35–65, March 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-004-0224-0.

French:2008:MPM

[Fre08] A. P. French. In memoriam: Philip Morrison. Physics in Perspective (PIP), 10(1):110–122, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-007-0343-5.

Friedlander:1989:CR

[Fri89] Michael W. Friedlander. Cosmic rays. Harvard University Press, Cambridge, MA, USA, 1989. ISBN 0-674-17458-5. 160 pp. LCCN QC485 .F75 1989.

Friedrich:1999:KXR

[Fri99] Bretislav Friedrich. The KLMN of X-ray spectroscopy: Dolejsek’s discovery of the N series. Physics in Perspective (PIP), 1(4):384–389, December 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050029.

Friedlander:2000:TCR

[Fri00] Michael W. Friedlander. A thin cosmic rain: particles from outer space. Harvard University Press, Cambridge, MA, USA, 2000. ISBN 0-674-00288-1 (hardcover). 241 pp. LCCN QC485 .F75 2000.

Friedman:2001:PEB

[Fri01] Robert Marc Friedman. The politics of excellence: behind the Nobel Prize in science. Times Books, New York, NY, USA, 2001. ISBN 0-7167-3103-7. xv + 379 pp. LCCN QC49 .F75 2001.

Friedlander:2008:BRA

[Fri08] Michael W. Friedlander. Book review: Arthur I. Miller, Empire of the Stars: Friendship, Obsession and Betrayal in the Quest for Black Holes. New York, Houghton Mifflin, 2005, xiv + 400 pages. $26.00 (cloth). Physics in Perspective (PIP), 10(1):132– 134, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Friedlander:2009:BRK

[Fri09] Michael W. Friedlander. Book review: Kelly Moore, Disrupting Science. Princeton: Princeton University Press, 2008, x + 311 pages. $35.00 (cloth). Physics in Perspective (PIP), 11(4): 465–466, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Friedlander:2011:BRJ

[Fri11] Michael W. Friedlander. Book review: John L. Heilbron, Galileo. New York: Oxford University Press, 2010, xiv + 508 pages, $34.95 (cloth). Physics in Perspective (PIP), 13(4):493–495, December 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Fuller:2000:TKP

[Ful00] Steve Fuller. Thomas Kuhn: a philosophical history of our time. University of Chicago Press, Chicago, IL, USA and London, UK, 2000. ISBN 0-226-26894-2 (hardcover). xvii + 472 pp. LCCN ????

Gablot:2000:PWA

[Gab00] Ginette Gablot. A Parisian walk along the landmarks of the discovery of radioactivity. Physics in Perspective (PIP), 2(1): 100–107, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s000160050038.

Gambassi:2003:EFP

[Gam03] Andrea Gambassi. Enrico Fermi in Pisa. Physics in Perspective (PIP), 5(4):384–397, December 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-003-0171-1;
 http://www.springerlink.com/content/542n12h2hmxk4qdp/

.

Gauvin:2018:PQT

[Gau18] Jean-François Gauvin. Playing with quantum toys: Julian Schwinger’s measurement algebra and the material culture of quantum mechanics pedagogy at Harvard in the 1960s. Physics in Perspective (PIP), 20(1):8–42, March 2018. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Gavroglu:2001:DYC

[Gav01] Kostas Gavroğlu. From defiant youth to conformist adulthood: The sad story of liquid helium. Physics in Perspective (PIP), 3(2):165–188, June 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/PL00000528.

Gerward:2006:BR

[GBS^{+}06] Leif Gerward, Laurie M. Brown, J. B. Shank, John L. Hubisz, Klaus Staubermann, William E. Evenson, David Park, Randall D. Knight, and Ronald E. Mickens. Book reviews.

Physics in Perspective (PIP), 8(4):470–488, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-006-0289-z.

Gearhart:1999:BRC

[Gea99] Clayton A. Gearhart. Book review: Crosbie Smith, The Science of Energy: A Cultural History of Energy Physics in Victorian Britain. Chicago: University of Chicago Press, 1998. xii + 404 pages. $60 (cloth), $25 (paper). Physics in Perspective (PIP), 1(4):458–459, December 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Gearhart:2002:PQH

[Gea02] Clayton A. Gearhart. Planck, the quantum, and the historians. Physics in Perspective (PIP), 4(2):170–215, May 2002. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-

002-8363-7. See also [Bad09c].

Gearhart:2014:FHE

[Gea14] Clayton A. Gearhart. The Franck–Hertz experiments, 1911– 1914 experimentalists in search of a theory. Physics in Perspective (PIP), 16(3):293–343, September 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-014-0139-3.

Good:2009:BR

[GEBF09] Gregory A. Good, Guy Emery, Benjamin Bederson, and Michael W. Friedlander. Book reviews. Physics in Perspective (PIP), 11(4):457–468, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-008-0399-x.

Gerward:1999:BRB

[Ger99a] Leif Gerward. Book review: Bettyann Holtzmann Kevles, Naked to the Bone: Medical Imaging in the Twentieth Century. New Brunswick: Rutgers University Press, 1997, xiv + 378 pp. $35.95 (cloth); Reading: Addison-Wesley/Perseus Books, 1998. $18.00 (paper). Physics in Perspective (PIP), 1(3):337, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Gerward:1999:PVH

[Ger99b] Leif Gerward. Paul Villard and his discovery of gamma rays. Physics in Perspective (PIP), 1(4):367–383, December 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160050028.

Gerjuoy:2003:BRJ

[Ger03] Edward Gerjuoy. Book review: Jagdish Mehra and Kimball A. Milton, Climbing the Mountain: The Scientific Biography of Julian Schwinger. New York: Oxford University Press, 2000, 677 pages. $49.95 (cloth). Physics in Perspective (PIP), 5(1):124– 126, April 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Gerjuoy:2006:BRJ

[Ger06a] Edward Gerjuoy. Book review: Jeremy Bernstein, Oppenheimer: Portrait of an Enigma. Chicago: Ivan R. Dee, 2004, 223 pages. $25.00 (cloth). David C. Cassidy, J. Robert Oppenheimer and the American Century. New York: Pi Press, 2005, 462 pages. $27.95 (cloth). Physics in Perspective (PIP), 8(1):109–112, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Gerward:2006:BRD

[Ger06b] Leif Gerward. Book review: Denis Brian, The Curies: A Biography of the Most Controversial Family in Science. Hoboken, New Jersey, John Wiley & Sons, 2005, ix + 438 pages. $30.00 (cloth). Physics in Perspective (PIP), 8(4):470–473, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Greenberg:2005:CWFa

[GF05a] John Greenberg and William A. Fowler. A conversation with William A. Fowler —- Part I. Physics in Perspective (PIP), 7 (1):66–106, March 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-004-0216-0.

Greenberg:2005:CWFb

[GF05b] John Greenberg and William A. Fowler. A conversation with William A. Fowler — Part II. Physics in Perspective (PIP), 7(2):165–203, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-004-0217-z.

Goodstein:2000:RFH

[GG00] David Goodstein and Judith Goodstein. Richard Feynman and the history of superconductivity. Physics in Perspective (PIP), 2(1):30–47, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s000160050035.

Ghirardi:2005:SLG

[Ghi05] G. C. Ghirardi. Sneaking a Look at God’s Cards: Unraveling the Mysteries of Quantum Mechanics. Princeton University Press, Princeton, NJ, USA, revised edition, 2005. ISBN 0-691-04934-3, 0-691-12139-7, 0-691-13037-X. xix + 488 pp. LCCN Q173 .G4813 2004.

Giberson:2005:BRS

[Gib05] Karl Giberson. Book review: Stephen M. Barr, Modern Physics and Ancient Faith. Notre Dame: University of Notre Dame Press, 2003, xii + 313 pages. $30.00 (cloth). Physics in Perspective (PIP), 7(2):257–259, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Gingras:2010:TP

[Gin10] Yves Gingras. The transformation of physics from 1900 to 1945. Physics in Perspective (PIP), 12(3):248–265, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-010-0017-6.

Giudice:2010:ZOJ

[Giu10] Gian Francesco Giudice. A zeptospace odyssey: a journey into the physics of the LHC. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-958191-6 (hardcover). 276 pp. LCCN QC787.P73 G58 2010.

Giudice:2012:BSL

[Giu12] Gian Francesco Giudice. Big Science and the Large Hadron Collider. Physics in Perspective (PIP), 14(1):95–112, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-011-0078-1.

Guerra:2006:EFD

[GLR06] Francesco Guerra, Matteo Leone, and Nadia Robotti. Enrico Fermi’s discovery of neutron-induced artificial radioactivity: Neutrons and neutron sources. Physics in Perspective (PIP), 8(3):255–281, September 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-006-0296-0;
 http://www.springerlink.com/content/w233203272t22206/

.

Guerra:2012:DAR

[GLR12] Francesco Guerra, Matteo Leone, and Nadia Robotti. The discovery of artificial radioactivity. Physics in Perspective (PIP), 14(1):33–58, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-011-0064-7;   http://www.
springerlink.com/content/v1733673628jr058/.

Guerra:2020:EFD

[GLR20] Francesco Guerra, Matteo Leone, and Nadia Robotti. Enrico Fermi’s discovery of neutron-induced artificial radioactivity: A case of “emanation” from “Divine Providence”. Physics in Perspective (PIP), 22(3):129–161, September 2020. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Gould:2003:SLT

[GMS^{+}03] Stephen Jay Gould, Stephanie Marshall, Judith A. Scheppler, Michael J. Palmisano, et al., editors. Science Literacy for the Twenty-First Century. Prometheus Books, Amherst, NY, USA, 2003. ISBN 1-59102-020-4. 321 pp. LCCN Q183.3.A1

S3565 2002. URL http://www.loc.gov/catdir/toc/fy035/
2002036718.html.

Goenner:2003:AEF

[Goe03] Hubert Goenner. Albert Einstein and Friedrich Dessauer: political views and political practice. Physics in Perspective (PIP), 5(1):21–66, April 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160300002;   http:/
/www.springerlink.com/content/jkg195fabv2pjdcq/fulltext.
pdf.

Goldberger:1999:BRD

[Gol99a] Marvin L. Goldberger. Book review: Dan Cooper, Enrico Fermi and the Revolution in Modern Physics. New York: Oxford University Press, 1999, 117 pages. $21.00 (cloth). Physics in Perspective (PIP), 1(2):226–227, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Goldberger:1999:AEF

[Gol99b] Marvin L. Goldberger. In appreciation: Enrico Fermi (1901– 1954): The complete physicist. Physics in Perspective (PIP), 1(3):328–336, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050024;   http:/
/www.springerlink.com/content/y1h02gr3rv7wurrl/.

Gold:2010:TEV

[Gol10] Barri J. Gold. ThermoPoetics: energy in Victorian literature and science. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01372-X (hardcover). x + 343 pp. LCCN PR468.S34 G65 2010.

Gold:2012:TEV

[Gol12] Barri J. Gold. ThermoPoetics: energy in Victorian literature and science. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-51731-0 (paperback). ???? pp. LCCN ????

Goldhaber:2016:GSG

[Gol16] Michael H. Goldhaber. Gertrude Scharff-Goldhaber, 1911–1998: Nuclear physicist against the odds. Physics in Perspective (PIP), 18(2):182–208, August 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-016-0181-4.

Goodstein:1999:CHB

[Goo99] Judith Goodstein. A conversation with Hans Bethe. Physics in Perspective (PIP), 1(3):253–281, October 1999. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160050022;   http://www.springerlink.com/content/
k4f0pllveakla47q/.

Goodstein:2001:CFR

[Goo01] Judith R. Goodstein. A conversation with Franco Rasetti. Physics in Perspective (PIP), 3(3):271–313, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/

PL00000534. This interview was conducted 4 February 1982 but was unpublished for 19 years.

Goodstein:2002:BRB

[Goo02] David Goodstein. Book review: Einstein, Picasso. Space, time, and the beauty that causes havoc. Physics in Perspective (PIP), 4(2):247–248, May 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Goodstein:2003:RSN

[Goo03a] D. Goodstein. The road to Stockholm: Nobel prizes, science and scientists. Physics in Perspective (PIP), 5(4):473, December 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Goodstein:2003:CLAa

[Goo03b] Judith R. Goodstein. A conversation with Lee Alvin DuBridge — Part I. Physics in Perspective (PIP), 5(2):174–205, May 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-003-0144-4.

Goodstein:2003:CLAb

[Goo03c] Judith R. Goodstein. A conversation with Lee Alvin DuBridge — Part II. Physics in Perspective (PIP), 5(3):281–309, September 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-003-0145-3.

Goodstein:2004:CFP

[Goo04] Judith R. Goodstein. A conversation with Frank Press. Physics in Perspective (PIP), 6(2):184–196, June 2004. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
003-0204-9.

Goodstein:2008:BRH

[Goo08] David Goodstein. Book review: Helge S. Kragh, Conceptions of Cosmos — From Myths to the Accelerating Universe: History of Cosmology. Oxford: Oxford University Press, 2007, 276 pages. $70.00 (cloth). Physics in Perspective (PIP), 10(2):254, June 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Good:2009:BRS

[Goo09a] Gregory A. Good. Book review: Silvan S. Schweber, Einstein and Oppenheimer: The Meaning of Genius. Cambridge, Mass. and London: Harvard University Press, 2008, ix + 412 pages. $29.95 (cloth). Physics in Perspective (PIP), 11(4):467–468, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Goodstein:2009:BRA

[Goo09b] David Goodstein. Book review: Alan Sokal, Behind the Hoax: Science, Philosophy, and Culture. Oxford: Oxford University Press, 2008, xxi + 465 pages. $34.95 (cloth). Physics in Perspective (PIP), 11(2):237–238, June 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Goodstein:2010:FFC

[Goo10] David L. Goodstein. On fact and fraud: cautionary tales from the front lines of science. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-13966-0 (hardcover), 1-4008-3457-0 (e-book). xiv + 168 pp. LCCN Q175.37 .G66 2010.

Good:2012:BRM

[Goo12] Gregory A. Good. Book review: Margaret J. Osler, Reconfiguring the World: Nature, God, and Human Understanding from the Middle Ages to Early Modern Europe. Baltimore: Johns Hopkins University Press, 2010, 184 pages. $25.00 (paper). Physics in Perspective (PIP), 14(2):252–254, June 2012. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Gorelik:2009:PHB

[Gor09] Gennady Gorelik. The paternity of the H-bombs: Soviet-American perspectives. Physics in Perspective (PIP), 11(2): 169–197, June 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-007-0377-8.

Grasso:2009:NSH

[GORS09] Giacomo Grasso, Carlo Oppici, Federico Rocchi, and Marco Sumini. A neutronics study of the 1945 Haigerloch B-VIII nuclear reactor. Physics in Perspective (PIP), 11(3):318– 335, September 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-008-0396-0.

Gould:2001:BRB

[Gou01] Harvey Gould. Book review: I. Bernard Cohen, Howard Aiken: Portrait of a Computer Pioneer. Cambridge: MIT Press, 1999, xi + 329 pp. $34.95 (Hardbound). Physics in Perspective (PIP), 3(1):128–130, March 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Golub:2001:NSS

[GP01] L. (Leon) Golub and Jay M. Pasachoff. Nearest Star: the Surprising Science of our Sun. Harvard University Press, Cambridge, MA, USA, 2001. ISBN 0-674-00467-1 (hardcover). xii + 267 pp. LCCN QB521 .G65 2001.

Guerra:2008:EMA

[GR08a] Francesco Guerra and Nadia Robotti. Ettore Majorana: aspects of his scientific and academic activity, volume 6 of CRM Series. Edizioni della Normale, Pisa, Italia, 2008. ISBN 88-7642-331-1 (paperback), 88-7642-331-4 (paperback). ISSN 2239-4524. xii + 243 pp. LCCN ????

Guerra:2008:EMF

[GR08b] Francesco Guerra and Nadia Robotti. Ettore Majorana’s forgotten publication on the Thomas–Fermi model. Physics in Perspective (PIP), 10(1):56–76, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-007-0340-8;
 http://www.springerlink.com/content/53l26g8353l2421g/

.

Guerra:2009:EFD

[GR09] Francesco Guerra and Nadia Robotti. Enrico Fermi’s discovery of neutron-induced artificial radioactivity: The influence of his theory of beta decay. Physics in Perspective (PIP), 11(4): 379–404, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-008-0415-1;   http://www.
springerlink.com/content/65w213132524072n/.

Guerra:2013:DDE

[GR13] Francesco Guerra and Nadia Robotti. The disappearance and death of Ettore Majorana. Physics in Perspective (PIP), 15 (2):160–177, June 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-013-0111-7.

Graney:2008:SIM

[Gra08] Christopher M. Graney. But still, it moves: Tides, stellar parallax, and Galileo’s commitment to the Copernican theory. Physics in Perspective (PIP), 10(3):258–268, September 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-007-0345-3.

Graney:2010:STR

[Gra10] Christopher M. Graney. Seeds of a Tychonic revolution: Telescopic observations of the stars by Galileo Galilei and Simon Marius. Physics in Perspective (PIP), 12(1):4–24, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-009-0002-0.

Graney:2011:CGR

[Gra11] Christopher M. Graney. Contra Galileo: Riccioli’s “Coriolisforce” argument on the Earth’s diurnal rotation. Physics in Perspective (PIP), 13(4):387–400, December 2011. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
011-0058-5.

Greenslade:2000:PTS

[Gre00] Thomas B. Greenslade, Jr. The physical tourist: Scientific travels in the Irish countryside. Physics in Perspective (PIP), 2(3): 313–326, September 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s000160050048.

Greenberg:2001:SMP

[Gre01] Daniel S. Greenberg. Science, Money, and Politics: Political Triumph and Ethical Erosion. University of Chicago Press, Chicago, IL, USA and London, UK, 2001. ISBN 0-226-30634-8 (hardcover). x + 530 pp. LCCN Q180.55.G6 G74 2001.

URL http://www.loc.gov/catdir/bios/uchi051/00013226.
html;   http://www.loc.gov/catdir/description/uchi051/
00013226.html;   http://www.loc.gov/catdir/toc/uchi051/
00013226.html.

Greenberger:2005:BRJ

[Gre05] D. M. Greenberger. Book review: Jim Baggott, Beyond Measure: Modern Physics, Philosophy and the Meaning of Quantum Theory. Oxford: Oxford University Press, 2004, 379 pp. $74.50 (cloth), $34.50 (paper). Physics in Perspective (PIP), 7 (3):384–386, September 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Greenslade:2006:BRR

[Gre06] T. B. Greenslade, Jr. Book review: Robert Fox and Graeme Gooday, ed., Physics in Oxford 1839–1939: Laboratories, Learning, and College Life. Oxford: Oxford University Press, 2005, xxi + 363 pages. $134.50/£60.00 (cloth). Physics in Perspective (PIP), 8(2):229–231, May 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Greenler:2007:BRP

[Gre07a] Robert Greenler. Book review: Peter Pesic, Sky in a Bottle. Cambridge, Mass. and London: The MIT Press, 2005, 262 pages. $24.95 (cloth). Physics in Perspective (PIP), 9(1):122– 123, January 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Greenslade:2007:BRD

[Gre07b] Thomas B. Greenslade, Jr. Book review: Domenico Bertoloni Meli, Thinking with Objects: The Transformation of Mechanics in the Seventeenth Century. Baltimore: Johns Hopkins University Press, 2006, xii + 389 pages. $29.95 (paper). Physics in Perspective (PIP), 9(4):510–512, December 2007. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Greenslade:2009:BRL

[Gre09] Thomas B. Greenslade, Jr. Book review: Lissa Roberts, Simon Schaffer, and Peter Dear, ed., The Mindful Hand; Inquiry and Invention from the Late Renaissance to Early Industrialisation. Amsterdam: Koninkliijke Nederlandse Akademie van Wetenschappen, 2007, xxvii + 503 pages. $110.00 (cloth). Physics in Perspective (PIP), 11(1):111–112, March 2009. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Greenlee:2012:CNN

[Gre12] Thomas R. Greenlee. Cosmic numbers: The numbers that define our universe. Physics in Perspective (PIP), 14(4):517–518, December 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Greenslade:2020:PNP

[Gre20] Thomas B. Greenslade, Jr. Professors of natural philosophy in the nineteenth century. Physics in Perspective (PIP), 22(4): 226–238, December 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Greenslade:2021:ANC

[Gre21] Thomas B. Greenslade, Jr. American Nineteenth-Century manufacturers and importers of philosophical apparatus. Physics in Perspective (PIP), 23(4):202–230, December 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Griffiths:2006:BRA

[Gri06] David J. Griffiths. Book review: Alberto P. Guimarães, From Lodestone to Supermagnets: Understanding Magnetic Phenomena. Weinheim: Wiley-VCH, 2005, xii + 236 pages. $27.95 (cloth). Physics in Perspective (PIP), 8(3):353–354, September 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Guimaraes:2005:LSU

[Gui05] Alberto Passos Guimaräes. From lodestone to supermagnets: understanding magnetic phenomena. Wiley-VCH, Berlin, Germany; Weinheim, Germany; New York, NY, USA, 2005. ISBN 3-527-40557-7. xi + 236 pp. LCCN QC753.2 .G85

2005. URL http://catdir.loc.gov/catdir/enhancements/
fy0653/2006280145-b.html;   http://catdir.loc.gov/catdir/

enhancements/fy0653/2006280145-d.html;   http://catdir.
loc.gov/catdir/enhancements/fy0653/2006280145-t.html.

Goodstein:2003:BR

[GWL03] David Goodstein, William A. Wallace, and Peter Lindenfeld. Book reviews. Physics in Perspective (PIP), 5(4):473–476, December 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-003-0180-0.

Haack:2003:DSW

[Haa03] Susan Haack. Defending science — within reason: between scientism and cynicism. Prometheus Books, Amherst, NY, USA, 2003. ISBN 1-59102-117-0. 411 pp. LCCN Q172.5.C74

H33 2003. URL http://www.loc.gov/catdir/toc/ecip044/
2003012541.html.

Haack:2007:DSW

[Haa07] Susan Haack. Defending science — within reason. Prometheus Books, Amherst, NY, USA, 2007. ISBN 1-59102-458-7 (softcover). 411 pp. LCCN Q172.5.C74 H33 2007. URL http://

www.loc.gov/catdir/toc/ecip0620/2006027746.html.

Habfast:2010:DGJ

[Hab10] Claus Habfast. The DESY Golden Jubilee in Hamburg: Lessons from the past. Physics in Perspective (PIP), 12(2):219–230, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-010-0020-y.

Hagar:2012:PP

[Hag12] Amit Hagar. On physics and philosophy. Physics in Perspective (PIP), 14(4):512–514, December 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Halpern:2004:NER

[Hal04] Paul Halpern. Nordström, Ehrenfest, and the role of dimensionality in physics. Physics in Perspective (PIP), 6(4):390–400, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-004-0221-3.

Halpern:2005:PBE

[Hal05] Paul Halpern. Peter Bergmann: The education of a physicist. Physics in Perspective (PIP), 7(4):390–403, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-005-0237-3.

Halpern:2007:KEF

[Hal07] Paul Halpern. Klein, Einstein, and five-dimensional unification. Physics in Perspective (PIP), 9(4):390–405, December 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-006-0319-x;   http://www.springerlink.com/content/
j97p35877401w357/fulltext.pdf.

Hall:2009:BRL

[Hal09a] Donald E. Hall. Book review: Leo Beranek, Riding the Waves: A Life in Sound, Science, and Industry. Cambridge: MIT Press, 2008, x + 235 pages. $24.95 (cloth). Physics in Perspective (PIP), 11(2):232–235, June 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Halpern:2009:PLL

[Hal09b] Paul Halpern. Philadelphia: Life, liberty, and the pursuit of physics. Physics in Perspective (PIP), 11(2):209–227, June 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-008-0386-2.

Halpern:2010:WDC

[Hal10] Paul Halpern. Washington: a DC circuit tour. Physics in Perspective (PIP), 12(4):443–466, December 2010. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
010-0029-2.

Halpern:2012:QHP

[Hal12] Paul Halpern. Quantum humor: The playful side of physics at Bohr’s Institute for Theoretical Physics. Physics in Perspective (PIP), 14(3):279–299, September 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-011-0071-8.

Halpern:2018:CPH

[Hal18] Paul Halpern. Celebrity physicist: How the press sensationalized Einstein’s search for a unified field theory. Physics in Perspective (PIP), 20(3):254–271, September 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hamblin:2000:SIA

[Ham00] Jacob Darwin Hamblin. Science in isolation: American marine geophysics research, 1950–1968. Physics in Perspective (PIP), 2(3):293–312, September 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050047.

Hamblin:2011:BRB

[Ham11] Jacob Darwin Hamblin. Book review: Bruce J. Hunt, Pursuing Power and Light: Technology and Physics from James Watt to Albert Einstein. Baltimore: Johns Hopkins University Press, 2010, 182 pages. $20.00 (paper). Physics in Perspective (PIP), 13(1):117–118, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hankins:2004:BRR

[Han04] Thomas L. Hankins. Book review: Roy Porter, ed., Eighteenth-Century Science. Vol. 4. The Cambridge History of Science. Cambridge University Press, 2003, xxx + 912 pages. $95.00 (cloth). Physics in Perspective (PIP), 6(3):364–365, September 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Harman:1985:WPS

[Har85] P. M. (Peter Michael) Harman, editor. Wranglers and physicists: studies on Cambridge physics in the nineteenth century. Manchester University Press, Manchester, UK, 1985. ISBN 0-7190-1756-4, 0-226-12017-1 (hardcover), 0-226-12019-8 (paperback). LCCN QC9.G7 W73 1985. US$27.00.

Harper:2001:AGG

[Har01] Eamon Harper. In appreciation: George Gamow: Scientific amateur and polymath. Physics in Perspective (PIP), 3(3):335–372, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://adsabs.harvard.edu/
abs/2001PhP.....3..335H;   http://link.springer.com/

article/10.1007/PL00000536;   http://www.springerlink.
com/content/m0htmt9ww9b9jjv1/;   http://www.springerlink.
com/openurl.asp?genre=journal&issn=1422-6944.

Hargittai:2002:RSN

[Har02a] István Hargittai. The road to Stockholm: Nobel Prizes, science, and scientists. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-850912-X. xvii + 342 + 24 pp.

LCCN Q141 .H267 2002. URL http://www.loc.gov/catdir/
enhancements/fy0614/2002283888-d.html;   http://www.
loc.gov/catdir/enhancements/fy0724/2002283888-b.html;
 http://www.loc.gov/catdir/toc/fy033/2002283888.html.

Harkness:2002:ALC

[Har02b] Jon M. Harkness. In appreciation: a lifetime of connections: Otto Herbert Schmitt, 1913–1998. Physics in Perspective (PIP), 4(4):456–490, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160200005.

Harrison:2005:BRA

[Har05] M. J. Harrison. Book review: Andrew Warwick, Masters of Theory: Cambridge and the Rise of Mathematical Physics. Chicago: University of Chicago Press, 2003, xiv + 572 pages. $85.00 (cloth), $29.00 (paper). Physics in Perspective (PIP), 7(1):130– 132, March 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hargittai:2006:MSF

[Har06] István Hargittai. The Martians of Science: five physicists who changed the twentieth century. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-517845-9 (hardcover). xxiv + 313 + 32 pp.

LCCN QC15 .H27 2006. URL http://www.loc.gov/catdir/
enhancements/fy0636/2005029427-d.html;   http://www.
loc.gov/catdir/enhancements/fy0724/2005029427-b.html;
 http://www.loc.gov/catdir/toc/ecip061/2005029427.html.

Harrison:2009:FEB

[Har09] Walter A. Harrison. Finding the energy bands of silicon. Physics in Perspective (PIP), 11(2):198–208, June 2009. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
008-0397-z.

Haussecker:2011:IAS

[HC11] Enzo F. Haussecker and Alexander W. Chao. The influence of accelerator science on physics research. Physics in Perspective (PIP), 13(2):146–160, June 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-010-0049-y.

Hoskin:2012:CHW

[HDS12] Michael A. Hoskin, David Dewirst, and Wolfgang Steinicke. The construction of the heavens: William Herschel’s cosmology. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-01838-2 (hardback). viii + 205 pp. LCCN QB36.H6 H719

2012. URL http://assets.cambridge.org/97811070/18389/
cover/9781107018389.jpg.

Hecht:2004:CLS

[Hec04] Jeff Hecht. City of Light: the Story of Fiber Optics. The Sloan technology series. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, revised and expanded edition, 2004. ISBN 0-19-516255-2. xii + 340 pp. LCCN TA1800 .H42 2004.

Hecht:2007:BRS

[Hec07] Jeff Hecht. Book review: Sean F. Johnston, Holographic Visions: A History of New Science. Oxford: Oxford University Press, 2006, 568 pages. $124.50 (cloth). Physics in Perspective (PIP), 9(3):383–385, September 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hecht:2010:BRM

[Hec10] Jeff Hecht. Beam: The Race to Make the Laser. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-514210-1 (hardcover), 0-19-973871-8 (paperback). x + 274 pp. LCCN TA1677 .H42 2010.

Heering:2006:RTR

[Hee06] Peter Heering. Regular twists: Replicating Coulomb’s wiretorsion experiments. Physics in Perspective (PIP), 8(1):52–63, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-005-0262-2.

Heilbron:2010:G

[Hei10a] J. L. Heilbron. Galileo. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-958352-8, 0-19-161295-2 (e-book). xiv + 508 + 16 pp. LCCN QB36.G2 H445 2010.

Hein:2010:OH

[Hei10b] Piet Hein. Occupational hazard. Physics in Perspective (PIP), 12(3):368, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-010-0033-6.pdf.

Hein:2010:WR

[Hei10c] Piet Hein. Wide road. Physics in Perspective (PIP), 12(2):244, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00016-010-0026-5.pdf.

Hein:2011:WI

[Hei11] Piet Hein. We have to have it. Physics in Perspective (PIP), 13(4):498, December 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-011-0076-3.pdf.

Hellyer:2005:CPJ

[Hel05] Marcus Hellyer. Catholic physics: Jesuit natural philosophy in early modern Germany. University of Notre Dame Press, Notre Dame, IN, USA, 2005. ISBN 0-268-03071-5 (hardcover). xii + 336 pp. LCCN QC47.G3 H45 2004. URL http://www.loc.

gov/catdir/toc/ecip053/2004024688.html.

Hentschel:1999:CVR

[Hen99a] Klaus Hentschel. The culture of visual representations in spectroscopic education and laboratory instruction. Physics in Perspective (PIP), 1(3):282–327, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s000160050023.

Hentschel:1999:SHP

[Hen99b] Klaus Hentschel. Some historical points of interest in Göttingen. Physics in Perspective (PIP), 1(1):110–117, March 1999. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160050009.

Hentschel:2005:PHB

[Hen05a] Ann M. Hentschel. Peripatetic highlights in Bern. Physics in Perspective (PIP), 7(1):107–129, March 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
004-0232-0.

Hentschel:2005:GUH

[Hen05b] Klaus Hentschel. Gaußens unsichtbare Hand: der Universitäts-Mechanicus und Maschinen-Inspector Moritz Meyerstein: ein Instrumentenbauer im 19. Jahrhundert. (German) [Gauss’s invisible hand: the university mechanics and machine inspector Moritz Meyerstein: an instrument maker in the 19th Century], volume 52 (folge 3) of Abhandlungen der Akademie der Wissenschaften zu Göttingen, Mathematisch-Physikalische Klasse. Vandenhoeck and Ruprecht, Göttingen, Germany, 2005. ISBN 3-525-82126-3. 321 pp. LCCN T40.M49. EUR 79.00. URL http:

//www.gbv.de/dms/hebis-darmstadt/toc/127677550.pdf.

Henry:2007:PTP

[Hen07a] John Henry. The physical tourist: Physics in Edinburgh: From Napier’s bones to Higgs’s boson. Physics in Perspective (PIP), 9(4):468–501, December 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-007-0348-0.

Hentschel:2007:MAM

[Hen07b] Klaus Hentschel. The mental aftermath: the mentality of German physicists 1945–1949. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-920566-3 (hardcover). 205 pp. LCCN QC9.G3 H46 2007. URL http://www.

loc.gov/catdir/toc/fy0805/2007280005.html.

Herken:2000:CCP

[Her00] Gregg Herken. Cardinal Choices: Presidential Science Advising from the Atomic Bomb to SDI. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, revised and expanded edition, 2000. ISBN 0-8047-3966-8, 0-8047-3770-3 (paperback). xv + 358 pp. LCCN Q127.U6 H394

2000. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/description/
cam021/00026546.html;   http://www.loc.gov/catdir/toc/
cam027/00026546.html.

Hertz:2003:PMP

[Her03] Heinrich Hertz. The principles of mechanics presented in a new form. Dover phoenix editions. Dover Publications, Inc., New York, NY, USA, 2003. ISBN 0-486-49557-4. 271 pp. LCCN

QA805 .H5713 2003. URL http://www.loc.gov/catdir/
enhancements/fy0615/2003062517-d.html;   http://www.
openisbn.com/preview/1602062943/. Preface by H. von

Helmholtz. Authorized English translation by D. E. Jones and J. T. Walley. Introduction by Robert S. Cohen.

Hey:1999:FCE

[Hey99] Anthony J. G. Hey, editor. Feynman and Computation: Exploring the Limits of Computers. Perseus Publishers, Cambridge, MA, USA, 1999. ISBN 0-7382-0057-3. xxiii + 438 pp. LCCN QA76 .F46 1999, QC52 .F49 199. US$50.00.

Heisenberg:2003:LEB

[HHH03] Werner Heisenberg and Anna M. Hirsch-Heisenberg. Liebe Eltern! Briefe aus kritischer Zeit; 1918 bis 1945. (German) [Dear parents! Letters from a critical period, 1918 to 1945]. Langen Müller, Munich, Germany, 2003. ISBN 3-7844-2900-9. 400 pp. LCCN ????

Heinze:2017:TST

[HHH17] Thomas Heinze, Olof Hallonsten, and Steffi Heinecke. Turning the ship: The transformation of DESY, 1993–2009. Physics in Perspective (PIP), 19(4):424–451, December 2017. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00016-017-0209-4.pdf.

Hiebert:1999:BRR

[Hie99] Erwin N. Hiebert. Book review: Robert D. Purrington, Physics in the Nineteenth Century. New Brunswick, NJ: Rutgers University Press, 1997, xx, 251 pages. $55.00 (cloth), $25.00 (paper). Physics in Perspective (PIP), 1(2):225–226, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hiebert:2000:CFE

[Hie00] Erwin N. Hiebert. Common frontiers of the exact sciences and the humanities. Physics in Perspective (PIP), 2(1):6–29, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s000160050034.

Hilborn:2007:BRR

[Hil07] Robert C. Hilborn. Book review: Robert B. Laughlin, A Different Universe: Reinventing Physics from the Bottom Down. New York: Basic Books, 2006, xviii + 254 pages. $16.00 (paper). Physics in Perspective (PIP), 9(1):118–120, January 2007. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hilborn:2011:BRR

[Hil11] Robert C. Hilborn. Book review: Richard A. Muller, Physics and Technology for Future Presidents: An Introduction to the Essential Physics Every World Leader Needs to Know. Princeton: Princeton University Press, 2010, xi + 517 pages. $49.50 (cloth). Physics in Perspective (PIP), 13(2):247–249, June 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hintz:2012:MLN

[Hin12] Norton M. Hintz. My life in nuclear physics, photography, and opera. Physics in Perspective (PIP), 14(2):196–238, June 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-012-0086-9.

Hirshfeld:2009:EML

[Hir10] Alan Hirshfeld. Eureka man: the life and legacy of Archimedes. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, pbk. edition, 2009–2010. ISBN 0-8027-1618-0 (hardcover), 0-8027-7766-X (paperback), 0-8027-1979-1 (e-book). viii + 242 + 8 (hardcover), vi + 247 + 8 (paperback) pp. LCCN Q143.A62 H57 2010.

Hoddeson:2003:VRR

[HK03] Lilian Hoddeson and Adrienne Kolb. Vision to reality: from Robert R. Wilson’s frontier to Leon M. Lederman’s Fermilab. Physics in Perspective (PIP), 5(1):67–86, April 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160300003.

Huijnen:2007:PER

[HK07] Pim Huijnen and A. J. Kox. Paul Ehrenfest’s rough road to Leiden: a physicist’s search for a position, 1904–1912. Physics in Perspective (PIP), 9(2):186–211, June 2007. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00016-006-0287-1.pdf.

Hirsch:2002:NDC

[HKT02] E. D. (Eric Donald) Hirsch, Jr., Joseph F. Kett, and James S. Trefil. The New Dictionary of Cultural Literacy. Houghton Mifflin, Boston, MA, USA, third edition, 2002. ISBN 0-618-22647-8. xix + 647 pp. LCCN E169.1 .H614 2002. URL http:

//www.bartelby.com/59/;   http://www.loc.gov/catdir/
description/hm022/2002027609.html;   http://www.loc.
gov/catdir/enhancements/fy0736/2002027609-b.html;   http:
//www.loc.gov/catdir/samples/hm051/2002027609.html;
http://www.loc.gov/catdir/toc/hm021/2002027609.html.

Hunt:2001:BR

[HMHB01] J. Christopher Hunt, Eugen Merzbacher, Charles H. Holbrow, and Benjamin Bederson. Book reviews. Physics in Perspective (PIP), 3(2):249–254, June 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/PL00000532.

Hobson:2002:BRG

[Hob02] Art Hobson. Book review: George Gamow and Russell Stannard, The New World of Mr Tompkins. Cambridge, United Kingdom: Cambridge University Press, 1999, ix + 258 pages, $24.95 (cloth), $16.95 (paper). Physics in Perspective (PIP), 4 (4):494–495, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hobson:2004:BRS

[Hob04] Art Hobson. Book review: Stephanie Pace Marshall, Judith A. Scheppler, and Michael J. Palmisano, ed., Science Literacy for the Twenty-First Century. Amherst and New York: Prometheus Books, 2003, 321 pages. $29.00 (cloth). Physics in Perspective (PIP), 6(3):365–367, September 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hoffmann:1999:PBW

[Hof99] Dieter Hoffmann. Physics in Berlin: a walk through the historical city center. Physics in Perspective (PIP), 1(4):445–454, December 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s000160050031.

Hoffmann:2000:PTP

[Hof00] Dieter Hoffmann. The physical tourist: Physics in Berlin: Walking tours in Charlottenburg and Dahlem and excursions in the vicinity of Berlin. Physics in Perspective (PIP), 2(4): 426–445, December 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s000160050054.

Hoffleit:2002:PWS

[Hof02] E. Dorrit Hoffleit. Pioneering women in the spectral classification of stars. Physics in Perspective (PIP), 4(4):370–398, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s000160200001.

Hoffmann:2005:BAA

[Hof05] Dieter Hoffmann. Between autonomy and accommodation: The German Physical Society during the Third Reich. Physics in Perspective (PIP), 7(3):293–329, September 2005. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
004-0235-x.

Hoffmann:2009:FLK

[Hof09] Dieter Hoffmann. Fritz Lange, Klaus Fuchs, and the remigration of scientists to East Germany. Physics in Perspective (PIP), 11(4):405–425, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-009-0427-5.

Hofmann:2017:ERF

[Hof17] James R. Hofmann. Essay review: Friedrich Steinle, Exploratory Experiments: Ampère, Faraday, and the Origins of Electrodynamics, translated by Alex Levine, Pittsburgh: University of Pittsburgh Press, 2016. Physics in Perspective (PIP), 19(3):

307–318, September 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). See reply [Ste17].

Holton:1999:RMS

[Hol99] Gerald Holton. R. A. Millikan’s struggle with the meaning of Planck’s constant. Physics in Perspective (PIP), 1(3):231–237, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s000160050020.

Holbrow:2001:BRA

[Hol01] Charles H. Holbrow. Book review: Alan Isaacs, ed., A Dictionary of Physics, fourth edition. Oxford: Oxford University Press, 2000, vi + 546 pages. $14.95 (paper). Physics in Perspective (PIP), 3(2):251–253, June 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Holbrow:2003:ACC

[Hol03] Charles H. Holbrow. In appreciation: Charles C. Lauritsen: a reasonable man in an unreasonable world. Physics in Perspective (PIP), 5(4):419–472, December 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-003-0178-7.

Holmberg:2004:PTP

[Hol04] Peter Holmberg. The physical tourist: Physics in Helsinki. Physics in Perspective (PIP), 6(1):76–113, April 2004. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
003-0194-7.

Holton:2005:VVS

[Hol05] Gerald James Holton. Victory and vexation in science: Einstein, Bohr, Heisenberg, and others. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01519-3. xi + 229 pp. LCCN Q180.A3 H65 2005.

Holton:2007:BRJ

[Hol07] Gerald Holton. Book review: Jeffrey Crelinsten, Einstein’s Jury: The Race to Test Relativity. Princeton: Princeton University Press, 2006, xxix + 397 pages. $35.00 (cloth). Physics in Perspective (PIP), 9(2):257–259, June 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Holmes:2008:AWHa

[Hol08a] Richard Holmes. The age of wonder: how the romantic generation discovered the beauty and terror of science. HarperCollins, London, UK, 2008. ISBN 0-00-714953-0 (paperback). xxi + 554 pp. LCCN Q127.G4 H65 2009.

Holmes:2008:AWHb

[Hol08b] Richard Holmes. The age of wonder: how the romantic generation discovered the beauty and terror of science. Pantheon Books, New York, NY, USA, 2008. ISBN 0-375-42222-6. xxi + 552 + 16 pp. LCCN Q127.G4 H65 2008.

Holmes:2008:AWHc

[Hol08c] Richard Holmes. The age of wonder: how the romantic generation discovered the beauty and terror of science. Vintage Books, New York, NY, USA, 2008. ISBN 1-4000-3187-7 (paperback), 0-307-37832-2 (e-book). xxi + 552 + 24 pp. LCCN Q127.G4 H65 2010.

Holbrow:2009:BRD

[Hol09] Charles H. Holbrow. Book review: David E. Rowe and Robert Schulmann, ed., Einstein on Politics: His Private Thoughts and Public Stands on Nationalism, Zionism, War, Peace, and the Bomb. Princeton: Princeton University Press, 2007, xxxvi + 524 pages. $29.95 (cloth). Physics in Perspective (PIP), 11(1):112– 115, March 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Holbrow:2011:DCC

[Hol11] Charles H. Holbrow. Dick Crane’s California days. Physics in Perspective (PIP), 13(1):36–57, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-010-0041-6.

Holton:2019:BRA

[Hol19] Gerald Holton. Book review: Allen Esterson and David C. Cassidy, Einstein;s Wife: The Real Story of Mileva Einstein-Marić, Cambridge, MA: MIT Press, 2019, 336 pages, $29.99 (hardcover). Physics in Perspective (PIP), 21(3):253–254, September 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Home:2005:PTP

[Hom05] R. W. Home. The physical tourist: Physics in Melbourne. Physics in Perspective (PIP), 7(4):473–490, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-005-0222-x.

Hong:2000:OUT

[Hon00] Sungook Hong. Once upon a time in physics when both mathematics and experiment were helpless: a strange life of Voltaic contact potential. Physics in Perspective (PIP), 2(3): 269–292, September 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s000160050046.

Hon:2003:PSE

[Hon03] Giora Hon. From propagation to structure: The experimental technique of bombardment as a contributing factor to the emerging quantum physics. Physics in Perspective (PIP), 5 (2):150–173, May 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-003-0124-8.

Hooper:2008:NBS

[Hoo08] Dan Hooper. Nature’s blueprint: supersymmetry and the search for a unified theory of matter and force. Smithsonian Books, Washington, DC, USA, 2008. ISBN 0-06-155836-

2. 230 pp. LCCN QC174.17.S9 H66 2008. URL http://

www.loc.gov/catdir/enhancements/fy0911/2008013392-b.
html;   http://www.loc.gov/catdir/enhancements/fy0911/
2008013392-d.html;   http://www.loc.gov/catdir/toc/ecip0814/
2008013392.html.

Hore:2003:PBS

[Hor03] Peter Hore. Patrick Blackett: sailor, scientist, and socialist. Frank Cass, London, UK; Portland, OR, USA, 2003. ISBN 0-7146-5317-9 (hardcover). xiii + 330 + 16 pp. LCCN

QC16.B59 P37 2003. URL http://www.loc.gov/catdir/
enhancements/fy0652/2002067626-d.html;   http://www.
loc.gov/catdir/toc/fy034/2002067626.html. With a fore-

word by Tam Dalyell.

Hentschel:2001:PUG

[HR01] Klaus Hentschel and Gerhard Rammer. Physicists at the University of Göttingen, 1945–1955. Physics in Perspective (PIP), 3(2):189–209, June 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/PL00000529.

Heering:2000:ILP

[HRS00] Peter Heering, Falk Rieß, and Christian Sichau. Im Labor der Physikgeschichte: zur Untersuchung historischer Experimentalpraxis. (German) [In the laboratory of the history of physics — Research on historical experiments]. Bibliotheks- und Informationssystem der Univ., Oldenburg, Germany, 2000. ISBN 3-8142-0735-1. 220 pp. LCCN ????

Harrison:2005:BR

[HTR05] Michael J. Harrison, Edwin F. Taylor, and Peter J. Ramberg. Book reviews. Physics in Perspective (PIP), 7(1):130–134, March 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-003-0220-4.

Hu:2012:MJK

[Hu12] Danian Hu. Martin J. Klein: From physicist to historian. Physics in Perspective (PIP), 14(4):498–507, December 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-012-0095-8.

Hu:2016:AIC

[Hu16] Danian Hu. American influence on Chinese physics study in the early Twentieth Century. Physics in Perspective (PIP), 17 (4):268–297, January 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-015-0174-8.

Hu:2019:CLB

[Hu19a] Danian Hu. Correction to: From Liverpool to Beijing and Chongqing: William Band’s Adventure in Wartime China. Physics in Perspective (PIP), 21(3):252, September 2019. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00016-019-00244-x.pdf. See [Hu19b].

Hu:2019:LBC

[Hu19b] Danian Hu. From Liverpool to Beijing and Chongqing: William Band’s adventure in wartime China. Physics in Perspective (PIP), 21(3):222–251, September 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). See correction [Hu19a].

Hubisz:2006:BRM

[Hub06] John L. Hubisz. Book review: Marcus Hellyer, Catholic Physics: Jesuit Natural Philosophy in Early Modern Germany. Notre Dame: University of Notre Dame Press, 2005, xii + 337 pages. $50.00 (cloth). Physics in Perspective (PIP), 8(4):477–478, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Huff:2011:ICS

[Huf11] Toby E. Huff. Intellectual curiosity and the scientific revolution: a global perspective. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-107-00082-3, 0-521-17052-4 (paperback). xiii + 354 pp. LCCN Q127.E8

H84 2011. URL http://assets.cambridge.org/97805211/
70529/cover/9780521170529.jpg;   http://www.loc.gov/
catdir/enhancements/fy1010/2010021876-b.html;   http:
//www.loc.gov/catdir/enhancements/fy1010/2010021876-
d.html;   http://www.loc.gov/catdir/enhancements/fy1010/
2010021876-t.html.

Hufbauer:2012:SPH

[Huf12] Karl Hufbauer. From student of physics to historian of science: T. S. Kuhn’s education and early career, 1940–1958. Physics in Perspective (PIP), 14(4):421–470, December 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-012-0098-5.

Huggett:1999:SZE

[Hug99] Nick Huggett. Space from Zeno to Einstein: classic readings with a contemporary commentary. MIT Press, Cambridge, MA, USA, 1999. ISBN 0-262-08271-3 (hardcover), 0-262-58169-8 (paperback). xi + 274 pp. LCCN QC6 .S6625 1997.

Hughes:2009:BRM

[Hug09] Jeff Hughes. Book review: Maria Rentetzi, Trafficking Materials and Gendered Experimental Practices: Radium Research in Early 20th Century Vienna. 0-231-50959-6. Gutenberg, Columbia University Press. 2007. Physics in Perspective (PIP), 11(3):348–350, September 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://www.

gutenberge.org/rentetzi.

Huggett:2010:EEA

[Hug10] Nick Huggett. Everywhere and everywhen: adventures in physics and philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-537951-9 (hardcover), 0-19-537950-0 (paperback). xiii + 217 pp. LCCN QC6 .H824 2010.

Hunt:2001:BRH

[Hun01] J. Christopher Hunt. Book review: Helge Kragh, Cosmology and Controversy: The Historical Development of Two Theories of the Universe. Princeton: Princeton University Press, 1996, xiii + 500 pages $77.50 (cloth), $19.95 (paper). Physics in Perspective (PIP), 3(2):249–250, June 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hunter:2004:LML

[Hun04] Graeme K. Hunter. Light is a messenger: the life and science of William Lawrence Bragg. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-852921-x. xxi + 301 pp. LCCN ????

Hunt:2010:PPL

[Hun10] Bruce J. Hunt. Pursuing power and light: technology and physics from James Watt to Albert Einstein. Johns Hopkins introductory studies in the history of science. The Johns Hopkins University Press, Baltimore, MD, USA, 2010. ISBN 0-8018-9358-5 (hardcover), 0-8018-9359-3 (paperback). 182 pp. LCCN T173.8 .H92 2010.

Hunziker:2015:PTA

[Hun15] Herbert Hunziker. The physical tourist: Albert Einstein’s magic mountain: An Aarau education. Physics in Perspective (PIP), 17(1):55–69, March 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-014-0153-5.

Hey:1987:QU

[HW87] Anthony J. G. Hey and Patrick Walters. The Quantum Universe. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-26744-7, 0-521-31845-9 (paperback). vii + 180 pp. LCCN QC174.12 .H48 1987. URL http://www.loc.

gov/catdir/description/cam023/86006830.html;   http:
//www.loc.gov/catdir/toc/cam028/86006830.html.

Hey:2003:NQU

[HW03] Anthony J. G. Hey and Patrick Walters. The New Quantum Universe. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-56418-2, 0-521-56457-3 (paperback). xv + 357 pp. LCCN QC174.12 .H478 2003. URL http://www.loc.gov/

catdir/description/cam031/2002074047.html;   http://
www.loc.gov/catdir/enhancements/fy0731/2002074047-b.
html;   http://www.loc.gov/catdir/toc/cam031/2002074047.
html.

Hoffmann:2008:PZA

[HW08] Dieter Hoffmann and Mark Walker. Physiker zwischen Autonomie und Anpassung: Die Deutsche Physikalische Gesellschaft im Dritten Reich. (German) [Physicists between autonomy and adjustment: The German Physical Society in the Third Reich]. Wiley-VCH, Berlin, Germany; Weinheim, Germany; New York, NY, USA, 2008. ISBN 3-527-40585-2. xii + 675 pp. LCCN QC9.G3 .P49 2007.

URL http://ebooks.ciando.com/book/index.cfm/bok_id/
477257;   http://gbv.eblib.com/patron/FullRecord.aspx?
p=481854;   http://site.ebrary.com/lib/alltitles/docDetail.
action?docID=10518774;   http://www.ciando.com/img/books/
width167/3527622217_k.jpg;   http://www.ciando.com/pictures/
bib/3527622217bib_t\_1.jpg.

Hoffmann:2011:BFS

[HW11] Ronald Hoffmann and Iain Boyd Whyte, editors. Beyond the finite: the sublime in art and science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-973769-X (hardcover). xvii + 173 pp. LCCN Q175.32.A47 B49 2011.

Illy:2012:PEE

[Ill12] József Illy. The practical Einstein: experiments, patents, inventions. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0457-5 (hardcover), 1-4214-0533-4 (ebook). xi + 202 pp. LCCN QC16.E5 I45 2012.

Illy:2019:EG

[Ill19] József Illy. Einstein’s gyros. Physics in Perspective (PIP), 21(4):274–295, December 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00016-019-00248-7.

pdf. See correction [Ill20].

Illy:2020:CEG

[Ill20] József Illy. Correction to: Einstein’s gyros. Physics in Perspective (PIP), 22(1):52, March 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00016-020-00250-4.

pdf. See [Ill19].

Inwood:2002:MWK

[Inw02] Stephen Inwood. The man who knew too much: the strange and inventive life of Robert Hooke, 1635–1703. MacMillan Publishing Company, New York, NY, USA, 2002. ISBN 0-333-78286-0. xxix + 497 + 16 pp. LCCN Q143.H7 I58 2002.

Isaacs:2000:DP

[Isa00] Alan Isaacs. A dictionary of physics. Oxford paperback reference. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, fourth edition, 2000. ISBN 0-19-280103-1. 546 pp.

LCCN QC5 .C56 2000. URL http://www.loc.gov/catdir/
enhancements/fy0636/00501864-d.html.

Ienna:2023:JEC

[IT23] Gerardo Ienna and Simone Turchetti. JASON in Europe: Contestation and the physicists’ dilemma about the Vietnam War. Physics in Perspective (PIP), 25(3):??, November 2023. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL https://link.springer.com/article/10.
1007/s00016-023-00302-5.

Jackson:2000:SBJ

[Jac00] Myles W. Jackson. Spectrum of Belief: Joseph von Fraunhofer and the Craft of Precision Optics. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-10084-3, 0-226-38932-4 (hardcover), 0-226-38933-2 (paperback). x + 284 pp. LCCN QC450.5 .J33 2000.

Jacob:2003:BRD

[Jac03] Richard Jacob. Book review: D. Allan Bromley, A Century of Physics. New York: Springer-Verlag, 2001, x + 114 pages. $59.95 (cloth). Physics in Perspective (PIP), 5(1):126, April 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Jacob:2004:BRK

[Jac04] Richard J. Jacob. Book review: Kenneth D. Bergeron, Tritium on Ice. Cambridge, Mass.: The MIT Press, 2002, xii + 234 pages. $24.95 (cloth). Physics in Perspective (PIP), 6(1):121– 122, April 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Jackson:2006:HTP

[Jac06] Myles W. Jackson. Harmonious triads: physicists, musicians, and instrument makers in Nineteenth-Century Germany. Transformations. MIT Press, Cambridge, MA, USA, 2006. ISBN 0-262-10116-5. x + 395 pp. LCCN ML3805 .J33 2006. URL

http://www.loc.gov/catdir/toc/fy0710/2006044991.html.

Jackiw:2009:HBM

[Jac09] Roman Jackiw. Hans Bethe, my teacher. Physics in Perspective (PIP), 11(1):98–103, March 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-008-0412-4;
 http://www.springerlink.com/content/707k21m7328t0860/

.

Jackson:2010:PAM

[Jac10] John David Jackson. A personal adventure in muon-catalyzed fusion. Physics in Perspective (PIP), 12(1):74–88, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00016-009-0006-9.pdf.

Jackiw:2011:CG

[Jac11] Roman Jackiw. Celebration of Gerry. Physics in Perspective (PIP), 13(1):104–109, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-010-0036-3.

Jacquette:2014:NMS

[Jac14] Dale Jacquette. Newton’s metaphysics of space as God’s emanative effect. Physics in Perspective (PIP), 16(3):344–370, September 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-014-0142-8.

James:2000:BRW

[Jam00] Philip B. James. Book review: William Poundstone, Carl Sagan: A Life in the Cosmos. New York, Henry Holt and Company, LLC, 1999, xii + 473 pages. $30.00 (cloth). Physics in Perspective (PIP), 2(2):219–220, June 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Jammer:2007:CTP

[Jam07] Max Jammer. Concepts of time in physics: a synopsis. Physics in Perspective (PIP), 9(3):266–280, September 2007. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
007-0349-z.

Janssen:2002:RSR

[Jan02] Michel Janssen. Reconsidering a scientific revolution: The case of Einstein versus Lorentz. Physics in Perspective (PIP), 4(4):421–446, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160200003;   http:/
/www.springerlink.com/content/53vxdwrfnm99h8l3/fulltext.
pdf.

Januszajtis:2011:WAG

[Jan11] Andrzej Januszajtis. A walk around Gdańsk for physicists. Physics in Perspective (PIP), 13(4):456–480, December 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00016-011-0066-5.pdf.

Jayawardhana:2011:SNW

[Jay11] Ray Jayawardhana. Strange new worlds: the search for alien planets and life beyond our solar system. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14254-8 (paperback). 255 pp. LCCN QB820 .J39 2011.

Jenkin:2002:GMJ

[Jen02] John Jenkin. G. E. M. Jauncey and the Compton Effect. Physics in Perspective (PIP), 4(3):320–332, August 2002. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
002-8371-7.

Jenkin:2008:WLB

[Jen08] John (John Grenfell) Jenkin. William and Lawrence Bragg, father and son: the most extraordinary collaboration in science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-923520-1 (hardcover), 0-19-960670-6 (paperback). xiv + 458 pp. LCCN QC15. £35.00£35.00. URL http:

//www.gbv.de/dms/bowker/toc/9780199235209.pdf.

Jenkin:2011:AEM

[Jen11] John G. Jenkin. Atomic energy is “moonshine”: What did Rutherford really mean? Physics in Perspective (PIP), 13 (2):128–145, June 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-010-0038-1.

Joaquim:2015:QEB

[JFEH15] Leyla Joaquim, Olival Freire, Jr., and Charbel N. El-Hani. Quantum explorers: Bohr, Jordan, and Delbrück venturing into biology. Physics in Perspective (PIP), 17(3):236–250, September 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-015-0167-7.

Jha:2011:WPE

[Jha11] Stefania Jha. Wigner’s “Polanyian” epistemology and the measurement problem: The Wigner–Polanyi dialog on tacit knowledge. Physics in Perspective (PIP), 13(3):329–358, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-010-0050-5.

Johnson:1999:SBM

[Joh99a] George Johnson. Strange Beauty: Murray Gell-Mann and the Revolution in Twentieth-Century Physics. Vintage Books, New York, NY, USA, 1999. ISBN 0-679-75688-4. xii + 434 pp. LCCN QC774.G45 J65 2000.

Johnson:1999:SBT

[Joh99b] K. E. Johnson. Science at the breakfast table. Physics in Perspective (PIP), 1(1):22–34, March 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http:/

/link.springer.com/article/10.1007/s000160050003.

Johnson:2004:NHN

[Joh04] Karen E. Johnson. From natural history to the nuclear shell model: Chemical thinking in the work of Mayer, Haxel, Jensen, and Suess. Physics in Perspective (PIP), 6(3):295–309, September 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-003-0203-x.

Johnston:2006:ANS

[Joh06a] Sean F. Johnston. Absorbing new subjects: Holography as an analog of photography. Physics in Perspective (PIP), 8 (2):164–188, May 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0264-8.

Johnston:2006:PTP

[Joh06b] Sean F. Johnston. The physical tourist: Physics in Glasgow: a heritage tour. Physics in Perspective (PIP), 8(4):451–465, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-006-0310-6.

Johnston:2009:HSB

[Joh09] Sean Johnston. History of science: a beginner’s guide. Oneworld beginners guides. Oneworld, Oxford, UK, 2009. ISBN 1-85168-681-9 (paperback). viii + 218 pp. LCCN Q126 .J64 2009.

Jones:2007:BRJ

[Jon07] Stanley T. Jones. Book review: Jurgen Audretsch, ed., Entangled World: The Fascination of Quantum Information and Computation, translated from the German by Rudolph Ehrlich. Weinheim: Wiley-VCH, 2006, xii + 300 pages. $35.95 (cloth). Physics in Perspective (PIP), 9(3):381–383, September 2007.

CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Jones:2008:QTS

[Jon08] Sheilla Jones. The Quantum Ten: a Story of Passion, Tragedy, Ambition and Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-536909-2. xii + 323 + 8 pp. LCCN QC174.12 .J66 2008.

Josephson:2010:LLZ

[Jos10] Paul R. Josephson. Lenin’s laureate: Zhores Alferov’s life in Communist science. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01458-0. 307 pp. LCCN QC16.A3417.

Johnson:1999:MAR

[JP99] Karen E. Johnson and Donald C. Peckham. In memoriam: Alfred Romer (1906–1998). Physics in Perspective (PIP), 1 (2):215–218, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s000160050016.

James:2007:CSS

[JP07] Frank A. J. L. James and Anthony Peers. Constructing space for science at the Royal Institution of Great Britain. Physics in Perspective (PIP), 9(2):130–185, June 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-006-0303-5.

Jackiw:2002:ADB

[JS02] Roman Jackiw and Abner Shimony. In appreciation: The depth and breadth of John Bell’s physics. Physics in Perspective (PIP), 4(1):78–116, February 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-002-8359-3.

Kupershtokh:2014:PNA

[KA14] Natalia Kupershtokh and Alexander Apolonskiy. Physics in Novosibirsk and Akademgorodok. Physics in Perspective (PIP), 16(2):250–276, June 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-014-0138-4;

 http://link.springer.com/content/pdf/10.1007/s00016-

014-0138-4.pdf. See erratum [KA15].

Kupershtokh:2015:EPN

[KA15] Natalia Kupershtokh and Alexander Apolonskiy. Erratum to: Physics in Novosibirsk and Akademgorodok. Physics in Perspective (PIP), 17(2):169–170, June 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-014-0145-5;
 http://link.springer.com/content/pdf/10.1007/s00016-
014-0145-5.pdf. See [KA14].

Kahn:2003:BRM

[Kah03a] Peter B. Kahn. Book review: Michael W. Friedlander, A Thin Cosmic Rain: Particles from Outer Space. Cambridge, Mass.: Harvard University Press, 2000, 241 pages. $29.95 (cloth), $17.95 (paper). Physics in Perspective (PIP), 5(2):234–235, May 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Kahn:2003:ARM

[Kah03b] Peter B. Kahn. In appreciation: Remembering Max Dresden (1918–1997). Physics in Perspective (PIP), 5(2):206–233, May 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-003-0167-x.

Kapusta:2008:ADS

[Kap08] Joseph I. Kapusta. Accelerator disaster scenarios, the Unabomber, and scientific risks. Physics in Perspective (PIP), 10 (2):163–181, June 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-007-0366-y.

Katzir:2005:PRP

[Kat05] Shaul Katzir. Poincaré’s relativistic physics: Its origins and nature. Physics in Perspective (PIP), 7(3):268–292, September 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-004-0234-y.

Kahn:2003:BR

[KBS03] Peter B. Kahn, Stephen G. Brush, and David P. Stern. Book reviews. Physics in Perspective (PIP), 5(2):234–239, May 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-003-0159-x.

Kelly:2007:MPB

[Kel07] Cynthia C. Kelly, editor. The Manhattan Project: the birth of the atomic bomb by its creators, eyewitnesses, and historians. Black Dog and Leventhal Publishers, New York, NY, USA, 2007. ISBN 1-57912-747-9, 1-57912-808-4 (paperback), 1-60376-206-X (e-book). xiv + 495 pp. LCCN QC773.3.U5

M27 2007. URL http://www.loc.gov/catdir/toc/ecip0720/
2007023984.html.

Kelly:2022:MTE

[Kel22] Nadya D. Kelly. (Mis)Translating entropy?: Camille Flammarion and the multiple theologies of the death of the Universe. Physics in Perspective (PIP), 24(4):208–222, December 2022. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL https://link.springer.com/article/10.
1007/s00016-022-00293-9.

Kennefick:2007:TST

[Ken07] Daniel Kennefick. Traveling at the speed of thought: Einstein and the quest for gravitational waves. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-11727-6 (hardcover). xii + 319 pp. LCCN QC179 .K46 2007. URL http://

www.loc.gov/catdir/enhancements/fy0834/2006938366-b.
html;   http://www.loc.gov/catdir/enhancements/fy0834/
2006938366-d.html;   http://www.loc.gov/catdir/enhancements/
fy0834/2006938366-t.html.

Kevles:1997:NBM

[Kev97] Bettyann Kevles. Naked to the Bone: Medical Imaging in the Twentieth Century. The Sloan technology series. Rutgers University Press, New Brunswick, NJ, USA, 1997. ISBN 0-8135-2358-3 (hardcover), 0-201-32833-X (paperback), 0-585-02803-6 (e-book). xiv + 378 pp. LCCN RC78.7.D53 K48 1997.

Khelfaoui:2019:PRP

[KG19] Mahdi Khelfaoui and Yves Gingras. Physical Review: From the periphery to the center of physics. Physics in Perspective (PIP), 21(1):23–42, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Kleppner:2011:BR

[KHC11] Daniel Kleppner, Robert C. Hilborn, and David C. Cassidy. Book reviews. Physics in Perspective (PIP), 13(2):244–255, June 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-011-0061-x.

Kipnis:2000:WOL

[Kip00] Nahum S. Kipnis. The window of opportunity: Logic and chance in Becquerel’s discovery of radioactivity. Physics in Perspective (PIP), 2(1):63–99, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050037.

Kirshner:2002:EUE

[Kir02] Robert P. Kirshner. The Extravagant Universe: Exploding Stars, Dark Energy, and the Accelerating Cosmos. Princeton University Press, Princeton, NJ, USA, 2002. ISBN 0-691-05862-

8. xii + 282 pp. LCCN QB843.S95 K57 2002. URL http:

//www.loc.gov/catdir/description/prin022/2002029268.
html;   http://www.loc.gov/catdir/samples/prin031/2002029268.
html;   http://www.loc.gov/catdir/toc/prin031/2002029268.
html.

Kuliev:2019:ICS

[KKG^{+}19] A. M. Kuliev, V. I. Kukharenko, K. N. Grinberg, V. V. Terskikh, A. D. Tamarkina, E. A. Bogomazov, P. S. Redkin, and S. S. Vasileysky. Investigation of a cell strain with trisomy 14 from a spontaneously aborted human fetus. Physics in Perspective (PIP), 21(1):1–12, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/content/pdf/10.1007/BF00278558.pdf.

Kleppner:2007:BRA

[Kle07] Daniel Kleppner. Book review: Abraham Pais, J. Robert Oppenheimer: A Life, with supplemental material by Robert P.

Crease. Oxford: Oxford University Press, 2006, xxii + 353 pages. $30.00 (cloth). Physics in Perspective (PIP), 9(4):505– 508, December 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Klein:2010:PEN

[Kle10] Martin J. Klein. Paul Ehrenfest, Niels Bohr, and Albert Einstein: Colleagues and friends. Physics in Perspective (PIP), 12(3):307–337, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-010-0025-6;
 http://www.springerlink.com/content/g08710644434v387/
fulltext.pdf.

Kleppner:2011:BRD

[Kle11] Daniel Kleppner. Book review: David Goodstein, Fact and Fraud in Science: Cautionary Tales from the Front Lines of Science. Princeton: Princeton University Press, 2010, xiv + 168 pages. $22.95 (cloth). Physics in Perspective (PIP), 13(2):244– 247, June 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Kleemans:2023:DDE

[Kle23] Machiel Kleemans. Dreams of declassification: The early Cold War quest for nuclear knowledge in The Netherlands and Norway. Physics in Perspective (PIP), 25(3):??, November 2023. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL https://link.springer.com/article/10.
1007/s00016-023-00305-2.

Kleppner:2007:BR

[KMG^{+}07] Daniel Kleppner, Barry R. Masters, Thomas B. Greenslade Jr., Allan Franklin, and Anders Carlsson. Book reviews. Physics in Perspective (PIP), 9(4):505–515, December 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
006-0331-1.

Knight:2006:ITM

[Kni06] Randall D. Knight. Inventing temperature: Measurement and scientific progress. Physics in Perspective (PIP), 8(4):483–486, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Kobe:2002:BRR

[Kob02] Donald Kobe. Book review: Roland Omnès, Understanding Quantum Mechanics. Princeton: Princeton University Press, 1999, xiii + 307 pages. $35.00 (Hardbound). Physics in Perspective (PIP), 4(3):360–361, August 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Kosso:2013:VPR

[Kos13] Peter Kosso. Void points, rosettes, and a brief history of planetary astronomy. Physics in Perspective (PIP), 15(4):373– 390, December 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-013-0112-6.

Kovacs:2003:PTB

[Kov03] László Kovács, Sr. The physical tourist: Budapest: a random walk in science and culture. Physics in Perspective (PIP), 5(3): 310–348, September 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-003-0174-y.

Kalos:2002:BRM

[KP02] Malvin H. Kalos and Douglass E. Post. Book review: Martin Davis, The Universal Computer: The Road From Leibniz to Turing. New York: W. W. Norton and Company, 2000, xii + 237 pages. $25.95 (cloth). Physics in Perspective (PIP), 4(1): 118–119, February 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Kronz:2010:BR

[KPP^{+}10] Frederick M. Kronz, Jay M. Pasachoff, Naomi Pasachoff, John L. Roeder, Hans Christian von Baeyer, Peter Lindenfeld, and Ruth Lewin Sime. Book reviews. Physics in Perspective (PIP), 12(1):102–117, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-009-0007-8.

Kragh:1996:CCH

[Kra96] Helge Kragh. Cosmology and Controversy: the Historical Development of Two Theories of the Universe. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-02623-8 (hardcover). xiii + 500 pp. LCCN QB981 .K73 1996.

URL http://www.loc.gov/catdir/description/prin031/
96005612.html;   http://www.loc.gov/catdir/toc/prin031/
96005612.html.

Kragh:1999:QGH

[Kra99] Helge Kragh. Quantum generations: a history of physics in the twentieth century. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-01206-7 (hardcover), 0-691-09552-3 (paperback). xiv + 494 pp. LCCN QC7 .K7 2002. URL http:

//press.princeton.edu/titles/6683.html.

Kragh:2000:UCG

[Kra00] Helge Kragh. An unlikely connection: Geochemistry and nuclear structure. Physics in Perspective (PIP), 2(4):381–397, December 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s000160050051.

Kragh:2007:CCM

[Kra07] Helge Kragh. Conceptions of cosmos: from myths to the accelerating universe: a history of cosmology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-920916-2. 276 pp. LCCN QB981 .K729 2007. URL http://

www.loc.gov/catdir/enhancements/fy0723/2006027692-b.
html;   http://www.loc.gov/catdir/toc/ecip0619/2006027692.
html.

Kragh:2008:MWS

[Kra08a] Helge Kragh. The moon that wasn’t: the saga of Venus’ spurious satellite, volume 37 of Science networks. Historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2008. ISBN 3-7643-8908-7 (hardcover), 3-7643-8909-5 (e-book). xii + 199 pp. LCCN QB401 .K73 2008. With the assistance of Kurt Møller Pedersen.

Kragh:2008:PDE

[Kra08b] Helge Kragh. Pierre Duhem, entropy, and Christian faith. Physics in Perspective (PIP), 10(4):379–395, December 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-007-0365-z.

Kragh:2011:RBA

[Kra11a] Helge Kragh. Resisting the Bohr atom: The early British opposition. Physics in Perspective (PIP), 13(1):4–35, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-010-0048-z.

Krauss:2011:QMR

[Kra11b] Lawrence Maxwell Krauss. Quantum Man: Richard Feynman’s Life in Science. Great discoveries. W. W. Norton & Co., New York, NY, USA, 2011. ISBN 0-393-06471-9. xvii + 350 pp. LCCN QC16.F49 K73 2011. URL http://www.

scientificcomputing.com/news-DS-Illuminating-the-Life-
and-Legacy-of-Richard-Feynman-032211.aspx.

Kragh:2012:ZU

[Kra12] Helge Kragh. Zöllner’s Universe. Physics in Perspective (PIP), 14(4):392–420, December 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-012-0099-4.

Kragh:2015:PJV

[Kra15] Helge Kragh. Pascual Jordan, varying gravity, and the expanding Earth. Physics in Perspective (PIP), 17(2):107–134, June 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-015-0157-9.

Kragh:2016:BRF

[Kra16] Helge Kragh. Book review: Finn Aaserud and John Heilbron, Love, Literature, and the Quantum Atom: Bohr’s 1913 Trilogy Revisited. Oxford: Oxford University Press, 2013, x + 284 pages, $61 (hardcover). Physics in Perspective (PIP), 18(3): 351–354, September 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-016-0188-x.

Kragh:2018:LLH

[Kra18] Helge Kragh. Ludvig Lorenz and his non–Maxwellian electrical theory of light. Physics in Perspective (PIP), 20(3):221–253, September 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Kragh:2019:VCN

[Kra19] Helge Kragh. Varying constants of nature: Fragments of a history. Physics in Perspective (PIP), 21(4):257–273, December 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Kragh:2022:BGM

[Kra22] Helge Kragh. Brecht, Galileo, and Møller: a view from Copenhagen, 1938–1939. Physics in Perspective (PIP), 24(2–3):99– 124, October 2022. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL https://link.springer.com/
article/10.1007/s00016-022-00289-5.

Krige:2005:IRC

[Kri05] John Krige. Isidor I. Rabi and CERN. Physics in Perspective (PIP), 7(2):150–164, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-004-0225-z.

Kronz:2010:BRA

[Kro10] Frederick M. Kronz. Book review: Alisa Bokulich, Reexamining the Quantum–Classical Relation: Beyond Reductionism and Pluralism. Cambridge: Cambridge University Press, 2008, x + 195 pages. $80.00 (cloth). Physics in Perspective (PIP), 12 (1):102–105, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Kobe:2002:BR

[KSAF02] Donald Kobe, Alan E. Shapiro, John F. Ahearne, and Kenneth W. Ford. Book reviews. Physics in Perspective (PIP), 4 (3):360–366, August 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-002-8373-5.

Kurian:2002:NSB

[Kur02] George Thomas Kurian. The Nobel scientists: a biographical encyclopedia. Prometheus Books, Amherst, NY, USA, 2002. ISBN 1-57392-927-1. 420 pp. LCCN Q141 .K78 2002.

Kolb:2008:FPF

[KW08] Adrienne. Kolb and Catherine Westfall. Fermilab: Physics, the Frontier, and Megascience. University of Chicago Press, Chicago, IL, USA and London, UK, 2008. ISBN 0-226-34623-4 (hardcover), 0-226-34625-0 (e-book), 1-282-23966-X. xi + 497

pp. LCCN HF1008. URL http://www.myilibrary.com?id=
223966;   http://www.myilibrary.com?id=223966\%26ref=
toc.

Kortemeyer:2010:PT

[KW10] Gerd Kortemeyer and Catherine Westfall. The physical tourist. Physics in Perspective (PIP), 12(1):89–99, March 2010. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
009-0005-x.

Lacki:2007:PTG

[Lac07] Jan Lacki. The physical tourist: Geneva: From the science of the Enlightenment to CERN. Physics in Perspective (PIP), 9 (2):231–252, June 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-007-0327-5.

Laidler:2004:SSE

[Lai04] Keith James Laidler. Science and sensibility: the elegant logic of the universe. Prometheus Books, Amherst, NY, USA, 2004. ISBN 1-59102-138-3. 233 pp. LCCN Q162

.L315 2004. URL http://www.loc.gov/catdir/toc/ecip048/
2003018605.html.

Lalli:2019:BRD

[Lal19] Roberto Lalli. Book review: Dominique Lambert, The Atom of the Universe: The Life and Work of Georges Lemaître, Krakow: Copernicus Center Press, 2015, 464 + xix pages, EUR 59.90 (hardcover). Physics in Perspective (PIP), 21(1):87–90, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Lambert:2014:HPK

[Lam14] Kevin Lambert. Hearing Pygmalion’s kiss: A scientific object at the Paris Opéra. Physics in Perspective (PIP), 16(4): 417–439, December 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-014-0148-2.

Landolt:2010:BRC

[Lan10] Arlo U. Landolt. Book review: Carlos I. Calle, The Universe: Order Without Design. Amherst, NY: Prometheus Books, 2009, 304 pages. $27.98 (cloth). Physics in Perspective (PIP), 12(2): 242–243, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Landolt:2012:HWS

[Lan12] Arlo U. Landolt. How we see the sky: a naked-eye tour of day and night. Physics in Perspective (PIP), 14(4):518–519, December 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Larsen:2005:SHB

[Lar05a] Kristine M. Larsen. Stephen Hawking: a biography. Greenwood biographies. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2005. ISBN 0-313-32392-5. ISSN 1540-4900. xix + 165 pp. LCCN QC16.H33 L37 2005. URL http://www.

loc.gov/catdir/toc/ecip0511/2005011150.html.

Larson:2005:BR

[Lar05b] Michelle B. Larson. Book review. Physics in Perspective (PIP), 7(4):491–502, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-005-0268-9.

Larson:2005:BRK

[Lar05c] Michelle B. Larson. Book review: Keith J. Laidler, Science and Sensibility: The Elegant Logic of the Universe. Amherst, New York: Prometheus Books, 2004, 233 pages. $28.00 (cloth). Physics in Perspective (PIP), 7(4):501–502, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Larson:2007:PCI

[Lar07] Ronald G. Larson. Perspectives on current issues is “anthropic selection” science? Physics in Perspective (PIP), 9(1):58–69, January 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-007-0333-7.

Labinger:2001:OCC

[LC01] Jay A. Labinger and H. M. (Harry M.) Collins, editors. The one culture?: a conversation about science. University of Chicago Press, Chicago, IL, USA and London, UK, 2001. ISBN 0-226-46722-8 (hardcover), 0-226-46723-6 (paper). xi + 329 pp. LCCN Q175.55.O54 2001.

Leibowitz:2008:HHC

[Lei08] J. R. (Jack R.) Leibowitz. Hidden harmony: the connected worlds of physics and art. The Johns Hopkins University Press, Baltimore, MD, USA, 2008. ISBN 0-8018-8866-2 (hardcover). xi + 148 pp. LCCN N70 .L454 2008. URL http://

www.loc.gov/catdir/enhancements/fy0805/2007043958-b.
html;   http://www.loc.gov/catdir/enhancements/fy0805/
2007043958-d.html;   http://www.loc.gov/catdir/toc/ecip083/
2007043958.html.

Lustig:2004:BR

[LEJ04] Harry Lustig, William E. Evenson, and Richard J. Jacob. Book reviews. Physics in Perspective (PIP), 6(1):114–122, April 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-003-0189-4.

Lemmerich:2007:AIS

[Lem07] Jost Lemmerich. Aufrecht im Sturm der Zeit: Der Physiker James Franck: 1882–1964. (German) [Upright in the storm of time: the physicist James Franck]. Verlag für Geschichte der Naturwissenschaften und der Technik, Diepholz, Germany, 2007. ISBN 3-928186-83-3. 362 pp. LCCN QC16.F67

2007. URL http://paperc.de/4167-aufrecht-im-sturm-
der-zeit-9783928186834;   http://www.gnt-verlag.de/programm/

83/. See also English translation [LH11].

Lepage:2007:BRD

[Lep07] G. Peter Lepage. Book review: David Kaiser, Drawing Theories Apart: The Dispersion of Feynman Diagrams in Postwar Physics. Chicago: University of Chicago Press, 2005, xix + 469 pages. $30.00 (paper). Physics in Perspective (PIP), 9(1):120– 122, January 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Levine:2004:EGW

[Lev04] James L. Levine. Early gravity-wave detection experiments, 1960–1975. Physics in Perspective (PIP), 6(1):42–75, April 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-003-0179-6.

Levitt:2009:SEO

[Lev09] Theresa Levitt. The shadow of enlightenment: optical and political transparency in France, 1789–1848. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954470-0. 192 pp. LCCN TA1522 .L48 2009.

Lippincott:2008:CVT

[LG08] Sara Lippincott and Judith R. Goodstein. A conversation with Valentine L. Telegdi — Part II. Physics in Perspective (PIP), 10(1):77–109, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0307-1.

Lemmerich:2011:SCL

[LH11] Jost Lemmerich and Ann Hentschel. Science and conscience: the life of James Franck. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 2011. ISBN 0-8047-6310-0. xviii + 369 pp. LCCN QC16.F67 L45513 2011. Translation of German original [Lem07].

Lindenfeld:2003:HUG

[Lin03] P. Lindenfeld. How the universe got its spots: Diary of a finite time in a finite space. Physics in Perspective (PIP), 5(4): 475–476, December 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Lindley:2004:DKT

[Lin04] David Lindley. Degrees Kelvin: a tale of genius, invention, and tragedy. Joseph Henry Press, Washington, DC, USA, 2004. ISBN 0-309-09073-3. viii + 366 pp. LCCN QC16.K3 L56 2004. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/toc/ecip0410/2003022885.html.

Lincoln:2009:QFL

[Lin09] Don Lincoln. The Quantum Frontier: the Large Hadron Collider. The Johns Hopkins University Press, Baltimore, MD, USA, 2009. ISBN 0-8018-9144-2 (hardcover). xiv + 172 pp. LCCN QC793.5.B62 L56 2009. URL http://www.loc.gov/

catdir/toc/ecip0819/2008022647.html.

Lindenfeld:2010:BRD

[Lin10] Peter Lindenfeld. Book review: Deborah R. Coen, Vienna in the Age of Uncertainty: Science, Liberalism, and Private Life. Chicago and London: The University of Chicago Press, 2007, xi + 380 pages. $45.00 (cloth). Physics in Perspective (PIP), 12 (1):112–115, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Lippincott:2006:CRFa

[Lip06a] Sara Lippincott. A conversation with Robert F. Christy — Part I. Physics in Perspective (PIP), 8(3):282–317, September 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-006-0270-x.

Lippincott:2006:CRFb

[Lip06b] Sara Lippincott. A conversation with Robert F. Christy — Part II. Physics in Perspective (PIP), 8(4):408–450, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-006-0269-3.

Lippincott:2007:CVT

[Lip07] Sara Lippincott. A conversation with Valentine L. Telegdi — Part I. Physics in Perspective (PIP), 9(4):434–467, December 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-006-0306-2.

Lisle:2016:ES

[Lis16] John Lisle. Einstein up in smoke. Physics in Perspective (PIP), 17(4):354–360, January 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-015-0171-y.

Liu:2020:WMA

[Liu20] Liyuan Liu. When missionary astronomy encountered Chinese astrology: Johann Adam Schall von Bell and Chinese calendar reform in the Seventeenth Century. Physics in Perspective (PIP), 22(2):110–126, June 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Lykknes:2004:EGP

[LKK04] Annette Lykknes, Helge Kragh, and Lise Kvittingen. Ellen Gleditsch: Pioneer woman in radiochemistry. Physics in Perspective (PIP), 6(2):126–155, June 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-003-0187-6.

Lowe-Ma:2006:BRG

[LM06] Charlotte Lowe-Ma. Book review: Graeme K. Hunter, Light Is a Messenger: The Life and Science of William Lawrence Bragg. New York: Oxford University Press, 2004, xxi + 301 pages. $59.50 (cloth). Physics in Perspective (PIP), 8(1):106– 107, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Loettgers:2003:SPL

[Loe03] Andrea Loettgers. Samuel Pierpont Langley and his contributions to the empirical basis of black-body radiation. Physics in Perspective (PIP), 5(3):262–280, September 2003. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
003-0143-5.

LoSecco:2016:HAA

[LoS16] John M. LoSecco. The history of “anomalous” atmospheric neutrino events: A first person account. Physics in Perspective (PIP), 18(2):209–241, August 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-016-0185-0.

Leone:2004:SDC

[LPR04] Matteo Leone, Alessandro Paoletti, and Nadia Robotti. A simultaneous discovery: The case of Johannes Stark and Antonino Lo Surdo. Physics in Perspective (PIP), 6(3):271–294, September 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-003-0170-2.

Leone:2003:EEN

[LR03] Matteo Leone and Nadia Robotti. Are the elements elementary? Nineteenth-Century chemical and spectroscopical answers. Physics in Perspective (PIP), 5(4):360–383, December 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-003-0176-9.

Lustig:2004:BRW

[Lus04] Harry Lustig. Book review: William H. Cropper, Great Physicists: The Life and Times of Leading Physicists from Galileo to Hawking. New York: Oxford University Press, 2001, xii + 500 pages. $35.00 (cloth). Physics in Perspective (PIP), 6(1):114– 119, April 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Lustig:2010:ERL

[Lus10] Harry Lustig. Essay review: The life and times of Werner Heisenberg. David C. Cassidy, Beyond Uncertainty: Heisenberg, Quantum Physics, and the Bomb. New York: Bellevue Literary Press, 2009, 480 pages. $27.00 (cloth). Physics in Perspective (PIP), 12(4):470–496, December 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-010-0034-5;
 http://www.springerlink.com/content/h348l1l5x5955377/

.

Littmann:2009:TES

[LWE09] Mark Littmann, Ken Willcox, and Fred Espenak. Totality: eclipses of sun: updated with guides to total eclipses from 2009 through 2017. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, third edition, 2009. ISBN 0-19-956552-X (paperback). xv + 343 pp. LCCN QB541 .L69 2009.

Lyonhart:2020:BR

[Lyo20] Jonathan Lyonhart. Book review. Physics in Perspective (PIP), 22(4):243–245, December 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Maas:2007:EEC

[Maa07] Ad Maas. Einstein as engineer: The case of the little machine. Physics in Perspective (PIP), 9(3):305–328, September 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0308-0;   http://www.
springerlink.com/content/3x7341xrt45047gx/fulltext.pdf.

Magueijo:2009:BDE

[Mag09] João Magueijo. A brilliant darkness: the extraordinary life and disappearance of Ettore Majorana, the troubled genius of the nuclear age. Basic Books, New York, NY, USA, 2009. ISBN 0-465-00903-4. xxi + 280 pp. LCCN QC774.M34 M35 2009.

Majid:2008:ST

[Maj08] Shahn Majid, editor. On space and time. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-88926-X (hardcover). xx + 287 pp. LCCN QC173.59.S65 O49 2008. URL http://

www.loc.gov/catdir/enhancements/fy0834/2008013428-b.
html;   http://www.loc.gov/catdir/enhancements/fy0834/
2008013428-d.html;   http://www.loc.gov/catdir/enhancements/
fy0834/2008013428-t.html.

Majid:2012:ST

[Maj12] Shahn Majid, editor. On space and time. Canto classics. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-64168-3. xx + 320 pp. LCCN QC173.59.S65 2008.

Malley:2011:RHM

[Mal11] Marjorie Caroline Malley. Radioactivity: a history of a mysterious science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-976641-X (hardcover). xxi + 267 pp. LCCN QC794.6.R3 M35 2011.

Malin:2012:NLH

[Mal12] Shimon Malin. Nature Loves to Hide: Quantum Physics and the Nature of Reality, a Western Perspective (Revised Edition). World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2012. ISBN 981-4324-56-6, 981-4324-60-4 (e-book). xv + 290 pp. LCCN

QC174.12 .M34 2012. URL http://gbv.eblib.com/patron/
FullRecord.aspx?p=919130.

Mantovani:2022:MCI

[Man22] Roberto Mantovani. Making and collecting instruments in fair Verona: The case of the Italian amateur scientist Gaetano Spandri (1796–1859). Physics in Perspective (PIP), 24(1):3– 34, March 2022. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL https://link.springer.com/
article/10.1007/s00016-021-00283-3.

Mantovani:2023:BON

[Man23] Roberto Mantovani. Between old and new interpretations of life: Animal electricity at the First Congress of Italian Scientists. Physics in Perspective (PIP), 25(1–2):3–40, June 2023. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL https://link.springer.com/article/10.
1007/s00016-023-00296-0.

March:2003:PUW

[Mar03a] Robert H. March. Physics at the University of Wisconsin: a history. Physics in Perspective (PIP), 5(2):130–149, May 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-003-0142-6.

Martinez:2003:BRJ

[Mar03b] Alberto A. Martínez. Book review: John Stachel, Einstein from ‘B’ to ‘Z’. Boston/Basel… Berlin: Birkhäuser, 2002, xi + 556 pages. $69.95 (cloth). Physics in Perspective (PIP), 5(3): 352–354, September 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Martinez:2004:MHI

[Mar04a] Alberto A. Martínez. Material history and imaginary clocks: Poincaré, Einstein, and Galison on simultaneity. Physics in Perspective (PIP), 6(2):224–240, June 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-004-0209-z;
 http://www.springerlink.com/content/qd3k029w8d2lfp1m/
fulltext.pdf.

Martinez:2004:REE

[Mar04b] Alberto A. Martínez. Ritz, Einstein, and the Emission Hypothesis. Physics in Perspective (PIP), 6(1):4–28, April 2004. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
003-0195-6;   http://www.springerlink.com/content/jq5dekc9hntg2qte/

.

Martinez:2009:KLO

[Mar09] Alberto A. Martínez. Kinematics: the lost origins of Einstein’s relativity. The Johns Hopkins University Press, Baltimore, MD, USA, 2009. ISBN 0-8018-9135-3 (hardcover). xix + 464 pp. LCCN QA841 .M37 2009.

Martinez:2011:SST

[Mar11] Alberto A. Martínez. Science secrets: the truth about Darwin’s finches, Einstein’s wife, and other myths. University of Pittsburgh Press, Pittsburgh, PA, USA, 2011. ISBN 0-8229-4407-3 (hardcover). xviii + 324 pp. LCCN Q173 .M316 2011.

Marburger:2014:SSU

[Mar14] John H. Marburger III. The Superconducting Supercollider and US science policy. Physics in Perspective (PIP), 16(2):218–249, June 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-014-0133-9.

Martin:2015:WNC

[Mar15] Joseph D. Martin. What’s in a name change? Physics in Perspective (PIP), 17(1):3–32, March 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-014-0151-7.

Martin:2020:E

[Mar20] Joseph D. Martin. Editorial. Physics in Perspective (PIP), 22(3):127–128, September 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00016-020-00262-0.
pdf.

Martin:2021:RS

[Mar21] Joseph D. Martin. Reporting science. Physics in Perspective (PIP), 23(2–3):83–84, October 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00016-021-00279-z.
pdf.

Masters:2007:BRM

[Mas07] Barry R. Masters. Book review: Mario Biagioli, Galileo’s Instruments of Credit: Telescopes, Images, Secrecy. Chicago: The University of Chicago Press, 2006, 302 pages. $35.00 (cloth). Physics in Perspective (PIP), 9(4):508–510, December 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

McCahey:2022:BR

[McC22] Daniella McCahey. Book review. Physics in Perspective (PIP), 24(1):93–96, March 2022. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL https://link.springer.

com/article/10.1007/s00016-022-00286-8.

McMullin:2002:OFC

[McM02] Ernan McMullin. The origins of the field concept in physics. Physics in Perspective (PIP), 4(1):13–39, February 2002. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
002-8357-5.

McMillan:2005:RJR

[McM05] Priscilla Johnson McMillan. The ruin of J. Robert Oppenheimer, and the birth of the modern arms race. Viking, New York, NY, USA, 2005. ISBN 0-670-03422-3. viii + 373 + 16 pp. LCCN QC16.O62 M36 2005.

Mukherjee:2019:SSS

[MDK19] B. N. Mukherjee, S. K. Das, and G. Kellermann. Study of some serum group systems in the Mahishyas and the Muslims in 24-Parganas district, West Bengal. Physics in Perspective (PIP), 21(1):27–32, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/BF00278561.pdf.

Melia:2009:CEC

[Mel09] Fulvio Melia. Cracking the Einstein code: Relativity and the birth of black hole physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2009. ISBN 0-226-51951-1. xi + 137 pp. LCCN QC173.6 .M434 2009. With an afterword by Roy Kerr.

Merzbacher:2001:BRR

[Mer01] Eugen Merzbacher. Book review: Roland Omnès, Quantum Philosophy: Understanding and Interpreting Contemporary Science, translated from the French by Arturo Sangalli, Princeton: Princeton University Press, 1999, xxiii + 296 pages. $29.95 (cloth). Physics in Perspective (PIP), 3(2):250–251, June 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Munoz-Gargante:2021:BRA

[MG21] Núria Muñoz-Garganté. Book review: Andrew Zangwill, A Mind Over Matter: Philip Anderson and the Physics of the Very Many, Oxford University Press, 2021, 412 Pages, $25.00 (Hardcover). Physics in Perspective (PIP), 23(2–3):170–175, October 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00016-021-00276-2.pdf.

Mickens:2006:CRT

[Mic06] Ronald E. Mickens. Converging realities: Toward a common philosophy of physics and mathematics. Physics in Perspective (PIP), 8(4):486–488, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Miller:1999:EFS

[Mil99] Arthur I. Miller. Einstein’s first steps toward General Relativity: Gedanken experiments and axiomatics. Physics in Perspective (PIP), 1(1):85–104, March 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050007;   http:/
/www.springerlink.com/content/qx2xj7m80nx0ph79/fulltext.
pdf.

Miller:2001:EPS

[Mil01] Arthur I. Miller. Einstein, Picasso: space, time, and beauty that causes havoc. Basic Books, New York, NY, USA, 2001. ISBN 0-465-01859-9 (paperback). x + 357 pp. LCCN N72.S3

M55 2001. URL http://www.loc.gov/catdir/enhancements/
fy0831/00065130-d.html.

Miller:2005:ESO

[Mil05] Arthur I. Miller. Empire of the stars: obsession, friendship, and betrayal in the quest for black holes. Houghton-Mifflin, Boston, MA, USA, 2005. ISBN 0-618-34151-X. xx + 364 pp.

LCCN QB35 .M55 2005. URL http://www.loc.gov/catdir/
enhancements/fy0618/2004060909-d.html;   http://www.
loc.gov/catdir/enhancements/fy0618/2004060909-s.html;
 http://www.loc.gov/catdir/enhancements/fy0618/2004060909-
t.html.

Milton:2007:AJS

[Mil07] Kimball A. Milton. In appreciation: Julian Schwinger: From nuclear physics and quantum electrodynamics to source theory and beyond. Physics in Perspective (PIP), 9(1):70–114, January 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-007-0326-6.

Miller:2009:DCN

[Mil09] Arthur I. Miller. Deciphering the Cosmic Number: the Strange Friendship of Wolfgang Pauli and Carl Jung. W. W. Norton & Co., New York, NY, USA, 2009. ISBN 0-393-06532-4 (hardcover). xxv + 336 + 8 pp. LCCN QC16.P37 M55 2009.

Milton:2010:BRS

[Mil10] Kimball A. Milton. Book review: Steven Weinberg, Lake Views. Cambridge, Mass.: Harvard University Press, 2009, x + 259 pages. $26.95 (cloth). Physics in Perspective (PIP), 12(4): 503–505, December 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Mehra:2000:CMS

[MM00] Jagdish Mehra and Kimball A. Milton. Climbing the mountain: the scientific biography of Julian Schwinger. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850658-9 (hardcover), 0-19-852745-4 (paperback). xii + 677 pp. LCCN QC16.S29 M45

2000. URL http://www.loc.gov/catdir/enhancements/
fy0606/99087323-d.html;   http://www.loc.gov/catdir/
enhancements/fy0606/99087323-t.html.

McGuire:2023:FRR

[MM23] Coreen McGuire and Joseph D. Martin. Film review: Revenge of the Nerds. Physics in Perspective (PIP), 25(3):??, November 2023. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL https://link.springer.com/article/10.
1007/s00016-023-00304-3.

Martin:2022:SP

[MNS22] Joseph D. Martin, Robert L. Naylor, and Richard Staley. Small physics. Physics in Perspective (PIP), 24(4):179–180, December 2022. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL https://link.springer.com/article/10.
1007/s00016-022-00295-7.

Martin:2023:SS

[MNS23] Joseph D. Martin, Robert L. Naylor, and Richard Staley. Secret science. Physics in Perspective (PIP), 25(3):??, November 2023. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL https://link.springer.com/article/10.
1007/s00016-023-00303-4.

Mody:2018:BRL

[Mod18] Cyrus C. M. Mody. Book reviews: Lillian Hoddeson and Peter Garrett, The Man Who Saw Tomorrow: The Life and Inventions of Stanford R. Ovshinsky, Cambridge, MA: MIT Press, 2018, 400 pages, $29.95 (hardcover). Physics in Perspective (PIP), 20(3):307–310, September 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Montgomery:2000:STM

[Mon00] Scott L. Montgomery. Science in Translation: Movements of Knowledge Through Cultures and Time. University of Chicago Press, Chicago, IL, USA and London, UK, 2000. ISBN 0-226-53480-4 (hardcover). xi + 325 pp. LCCN Q124 .M66

2000. URL http://www.h-net.org/review/hrev-a0c0g0-
aa;   http://www.loc.gov/catdir/description/uchi051/
99053389.html;   http://www.loc.gov/catdir/toc/uchi052/
99053389.html.

Moore:2008:DSS

[Moo08] Kelly Moore. Disrupting Science: Social Movements, American Scientists, and the Politics of the Military, 1945–1975. Princeton studies in cultural sociology. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-11352-1 (hardcover). x + 311 pp. LCCN Q127.U6. alk. paper. URL http://

www.gbv.de/dms/bowker/toc/9780691113524.pdf;   http:
//www.loc.gov/catdir/toc/ecip0717/2007019974.html.

Morrison:1990:PML

[Mor90] Philip Morrison. Philip Morrison’s long look at the literature: his reviews of a hundred memorable science books. W. H. Freeman, New York, NY, USA, 1990. ISBN 0-7167-2107-4. xi + 351 pp. LCCN Q158.5 .M67 1990. US$24.95.

Morris:2003:LSP

[Mor03] Richard Morris. The last sorcerers: the path from alchemy to the periodic table. Joseph Henry Press, Washington, DC, USA, 2003. ISBN 0-309-08905-0 (hardcover), 0-309-50593-3 (PDF). xi + 282 pp. LCCN QD11 .M86 2003. URL http://www.loc.

gov/catdir/toc/ecip046/2003014790.html.

Morus:2005:WPB

[Mor05] Iwan Rhys Morus. When physics became king. University of Chicago Press, Chicago, IL, USA and London, UK, 2005. ISBN 0-226-54201-7, 0-226-54202-5 (paperback). xii + 303 pp. LCCN QC9.E89 M67 2005. URL http://www.loc.gov/

catdir/description/uchi051/2004015207.html;   http:/
/www.loc.gov/catdir/enhancements/fy0616/2004015207-b.
html;   http://www.loc.gov/catdir/toc/ecip0419/2004015207.
html.

Moura:2020:PRE

[Mou20] Breno Arsioli Moura. The problem of reflection in eighteenthcentury projectile theories of light. Physics in Perspective (PIP), 22(4):191–214, December 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Mauff:2019:VUP

[MPFP19] G. Mauff, B. G. Potrafki, H. Freis, and G. Pulverer. Vergleichende Untersuchungen zum Polymorphismus des Posttransferrins (Pt) und der dritten Komponente des Humankomplements (C3). (German) [Comparative studies on the polymorphism of the Posttransferrin (Pt) and the third component of human complexity (C3)]. Physics in Perspective (PIP), 21(1):75–80, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/content/pdf/
10.1007/BF00278569.pdf.

Mulligan:1999:HHP

[Mul99] Joseph F. Mulligan. Heinrich Hertz and Philipp Lenard: two distinguished physicists, two disparate men. Physics in Perspective (PIP), 1(4):345–366, December 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050027.

Mulligan:2001:AHH

[Mul01] Joseph F. Mulligan. The aether and Heinrich Hertz’s The Principles of Mechanics presented in a new form. Physics in Perspective (PIP), 3(2):136–164, June 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/PL00000527.

Muller:2010:PTF

[Mul10] R. (Richard) Muller. Physics and technology for future presidents: an introduction to the essential physics every world leader needs to know. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-13504-5 (hardcover). xi + 517 pp. LCCN QC24.5 .M85 2010.

Marage:1999:SCB

[MW99] Pierre Marage and Grégoire Wallenborn, editors. The Solvay councils and the birth of modern physics, volume 22 of Science networks, historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 3-7643-5705-3 (Basel), 0-8176-5705-3 (Boston). xiii + 224 pp. LCCN QC7.

Matricon:2003:CWH

[MW03] Jean Matricon and Georges Waysand. The cold wars: a history of superconductivity. Rutgers University Press, New Brunswick, NJ, USA, 2003. ISBN 0-8135-3294-9, 0-8135-3295-7. xiii + 271 pp. LCCN ????

Nall:2020:BRD

[Nal20] Joshua Nall. Book review Daniel Kennefick, No Shadow of a Doubt: The 1919 Eclipse That Confirmed Einstein s Theory of Relativity, Princeton, NJ: Princeton University Press, 2019, 416 pages, $29.95 (hardcover). Physics in Perspective (PIP), 22(4): 239–242, December 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Nauenberg:2005:RHS

[Nau05] Michael Nauenberg. Robert Hooke’s seminal contribution to orbital dynamics. Physics in Perspective (PIP), 7(1):4–34, March 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-004-0226-y.

Navarro:2006:EAD

[Nav06] Jaume Navarro. Early attempts to detect the neutrino at the Cavendish Laboratory. Physics in Perspective (PIP), 8 (1):64–82, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-005-0249-z.

Nussbaumer:2009:DEU

[NB09] Harry Nussbaumer and Lydia Bieri. Discovering the Expanding Universe. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-51484-3 (hardcover). xvii + 226 pp. LCCN

QB981 .N865 2009. URL http://www.gbv.de/dms/bowker/
toc/9780521514842.pdf.

Neufeld:1999:BRP

[Neu99] Michael J. Neufeld. Book review: Paul Lawrence Rose, Heisenberg and the Nazi Atomic Bomb Project: A Study in German Culture. Berkeley/Los Angeles/London: University of California Press, 1998, xx, 352 pages. $35.00 (cloth). Physics in Perspective (PIP), 1(2):227–228, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Newton:1999:BRG

[New99] Roger G. Newton. Book review: Gary William Flake, The Computational Beauty of Nature: Computer Explorations of Fractals, Chaos, Complex Systems, and Adaptation. Cambridge, Mass.: The MIT Press, 1998, xviii + 493 pp. $45.00 (cloth). Physics in Perspective (PIP), 1(1), March 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Newton:2004:BRP

[New04] Roger G. Newton. Book review: Peter Pesic, Abel’s Proof: An Essay on the Sources and Meaning of Mathematical Unsolvability. Cambridge, Mass: The MIT Press, 2003, viii + 213 pages. $24.95 (cloth). Physics in Perspective (PIP), 6(4):482–484, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Newton:2007:CCH

[New07] Roger G. Newton. From clockwork to crapshoot: a history of physics. Belknap Press of Harvard University Press, Cambridge, MA, USA; London, UK, 2007. ISBN 0-674-02337-4 (paperback). viii + 340 pp. LCCN QC7 .N398 2007.

Niu:2023:CMD

[NGZ23] Weixing Niu, Yejing Ge, and Haohao Zhu. Calculating and mapping the disposition of future heavens: Ferdinand Verbiest’s weather forecasting and its termination during the Qing Court (1669–1680). Physics in Perspective (PIP), 25(3):??, November 2023. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL https://link.springer.com/article/10.
1007/s00016-023-00301-6.

Nimmo:2014:PDC

[NHKL14] John R. Nimmo, Paula M. Hermann, M. B. Kirkham, and Edward R. Landa. Pollen dispersal by catapult: Experiments of Lyman J. Briggs on the flower of mountain laurel. Physics in Perspective (PIP), 16(3):371–389, September 2014. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
014-0141-9.

Nimtz:2008:ZTS

[NHW08] Günter Nimtz, Astrid Haibel, and Ulrich Vorr Walter. Zero time space: how quantum tunneling broke the light speed barrier. Wiley-VCH, Berlin, Germany; Weinheim, Germany; New York, NY, USA, 2008. ISBN 3-527-40735-9 (paperback). xvi + 150 pp. LCCN QC176.8.T8. ca. EUR 24.90ca. EUR 24.90.

URL http://www.gbv.de/dms/hbz/toc/ht015424128.pdf;
http://www.loc.gov/catdir/enhancements/fy0838/2008447837-
b.html;   http://www.loc.gov/catdir/enhancements/fy0838/
2008447837-d.html;   http://www.loc.gov/catdir/enhancements/
fy0838/2008447837-t.html;   http://www.zentralblatt-
math.org/zmath/en/search/?an=1152.81002.

Naylor:2022:PPU

[NMS22] Robert L. Naylor, Joseph D. Martin, and Richard Staley. Physics and philosophy — uneasy bedfellows? Physics in Perspective (PIP), 24(2–3):97–98, October 2022. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL https://

link.springer.com/article/10.1007/s00016-022-00292-w.

Naylor:2023:VP

[NMS23] Robert L. Naylor, Joseph D. Martin, and Richard Staley. Vital physics. Physics in Perspective (PIP), 25(1–2):1–2, June 2023. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL https://link.springer.com/article/10.
1007/s00016-023-00297-z.

Noer:2009:BR

[NNT^{+}09] Richard Noer, Paul G. Nyce, Virginia Trimble, Thomas B. Greenslade Jr., and Charles H. Holbrow. Book reviews. Physics in Perspective (PIP), 11(1):106–116, March 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
008-0388-0.

Noer:2009:BRM

[Noe09] Richard Noer. Book review: Michael J. Crowe, Mechanics from Aristotle to Einstein. Santa Fe: Green Lion Press, 2007, xxii + 332 pages. $17.95 (paper). Physics in Perspective (PIP), 11 (1):106–107, March 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Noer:2011:BRJ

[Noe11] Richard Noer. Book review: Jennifer Coopersmith, Energy, the Subtle Concept: The Discovery of Feynman’s Blocks from Leibniz to Einstein. Oxford: Oxford University Press, 2010, xiv + 400 pages. $55.00 (cloth). Physics in Perspective (PIP), 13(3): 379–380, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Norton:2016:HED

[Nor16] John D. Norton. How Einstein did not discover. Physics in Perspective (PIP), 18(3):249–282, September 2016. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
016-0186-z.

Neuenschwander:2008:APP

[NW08] Dwight E. Neuenschwander and Sallie A. Watkins. In appreciation: Professional and personal coherence: The life and work of Melba Newell Phillips. Physics in Perspective (PIP), 10(3): 295–364, September 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-007-0373-z.

Nyce:2009:BRM

[Nyc09] Paul G. Nyce. Book review: Matthew Stanley, Practical Mystic: Religion, Science, and A. S. Eddington. Chicago: The University of Chicago Press, 2007, x + 313 pages. $37.50 (cloth). Physics in Perspective (PIP), 11(1):107–109, March 2009. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Nye:1999:PCP

[Nye99] M. J. Nye. A physicist in the corridors of power: P. M. S. Blackett’s opposition to atomic weapons following the War. Physics in Perspective (PIP), 1(2):136–156, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s000160050013.

Nye:2008:CHS

[Nye08] Mary Jo Nye, editor. The Cambridge history of science. Vol. 5, The modern physical and mathematical sciences. Cambridge histories online. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-511-46834-2 (e-book). xxix + 650 pp. LCCN Q125.

Nye:2019:STM

[Nye19] Mary Jo Nye. Shifting trends in modern physics, Nobel recognition, and the histories that we write. Physics in Perspective (PIP), 21(1):3–22, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

OConnor:2010:SWJ

[O’C10] Thomas O’Connor. The scientific work of John A. McClelland: a recently discovered manuscript. Physics in Perspective (PIP), 12(3):266–306, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-010-0023-8.

OConnor:2014:DDP

[O’C14] Thomas C. O’Connor. Daedalus in Dublin: A physicist’s labyrinth. Physics in Perspective (PIP), 16(1):98–128, March 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-014-0131-y.

Ohanian:2008:BRJ

[Oha08a] Hans C. Ohanian. Book review: Jean Eisenstaedt, The Curious History of Relativity: How Einstein’s Theory was Lost and Found Again. Translated from the French by Arturo Sangalli. Princeton: Princeton University Press, 2006, ix + 363 pages. $29.95 (cloth). Physics in Perspective (PIP), 10(1):126–127, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Ohanian:2008:BRS

[Oha08b] Hans C. Ohanian. Book review: Sander Bais, Very Special Relativity: An Illustrated Guide. Cambridge, Mass.: Harvard University Press, 2007, 120 pages. $20.95 (cloth). Physics in Perspective (PIP), 10(3):375–376, September 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Ohanian:2010:BRA

[Oha10] Hans C. Ohanian. Book review: Alberto A. Martinez, Kinematics: The Lost Origins of Einstein’s Relativity. Baltimore: The Johns Hopkins University Press, 2009, xix + 464 pages. $65.00 (cloth). Physics in Perspective (PIP), 12(2):236–238, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

ORaifeartaigh:2018:ILE

[OM18] Cormac O’Raifeartaigh and Simon Mitton. Interrogating the legend of Einstein’s ”biggest blunder”. Physics in Perspective (PIP), 20(4):318–341, December 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Omnes:1999:QPU

[Omn99a] Roland Omnès. Quantum Philosophy: Understanding and Interpreting Contemporary Science. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-02787-0 (hardcover). xxiii + 296 pp. LCCN QC6 .O55 1999.

URL http://www.loc.gov/catdir/description/prin021/
98042445.html;   http://www.loc.gov/catdir/toc/prin032/

98042445.html. Translated by Arturo Sangalli.

Omnes:1999:UQM

[Omn99b] Roland Omnès. Understanding Quantum Mechanics. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-00435-8 (hardcover). xiii + 307 pp. LCCN QC174.12 .O465

1999. URL http://catdir.loc.gov/catdir/description/
prin021/98042442.html;   http://catdir.loc.gov/catdir/
toc/prin032/98042442.html.

ORaifeartaigh:2020:RVP

[OO20] Cormac O’Raifeartaigh and Michael O’Keeffe. Redshifts versus paradigm shifts: Against renaming Hubble’s Law. Physics in Perspective (PIP), 22(4):215–225, December 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Orrman-Rossiter:2021:OAD

[OR21] Kevin Orrman-Rossiter. Observation and annihilation: The discovery of the antiproton. Physics in Perspective (PIP), 23(1): 3–24, March 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Ortega-Rodriguez:2017:ESC

[ORSSBO^{+}17] M. Ortega-Rodríguez, H. Solís-Sánchez, E. Boza-Oviedo, K. Chaves-Cruz, M. Guevara-Bertsch, M. Quirós-Rojas, S. Vargas-Hernández, and A. Venegas-Li. The early scientific contributions of J. Robert Oppenheimer: Why did the scientific community miss the black hole opportunity? Physics in Perspective (PIP), 19(1):60–75, April 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-017-0195-6.

Osler:2010:RWN

[Osl10] Margaret J. Osler. Reconfiguring the world: nature, God, and human understanding from the Middle Ages to early modern Europe. The Johns Hopkins University Press, Baltimore, MD, USA, 2010. ISBN 0-8018-9655-X (hardcover), 0-8018-9656-8 (paperback). x + 184 pp. LCCN Q124.97 .O85 2010.

Olivotto:2010:GM

[OT10] Cristina Olivotto and Antonella Testa. Galileo and the movies. Physics in Perspective (PIP), 12(4):372–395, December 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-010-0027-4.

Ohanian:2008:BR

[OTRF08] Hans C. Ohanian, Virginia Trimble, Thomas D. Rossing, and Michael W. Friedlander. Book reviews. Physics in Perspective (PIP), 10(1):126–134, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-006-0355-6.

Olano:2022:WSH

[OV22] Juan A. Queijo Olano and Antonio A. P. Videira. Walter Scott Hill and Uruguayan physics. Physics in Perspective (PIP), 24 (1):35–71, March 2022. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL https://link.springer.

com/article/10.1007/s00016-021-00284-2.

Overbye:2000:ELS

[Ove00] Dennis Overbye. Einstein in love: a scientific romance. Viking, New York, NY, USA, 2000. ISBN 0-670-89430-3. xv + 416 pp. LCCN QC16.E5 O9 2000.

Overbye:2001:ELS

[Ove01] Dennis Overbye. Einstein in love: a scientific romance. Penguin, London, UK and New York, NY, USA, 2001. ISBN 0-670-89430-3 (hardcover), 0-14-100221-2 (paperback). xv + 416 pp. LCCN QC16.E5 O9 2001. US$15.00.

Oppenheimer:2013:PASa

[OW13a] Frank Oppenheimer and Charles Weiner. A physicist for all seasons: Part I. Physics in Perspective (PIP), 15(1):33–91, March 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-009-0009-6.

Oppenheimer:2013:PASb

[OW13b] Frank Oppenheimer and Charles Weiner. A physicist for all seasons: Part II. Physics in Perspective (PIP), 15(2):178–240, June 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-009-0010-0.

Pors:2001:PTH

[PA01] Felicity Pors and Finn Aaserud. The physical tourist: Historical sites of physical science in Copenhagen. Physics in Perspective (PIP), 3(2):230–248, June 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/PL00000531.

Panoutsopoulos:2021:CBA

[PA21] Grigoris Panoutsopoulos and Theodore Arabatzis. CERN’s balancing act between unity and disunity: The ”Sister Experiments” UA1 and UA2 and CERN’s first Nobel Prize. Physics in Perspective (PIP), 23(4):181–201, December 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pais:1999:MRS

[Pai99] Abraham Pais. In memoriam: Robert Serber (1909–1997). Physics in Perspective (PIP), 1(1):105–109, March 1999. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160050008.

Palevsky:2000:AFD

[Pal00] Mary Palevsky. Atomic Fragments: a Daughter’s Questions. University of California Press, Berkeley, CA, USA, 2000. ISBN 0-520-22055-2 (hardcover), 0-250-22055-2 (hardcover). xiv + 289 pp. LCCN QC773.3.U5 P35 2000. URL http://catdir.

loc.gov/catdir/bios/ucal052/99087422.html;   http://
catdir.loc.gov/catdir/description/ucal042/99087422.html.

Pancaldi:2003:VSC

[Pan03] Giuliano Pancaldi. Volta: science and culture in the age of Enlightenment. Princeton University Press, Princeton, NJ, USA, 2003. ISBN 0-691-09685-6. xv + 381 pp. LCCN QC515.V8 P36 2003.

Park:2000:BRM

[Par00] David Park. Book review: Mara Beller, Quantum Dialogue: The Making of a Revolution. Chicago: University of Chicago Press, 1999, xv + 365 pages. $35.00 (cloth). Physics in Perspective (PIP), 2(3):327–330, September 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Park:2006:BRG

[Par06] David Park. Book review: Gerald Holton, Victory and Vexation in Science: Einstein, Bohr, Heisenberg, and Others. Cambridge, Mass. and London: Harvard University Press, 2005. xiv + 229 pages. $35.00 (cloth). Physics in Perspective (PIP), 8(4): 481–483, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Park:2007:GCW

[Par07] David Park. The grand contraption: the world as myth, number, and chance. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-13053-1 (paperback). xiv + 331 pp. LCCN Q125 .P323 2007.

Paraoanu:2011:PCI

[Par11] G. S. Paraoanu. Perspectives on current issues: Quantum computing: Theoretical versus practical possibility. Physics in Perspective (PIP), 13(3):359–372, September 2011. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
011-0057-6.

Pasachoff:2004:BRR

[Pas04] Jay M. Pasachoff. Book review: Robert P. Kirshner, The Extravagant Universe: Exploding Stars, Dark Energy, and the Accelerating Cosmos. Princeton: Princeton University Press, 2002, xii + 282 pages. $29.95 (cloth). Physics in Perspective (PIP), 6(2):241–243, June 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pasachoff:2010:BRHb

[Pas10] Naomi Pasachoff. Book review: Harry Nussbaumer and Lydia Bieri, Discovering the Expanding Universe. Cambridge: Cambridge University Press, 2009, 226 pages. $59.00 (cloth). Physics in Perspective (PIP), 12(3):353–355, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pasachoff:2012:BRR

[Pas12a] Naomi Pasachoff. Book review: Ray Jayawardhana, Strange New Worlds: The Search for Alien Planets and Life Beyond Our Solar System. Princeton: Princeton University Press, 2011, 255 pages. $24.95 (cloth). Physics in Perspective (PIP), 14(1):117– 119, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pasachoff:2012:SHP

[Pas12b] Naomi Pasachoff. A short history of physics in the American Century. Physics in Perspective (PIP), 14(3):384–387, September 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pavlish:2010:RVP

[Pav10] Ursula Pavlish. Robert Vivian Pound and the discovery of nuclear magnetic resonance in condensed matter. Physics in Perspective (PIP), 12(2):180–189, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-009-0011-z.

Pavlish:2011:GGL

[Pav11] Ursula Pavlish. Gerson Goldhaber: a life in science. Physics in Perspective (PIP), 13(2):189–214, June 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
010-0040-7.

Pasachoff:2002:BRR

[PCGD02] Jay M. Pasachoff, James T. Cushing, David Goodstein, and Michael A. Day. Book reviews: Ralph A. Alpher and Robert Herman, The Genesis of the Big Bang. New York: Oxford University Press, 2001, viii + 224 pages. $29.95 (hardbound). Shimon Malin, Nature Loves to Hide: Quantum Physics and Reality, a Western Perspective. New York: Oxford University Press, 2001, xv + 288 pages. $27.50 (hardbound). Arthur I. Miller, Einstein, Picasso. Space, Time, and the Beauty That Causes Havoc. New York: Basic Books, 2001, x + 357 pages. $30.00 (cloth). Jay A. Labinger and Harry Collins, ed., The One Culture? A Conversation about Science. Chicago University of Chicago Press, 2001, 321 pages. $18.00 (paper), $65.00 (cloth). Physics in Perspective (PIP), 4(2):243–250, May 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http:/

/link.springer.com/article/10.1007/s00016-002-8367-3.

Pedersen:2011:PSO

[Ped11] Bjørn Pedersen. Physical science in Oslo. Physics in Perspective (PIP), 13(2):215–238, June 2011. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s00016-
011-0055-8.pdf.

Perkovich:1999:INB

[Per99] George Perkovich. India’s nuclear bomb: the impact on global proliferation. University of California Press, Berkeley, CA, USA, 1999. ISBN 0-520-21772-1. xii + 597 pp. LCCN UA840 .P47

Perl:2001:BRG

[Per01] Martin L. Perl. Book review: George Johnson, Strange Beauty: Murray Gell-Mann and the Revolution in Twentieth Century Physics. New York: Knopf, 1999, x + 434 pages. $30.00 (Hardbound). Physics in Perspective (PIP), 3(1):130–131, March 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Perl:2004:DTL

[Per04] Martin L. Perl. The discovery of the tau lepton and the changes in elementary-particle physics in forty years. Physics in Perspective (PIP), 6(4):401–427, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-003-0218-3.

Perkowitz:2010:BRT

[Per10] Sidney Perkowitz. Book review: Theresa Levitt, The Shadow of Enlightenment: Optical and Political Transparency in France 1789–1848. New York: Oxford University Press, 2009, 304 pages. $70.00 (cloth). Physics in Perspective (PIP), 12(2):234– 236, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pesic:2003:APE

[Pes03] Peter Pesic. Abel’s proof: an essay on the sources and meaning of mathematical unsolvability. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-16216-4. viii + 213 pp. LCCN

QA212 .P47 2003. URL http://www.loc.gov/catdir/toc/
fy037/2002031991.html.

Pesic:2009:HWM

[Pes09] Peter Pesic, editor. Hermann Weyl: Mind and nature: selected writings on philosophy, mathematics, and physics. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-13545-2 (hardcover), 1-4008-3332-9. 261 pp. LCCN QA8.4 .W49 2009.

Pesic:2011:BRN

[Pes11] Peter Pesic. Book review: Nick Huggett, Everywhere and Everywhen: Adventures in Physics and Philosophy. New York: Oxford University Press, 2010, 217 pages. $24.95 (cloth). Physics in Perspective (PIP), 13(3):376–378, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pesic:2012:IPQ

[Pes12] Peter Pesic. The infinity puzzle: Quantum field theory and the hunt for an orderly universe. Physics in Perspective (PIP), 14 (3):373–375, September 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pesic:2013:HRS

[Pes13] Peter Pesic. Helmholtz, Riemann, and the sirens: Sound, color, and the “problem of space”. Physics in Perspective (PIP), 15(3): 256–294, September 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-013-0109-1.

Pesic:2015:BRS

[Pes15] Peter Pesic. Book review: Steven Weinberg, To Explain the World: The Discovery of Modern Science. HarperCollins, 2015, xiv + 416 pages, $28.99. Physics in Perspective (PIP), 17 (2):156–160, June 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-015-0162-z.

Peterson:2011:GMR

[Pet11] Mark A. Peterson. Galileo’s muse: Renaissance mathematics and the arts. Harvard University Press, Cambridge, MA, USA, 2011. ISBN 0-674-05972-7. vi + 336 pp. LCCN QB36.G2 P48 2011.

Pasachoff:2010:BR

[PFC^{+}10] Naomi Pasachoff, Allan Franklin, David C. Cassidy, Hans Christian von Baeyer, Catherine Westfall, and Gino Segrè. Book reviews. Physics in Perspective (PIP), 12(3):353–367, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-010-0030-9.

Pessoa:2008:TCF

[PFD08] Osvaldo Pessoa, Jr., Olival Freire, Jr., and Alexis De Greiff. The Tausk controversy on the foundations of quantum mechanics: Physics, philosophy, and politics. Physics in Perspective (PIP), 10(2):138–162, June 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-007-0347-1.

Pasachoff:2004:BR

[PFW04] Jay M. Pasachoff, A. P. French, and Robert Weinstock. Book reviews. Physics in Perspective (PIP), 6(2):241–248, June 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-003-0190-y.

Phillips:2013:BRP

[Phi13] Philip Phillips. Book review: Philip W. Anderson, More and Different Notes from a Thoughtful Curmudgeon. Singapore: World Scientific, 2011, ix + 412 pages. $38.00 (paper). Physics in Perspective (PIP), 15(1):118–121, March 2013. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pippard:1999:WMC

[Pip99a] B. Pippard. The Whipple Museum and Cavendish Laboratory, Cambridge. Physics in Perspective (PIP), 1(2):219–223, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s000160050017.

Pippard:1999:BRP

[Pip99b] Sir Brian Pippard. Book review: Peter Day, ed., The Philosopher’s Tree: Michael Faraday’s Life and work in his own words. London: Institute of Physics, 1999. xv + 211 pages. $75.00 (cloth), $29.00 (paper). Physics in Perspective (PIP), 1(3):338, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pippard:2001:DEL

[Pip01] Sir Brian Pippard. Dispersion in the ether: Light over the water. Physics in Perspective (PIP), 3(3):258–270, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
PL00000533.

Pippard:2002:VEH

[Pip02] Sir Brian Pippard. Vignette — Elisabeth Hertz (née Doll) 1864– 1941, widow of Heinrich Rudolf Hertz 1857–1894, Professor of Physics, University of Bonn. Physics in Perspective (PIP), 4 (2):241–242, May 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-002-8366-4.

Priscu:2019:TAA

[PMAS19] R. Priscu, I. Moraru, A. Albu, and S. Sichitiu. Les taux des amino-acides libres dans le liquide cérébro-spinal chez les nourrissons trisomiques 21. Physics in Perspective (PIP), 21 (1):63–68, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/BF00278567.pdf.

Porter:2003:CHS

[Por09] Roy Porter, editor. The Cambridge history of science. Cambridge University Press, Cambridge, UK, 2003–2009. ISBN 0-521-57244-4 (vol. 3: hardcover), 0-521-57243-6 (vol. 4), 0-521-57199-5 (vol. 5), 0-521-59442-1 (vol. 7). ???? pp. LCCN

Q125 .C32 2003. URL http://www.h-net.org/review/hrev-
a0d3l3-aa;   http://www.h-net.org/review/hrev-a0d7k4-
aa;   http://www.h-net.org/review/hrev-a0f4e9-aa;   http:
//www.loc.gov/catdir/description/cam021/2001025311.html;
 http://www.loc.gov/catdir/enhancements/fy0731/2001025311-
b.html;   http://www.loc.gov/catdir/samples/cam033/2001025311.
html;   http://www.loc.gov/catdir/toc/cam026/2001025311.
html.

Perkowitz:2010:BR

[PORL10] Sidney Perkowitz, Hans C. Ohanian, Thomas Ryckman, and Arlo U. Landolt. Book reviews. Physics in Perspective (PIP), 12(2):234–243, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-010-0024-7.

Poundstone:1999:CSL

[Pou99] William Poundstone. Carl Sagan: a life in the cosmos. Henry Holt, New York, NY, USA, 1999. ISBN 0-8050-5766-8. xvii + 473 pp. LCCN QB36.S15 P68

1999. URL http://www.loc.gov/catdir/enhancements/
fy0665/99014615-b.html;   http://www.loc.gov/catdir/
enhancements/fy0665/99014615-d.html.

Pound:2000:WP

[Pou00] Robert V. Pound. Weighing photons, I. Physics in Perspective (PIP), 2(3):224–268, September 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050045.

Pound:2001:WPI

[Pou01] Robert V. Pound. Weighing photons, II. Physics in Perspective (PIP), 3(1):4–51, March 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050055.

Pasachoff:2008:BRK

[PP08] Jay M. Pasachoff and Naomi Pasachoff. Book review: Kristine Larsen, Stephen Hawking: A Biography. Amherst, New York: Prometheus Books, 2007, 215 pages. $16.95 (paper). Physics in Perspective (PIP), 10(4):489–491, December 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pasachoff:2010:BRHa

[PP10] Jay M. Pasachoff and Naomi Pasachoff. Book review: Helge Kragh, The Moon that Wasn’t: The Saga of Venus’ Spurious Satellite. Basel: Birkhäuser Verlag, 2008, 199 pages. $59.95 (cloth). Physics in Perspective (PIP), 12(1):105–108, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Price:1999:BRN

[Pri99] Richard H. Price. Book review: Nick Huggett, Space from Zeno to Einstein. Cambridge, Mass.: MIT Press, 1999. xi + 274 pages. $45.00 (cloth), $22.50 (paper). Physics in Perspective (PIP), 1 (4):457–458, December 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pronskikh:2016:FPM

[Pro16] Vitaly S. Pronskikh. E-36: The first proto-megascience experiment at NAL. Physics in Perspective (PIP), 18(4):357–378, December 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-016-0192-1.

Pera:2019:PVF

[PS19] F. Pera and P. Scholz. Polyploidization in vitro: Formation of a predominantly triploid cell population in an originally diploid tissue culture of Microtus agrestis. Physics in Perspective (PIP), 21(1):17–26, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/BF00278560.pdf.

Phillips:2013:BR

[PSTR13] Philip Phillips, Gino Segrè, Mason Tattersall, and Barbara Ryden. Book reviews. Physics in Perspective (PIP), 15(1):118–126, March 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-013-0105-5.

Pullman:1998:AHH

[Pul98] Bernard Pullman. The atom in the history of human thought. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-511447-7 (hardcover). x + 403 pp. LCCN QC171.2 .P8513 1998. URL http://www.loc.gov/

catdir/enhancements/fy0635/97036040-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy0723/97036040-b.html.

Translated by Axel Reisinger from the French original.

Purrington:1997:PNC

[Pur97] Robert D. Purrington. Physics in the Nineteenth Century. Rutgers University Press, New Brunswick, NJ, USA, 1997. ISBN 0-8135-2441-5 (hardcover), 0-8135-2442-3 (paperback). xvii + 249 pp. LCCN QC7 .P84 1997.

Papiha:2019:TVI

[PW19] S. S. Papiha and H. J. Wastell. Transferrin variants in the Indian subcontinent. Physics in Perspective (PIP), 21(1):69–73, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/content/pdf/
10.1007/BF00278568.pdf.

Quinn:2008:MMA

[QN08] Helen R. Quinn and Yossi Nir. The Mystery of the Missing Antimatter. Science essentials. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-13309-3. xii + 278 pp. LCCN

QC173.3 .Q856 2008. URL http://www.loc.gov/catdir/
enhancements/fy0835/2007934402-b.html;   http://www.
loc.gov/catdir/enhancements/fy0835/2007934402-d.html;
 http://www.loc.gov/catdir/enhancements/fy0835/2007934402-
t.html.

Reif-Acherman:2004:HKO

[RA04] Simón Reif-Acherman. Heike Kamerlingh Onnes: Master of experimental technique and quantitative research. Physics in Perspective (PIP), 6(2):197–223, June 2004. CODEN PHPEF2.

ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-003-0193-8.

Reif-Acherman:2010:HVR

[RA10] Simón Reif-Acherman. Henri Victor Regnault: Experimentalist of the science of heat. Physics in Perspective (PIP), 12(4): 396–442, December 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-010-0028-3.

Reif-Acherman:2013:HKO

[RA13] Simón Reif-Acherman. Heike Kamerlingh Onnes and the Nobel Prize in Physics for 1913: The highest honor for the lowest temperatures. Physics in Perspective (PIP), 15(4):415–450, December 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-013-0118-0.

Reif-Acherman:2015:BTT

[RA15] Simón Reif-Acherman. Between thermodynamics and thermometry: The life and scientific achievements of Hugh Longbourne Callendar. Physics in Perspective (PIP), 17(3):198– 235, September 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-015-0166-8.

Ramsey:1999:EHM

[Ram99] N. F. Ramsey. Early history of magnetic resonance. Physics in Perspective (PIP), 1(2):123–135, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s000160050012.

Ramberg:2005:BRR

[Ram05] Peter J. Ramberg. Book review: Richard Morris, The Last Sorcerers: The Path from Alchemy to the Periodic Table. Washington, DC: Joseph Henry Press, 2003, xi + 225, $24.95 (cloth). Physics in Perspective (PIP), 7(1):134, March 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Renn:2005:DLA

[RD05] Jürgen Renn and Peter Damerow, editors. Dokumente eines Lebensweges: zur Ausstellung Albert Einstein. (German) [Documents of a life’s pathway: for the Albert Einstein exhibition].

Wiley-VCH, Berlin, Germany; Weinheim, Germany; New York, NY, USA, 2005. ISBN 3-527-40571-2. 560 pp. LCCN ????

Reed:2009:CDM

[Ree09] B. Cameron Reed. Centrifugation during the Manhattan Project. Physics in Perspective (PIP), 11(4):426–441, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-009-0429-3.

Reed:2011:LTD

[Ree11] B. Cameron Reed. Liquid thermal diffusion during the Manhattan Project. Physics in Perspective (PIP), 13(2):161–188, June 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-010-0039-0.

Reed:2014:FMP

[Ree14] B. Cameron Reed. The feed materials program of the Manhattan Project: A foundational component of the nuclear weapons complex. Physics in Perspective (PIP), 16(4):461–479, December 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-014-0146-4.

Reiter:2001:ASM

[Rei01a] Wolfgang L. Reiter. In appreciation: Stefan Meyer: Pioneer of radioactivity. Physics in Perspective (PIP), 3(1):106–127, March 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s000160050058.

Reiter:2001:PTV

[Rei01b] Wolfgang L. Reiter. The physical tourist: Vienna: a random walk in science. Physics in Perspective (PIP), 3(4):462– 489, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/PL00000541.

Reiter:2007:MLB

[Rei07] Wolfgang L. Reiter. In memoriam: Ludwig Boltzmann: a life of passion. Physics in Perspective (PIP), 9(3):357–374, September 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-007-0339-1.

Reidy:2010:JTV

[Rei10] Michael S. Reidy. John Tyndall’s vertical physics: From rock quarries to icy peaks. Physics in Perspective (PIP), 12(2): 122–145, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-009-0012-y.

Reiter:2019:KPR

[Rei19] Wolfgang L. Reiter. Karl Przibram: Radioactivity, crystals, and colors. Physics in Perspective (PIP), 21(3):163–193, September 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00016-019-00242-z.pdf.

Renn:2005:AEIa

[Ren05a] Jürgen Renn, editor. Albert Einstein: Ingenieur des Universums. Hundert Autoren für Einstein. (German) [Albert Einstein: Chief Engineer of the Universe. One hundred authors for Einstein]. Wiley-VCH, Berlin, Germany; Weinheim, Germany; New York, NY, USA, 2005. ISBN 3-527-40569-0, 3-527-40571-2. 254 pp. LCCN QC16.E5 A6759 2005.

Renn:2005:AEI

[Ren05b] Jürgen Renn, editor. Albert Einstein: Ingenieur des Universums. Hundert Autoren für Einstein. (German) [Albert Einstein: Chief Engineer of the Universe. One hundred authors for Einstein]. Wiley-VCH, Berlin, Germany; Weinheim, Germany; New York, NY, USA, 2005. ISBN 3-527-40569-0, 3-527-40571-2. 254 pp. LCCN QC16.E5 A6759 2005.

Renn:2005:ELW

[Ren05c] Jürgen Renn, editor. Einsteins Leben und Werk im Kontext. (German) [Einstein’s Life and Work in Context]. Wiley-VCH, Berlin, Germany; Weinheim, Germany; New York, NY, USA, 2005. ISBN 3-527-40573-9. 254 pp. LCCN ????

Rentetzi:2008:TMG

[Ren08] Maria Rentetzi. Trafficking materials and gendered experimental practices: radium research in early 20th Century Vienna. Columbia University Press, New York, NY, USA, 2008. ISBN 0-231-13558-0 (hardcover), 0-231-50959-6 (e-book). xxiii + 279 pp. LCCN QC795.34 .R46 2008. URL http://catdir.loc.

gov/catdir/toc/ecip0828/2008040581.html.

Rhees:2002:PTP

[Rhe02] David J. Rhees. The physical tourist: Physics in “Lake Wobegon”: a tour of three Minnesota museums of science and technology. Physics in Perspective (PIP), 4(2):230–240, May 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-002-8365-5.

Rigden:2007:BRV

[Rig07] John S. Rigden. Book review: Valerio Scarani, Quantum Physics: A First Encounter, Interference, Entanglement, and Reality. Oxford: Oxford University Press, 2006, xvii + 125 pages. $34.50 (cloth). Physics in Perspective (PIP), 9(3):379– 381, September 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Rigden:2011:EMP

[Rig11] John S. Rigden. Edward Mills Purcell, August 30, 1912–March 7, 1997. Physics in Perspective (PIP), 13(1):91–103, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-010-0042-5.

Riordan:2000:BRR

[Rio00a] Michael Riordan. Book review: Robert P. Crease, Making Physics: A Biography of Brookhaven National Laboratory, 1946– 1972. Chicago: University of Chicago Press, 1999, xii + 434 pages, $38.00 (cloth), $22.50 (paper). Physics in Perspective (PIP), 2(2):218–219, June 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Riordan:2000:DSS

[Rio00b] Michael Riordan. The demise of the Superconducting Super Collider. Physics in Perspective (PIP), 2(4):411–425, December 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s000160050053.

Rosenblum:2011:QEP

[RK11] Bruce Rosenblum and Fred Kuttner. Quantum enigma: physics encounters consciousness. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2011. ISBN 0-19-975381-4 (paperback). x + 287 pp. LCCN QC174.13 .R67 2011.

Robey:2018:BRD

[Rob18] Sarah Robey. Book reviews: David C. Cassidy, Farm Hall and the German Atomic Project of World War II: A Dramatic History, Cham: Springer, 2017, xiv + 125 pages. $29.99 (hardcover). Physics in Perspective (PIP), 20(3):305–307, September 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Roeder:2010:BRN

[Roe10] John L. Roeder. Book review: Nicolaas A. Rupke, Alexander von Humboldt: A Metabiography. Chicago: University of Chicago Press, 2008, 316 pages. $21.00 (paper). Physics in Perspective (PIP), 12(1):108–110, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Roeder:2012:BRD

[Roe12] John L. Roeder. Book review: David A. Weintraub, How Old Is the Universe? Princeton: Princeton University Press, 2011, 370 pages. $29.95 (cloth). Physics in Perspective (PIP), 14(1):116– 117, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Romer:1999:WCB

[Rom99] A. Romer. The welcoming of Copernicus’s De revolutionibus: The Commentariolus and its reception. Physics in Perspective (PIP), 1(2):157–183, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050014.

Rose:1998:HNA

[Ros98] Paul Lawrence Rose. Heisenberg and the Nazi atomic bomb project: a study in German culture. University of California Press, Berkeley, CA, USA, 1998. ISBN 0-520-21077-8, 0-585-32190-6 (e-book). xx + 352 pp. LCCN

QC16.H35 R67 1998.  URL http://preterhuman.net/
texts/religion.occult.new_age/occult.conspiracy.and.
related/Rose,%20Paul%20Lawrence%20-%20Heisenberg%20and%
20the%20Nazi%20Atomic%20Bomb%20Project.pdf.

Rossing:2001:BRJ

[Ros01] Thomas D. Rossing. Book review: Jeff Hecht, City of Light: The Story of Fiber Optics. New York: Oxford University Press, 1999, xii + 316 pages. $29.95 (Cloth). Physics in Perspective (PIP), 3(4):496–497, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Rossing:2008:BRM

[Ros08] Thomas D. Rossing. Book review: Myles W. Jackson, Harmonious Triads: Physicists, Musicians, and Instrument Makers in Nineteenth-Century Germany. Cambridge, Mass.: The MIT Press, 2006, x + 395 pages. $40.00 (cloth). Physics in Perspective (PIP), 10(1):130–131, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Rosa:2012:MMP

[Ros12] Rodolfo Rosa. The Merli–Missiroli–Pozzi two-slit electroninterference experiment. Physics in Perspective (PIP), 14(2): 178–195, June 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-011-0079-0.pdf.

Rowe:2001:EMH

[Row01] David E. Rowe. Einstein meets Hilbert: At the crossroads of physics and mathematics. Physics in Perspective (PIP), 3(4):379–424, November 2001. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://link.springer.com/article/10.1007/PL00000538;
 http://www.springerlink.com/content/5byc4pv8q4gq5qy1/
fulltext.pdf.

Roeder:2012:BR

[RPF12] John L. Roeder, Naomi Pasachoff, and Allan Franklin. Book reviews. Physics in Perspective (PIP), 14(1):116–122, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-012-0080-2.

Roca-Rosell:2013:PSB

[RRR13] Antoni Roca-Rosell and Xavier Roqué. Physical science in Barcelona. Physics in Perspective (PIP), 15(4):470–498, December 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-013-0122-4.

Rigden:2001:CF

[RS01a] J. S. Rigden and R. H. Stuewer. Celebrate facts. Physics in Perspective (PIP), 3(3):255–257, September 2001. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Rigden:2001:GTM

[RS01b] J. S. Rigden and R. H. Stuewer. Good theories make for good experiments. Physics in Perspective (PIP), 3(1):1–3, March 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Rigden:2001:PNE

[RS01c] J. S. Rigden and R. H. Stuewer. Physics in a new era. Physics in Perspective (PIP), 3(4):377–378, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Rigden:2001:DWB

[RS01d] J. S. Rigden and R. H. Stuewer. “with these dark words begins my tale… ”. Physics in Perspective (PIP), 3(2):133–135, June 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Rigden:2002:LSS

[RS02a] J. S. Rigden and R. H. Stuewer. The lure of a simple story. Physics in Perspective (PIP), 4(2):125–126, May 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Rigden:2002:RCP

[RS02b] J. S. Rigden and R. H. Stuewer. Realism and the contraction of “pure” physics. Physics in Perspective (PIP), 4(1):1–2, February 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Rigden:2002:EPC

[RS02c] John S. Rigden and Roger H. Stuewer. Editorial — physics in crisis? Physics in Perspective (PIP), 4(4):367–369, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s000160200000.

Rigden:2002:EVF

[RS02d] John S. Rigden and Roger H. Stuewer. Editorial — Victor Frederick Weisskopf, September 19, 1908–April 21, 2002. Physics in Perspective (PIP), 4(3):251–252, August 2002. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
002-8368-2.

Rigden:2003:EGN

[RS03a] John S. Rigden and Roger H. Stuewer. Editorial — globalization of nuclear weapons? Physics in Perspective (PIP), 5(2):127–129, May 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-003-0168-9.

Rigden:2003:ESP

[RS03b] John S. Rigden and Roger H. Stuewer. Editorial — stop the physics centrifuge. Physics in Perspective (PIP), 5(1):1–3, April 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s000160300000.

Rigden:2003:ETT

[RS03c] John S. Rigden and Roger H. Stuewer. Editorial: Two twin paradoxes. Physics in Perspective (PIP), 5(3):241–242, September 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-003-0188-5.

Rigden:2003:FRB

[RS03d] John S. Rigden and Roger H. Stuewer. Faith, reason, and Physics in Perspective. Physics in Perspective (PIP), 5(4): 357–359, December 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-003-0196-5.

Rosner:2003:MBS

[RS03e] Robert W. Rosner and Brigitte Strohmaier, editors. Marietta Blau, Sterne der Zertrümmerung: Biographie einer Wegbereiterin der modernen Teilchenphysik. (German) [Marietta Blau, stars of disintegration: biography of a pioneer of particle physics], volume 3 of Beiträge zur Wissenschaftsgeschichte und Wissenschaftsforschung. Böhlau, Wien, Austria, 2003. ISBN 3-205-77088-9. 229 pp. LCCN QC774.B53 M37 2003.

Rigden:2004:ISS

[RS04a] John S. Rigden and Roger H. Stuewer. Ideology and science; science and ideology. Physics in Perspective (PIP), 6(3):249– 251, September 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-004-0213-3.

Rigden:2004:PSY

[RS04b] John S. Rigden and Roger H. Stuewer. Is physics simple? Yes and no. Physics in Perspective (PIP), 6(1):1–3, April 2004. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
004-0207-1.

Rigden:2004:JRO

[RS04c] John S. Rigden and Roger H. Stuewer. J. Robert Oppenheimer — brilliance and charisma. Physics in Perspective (PIP), 6(4): 369–371, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-004-0231-1.

Rigden:2004:SQN

[RS04d] John S. Rigden and Roger H. Stuewer. Science questions nature and nature answers. Physics in Perspective (PIP), 6(2):123–125, June 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-004-0210-6.

Rigden:2005:WYP

[RS05a] John S. Rigden and Roger H. Stuewer. 2005: The World Year of Physics. Physics in Perspective (PIP), 7(1):1–3, March 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-004-0250-y.

Rigden:2005:EDP

[RS05b] John S. Rigden and Roger H. Stuewer. Editorial: Do physicists understand physics? Physics in Perspective (PIP), 7(4): 387–389, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-005-0269-8.

Rigden:2005:FOF

[RS05c] John S. Rigden and Roger H. Stuewer. Faith of our fathers. Physics in Perspective (PIP), 7(3):265–267, September 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-005-0267-x.

Rigden:2005:LWF

[RS05d] John S. Rigden and Roger H. Stuewer. “Lest we forget … ”. Physics in Perspective (PIP), 7(2):135–137, June 2005. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
005-0255-1.

Rigden:2006:BNA

[RS06a] John S. Rigden and Roger H. Stuewer. Book notes [Albert Einstein, The Meaning of Relativity (Princeton University Press, 2005); Jürgen Renn, Albert Einstein: Chief Engineer of the Universe: Einstein’s Life and Work In Context (vol. 1), Documents of a Life’s Pathway (vol. 2), One Hundred Authors For Einstein (vol. 3) (Wiley-VCH 2005)]. Physics in Perspective (PIP), 8(2):222–224, May 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0284-4.

Rigden:2006:BNB

[RS06b] John S. Rigden and Roger H. Stuewer. Book notes [Classic Feynman: All the Adventures of a Curious Character (New York: W. W. Norton, 2006) [combines Surely You’re Joking Mr. Feynman and What Do You Care What Other People Think?]; Tips on Physics: A Problem-Solving Supplement to the Feynman Lectures on Physics (San Francisco: Pearson Addison Wesley, 2006)]. Physics in Perspective (PIP), 8(1):102–103, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-005-0281-z.

Rigden:2006:BNL

[RS06c] John S. Rigden and Roger H. Stuewer. Book notes [Laurie Brown, ed., Feynman’s Thesis: A New Approach To Quantum Theory (World Scientific, 2005); Gordon Fraser, ed., The New Physics of the Twenty-first Century (Cambridge University Press, 2006)]. Physics in Perspective (PIP), 8(3):344– 346, September 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-005-0295-1.

Rigden:2006:BNR

[RS06d] John S. Rigden and Roger H. Stuewer. Book notes [Robyn Arianrhod, Einstein’s Heroes: Imagining the World Through the Language of Mathematics (New York: Oxford University Press, 2006), $16.95 (paper); Priscilla J. McMillan, The Ruin of J. Robert Oppenheimer and the Birth of the Modern Arms Race (New York: Penguin Books, 2005), $16.00 (paper)]. Physics in Perspective (PIP), 8(4):466–469, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-006-0324-0.

Rigden:2006:BPN

[RS06e] John S. Rigden and Roger H. Stuewer. Physics in Perspective: New features to begin the eighth year. Physics in Perspective (PIP), 8(1):1–2, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-005-0280-0.

Rigden:2006:ERB

[RS06f] John S. Rigden and Roger H. Stuewer. Editorial: Remember the basics. Physics in Perspective (PIP), 8(3):233–235, September 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-006-0294-2.

Rigden:2006:EMW

[RS06g] John S. Rigden and Roger H. Stuewer. Editorial: The meek weak. Physics in Perspective (PIP), 8(4):357–359, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-006-0325-z.

Rigden:2006:PPE

[RS06h] John S. Rigden and Roger H. Stuewer. PSSC physics, Eisenhower, and today, 50 years later. Physics in Perspective (PIP), 8(2):113–115, May 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0283-5.

Rigden:2007:BNB

[RS07a] John S. Rigden and Roger H. Stuewer. Book notes [Undermining Science: Suppression and Distortion in the Bush Administration by Seth Shulman (University of California Press, 2006, $24.95); Gino Segrè, Faust in Copenhagen: A Struggle for the Soul of Physics (Viking, 2007, $25.95)]. Physics in Perspective (PIP), 9(4):502–504, December 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-007-0369-8.

Rigden:2007:BND

[RS07b] John S. Rigden and Roger H. Stuewer. Book notes [David Lindley, Uncertainty: Einstein, Heisenberg, Bohr and the Struggle for the Soul of Science (New York: Doubleday, 2007, 257 pages, $32.00); Stuart Clark, The Sun Kings: The Unexpected Tragedy of Richard Carrington and the Tale of How Modern Astronomy Began (Princeton: Princeton University Press, 2007, xii + 211 pages, $24.95)]. Physics in Perspective (PIP), 9(3): 375–377, September 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-007-0360-4.

Rigden:2007:BNM

[RS07c] John S. Rigden and Roger H. Stuewer. Book notes [Mark Haw, Middle World: The Restless Heart of Matter and Life (New York: Macmillian, 2007, $24.95); Committee on Elementary Particle Physics in the 21st Century, Revealing the Hidden Nature of Space and Time: Charting the Course for Elementary Particle Physics]. Physics in Perspective (PIP), 9(2):253–255, June 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-007-0341-7.

Rigden:2007:BNP

[RS07d] John S. Rigden and Roger H. Stuewer. Book notes [Peter Dear, The Intelligibility of Science: How Science Makes Sense of the World, (Chicago: University of Chicago Press, 2006); Bruce Rosenblum and Fred Kuttner, Quantum Enigma: Physics Encounters Consciousness (Oxford: Oxford University Press, 2006)]. Physics in Perspective (PIP), 9(1):115–117, January 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-007-0335-5.

Rigden:2007:Ec

[RS07e] John S. Rigden and Roger H. Stuewer. Editorial. Physics in Perspective (PIP), 9(3):263–265, September 2007. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
007-0359-x.

Rigden:2007:EF

[RS07f] John S. Rigden and Roger H. Stuewer. Editorial: FAPP. Physics in Perspective (PIP), 9(1):1–3, January 2007. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
007-0334-6.

Rigden:2007:EPH

[RS07g] John S. Rigden and Roger H. Stuewer. Editorial: Physicists: Head and heart. Physics in Perspective (PIP), 9(4):387–389, December 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-007-0368-9.

Rigden:2007:EWU

[RS07h] John S. Rigden and Roger H. Stuewer. Editorial: Will the US lead or follow? Physics in Perspective (PIP), 9(2):127–129, June 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-007-0342-6.

Rowe:2007:EPH

[RS07i] David E. Rowe and Robert J. Schulmann, editors. Einstein on politics: his private thoughts and public stands on nationalism, Zionism, war, peace, and the bomb. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-12094-3. xxxiv + 523 pp. LCCN QC16.E5 E5157 2007. URL http://

www.loc.gov/catdir/enhancements/fy0704/2006100303-b.
html;   http://www.loc.gov/catdir/enhancements/fy0704/

2006100303-d.html;   http://www.loc.gov/catdir/toc/ecip076/
2006100303.html.

Rigden:2008:BND

[RS08a] John S. Rigden and Roger H. Stuewer. Book notes [E. D. Hirsch, Jr., Joseph F. Kett, and James S. Trefil, The New Dictionary of Cultural Literacy; James Trefil Why Science? (Teachers College Press and NSTA Press, 2008)]. Physics in Perspective (PIP), 10(2):244–245, June 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-008-0392-4.

Rigden:2008:BNG

[RS08b] John S. Rigden and Roger H. Stuewer. Book notes [Gerald E. Brown and Chang-Hwan Lee, Editors, Hans Bethe and His Physics, (Singapore: World Scientific, 2006)]. Physics in Perspective (PIP), 10(1):123–125, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-007-0379-6.

Rigden:2008:BNJ

[RS08c] John S. Rigden and Roger H. Stuewer. Book notes [J. R. Leibowitz, Hidden Harmony: The Connected Worlds of Physics and Art (Baltimore: The Johns Hopkins University Press, 2008, $24.95)]. Physics in Perspective (PIP), 10(4):482–483, December 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-008-0414-2.

Rigden:2008:BNR

[RS08d] John S. Rigden and Roger H. Stuewer. Book notes [Richard Muller, Physics for Future Presidents (W. W. Norton, 2008, $24.95); George Saliba Islamic Science and the Making of the European Renaissance (MIT Press, 2007, $40.00)]. Physics in Perspective (PIP), 10(3):365–367, September 2008. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
008-0404-4.

Rigden:2008:Ea

[RS08e] John S. Rigden and Roger H. Stuewer. Editorial. Physics in Perspective (PIP), 10(1):1–3, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-007-0378-7.

Rigden:2008:Ec

[RS08f] John S. Rigden and Roger H. Stuewer. Editorial. Physics in Perspective (PIP), 10(3):255–257, September 2008. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
008-0403-5.

Rigden:2008:EPI

[RS08g] John S. Rigden and Roger H. Stuewer. Editorial: Public intellectuals needed. Physics in Perspective (PIP), 10(4):377–378, December 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-008-0413-3.

Rigden:2008:EWE

[RS08h] John S. Rigden and Roger H. Stuewer. Editorial: Will economics end the physics that captivates the public? Physics in Perspective (PIP), 10(2):135–137, June 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-008-0391-5.

Rigden:2009:BNA

[RS09a] John S. Rigden and Roger H. Stuewer. Book notes [Alan Hirshfeld, Eureka Man: The Life and Legacy of Archimedes (New York: Walker and Company, 2009, viii + 228 pages); Francesco Guerra and Nadia Robotti, Ettore Majorana: Aspects of His Scientific and Academic Activity (Pisa: Edizioni Della Normale, 2008, xii + 243 pages)]. Physics in Perspective (PIP), 11(4): 454–456, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-009-0435-5.

Rigden:2009:BND

[RS09b] John S. Rigden and Roger H. Stuewer. Book notes [Dan Hooper, Nature’s Blueprint: Supersymmetry and the Search For a Unified Theory of Matter and Force (HarperCollins 2008, 230 pages, $24.95)]. Physics in Perspective (PIP), 11(3):336–337, September 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-009-0433-7.

Rigden:2009:BNM

[RS09c] John S. Rigden and Roger H. Stuewer. Book notes [Mark Littmann, Fred Espenak, and Ken Willcox, Totality: Eclipses of the Sun (Third Edition, Oxford University Press, 2008, $34.95); Don Lincoln, The Quantum Frontier: The Large Hadron Collider (Johns Hopkins University Press, 2009, $25.00)]. Physics in Perspective (PIP), 11(2):228–230, June 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
009-0431-9.

Rigden:2009:BNS

[RS09d] John S. Rigden and Roger H. Stuewer. Book notes [Shahn Majid, ed., On Space and Time (Cambridge University Press, 2008, xx + 287 pages, $26.00); Sébastíen Balíbar, The Atom and the Apple: Twelve Tales from Contemporary Physics (Princeton University Press, 2008, x + 190 pages, $24.95)]. Physics in Perspective (PIP), 11(1):104–105, March 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-009-0424-8.

Rigden:2009:Eb

[RS09e] John S. Rigden and Roger H. Stuewer. Editorial. Physics in Perspective (PIP), 11(2):117–119, June 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-009-0430-x.

Rigden:2009:Ed

[RS09f] John S. Rigden and Roger H. Stuewer. Editorial. Physics in Perspective (PIP), 11(4):355–356, December 2009. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
009-0434-6.

Rigden:2009:ESH

[RS09g] John S. Rigden and Roger H. Stuewer. Editorial: The SSC and the Hubble: a predictable lesson for physicists. Physics in Perspective (PIP), 11(3):241–243, September 2009. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
009-0432-8.

Rigden:2009:EWI

[RS09h] John S. Rigden and Roger H. Stuewer. Editorial: Will it be a bang or a whimper? Physics in Perspective (PIP), 11(1):1–3, March 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-009-0423-9.

Rigden:2009:PTS

[RS09i] John S. Rigden and Roger H. Stuewer, editors. The Physical Tourist: a Science Guide for the Traveler. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2009. ISBN 3-7643-8932-X. viii + 249 pp. LCCN QC7 .P473 2009.

URL http://www.loc.gov/catdir/enhancements/fy1002/
2008936305-d.html;   http://www.loc.gov/catdir/enhancements/
fy1002/2008936305-t.html.

Rigden:2010:TPW

[RS10a] John S. Rigden and Roger H. Stuewer. Are talks by physicists weakened by their props? Physics in Perspective (PIP), 12 (2):119–121, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-010-0021-x.

Rigden:2010:BNBa

[RS10b] John S. Rigden and Roger H. Stuewer. Book notes [E. Brian Davies, Why Beliefs Matter: Reflections On the Nature of Science, (Oxford University Press, 2010, vii + 250 pages, $45.00)]. Physics in Perspective (PIP), 12(4):467–469, December 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-010-0046-1.

Rigden:2010:BNG

[RS10c] John S. Rigden and Roger H. Stuewer. Book notes [Graham Farmelo, The Strangest Man: The Hidden Life of Paul Dirac, Mystic of the Atom (Basic Books, 2009, 539 pages, $29.95); Vlatko Vedral, Decoding Reality: The Universe as Quantum Information (Oxford University Press, 2010, x + 229 pages, $29.95)]. Physics in Perspective (PIP), 12(3):350–352, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-010-0031-8.

Rigden:2010:BNJ

[RS10d] John S. Rigden and Roger H. Stuewer. Book notes [Jeremy Bernstein, Quantum Leaps (Harvard University Press 2009, vi + 230 pages, $18.95); John Close, Nothing: a very short introduction (Oxford University Press, 2009, 157 pages, $11.95)]. Physics in Perspective (PIP), 12(1):100–101, March 2010. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
010-0014-9.

Rigden:2010:BNR

[RS10e] John S. Rigden and Roger H. Stuewer. Book notes [Richard Holmes, The Age of Wonder: How the Romantic Generation Discovered the Beauty and Terror of Science (Pantheon Books, 2008, xxi + 552 pages, $40); Curt Suplee, The Plasma Universe (Cambridge University Press, 2009, xii + 76 pages, $20.99)]. Physics in Perspective (PIP), 12(2):231–233, June 2010. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
010-0022-9.

Rigden:2010:CHH

[RS10f] John S. Rigden and Roger H. Stuewer. Can the humanities help science? Physics in Perspective (PIP), 12(4):369–371, December 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00016-010-0047-0.pdf.

Rigden:2010:E

[RS10g] John S. Rigden and Roger H. Stuewer. Editorial. Physics in Perspective (PIP), 12(1):1–3, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-010-0015-8.

Rigden:2010:MFP

[RS10h] John S. Rigden and Roger H. Stuewer. Has the media forgotten physics? Physics in Perspective (PIP), 12(3):245–247, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-010-0032-7.

Rigden:2011:AD

[RS11a] John S. Rigden and Roger H. Stuewer. America in decline. Physics in Perspective (PIP), 13(4):385–386, December 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-011-0075-4.

Rigden:2011:BNBa

[RS11b] John S. Rigden and Roger H. Stuewer. Book notes [Information and the Nature of Reality: From Physics to Metaphysics (edited by Paul Davies and Niels Henrik Gregersen, Cambridge University Press, 2010, xvi + 382 pages, $30)]. Physics in Perspective (PIP), 13(4):481–483, December 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-011-0074-5.

Rigden:2011:BNG

[RS11c] John S. Rigden and Roger H. Stuewer. Book notes [Galileo’s Two Chief World Systems and Two New Sciences; Harald Fritzsch, You Are Wrong, Mr Einstein!]. Physics in Perspective (PIP), 13(3):373–375, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-011-0067-4.

Rigden:2011:BNJ

[RS11d] John S. Rigden and Roger H. Stuewer. Book notes [John Heilbron, Galileo; Galileo, de Revolutionibus; Galileo, Dialogue Concerning the Two Chief World Systems]. Physics in Perspective (PIP), 13(2):239–243, June 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-011-0060-y.

Rigden:2011:BNT

[RS11e] John S. Rigden and Roger H. Stuewer. Book notes [Timothy Ferris, The Science of Liberty: Democracy, Reason, and the Laws of Nature (HarperCollins 2010, 368 pages, $26.99); Frank Close, Neutrino (Oxford 2010, x + 181 pages, $18.95)]. Physics in Perspective (PIP), 13(1):110–113, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-011-0051-z.

Rigden:2011:CDC

[RS11f] John S. Rigden and Roger H. Stuewer. Constancy directs change. Physics in Perspective (PIP), 13(1):1–3, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-011-0053-x.

Rigden:2011:ESR

[RS11g] John S. Rigden and Roger H. Stuewer. Editorial: Scientific research: Reason and emotion. Physics in Perspective (PIP), 13(2):125–127, June 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-011-0062-9.

Rigden:2011:IPI

[RS11h] John S. Rigden and Roger H. Stuewer. Is it the public? Is it the scientists? Physics in Perspective (PIP), 13(3):257– 259, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-011-0068-3.

Rigden:2012:ASI

[RS12a] John S. Rigden and Roger H. Stuewer. Anti-science: It comes from the top. Physics in Perspective (PIP), 14(4):389–391, December 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-012-0103-z.

Rigden:2012:BNc

[RS12b] John S. Rigden and Roger H. Stuewer. Book notes. Physics in Perspective (PIP), 14(3):368–370, September 2012. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
012-0092-y.

Rigden:2012:BNd

[RS12c] John S. Rigden and Roger H. Stuewer. Book notes. Physics in Perspective (PIP), 14(4):508–511, December 2012. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
012-0102-0.

Rigden:2012:BNR

[RS12d] John S. Rigden and Roger H. Stuewer. Book notes [Roald Hoffmann and Iain Boyd Whyte, eds. Beyond the Finite: The Sublime in Art and Science (Oxford: Oxford University Press, 2011, xvii + 173 pages, $24.95)]. Physics in Perspective (PIP), 14 (2):242–244, June 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-012-0090-0.

Rigden:2012:EMD

[RS12e] John S. Rigden and Roger H. Stuewer. Editorial: Manipulated data does not escape notice. Physics in Perspective (PIP), 14(1):1–3, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-012-0082-0.

Rigden:2012:EUQ

[RS12f] John S. Rigden and Roger H. Stuewer. Editorial: Understanding of quantum mechanics eludes physicists for eighty-six years. Physics in Perspective (PIP), 14(2):123–125, June 2012. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
012-0091-z.

Rigden:2012:KAB

[RS12g] John S. Rigden and Roger H. Stuewer. Knowledge and appreciation based on understanding. Physics in Perspective (PIP), 14(3):255–257, September 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-012-0093-x.

Rigden:2012:BNL

[RS12h] John S. Rigden and Roger H. H. Stuewer. Book notes [Lawrence Krauss, Quantum Man: Richard Feynman’s Life in Science (W. W. Norton, 2011, xvii + 350, $24.95)]. Physics in Perspective (PIP), 14(1):113–115, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-012-0081-1.

Rigden:2013:BNb

[RS13a] John S. Rigden and Roger H. Stuewer. Book notes. Physics in Perspective (PIP), 15(2):241–243, June 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
013-0115-3.

Rigden:2013:BNc

[RS13b] John S. Rigden and Roger H. Stuewer. Book notes. Physics in Perspective (PIP), 15(3):359–360, September 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s00016-013-0119-z;   http://link.springer.com/article/
10.1007/s00016-013-0119-z.

Rigden:2013:BN

[RS13c] John S. Rigden and Roger H. Stuewer. Book notes. Physics in Perspective (PIP), 15(4):499–501, December 2013. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
013-0123-3.

Rigden:2013:BNS

[RS13d] John S. Rigden and Roger H. Stuewer. Book notes [Stuart Firestein, Ignorance: How it Drives Science, Oxford: Oxford University Press, 2012, viii + 195 pages, ISBN 0-19-982807-5]. Physics in Perspective (PIP), 15(1):116–117, March 2013. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00016-013-0106-4;   http://link.springer.com/article/
10.1007/s00016-013-0106-4.

Rigden:2013:DFL

[RS13e] John S. Rigden and Roger H. Stuewer. Don’t forget the liberal arts. Physics in Perspective (PIP), 15(1):1–2, March 2013. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00016-013-0107-3;   http://link.springer.com/article/
10.1007/s00016-013-0107-3.

Rigden:2013:EDU

[RS13f] John S. Rigden and Roger H. Stuewer. Enlightenment is dimming in the United States. Physics in Perspective (PIP), 15 (2):127–129, June 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-013-0116-2.

Rigden:2013:SPO

[RS13g] John S. Rigden and Roger H. Stuewer. Seventeen particles and one molecule. Physics in Perspective (PIP), 15(3):253– 255, September 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-013-0120-6.

Rigden:2013:WSL

[RS13h] John S. Rigden and Roger H. Stuewer. We shall be like the trees. Physics in Perspective (PIP), 15(4):371–372, December 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00016-013-0124-2;   http://link.springer.
com/article/10.1007/s00016-013-0124-2.

Roberts:2007:MHI

[RSD07] Lissa Roberts, Simon Schaffer, and Peter Dear, editors. The Mindful Hand: Inquiry and Invention from the Late Renaissance to Early Industrialisation, volume 9 of History of science and scholarship in the Netherlands. Koninkliijke Nederlandse Akademie van Wetenschappen, Amsterdam, The Netherlands, 2007. ISBN 90-6984-483-4 (hardcover). xxvii + 503 pp. LCCN T173.8 .M565 2007.

Ruhl:2020:BFP

[Ruh20] Christian P. Ruhl. “It’s better to forget physics”: The idea of the tactical nuclear weapon in the Early Cold War. Physics in Perspective (PIP), 22(1):26–51, March 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Rupke:2005:AHM

[Rup05] Nicolaas A. Rupke. Alexander von Humboldt: a metabiography. Peter Lang, Frankfurt am Main, Germany, 2005. ISBN 3-631-53932-0 (Frankfurt am Main), 0-8204-7693-5 (New York). 320 pp. LCCN ????

Russell:2000:MFP

[Rus00] Colin Archibald Russell. Michael Faraday: physics and faith. Oxford portraits in science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-511763-8 (hardcover). 124 pp. LCCN QC16.F2 .R87 2000. URL http://www.

loc.gov/catdir/enhancements/fy0610/00027008-d.html.

Rush:2015:UNF

[Rus15] John J. Rush. US neutron facility development in the last half-century: A cautionary tale. Physics in Perspective (PIP), 17(2):135–155, June 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-015-0158-8;
 http://link.springer.com/content/pdf/10.1007/s00016-
015-0158-8.pdf.

Ryckman:2010:BRH

[Ryc10] Thomas Ryckman. Book review: Hermann Weyl, Mind and Nature: Selected Writings on Philosophy, Mathematics, and Physics. Edited and with an introduction by Peter Pesic. Princeton: Princeton University Press, 2009, 272 pages. $35.00 (cloth). Physics in Perspective (PIP), 12(2):238–242, June 2010. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Ryden:2013:BRM

[Ryd13] Barbara Ryden. Book review: Michael Hoskin, The Construction of the Heavens: William Herschel’s Cosmology. Cambridge: Cambridge University Press, 2012, viii + 205 pages. $105.00 (cloth). Physics in Perspective (PIP), 15(1):125–126, March 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Rynasiewicz:1999:BRP

[Ryn99] Robert A. Rynasiewicz. Book review: Per F. Dahl, Flash of the Cathode Rays: A History of J. J. Thomson’s Electron. Bristol and Philadelphia: Institute of Physics Publishing, 1997. xvii + 526 pages. $49.50 (cloth). Physics in Perspective (PIP), 1(4): 455–457, December 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Saliba:2007:ISM

[Sal07] George Saliba. Islamic science and the making of the European Renaissance. Transformations. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-19557-7 (hardcover). xi + 315 pp. LCCN Q127.I742 S35 2007. URL http://www.loc.gov/

catdir/toc/ecip0617/2006023618.html.

Salvia:2014:G

[Sal14] Stefano Salvia. “Galileo’s Machine”: Late notes on free fall, projectile motion, and the force of percussion (ca. 1638–1639).

Physics in Perspective (PIP), 16(4):440–460, December 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-014-0149-1.

Salvia:2017:AHP

[Sal17] Stefano Salvia. From Archimedean hydrostatics to post-Aristotelian mechanics: Galileo’s early manuscripts De motu antiquiora (ca. 1590). Physics in Perspective (PIP), 19(2):105– 150, June 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Salvia:2019:ECP

[Sal19] Stefano Salvia. Embattled cooperation(s): Peaceful atoms, pacifist physicists, and partisans of peace in the Early Cold War (1947–1957). Physics in Perspective (PIP), 21(1):43–62, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Salvia:2020:BAJ

[Sal20] Stefano Salvia. The battle of the astronomers: Johann Adam Schall von Bell and Ferdinand Verbiest at the Court of the Celestial Emperors (1660–1670). Physics in Perspective (PIP), 22(2): 81–109, June 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sample:2010:MHG

[Sam10] Ian Sample. Massive: the hunt for the God particle. Virgin Books, London, UK, 2010. ISBN 1-905264-95-X (hardcover), 0-7535-2211-X (paperback). xi + 307 pp. LCCN QC793.5.B62 S26 2010b.

Seth:2019:GPH

[SB19] Swadesh Seth and Helmut Berndt. Genetic polymorphism of the human complement component C’3 among North Germans. Physics in Perspective (PIP), 21(1):59–61, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/BF00278566.pdf.

Suplee:2009:PU

[SBD09] Curt Suplee, A. (Amitava) Bhattacharjee, and Kristi Donahue. The plasma universe. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-51927-6 (paperback). xii + 76 pp. LCCN QC718 .S85 2009. Designed by Kristi Donahue.

Stuewer:2006:BR

[SBLM^{+}06] Roger H. Stuewer, Stephen G. Brush, Charlotte Lowe-Ma, J. R. Dorfman, and Edward Gerjuoy. Book reviews. Physics in Perspective (PIP), 8(1):104–112, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-005-0251-5.

Scerri:2007:PTS

[Sce07] Eric R. Scerri. The Periodic Table: Its Story and Its Significance. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-530573-6. xxii + 346 pp.

LCCN QD467 .S345 2007. URL http://www.loc.gov/catdir/
enhancements/fy0635/2005037784-d.html;   http://www.
loc.gov/catdir/enhancements/fy0722/2005037784-b.html;
 http://www.loc.gov/catdir/toc/ecip065/2005037784.html.

Schofield:1997:EJP

[Sch97] Robert E. Schofield. The enlightenment of Joseph Priestley: a study of his life and work from 1733 to 1773. Pennsylvania State University Press, University Park, PA, USA, 1997. ISBN 0-271-01662-0. xii + 305 pp. LCCN BX9869.P8 S36 1997. URL http:

//www.psupress.org/books/titles/0-271-01662-0.html.

Schewe:2000:BRK

[Sch00a] Philip F. Schewe. Book review: Keay Davidson, Carl Sagan: A Life. New York: John Wiley, 1999, xx + 540 pages. $30.00 (cloth). Physics in Perspective (PIP), 2(4):446–447, December 2000.

Schweber:2000:SBB

[Sch00b] S. S. (Silvan S.) Schweber. In the shadow of the bomb: Bethe, Oppenheimer, and the moral responsibility of the scientist. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-04989-0 (hardcover). xviii + 260 + 8 pp. LCCN QC774.O56 S32 2000.

Schwartz:2001:BRR

[Sch01] A. Truman Schwartz. Book review: Robert E. Schofield, The Enlightenment of Joseph Priestley: A Study of His Life and Work from 1733 to 1773. University Park: Pennsylvania State University Press, 1997, xii+ 305 pages. $45.00 (Cloth). Physics in Perspective (PIP), 3(4):495–496, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Schirrmacher:2003:PHN

[Sch03] Arne Schirrmacher. Planting in his neighbor’s garden: David Hilbert and early Göttingen quantum physics. Physics in Perspective (PIP), 5(1):4–20, April 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160300001.

Schlote:2004:CNC

[Sch04] Karl-Heinz H. Schlote. Carl Neumann’s contributions to electrodynamics. Physics in Perspective (PIP), 6(3):252–270, September 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-003-0192-9.

Schwitters:2005:BRA

[Sch05] Roy F. Schwitters. Book review: Andrew Watson, The Quantum Quark. Cambridge: Cambridge University Press, 2004, x + 464 pages. $30.00 (cloth). Physics in Perspective (PIP), 7(4): 499–501, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Schemmel:2006:EGT

[Sch06] Matthias Schemmel. The English Galileo: Thomas Harriot and the force of shared knowledge in early modern mechanics. Physics in Perspective (PIP), 8(4):360–380, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-006-0275-5.

Schweber:2008:EOM

[Sch08] S. S. (Silvan S.) Schweber. Einstein and Oppenheimer: the meaning of genius. Harvard University Press, Cambridge, MA, USA, 2008. ISBN 0-674-02828-7. xiv + 412 pp. LCCN QC16.E5

S3285 2008. URL http://www.loc.gov/catdir/toc/ecip082/
2007043108.html.

Schettino:2009:PMU

[Sch09a] Edvige Schettino. The Physics Museum of the University of Naples “Federico II”. Physics in Perspective (PIP), 11(4):

442–453, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-008-0402-6.

Schweber:2009:WPS

[Sch09b] S. S. Schweber. Weimar physics: Sommerfeld’s Seminar and the Causality Principle. Physics in Perspective (PIP), 11(3): 261–301, September 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-008-0385-3.

Schwarz:2011:STN

[Sch11] Stephan Schwarz. Science, technology, and the Niels Bohr Institute in Occupied Denmark. Physics in Perspective (PIP), 13(4): 401–432, December 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-011-0059-4.

Schweber:2014:WBH

[Sch14] Silvan S. Schweber. Writing the biography of Hans Bethe: Contextual history and Paul Forman. Physics in Perspective (PIP), 16(2):179–217, June 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-014-0136-6.

Schweber:2016:SIC

[Sch16] S. S. Schweber. The Shelter Island conferences revisited: “fundamental” physics in the decade 1975–1985. Physics in Perspective (PIP), 18(1):58–147, April 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-016-0180-5.

Schwarz:2021:ONB

[Sch21] Stephan Schwarz. The occupation of Niels Bohr’s Institute: December 6, 1943–February 3, 1944. Physics in Perspective (PIP), 23(1):49–82, March 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Schwarz:2022:DAW

[Sch22] Stephan Schwarz. Drama around a wartime Heisenberg letter. Physics in Perspective (PIP), 24(1):72–92, March 2022. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL https://link.springer.com/article/10.
1007/s00016-021-00285-1.

Schwarz:2023:BR

[Sch23] Stephan Schwarz. Book review. Physics in Perspective (PIP), 25(1–2):79–82, June 2023. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL https://link.springer.

com/article/10.1007/s00016-023-00299-x.

Simoes:2012:PSL

[SDC12] Ana Simões, Maria Paula Diogo, and Ana Carneiro. Physical sciences in Lisbon. Physics in Perspective (PIP), 14(3): 335–367, September 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-012-0096-7.

Shiltsev:2014:SAP

[SE14] Vladimir Shiltsev and Marat Eseev. Scientific Arkhangelsk and Pomorie: A walk through centuries and thousands of miles. Physics in Perspective (PIP), 16(3):390–405, September 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-014-0140-x.

Schmidtke:2019:PRG

[SE19] Jörg Schmidtke and Wolfgang Engel. On the problem of regional gene duplication in diploid fish of the orders Ostariophysi and Isospondyli. Physics in Perspective (PIP), 21(1):39–45, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/content/pdf/
10.1007/BF00278563.pdf.

Sime:2004:BR

[SEA^{+}04] Ruth Lewin Sime, Robert Ehrlich, Antonio Aurilia, Roger G. Newton, Allan Franklin, and Laurie M. Brown. Book reviews. Physics in Perspective (PIP), 6(4):477–486, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-004-0230-2.

Segre:2007:FCS

[Seg07] Gino Segrè. Faust in Copenhagen: a struggle for the soul of physics. Viking, New York, NY, USA, 2007. ISBN 0-670-03858-X. x + 310 pp. LCCN QC15 .S427 2007. URL http://

www.loc.gov/catdir/enhancements/fy0743/2006052807-b.

html;   http://www.loc.gov/catdir/enhancements/fy0743/
2006052807-d.html.

Segre:2009:BRL

[Seg09] Gino Segrè. Book review: Lillian Hoddeson, Adrienne W. Kolb, and Catherine Westfall, Fermilab: Physics, the Frontier, and Megascience. Chicago, The University of Chicago Press, 2008, 497 pages. $45.00 (cloth). Physics in Perspective (PIP), 11(3): 350–351, September 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Segre:2010:BRG

[Seg10a] Gino Segrè. Book review: Guido Bacciagaluppi and Antony Valentini, Quantum Theory at the Crossroads: Reconsidering the 1927 Solvay Conference. New York: Cambridge University Press, 2009, xxv + 530 pages. $126.00 (cloth). Physics in Perspective (PIP), 12(4):499–501, December 2010. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Segre:2010:BRJ

[Seg10b] Gino Segrè. Book review: João Magueijo, A Brilliant Darkness: The Extraordinary Life and Mysterious Disappearance of Ettore Majorana, the Troubled Genius of the Nuclear Age. New York: Basic Books, 2009. Physics in Perspective (PIP), 12(3):365–367, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Segre:2011:BRA

[Seg11a] Gino Segrè. Book review: Alberto Martinez, Science Secrets: The Truth about Darwin’s Finches, Einstein’s Wife, and Other Myths. Pittsburgh: University of Pittsburgh Press, 2011, xviii + 324 pages. $24.95 (cloth). Physics in Perspective (PIP), 13(4): 495–497, December 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Segre:2011:OGM

[Seg11b] Gino Segrè. Ordinary geniuses: Max Delbrück, George Gamow, and the origins of genomics and Big Bang cosmology. Viking, New York, NY, USA, 2011. ISBN 0-670-02276-4. xxi + 330 pp. LCCN QH31.D434 S44 2011.

Segre:2013:BRS

[Seg13] Gino Segrè. Book review: Simone Turchetti, The Pontecorvo Affair: A Cold War Defection and Nuclear Physics. Chicago: University of Chicago Press, 2012, 292 pages. $45.00 (cloth). Physics in Perspective (PIP), 15(1):121–123, March 2013. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Segre:2015:BRF

[Seg15] Gino Segrè. Book review: Frank Close: Half-Life: The Divided Life of Bruno Pontecorvo, Physicist or Spy. New York: Basic Books, 2015, xix + 366 pages. $29.99 (cloth). Physics in Perspective (PIP), 17(1):70–73, March 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-015-0155-y.

Sehested:2019:SMR

[Seh19] Jens Sehested. A simple method for R banding of human chromosomes, showing a pH-dependent connection between R and G bands. Physics in Perspective (PIP), 21(1):55–58, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/BF00278565.pdf.

Seidel:2000:OBB

[Sei00] Robert W. Seidel. Opening the Black Box at Bradbury Science Museum, Los Alamos. Physics in Perspective (PIP), 2 (2):211–216, June 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s000160050043.

Seth:2010:CQA

[Set10] Suman Seth. Crafting the quantum: Arnold Sommerfeld and the practice of theory, 1890–1926. Transformations. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01373-8 (hardcover). viii + 378 pp. LCCN QC16.S76 S48 2010. URL http://

catalogue.bnf.fr/ark:/12148/cb42209714w.

Steinle:2003:BR

[SGJ03] Friedrich Steinle, Edward Gerjuoy, and Richard Jacob. Book reviews. Physics in Perspective (PIP), 5(1):122–126, April 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160300005.

Shimony:2006:BR

[SGvB06] Abner Shimony, David J. Griffiths, and Hans Christian von Baeyer. Book reviews. Physics in Perspective (PIP), 8(3): 347–355, September 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-005-0292-4.

Shapiro:2002:BRS

[Sha02] Alan E. Shapiro. Book review: Scott L. Montgomery, Science in Translation: Movements of Knowledge through Cultures and Time. Chicago: University of Chicago Press, 2000, xi + 325 pages. $28.00 (Hardbound). Physics in Perspective (PIP), 4 (3):361–362, August 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Shank:2006:BRC

[Sha06] J. B. Shank. Book review: C. D. Andriesse, Huygens: The Man Behind the Principle, translated from the Dutch by Sally Miedema. Cambridge: Cambridge University Press, 2005, xxvi + 440 pages. $90.00 (cloth). Physics in Perspective (PIP), 8(4): 474–477, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sheehan:2013:TVB

[She13] William Sheehan. From the Transits of Venus to the birth of experimental psychology. Physics in Perspective (PIP), 15 (2):130–159, June 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-012-0101-1.

Stuewer:2004:BR

[SHH04] Roger H. Stuewer, Thomas L. Hankins, and Art Hobson. Book reviews. Physics in Perspective (PIP), 6(3):362–367, September 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-003-0197-4.

Shimony:2006:BRG

[Shi06] Abner Shimony. Book review: GianCarlo Ghirardi, Sneaking a Look at God’s Cards: Unraveling the Mysteries of Quantum Mechanics, Revised Edition, translated from the Italian by Gerald Malsbary. Princeton and Oxford: Princeton University Press, 2005, xxii + 488 pages. $35.00 (cloth). Physics in Perspective (PIP), 8(3):347–352, September 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sime:2007:BR

[SHL^{+}07] Ruth Lewin Sime, Robert C. Hilborn, G. Peter Lepage, Robert Greenler, and Virginia Trimble. Book reviews. Physics in Perspective (PIP), 9(1):118–125, January 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-007-0314-x.

Sichau:2006:SCM

[Sic06] Christian Sichau. Storming a citadel: Mathematical theory and experimental practice. Physics in Perspective (PIP), 8(3): 236–254, September 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0272-8.

Silverman:2005:SRM

[Sil05] Mark P. Silverman. Schrödinger’s rabbits: The many worlds of quantum. Physics in Perspective (PIP), 7(4):496–498, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Silverman:2007:BRA

[Sil07] Mark P. Silverman. Book review: Al Gore, An Inconvenient Truth: The Planetary Emergence of Global Warming And What We Can Do About It. New York: Rodale, 2006, 328 pages. $21.95 (paper). Physics in Perspective (PIP), 9(2):259–262, June 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sime:2000:STE

[Sim00] Ruth Lewin Sime. The search for transuranium elements and the discovery of nuclear fission. Physics in Perspective (PIP), 2(1):48–62, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050036;   http:/
/www.springerlink.com/content/nr2t13tndn6t9t72/.

Simoes:2002:DCC

[Sim02] Ana Simões. Dirac’s claim and the chemists. Physics in Perspective (PIP), 4(3):253–266, August 2002. CODEN PHPEF2.

ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-002-8369-1.

Sime:2004:BRM

[Sim04] Ruth Lewin Sime. Book review: Mary Jo Nye, ed., The Modern Physical and Chemical Sciences. Vol. 5. The Cambridge History of Science. Cambridge: Cambridge University Press, 2003, xxvii + 678 pages. $95.00 (cloth). Physics in Perspective (PIP), 6 (4):477, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sime:2006:PMO

[Sim06] Ruth Lewin Sime. The politics of memory: Otto Hahn and the Third Reich. Physics in Perspective (PIP), 8(1):3–51, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-004-0248-5.

Sime:2007:BRB

[Sim07] Ruth Lewin Sime. Book review: Barbara Goldsmith, Obsessive Genius: The Inner World of Marie Curie. New York: W. W. Norton & Company, 2005, 257 pages + Reading Group Guide. $14.95 (paper). Physics in Perspective (PIP), 9(1):118, January 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sime:2008:BRD

[Sim08] Ruth Lewin Sime. Book review: Dieter Hoffmann and Mark Walker, ed., Physiker zwischen Autonomie und Anpassung: Die Deutsche Physikalische Gesellschaft im Dritten Reich. (German) [Physicists between autonomy and adjustment: The German Physical Society in the Third Reich], Weinheim: Wiley-VCH Verlag, 2007, xii + 675 pages. EUR99.00 (cloth). Physics in Perspective (PIP), 10(4):487–489, December 2008. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sime:2010:BRM

[Sim10a] Ruth Lewin Sime. Book review: Martijn Eickhoff, In the name of science? P. J. W. Debye and his career in Nazi Germany. Translated by Peter Mason. Amsterdam: Aksant, 2008, 184 pages. EUR24.90 (paper). Physics in Perspective (PIP), 12 (1):115–117, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sime:2010:IHN

[Sim10b] Ruth Lewin Sime. An inconvenient history: the nuclear-fission display in the deutsches museum. Physics in Perspective (PIP), 12(2):190–218, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-009-0013-x;   http://www.
springerlink.com/content/w716842562715257/.

Sime:2012:BRM

[Sim12a] Ruth Lewin Sime. Book review: Marjorie C. Malley, Radioactivity: A History of a Mysterious Science. Oxford and New York: Oxford University Press, 2011, xxi + 267 pages. $21.95 (cloth). Physics in Perspective (PIP), 14(2):245–246, June 2012. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sime:2012:PFO

[Sim12b] Ruth Lewin Sime. The politics of forgetting: Otto Hahn and the German Nuclear-Fission Project in World War II. Physics in Perspective (PIP), 14(1):59–94, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-011-0065-6;
 http://www.springerlink.com/content/k12202vg92147h68/

.

Sime:2013:MBP

[Sim13] Ruth Lewin Sime. Marietta Blau: Pioneer of photographic nuclear emulsions and particle physics. Physics in Perspective (PIP), 15(1):3–32, March 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-012-0097-6.

Singh:2002:CVR

[Sin02] Rajinder Singh. C. V. Raman and the discovery of the Raman Effect. Physics in Perspective (PIP), 4(4):399– 420, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160200002;   http:/
/www.nobelprize.org/nobel_prizes/physics/laureates/1930/

.

Steinle:2016:EEA

[SL16] Friedrich Steinle and Alex Levine. Exploratory experiments: Ampère, Faraday, and the origins of electrodynamics. University of Pittsburgh Press, Pittsburgh, PA, USA, 2016. ISBN 0-8229-4450-2 (hardcover). x + 494 pp. LCCN QC630.5 .S7413 2016. Translation from German by Alex Levine.

Smith:1998:SEC

[Smi98] Crosbie Smith. The science of energy: a cultural history of energy physics in Victorian Britain. University of Chicago Press, Chicago, IL, USA and London, UK, 1998. ISBN 0-226-76420-6, 0-226-76421-4 (paperback). xi + 404 pp.

LCCN QC72 .S58 1998. URL http://www.loc.gov/catdir/
description/uchi051/98024960.html;   http://www.loc.
gov/catdir/enhancements/fy0608/98024960-t.html;   http:
//www.loc.gov/catdir/enhancements/fy0609/98024960-b.
html.

Smith:2020:BRJ

[Smi20] Robert W. Smith. Book review: Joshua Nall, News from Mars: Mass Media and the Forging of a New Astronomy, 1860–1910, Pittsburgh: University of Pittsburgh Press, 2019, 287 pages, $50 (hardcover). Physics in Perspective (PIP), 22(3):182–184, September 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sobotka:2008:BRE

[Sob08] Lee G. Sobotka. Book review: Eric R. Scerri, The Periodic Table: Its Story and Significance. Oxford: Oxford University Press, 2007, xxii + 346 pages. $35.00 (cloth). Physics in Perspective (PIP), 10(3):374–375, September 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sokal:2008:BHS

[Sok08] Alan Sokal. Beyond the hoax: Science, philosophy and culture. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-956183-4, 0-19-923920-7. xxi + 465 pp. LCCN ????

Solway:2006:QLB

[Sol06] Andrew Solway. Quantum leaps and big bangs!: a history of astronomy. Stargazers’ guides. Heinemann Library, Chicago, IL, USA, 2006. ISBN 1-4034-7712-4 ( hardcover), 1-4034-7719-1 (paperback). 48 pp. LCCN QB28 .S65 2006. URL http://www.

loc.gov/catdir/toc/ecip061/2005029111.html.

Sanchez-Ron:2006:PTP

[SR06] José M. Sánchez-Ron. The physical tourist: Physics in Madrid: Where science competed with art. Physics in Perspective (PIP), 8(3):318–343, September 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-005-0261-3.

Strohmaier:2006:MBS

[SRD06] Brigitte Strohmaier, Robert W. Rosner, and Paul F. Dvorak, editors. Marietta Blau, stars of disintegration: biography of a pioneer of particle physics. Studies in Austrian literature, culture, and thought. Ariadne Press, Riverside, CA, USA, 2006. ISBN 1-57241-147-3. 220 pp. LCCN QC774.B53 S77 2006. URL http:

//www.loc.gov/catdir/toc/ecip066/2005037944.html.

Siegmund-Schultze:2007:PFR

[SS07] Reinhard Siegmund-Schultze. Philipp Frank, Richard von Mises, and the Frank–Mises. Physics in Perspective (PIP), 9(1):26–57, January 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-006-0288-0.

Schreier:2008:PTP

[SS08] Wolfgang Schreier and Karl-Heinz Schlote. The physical tourist: Physics in Leipzig: An amble through the centuries. Physics in Perspective (PIP), 10(2):224–243, June 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
007-0356-0.

Sopka:2010:BTS

[SS10] Katherine R. Sopka and Elisabeth M. Sopka. The Bonebrake Theological Seminary: Top-secret Manhattan Project site. Physics in Perspective (PIP), 12(3):338–349, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-010-0019-4.

Staley:2001:LOT

[Sta01] Kent W. Staley. Lost origins of the third generation of quarks: Theory, philosophy, and experiment. Physics in Perspective (PIP), 3(2):210–229, June 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/PL00000530.

Stachel:2002:EBZ

[Sta02] John J. Stachel. Einstein from “B” to “Z”, volume 9 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2002. ISBN 0-8176-4143-2, 3-7643-4143-2 (Basel). xi + 556 pp. LCCN QC16.E5 S69 2002.

Staubermann:2006:BRK

[Sta06a] Klaus Staubermann. Book review: Klaus Hentschel, Gaussens unsichtbare Hand: Der Universitäts-Mechanicus und Maschinen-Inspector Moritz Meyerstein: Ein Instrumentenbauer im 19. Jahrhundert. Göttingen: Vandenhoeck & Ruprecht, 2005, 321 pages. EUR 79.00 (cloth). [Gauss’s invisible hand: the university mechanics and machine inspector Moritz Meyerstein: an instrument maker in the 19th Century]. Physics in Perspective (PIP), 8(4):479, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Staubermann:2006:PTS

[Sta06b] Klaus Staubermann. The physical tourist: Science sights in Utrecht. Physics in Perspective (PIP), 8(2):214–221, May 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-005-0260-4.

Stanley:2007:PMR

[Sta07] Matthew Stanley. Practical mystic: religion, science, and A. S. Eddington. University of Chicago Press, Chicago, IL, USA and London, UK, 2007. ISBN 0-226-77097-4 (hardcover). x + 313 pp. LCCN BL240.3 .S725 2007. URL http://

www.loc.gov/catdir/enhancements/fy0809/2007005482-d.
html;   http://www.loc.gov/catdir/enhancements/fy0809/
2007005482-t.html;   http://www.loc.gov/catdir/enhancements/
fy0827/2007005482-b.html.

Stacey:2010:QFE

[Sta10] Weston M. Stacey, editor. The quest for a fusion energy reactor: an insider’s account of the INTOR Workshop. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-973384-8 (hardcover). LCCN ????

Steinle:2003:BRP

[Ste03a] Friedrich Steinle. Book review: Peter Heering, Falk Riess, and Christian Sichau, ed., Im Labor der Physikgeschichte: Zur Untersuchung historischer Experimentalpraxis. [(German) [In the laboratory of the history of physics — Research on historical experiments]]. Oldenburg: BIS-Verlag, 2000, 220 pages. 11.30 euros (soft cover). Physics in Perspective (PIP), 5(1):122–123, April 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Stern:2003:BRL

[Ste03b] David P. Stern. Book review: Leon Golub and Jay M. Pasachoff, Nearest Star: The Surprising Science of our Sun. Cambridge, Mass.: Harvard University Press, 2001, xii + 267 pages, $29.95 (cloth). Physics in Perspective (PIP), 5(2):238–239, May 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Steinle:2005:EEA

[Ste05] Friedrich Steinle. Explorative Experimente: Ampère, Faraday und die Ursprünge der Elektrodynamik. (German) [Exploratory experiments: Ampère, Faraday, and the origins of electrodynamics], volume 50 of Boethius. Steiner, Stuttgart, Germany, 2005. ISBN 3-515-08185-2 (hardcover). 450 pp. LCCN QC630.5 .S74 2005.

Steinle:2017:RJR

[Ste17] Friedrich Steinle. Reply to James R. Hofmann. Physics in Perspective (PIP), 19(4):452–457, December 2017. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). See [Hof17].

Sterligov:2023:BAF

[Ste23] Ivan Sterligov. Before and after the fall: Geography of Soviet and post-Soviet physics surveyed via leading journals. Physics in Perspective (PIP), 25(1–2):41–75, June 2023. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL https://link.springer.com/article/10.
1007/s00016-023-00298-y.

Stinner:2008:BRI

[Sti08] Arthur Stinner. Book review: István Hargittai, The Martians of Science: Five Physicists Who Changed the Twentieth Century.

Oxford: Oxford University Press, 2006, xxiv + 313 pages. $34.50 (cloth). Physics in Perspective (PIP), 10(2):246–252, June 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Stoltzner:2002:FSE

[Stö02] Michael Stöltzner. Franz Serafin Exner’s indeterminist theory of culture. Physics in Perspective (PIP), 4(3):267–319, August 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-002-8370-8.

Strzalkowski:2004:PTP

[Str04] Adam Strzalkowski. The physical tourist: Physics in Cracow. Physics in Perspective (PIP), 6(3):344–361, September 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-003-0212-9.

Stuewer:2004:BRR

[Stu04] Roger H. Stuewer. Book review: Robert Rosner and Brigitte Strohmaier, ed., Marietta Blau — Sterne der Zertrümmerung: Biographie einer Wegbereiterin der modernen Teilchenphysik. Wien, Köln, Weimar: Böhlau Verlag, 2003, 229 pages. 29.90 euros (cloth). Physics in Perspective (PIP), 6(3):362–364, September 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Stuewer:2005:GHL

[Stu05] Roger H. Stuewer. George de Hevesy/Loránd Eötvös: Scientist– teacher/Eugene P. Wigner and his Hungarian teachers. Physics in Perspective (PIP), 7(4):498–499, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Stuewer:2006:BRC

[Stu06] Roger H. Stuewer. Book review: Carlo Bernardini and Luisa Bonolis, ed., Enrico Fermi: His Work and Legacy. Bologna: Società Italiana di Fisica and Berlin, Heidelberg, New York: Springer Verlag, 2004, xii + 411 pages. $49.95 (cloth). Physics in Perspective (PIP), 8(1):104–105, March 2006. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Stuewer:2018:MJR

[Stu18] Roger H. Stuewer. In memoriam: John S. Rigden. Physics in Perspective (PIP), 20(1):4–7, March 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Styer:2005:BRT

[Sty05] D. F. Styer. Book review: Tony Hey and Patrick Walters, The New Quantum Universe. Cambridge: Cambridge University Press, 2003, xv + 357 pages. $35.00 (paper). Physics in Perspective (PIP), 7(3):381, September 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sime:2012:BR

[SvBC^{+}12] Ruth Lewin Sime, Hans Christian von Baeyer, David C. Cassidy, Robert P. Crease, and Gregory A. Good. Book reviews. Physics in Perspective (PIP), 14(2):245–254, June 2012. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
012-0089-6.

Stinner:2008:BR

[SvBG08] Arthur Stinner, Hans Christian von Baeyer, and David Goodstein. Book reviews. Physics in Perspective (PIP), 10(2): 246–254, June 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0362-7.

Tattersall:2013:BRJ

[Tat13] Mason Tattersall. Book review: József Illy, The Practical Einstein: Experiments, Patents, Inventions. Baltimore: The Johns Hopkins University Press, 2012, xi + 202 pages. $60.00 (cloth). Physics in Perspective (PIP), 15(1):123–125, March 2013. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Taylor:2005:BRR

[Tay05] Edwin F. Taylor. Book review: Richard Wolfson, Simply Einstein: Relativity Demystified. New York: W. W. Norton, 2003, 272 pages. $13.95 (paper). Physics in Perspective (PIP), 7 (1):132–133, March 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Taylor:2010:BRF

[Tay10] Edwin F. Taylor. Book review: Fulvio Melia, Cracking the Einstein Code: Relativity and the Birth of Black Hole Physics. With an Afterward by Roy Kerr. Chicago: University of Chicago Press, 2009, 152 pages. $25.00 (cloth). Physics in Perspective (PIP), 12(4):502–503, December 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Teichmann:2002:BRM

[Tei02] Jürgen Teichmann. Book review: Myles W. Jackson, Spectrum of Belief: Joseph von Fraunhofer and the Craft of Precision Optics. Cambridge, Mass. and London: MIT Press, 2000, x + 284 pages. $34.95 (cloth). Physics in Perspective (PIP), 4(1):117– 118, February 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Teichmann:2002:PPM

[TEW02] Jürgen Teichmann, Michael Eckert, and Stefan Wolff. Physicists and physics in Munich. Physics in Perspective (PIP), 4 (3):333–359, August 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-002-8372-6.

Taylor:2007:AMS

[TF07] Philip L. Taylor and William J. Fickinger. In appreciation: Multiple scattering: Leslie Foldy’s winding road through physics. Physics in Perspective (PIP), 9(3):346–356, September 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-007-0338-2.

Tisza:2009:ATPa

[Tis09a] Laszlo Tisza. Adventures of a Theoretical Physicist, Part I: Europe. Physics in Perspective (PIP), 11(1):46–97, March 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/

s00016-008-0405-3. Interview with Andor Frenkel.

Tisza:2009:ATPb

[Tis09b] Laszlo Tisza. Adventures of a theoretical physicist, Part II: America. Physics in Perspective (PIP), 11(2):120–168, June 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-008-0406-2.

Teichmann:2002:BR

[TKP^{+}02] Jürgen Teichmann, Malvin H. Kalos, Douglass E. Post, Hans Christian von Baeyer, and Roberto Torretti. Book reviews. Physics in Perspective (PIP), 4(1):117–123, February 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-002-8360-x.

Torretti:1999:PP

[Tor99] Roberto Torretti. The Philosophy of Physics. The evolution of modern philosophy. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-56259-7 (hardcover), 0-521-56571-5 (paperback), 1-139-17298-0 (e-book). xvi + 512 pp.

LCCN QC6 .T656 1999. URL http://www.loc.gov/catdir/
description/cam0210/99042504.html;   http://www.loc.
gov/catdir/samples/cam032/99042504.html;   http://www.
loc.gov/catdir/toc/cam025/99042504.html.

Torretti:2000:GSC

[Tor00] Roberto Torretti. Gravity as spacetime curvature. Physics in Perspective (PIP), 2(2):118–134, June 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s000160050039.

Torretti:2002:BRS

[Tor02] Roberto Torretti. Book review: Steve Fuller, Thomas Kuhn: A Philosophical History for Our Times. Chicago: University of Chicago Press, 2000, xviii + 472 pages. $35.00 (cloth). Physics in Perspective (PIP), 4(1):120–123, February 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Townes:1999:HLH

[Tow99] Charles H. Townes. How the Laser Happened: Adventures of a Scientist. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-512268-2 (hardcover), 0-19-515376-6, 1-280-47085-2. 200 pp. LCCN QC687.2

.T68 1999. URL http://www.loc.gov/catdir/enhancements/
fy0603/98022216-d.html;   http://www.loc.gov/catdir/
enhancements/fy0603/98022216-t.html;   http://www.loc.
gov/catdir/enhancements/fy0724/98022216-b.html.

Trainer:2008:MLK

[Tra08] Matthew Trainer. In memoriam: Lord Kelvin, recipient of the John Fritz Medal in 1905. Physics in Perspective (PIP), 10 (2):212–223, June 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-007-0344-4.

Trefil:2008:WS

[Tre08]  James S. Trefil. Why science?  Teachers College Press,

New York, NY, USA, 2008. ISBN 0-8077-4830-7 (paperback), 0-8077-4831-5 (hardcover). xii + 208 pp. LCCN

Q181 .T8195 2008. URL http://www.loc.gov/catdir/toc/
ecip0717/2007018130.html.

Trimble:2005:SLA

[Tri05] Virginia Trimble. The sky is not the limit: Adventures of an urban astrophysicist. Physics in Perspective (PIP), 7(4):492–494, December 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Trimble:2007:BRJ

[Tri07] Virginia Trimble. Book review: Jose Wudka, Space-Time, Relativity, and Cosmology. Cambridge: Cambridge University Press, 2005, ix + 320 pages. $55.00 (cloth). Physics in Perspective (PIP), 9(1):123–125, January 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Trimble:2008:BRB

[Tri08] Virginia Trimble. Book review: Brigitte Strohmaier and Robert Rosner, Marietta Blau — Stars of Disintegration: Biography of a Pioneer of Particle Physics. Translated and expanded from the German. Riverside, California: Ariadne Press, 2006, [iv] + 220 pages. $27.00 (paper); Robert Rosner and Brigitte Strohmaier, ed., Marietta Blau — Sterne der Zertrümmerung: Biographie einer Wegbereiterin der modernen Teilchenphysik. Wien, Köln, Weimar: Böhlau Verlag, 2003 [Reviewed by Roger H. Stuewer, Physics in Perspective 6 (2004), 362-364]. Physics in Perspective (PIP), 10(1):127–130, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Trimble:2009:BRH

[Tri09] Virginia Trimble. Book review: Helge S. Kragh, Conceptions of the Cosmos: From Myths to the Accelerating Universe: A History of Cosmology. Oxford: Oxford University Press, 2007, viii + 276 pages. $85.00 (cloth). Physics in Perspective (PIP), 11(1):109–111, March 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Teller:2001:MTC

[TS01] Edward Teller and Judith L. Shoolery. Memoirs: a Twentieth-Century Journey in Science and Politics. Perseus Publishers, Cambridge, MA, USA, 2001. ISBN 0-7382-0532-X, 1-903985-12-9. xii + 628 pp. LCCN QC16.T37 M55 2001. URL http://

alsos.wlu.edu/information.aspx?id=784;   http://www.
loc.gov/catdir/enhancements/fy0833/2001097880-d.html;
 http://www.loc.gov/catdir/toc/fy02/2001097880.html.

Talebian:2012:ATI

[TT12] Mohammad Talebian and Ehsan Talebian. Alenush Terian: The Iranian solar mother: In memoriam. Physics in Perspective (PIP), 14(2):239–241, June 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-012-0085-x.

Turchetti:2012:PAC

[Tur12] Simone Turchetti. The Pontecorvo affair: a cold war defection and nuclear physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2012. ISBN 0-226-81664-8 (hardcover). 292 pp. LCCN QC774.P66 T8713 2012.

Unna:2000:GPH

[Unn00] Issachar Unna. The genesis of physics at the Hebrew University of Jerusalem. Physics in Perspective (PIP), 2(4):336–380, December 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s000160050050.

Vandenbosch:2003:BRP

[Van03] Robert Vandenbosch. Book review: Per F. Dahl, From Nuclear Transmutation to Nuclear Fission, 1932–1939. Bristol and Philadelphia: Institute of Physics Publishing, 2002, xii + 304 pages. $75.00 (cloth). Physics in Perspective (PIP), 5(3):354– 355, September 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

VanDongen:2007:REF

[Van07] Jeroen Van Dongen. Reactionaries and Einstein’s fame: “German scientists for the preservation of pure science,” Relativity, and the Bad Nauheim Meeting. Physics in Perspective (PIP), 9(2):212–230, June 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0318-y;   http://www.
springerlink.com/content/l6518246k7367437/fulltext.pdf.

vanDongen:2012:MIM

[van12] Jeroen van Dongen. Mistaken identity and mirror images: Albert and Carl Einstein, Leiden and Berlin, Relativity and revolution. Physics in Perspective (PIP), 14(2):126–177, June 2012. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s00016-012-0084-y.pdf;   http://
www.springerlink.com/content/w55417550px35577/fulltext.
pdf.

vonBaeyer:1999:BRB

[vB99] Hans Christian von Baeyer. Book review: Bernard Pullman, The Atom in the History of Human Thought, translated from the French by Axel Reisinger. New York: Oxford University Press, 1998, xii + 403 pages. $32.50 (cloth). Physics in Perspective (PIP), 1(1):118–120, March 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBaeyer:2000:BRP

[vB00] Hans Christian von Baeyer. Book review: Pierre Marage and Gregoire Wallenborn, eds., The Solvay Councils and the Birth of Modern Physics. Basel: Birkhäuser Verlag, 1999, xiii + 226 pages. $79.95 (cloth). Physics in Perspective (PIP), 2(1):111– 112, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBaeyer:2001:BRD

[vB01] Hans Christian von Baeyer. Book review: Dennis Overbye, Einstein in Love: A Scientific Romance. New York: Viking, 2000, xv + 416 pages. $27.95 (cloth). Physics in Perspective (PIP), 3(3):375–376, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBaeyer:2002:BRG

[vB02a] Hans Christian von Baeyer. Book review: Gregg Herken, Cardinal Choices: Presidential Science Advising from the Atomic Bomb to SDI, revised and expanded edition. Stanford: Stanford University Press, 2000, xv + 358 pages. $22.95 (paper). Physics in Perspective (PIP), 4(1):119–120, February 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBaeyer:2002:BRS

[vB02b] Hans Christian von Baeyer. Book review: Sean F. Johnston, A history of light and colour measurements: Science in the shadows, Bristol and Philadelphia: Institute of Physics Publishing, 2001, xi + 281 pages. $75 (cloth). Physics in Perspective (PIP), 4(4):491–492, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBaeyer:2006:BRD

[vB06] Hans Christian von Baeyer. Book review: David Park, The Grand Contraption: The World as Myth, Number, and Chance. Princeton: Princeton University Press, 2005, xiv + 331 pages. $29.95 (cloth). Physics in Perspective (PIP), 8(3):354–355, September 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBaeyer:2008:BRD

[vB08a] Hans Christian von Baeyer. Book review: Daniel Kennefick, Traveling at the Speed of Thought: Einstein and the Quest for Gravitational Waves. Princeton: Princeton University Press, 2007, xii + 319 pages. $35.00 (cloth). Physics in Perspective (PIP), 10(3):369–371, September 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBaeyer:2008:BRR

[vB08b] Hans Christian von Baeyer. Book review: Roger G. Newton, From Clockwork to Crapshoot: A History of Physics. Cambridge, Mass. and London: The Belknap Press of Harvard University Press, 2007, 340 pages. $29.95 (cloth). Physics in Perspective (PIP), 10(2):252–254, June 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBaeyer:2010:BRA

[vB10a] Hans Christian von Baeyer. Book review: Arthur I. Miller, Deciphering the Cosmic Number: The Strange Friendship of Wolfgang Pauli and Carl Jung. New York: W. W. Norton & Company, 2009, xxv + 336 pages. $27.95 (cloth). Physics in Perspective (PIP), 12(4):497–499, December 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBaeyer:2010:BRL

[vB10b] Hans Christian von Baeyer. Book review: Laura Dassow Walls, The Passage to Cosmos: Alexander von Humboldt and the Shaping of America. Chicago: The University of Chicago Press, 2009, xv + 404 pages. $35.00 (cloth). Physics in Perspective (PIP), 12(3):361–364, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBaeyer:2010:BRM

[vB10c] Hans Christian von Baeyer. Book review: Maria Rentetzi, Trafficking Materials and Gendered Experimental Practices: Radium Research in Early 20th Century Vienna. New York: Columbia University Press, 2008, xxiii + 291 pages. $60.00 (cloth). Physics in Perspective (PIP), 12(1):110–112, March 2010. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://www.gutenberg-e.org/.

vonBaeyer:2011:BRI

[vB11] Hans Christian von Baeyer. Book review: Ian Sample, Massive: The Missing Particle that Sparked the Greatest Hunt in Science. New York: Basic Books, 2010, xi + 260 pages. $25.95 (cloth). Physics in Perspective (PIP), 13(3):380–382, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBaeyer:2012:BRJ

[vB12] Hans Christian von Baeyer. Book review: Jost Lemmerich, Science and Conscience: The Life of James Franck. Translated by Ann M. Hentschel. Stanford: Stanford University Press, 2011, xviii + 369 pages. $65.00 (cloth). Physics in Perspective (PIP), 14(2):246–248, June 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

vonBaeyer:2002:BR

[vBBH02] Hans Christian von Baeyer, Stephen G. Brush, and Art Hobson. Book reviews. Physics in Perspective (PIP), 4(4):491–495, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s000160200006.

vonBaeyer:2010:BR

[vBSTM10] Hans Christian von Baeyer, Gino Segrè, Edwin F. Taylor, and Kimball A. Milton. Book reviews. Physics in Perspective (PIP), 12(4):497–505, December 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-010-0043-4.

vanDongen:2020:E

[vD20] Jeroen van Dongen. In Europe. Physics in Perspective (PIP), 22(1):3–25, March 2020. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00016-020-00252-2.pdf.

Veys:2013:JRM

[Vey13] Lucy Veys. Joseph Rotblat: Moral dilemmas and the Manhattan Project. Physics in Perspective (PIP), 15(4):451–469, December 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-013-0125-1.

Vincze:2010:EGH

[VJ10] Ildikó J. Vincze and István Jankovics. Eugen von Gothard and his X-ray experiments. Physics in Perspective (PIP), 12 (1):25–35, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-009-0003-z.

Vandeville:2019:PBP

[VMR19] Danièle Vandeville, J. P. Martin, and C. Ropartz. α_{1−} Polymorphism of a Bantu population: Description of a mew allele PiL. Physics in Perspective (PIP), 21(1):33–38, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/BF00278562.pdf.

vonReichenbach:2009:RGF

[von09] Maria Cecilia von Reichenbach. Richard Gans: The first quantum physicist in Latin America. Physics in Perspective (PIP), 11(3):302–317, September 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-008-0416-0.

Vieira:2014:CHC

[VV14] Cássio Leite Vieira and Antonio Augusto Passos Videira. Carried by history: Cesar Lattes, nuclear emulsions, and the discovery of the pi-meson. Physics in Perspective (PIP), 16(1):3–36, March 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-014-0128-6.

Walker:2003:OHV

[Wal03a] Mark Walker. Otto Hahn: Verantwortung und Verdrängung. (German) [Otto Hahn: responsibility and repression]. Forschungsprogramm ,,Geschichte der Kaiser-Wilhelm-Gesellschaft im Nationalsozialismus” (Research Program “History of the Kaiser Wilhelm Society in the National Socialist Era”), Berlin, Germany,

2003. 64 pp. LCCN ???? URL www.mpiwg-berlin.mpg.de/
KWG/Ergebnisse/Ergebnisse10.pdf.

Wallace:2003:GC

[Wal03b] W. A. Wallace. Galileo in context. Physics in Perspective (PIP), 5(4):474–475, December 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Walker:2006:OHR

[Wal06] Mark Walker. Otto Hahn: Responsibility and repression. Physics in Perspective (PIP), 8(2):116–163, May 2006. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
006-0277-3.

Wald:2007:BRM

[Wal07] Robert M. Wald. Book review: Michael Lockwood, The Labyrinth of Time: Introducing the Universe. New York: Oxford University Press, 2005, vii + 405 pages. $34.50 (cloth). Physics in Perspective (PIP), 9(3):378–379, September 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Walls:2009:PCA

[Wal09] Laura Dassow Walls. The passage to Cosmos: Alexander von Humboldt and the shaping of America. University of Chicago Press, Chicago, IL, USA and London, UK, 2009. ISBN 0-226-87182-7. xv + 404 pp. LCCN ????

Walls:2011:PCA

[Wal11] Laura Dassow Walls. The passage to Cosmos: Alexander von Humboldt and the shaping of America. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-87183-5 (paperback). ???? pp. LCCN ????

Warwick:2003:MTC

[War03] Andrew Warwick. Masters of theory: Cambridge and the rise of mathematical physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2003. ISBN 0-226-87374-9 (cloth), 0-226-87375-7 (paperback). xiv + 572 pp. LCCN QC19.6 .W37 2003. URL http://www.

loc.gov/catdir/bios/uchi051/2002153732.html;   http:
//www.loc.gov/catdir/description/uchi051/2002153732.
html;   http://www.loc.gov/catdir/toc/fy037/2002153732.
html.

Wattenberg:2001:BRP

[Wat01] Albert Wattenberg. Book review: Per F. Dahl, Heavy Water and the Wartime Race for Nuclear Energy. Bristol and Philadelphia: Institute of Physics, 1999, xvi + 399 pages. $60.00 (Hardbound). Physics in Perspective (PIP), 3(1):131–132, March 2001. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Wattenberg:2003:BRE

[Wat03] Albert Wattenberg. Book review: Edward Teller with Judith Shoolery, Memoirs: A Twentieth-Century Journey in Science and Politics. Cambridge, Massachusetts: Perseus Publishing, 2001, xii + 628 pages. $35.00 (cloth). Physics in Perspective (PIP), 5(3):349–352, September 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Watson:2004:QQ

[Wat04] Andrew Watson. The Quantum Quark. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-82907-0 (hardcover). x + 464 pp. LCCN QC793.3.Q35 W38

2004.  URL http://www.loc.gov/catdir/description/
cam041/2004040660.html;   http://www.loc.gov/catdir/
enhancements/fy0732/2004040660-b.html;   http://www.
loc.gov/catdir/toc/cam041/2004040660.html.

Wazeck:2010:EFC

[Waz10] Milena Wazeck. The 1922 Einstein film: Cinematic innovation and public controversy. Physics in Perspective (PIP), 12(2):163– 179, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s00016-009-0008-7.pdf;   http://
www.springerlink.com/content/f625752811824q46/fulltext.
pdf.

Weekes:2010:NCS

[Wee10] Trevor C. Weekes. The Nineteenth-Century spiral nebula whodunit. Physics in Perspective (PIP), 12(2):146–162, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-010-0018-5.

Weinberg:1983:DSP

[Wei83] Steven Weinberg. The discovery of subatomic particles. Scientific American Library, New York, NY, USA, 1983. ISBN 0-7167-1488-4, 0-7167-1489-2 (paperback). xiii + 206 pp. LCCN QC793.2 .W44 1983.

Weinberg:1990:DSP

[Wei90] Steven Weinberg. The discovery of subatomic particles. W. H. Freeman, New York, NY, USA, 1990. ISBN 0-7167-2121-X. xix + 222 + 2 pp. LCCN QC793.2 .W44 1990. US$10.95.

Weinberg:2001:FSC

[Wei01] Steven Weinberg, editor. Facing up: science and its cultural adversaries. Harvard University Press, Cambridge, MA, USA, 2001. ISBN 0-674-00647-X. LCCN Q171 .W419 2001. URL

http://www.loc.gov/catdir/toc/fy02/2001024219.html.

Weinberg:2003:DSP

[Wei03] Steven Weinberg. The discovery of subatomic particles. Cambridge University Press, Cambridge, UK, revised edition, 2003. ISBN 0-521-82351-X (hardback). xvi + 206 pp. LCCN QC793.2 .W44 2003. URL http://www.

e-streams.com/es0707/es0707_3414.html;   http://www.
loc.gov/catdir/description/cam041/2003283983.html;
http://www.loc.gov/catdir/enhancements/fy0732/2003283983-
b.html;   http://www.loc.gov/catdir/samples/cam041/2003283983.
html;   http://www.loc.gov/catdir/toc/fy043/2003283983.
html.

Weinstock:2004:BRS

[Wei04] Robert Weinstock. Book reviews: Stephen Inwood, The Man Who Knew Too Much: The Strange and Inventive Life of Robert Hooke, 1635–1703. London: Macmillan, 2002, xxix + 497 pages. $37.50 (cloth). Jim Bennett, Michael Cooper, Michael Hunter, Lisa Jardine, London’s Leonardo: The Life and Work of Robert Hooke. Oxford: Oxford University Press, 2003, xii + 224 pages. $35.00 (cloth). Physics in Perspective (PIP), 6(2):246–248, June 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Weiss:2007:PR

[Wei07] Richard J. (Richard Jerome) Weiss. A physicist remembers. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2007. ISBN 981-270-058-7. ix + 261 pp. LCCN QC16.W42 A3 2007.

Weisel:2008:PPB

[Wei08] Gary J. Weisel. Properties and phenomena: Basic plasma physics and fusion research in Postwar America. Physics in Perspective (PIP), 10(4):396–437, December 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
007-0371-1.

Weinberg:2009:LVW

[Wei09] Steven Weinberg. Lake views: this world and the universe. Belknap Press, Cambridge, MA, USA, 2009. ISBN 0-674-03515-1. x + 259 pp. LCCN Q171 .W4194 2009.

Weintraub:2011:HOU

[Wei11] David A. (David Andrew) Weintraub. How Old Is the Universe? Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14731-0 (hardcover). 370 pp. LCCN QB501 .W45 2011.

Weinberg:2015:EWD

[Wei15] Steven Weinberg. To explain the world: the discovery of modern science. Harper, New York, NY, USA, 2015. ISBN 0-06-234665-2 (hardcover). xiv + 416 pp. LCCN Q124.95 .W45 2015.

Weisel:2017:PAP

[Wei17] Gary J. Weisel. The plasma archipelago: Plasma physics in the 1960s. Physics in Perspective (PIP), 19(3):183–226, September 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Westfall:2006:DLT

[Wes06] Catherine Westfall. A different laboratory tale: Fifty years of Mössbauer spectroscopy. Physics in Perspective (PIP), 8 (2):189–213, May 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0274-6.

Westfall:2010:BRS

[Wes10] Catherine Westfall. Book review: Sean Johnson, History of Science: A Beginner’s Guide. Oxford: Oxford University Press, 2009, viii + 218 pages. $14.95 (paper). Physics in Perspective (PIP), 12(3):364–365, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Westfall:2014:EBC

[Wes14] Catherine Westfall. Engineering the Big Chill: The story of JLab’s central helium liquefier. Physics in Perspective (PIP), 16(1):37–68, March 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-014-0127-7.

Westfall:2015:AVV

[Wes15] Catherine Westfall. Adventure in Ven: Visiting Tycho’s island. Physics in Perspective (PIP), 17(3):251–258, September 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-015-0165-9.

Westfall:2016:DDH

[Wes16] Catherine Westfall. From desire to data: How JLab’s experimental program evolved. Physics in Perspective (PIP), 18(3):

301–350, September 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-016-0189-9.

Westfall:2017:BLF

[Wes17] Catherine Westfall. Between the lines: A first-person account of Berkeley’s loss of Fermilab. Physics in Perspective (PIP), 19(2): 91–104, June 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Westfall:2018:DDH

[Wes18] Catherine Westfall. From desire to data: How JLab’s experimental program evolved. Part 2: The painstaking transition to concrete plans, mid-1980s to 1990. Physics in Perspective (PIP), 20(1):43–123, March 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Westfall:2019:DDH

[Wes19] Catherine Westfall. From desire to data: How JLab’s experimental program evolved. Part 3: From experimental plans to concrete reality, JLab gears up for research, mid-1990 through 1997. Physics in Perspective (PIP), 21(2):108–159, June 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Wheeler:1998:GBH

[WF98] John Archibald Wheeler and Kenneth William Ford. Geons, black holes, and quantum foam: a life in physics. W. W. Norton & Co., New York, NY, USA, 1998. ISBN 0-393-04642-7. 380 pp. LCCN QC16.W48 A3 1998.

Wiescher:2017:AHH

[Wie17a] Michael Wiescher. Arthur E. Haas, his life and cosmologies. Physics in Perspective (PIP), 19(1):3–59, April 2017. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
017-0197-4.

Wiescher:2017:FLN

[Wie17b] Michael Wiescher. The four lives of a nuclear accelerator. Physics in Perspective (PIP), 19(2):151–179, June 2017. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Wiescher:2018:HIC

[Wie18] Michael Wiescher. The history and impact of the CNO cycles in nuclear astrophysics. Physics in Perspective (PIP), 20(1):124– 158, March 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Wilson:1999:GRV

[Wil99] D. B. Wilson. Galileo’s religion versus the Church’s science? Rethinking the history of science and religion. Physics in Perspective (PIP), 1(1):65–84, March 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050006.

Wilson:2000:FP

[Wil00] Robert Rathbun Wilson. From frontiersman to physicist. Physics in Perspective (PIP), 2(2):141–203, June 2000. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160050041.

Willen:2017:BMB

[Wil17] Erich Willen. Building magnets at Brookhaven National Laboratory: A condensed account. Physics in Perspective (PIP), 19 (3):227–290, September 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Wittje:2007:NPN

[Wit07] Roland Wittje. Nuclear physics in Norway, 1933–1955. Physics in Perspective (PIP), 9(4):406–433, December 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
006-0317-z.

Wautier:2012:LWJ

[WJS12] Kristel Wautier, Alexander Jonckheere, and Danny Segers. The life and work of Joseph Plateau: Father of film and discoverer of surface tension. Physics in Perspective (PIP), 14(3): 258–278, September 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-012-0087-8.

Wattenberg:2003:BR

[WMV03] Albert Wattenberg, Alberto A. Martínez, and Robert Vandenbosch. Book reviews. Physics in Perspective (PIP), 5(3):349– 355, September 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-003-0183-X.

Wolfson:2003:SER

[Wol03] Richard Wolfson. Simply Einstein: Relativity demystified. W. W. Norton & Co., New York, NY, USA, 2003. ISBN 0-393-05154-4 (hardcover). x + 261 pp. LCCN QC173.57

.W65 2003. URL http://www.loc.gov/catdir/toc/fy035/
2002002984.html.

Wald:2007:BR

[WRJH07] Robert M. Wald, John S. Rigden, Stanley T. Jones, and Jeff Hecht. Book reviews. Physics in Perspective (PIP), 9(3):378– 385, September 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0311-5.

Wynn:2001:QLW

[WWH01] Charles M. Wynn, Arthur W. Wiggins, and Sidney Harris. Quantum leaps in the wrong direction: where real science ends — and pseudoscience begins. Joseph Henry Press, Washington, DC, USA, 2001. ISBN 0-309-07309-X (paperback). x + 226 pp. LCCN Q172.5.P77 W96 2001. With cartoons by Sidney Harris.

Yeh:2005:BRJ

[Yeh05] Nai-Chang Yeh. Book review: Jean Matricon and Georges Waysand, The Cold Wars: A History of Superconductivity, translated from the French by Charles Glashausser. New Brunswick, NJ: Rutgers University Press, 2003, xiii + 272 pages, $26.00 (paper). Physics in Perspective (PIP), 7(2):259–261, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Ziman:2002:RSW

[Zim02] John Ziman. Real science: what it is, and what it means. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-77229-X (hardcover), 0-521-89310-0 (paperback), 0-511-03450-4 (e-book). xii + 399 pp. LCCN Q175 .Z547 2002.

Zhang:2018:FFE

[ZZ18] Andrew Zhang and Andrew Zangwill. Four facts everyone ought to know about science: The two-culture concerns of Philip W. Anderson. Physics in Perspective (PIP), 20(4):342–369, December 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).